### v1

# Analysis of a class of Recursive Distributional Equations Including the Resistance of the Series-parallel graph

Abstract.
This paper analyzes a class of recursive distributional equations (RDE’s) proposed by Gurel-Gurevich [17] and involving a bias parameter $p$, which includes the logarithm of the resistance of the series-parallel graph. A discrete-time evolution equation resembling a nonlinear, fractional Fisher-KPP equation is derived to describe the CDF’s of solutions. When the bias parameter $p=\frac{1}{2}$, this equation is shown to have a PDE scaling limit, from which distributional limit theorems for the RDE are derived. Applied to the series-parallel graph, the results imply that $N^{-1/3}\log R^{(N)}$ has a nondegenerate limit when $p=\frac{1}{2}$, as conjectured by Addario-Berry, Cairns, Devroye, Kerriou, and Mitchell [1].
1. Introduction
1.1. Resistance of the Series-Parallel Graph
This paper seeks to address an old problem, namely, the asymptotic analysis of the resistance of the series-parallel graph. This is the random hierarchical lattice introduced by Hambly and Jordan [18] and constructed as follows: Start with a graph $G^{(0)}$ consisting of one edge connecting two nodes, $A$ and $B$. At step $n$ of the construction, obtain $G^{(n)}$ from $G^{(n-1)}$ by replacing every edge by two edges, either in series or in parallel, the result being decided by independent coin flips. Interpreting $G^{(n)}$ as a resistor network where each edge has resistance one, one can study the effective resistance $R^{(n)}$ between $A$ and $B$.
The asymptotic analysis of $R^{(n)}$ as $n\to\infty$ was initiated in [18]. Since $G^{(n)}$ can be realized by combining two independent copies $G^{(n-1)}_{1}$ and $G^{(n-1)}_{2}$ of $G^{(n-1)}$, say, in series with probability $p$ and in parallel with probability 1p1-p, the law of the resistance $R^{(n)}$ is a solution of the recursive distributional equation (RDE)
$R^{(n)}\overset{d}{=}\left\{\begin{array}[]{r l}R^{(n-1)}_{1}+R^{(n-1)}_{2},&\text{with probability}\,\,p,\\ \left(\frac{1}{R_{1}^{(n-1)}}+\frac{1}{R_{2}^{(n-1)}}\right)^{-1},&\text{with probability}\,\,1-p.\end{array}\right$.
Using this fact, a number of results were proved in [18]: When $p>\frac{1}{2}$, the logarithm $\log R^{(n)}$ grows linearly in $n$. More precisely, there is a constant $\lambda(p)>0$ such that
(1) $\mathbb{P}\left\{(2p-1)\log 2\leq\frac{1}{N}\log R^{(N)}\leq\lambda(p)\right\}\to 1\quad\text{as}\,\,N\to\infty.$
Due to symmetry, this implies that $\log R^{(n)}\to$-$\infty$ linearly when $p<\frac{1}{2}$.
At the critical point $p=\frac{1}{2}$, Hambly and Jordan instead proved that $\log R^{(n)}$ exhibits sublinear growth. The main results of this paper (see Theorem 1 and Corollary 1 below) show that, in fact, $\log R^{(N)}$ grows like $N^{\frac{1}{3}}$. Specifially,
$\frac{1}{(72\zeta(3))^{\frac{1}{3}}N^{\frac{1}{3}}}\log R^{(N)}+\frac{1}{2}\overset{d}{\to}\text{Beta}(2,2)\quad\text{as}\,\,N\to\infty,$
where $\zeta$ is the Riemann zeta function. This confirms a conjecture of Addario-Berry, Cairns, Devroye, Kerriou, and Mitchell [1].
The starting point for this work is an observation made by Gurel-Gurevich during a presentation at the open problem session at the Workshop for Disordered Media at the Rényi Institute in Budapest in January 2025 [17]. He noted that the transformation $X^{(n)}=\log R^{(n)}$ leads to an RDE of the form
(2) $X^{(n)}=\Phi(X^{(n-1)}_{1},X^{(n-1)}_{2})$,
where $\Phi$ is the random function defined by
(5) $\Phi(x,y)=\left\{\begin{array}[]{r l}\max\{x,y\}+f(|y-x|),&\text{with probability}\,\,p,\\ \min\{x,y\}-f(|y-x|),&\text{with probability}\,\,1-p,\end{array}\right$.
and $f(u)=\log(1+e^{-u})$. This led him to pose the following problem:
Problem.
Consider the RDE (2) with forcing $\Phi$ given by (5) under general assumptions on the nonincreasing function $f:[0,+\infty)\to[0,+\infty)$. What can be said about the long-time behavior of the law of solutions?
This paper endeavors to address this problem in some generality when $p=\frac{1}{2}$, focusing on the distributional limit of $n^{-1/3}X^{(n)}$. The general framework developed here applies not only to the resistance of the critical series-parallel graph, but also to some other RDE’s considered recently in the literature. For instance, as is shown in Section 4 below, the class of RDE’s determined by (2) and (5) includes the symmetric hipster random walk from [1] and two-player symmetric cooperative motion, one of the RDE’s analyzed very recently by Addario-Berry, Beckman, and Lin in [3].
While the analysis of the asymptotic behavior in the case $p\neq\frac{1}{2}$ is not touched on here, the framework developed in this paper suggests that it is related to the singular limit of certain fractional, quasilinear variants of the Fisher-KPP equation. This is made precise in Section 1.3.4 below.
1.2. General Setting
Consider the following general setting, which generalizes the forcing $\Phi$ in (2). Given a Bernoulli random variable $\Theta$ and random nonincreasing functions $f_{+},f_{-}:\mathbb{R}\to[0,+\infty)$, define $\Phi$ by
(8) $\Phi(x,y)=\left\{\begin{array}[]{r l}\max\{x,y\}+f_{+}(|x-y|),\quad\text{if}\,\,\Theta=1,\\ \min\{x,y\}-f_{-}(|x-y|),\quad\text{if}\,\,\Theta=0.\end{array}\right$.
For simplicity, the pair $(f_{+},f_{-})$ will be assumed to be independent of $\Theta$ henceforth.
The object of interest in this paper is the RDE determining the law of random variables $(X^{(n)})_{n\in\mathbb{N}}$ through the recursion
(9) $X^{(n)}\overset{d}{=}\Phi(X^{(n)}_{1},X^{(n)}_{2})$,
where $X^{(n)}_{1},X^{(n)}_{2}$ are i.i.d. copies of $X^{(n-1)}$, which are independent of $\Phi$. Throughout the paper, equality is only required to hold in law; questions of couplings or constructions of $\{X^{(n)}\}$ will not be treated.
The form (8) of $\Phi$ is not just a repackaging of (5). In the asymmetric setting where $f_{+}\neq f_{-}$, this class of RDE’s includes the one associated with Pemantle’s min-plus binary tree, the asymptotics of which were analyzed by Auffinger and Cable [5], as well as certain versions of the asymmetric hipster random walk from [1] and two-player totally asymmetric, $q$-lazy cooperative motion, one of the RDE’s analyzed by Addario-Berry, Beckman, and Lin in [2]. In all of these examples, as well as the symmetric models from [1, 3] already mentioned above, the asymptotic analysis carried out here generalizes the limit theorems previously obtained within a unified setting.
Henceforth define the bias parameter $p$ by
$p=\mathbb{P}\{\Theta=1\}.$
The joint law of the functions $(f_{+},f_{-})$ will be denoted by $\mathbf{P}$ in what follows. $\mathbf{P}$ is assumed to be a Borel probability measure on the space $BC([0,\infty))\times BC([0,\infty))$ of pairs of bounded continuous functions in $0,\infty)$ with the uniform norm topology. In most of the examples of interest here, however, $f_{+}$ and $f_{-}$ are deterministic; see the discussion in Section 4.
1.3. Evolution of the CDF
The main results of this paper are obtained by treating the evolution of the law of $X^{(n)}$ as if it were governed by a parabolic partial differential equation (PDE). Following [2, 3], this is done at the level of cumulative distribution functions (CDF’s) rather than probability measures.
In particular, in what follows, define the operator $T$ by taking a CDF $F$, letting $X_{1},X_{2}$ be i.i.d. random variables distributed according to $F$, and defining $TF$ to be the CDF of the random variable $Y$ determined by
(10) $Y=\Phi(X_{1},X_{2}),$
where $\Phi$ is sampled independently of $(X_{1},X_{2})$. With this notation, it follows that the random variables $\{X^{(n)}\}$ solve the RDE (9) if and only if the corresponding CDF’s $\{F_{n}\}$ solve the recursive equation
$F_{n+1}=TF_{n}.$
In the first step of this work, a discrete-time evolution equation is derived to describe the evolution of the CDF of $X^{(n)}$. Namely, there is a nonlinear operator $\mathscr{L}$ defined on the space of CDF’s such that the recursive equation $F_{n+1}=TF_{n}$ can be rewritten in the following suggestive form
(11) $F_{n+1}-F_{n}=\mathscr{L}F_{n}+(1-2p)F_{n}(1-F_{n}).$
As will be argued below, $\mathscr{L}$ has the character of an advection-diffusion operator, albeit a nonlinear one, while the “reaction term” F(1F)F(1-F) is familiar from the Fisher-KPP equation. The equation thus resembles a reaction-diffusion equation.
As far as convergence in distribution is concerned, deriving a limit for $n^{-1/\alpha}X^{(n)}$ for some exponent $\alpha$ is equivalent to proving a scaling limit for the equation (11). In the main result of this paper, which treats the critical point $p=\frac{1}{2}$, this is done by adapting the classical approach to numerical approximations of parabolic PDE’s developed by Barles and Souganidis [6].
To have a sense of why it is reasonable to expect a PDE scaling limit, consider what happens in (8) when $f_{+}=f_{-}=f_{\mathbb{Z}}$, where $f_{\mathbb{Z}}(u)=(1-u)_{+}$. In this case, if the initial datum $X^{(0)}$ is integer-valued, then that remains true of $X^{(n)}$ for all $n$, and thus the CDF $F_{n}$ can be regarded as a function on $\mathbb{Z}$. In fact, these $\mathbb{Z}$-valued solutions of the RDE coincide with two-player symmetric cooperative motion, one of the RDE’s introduced in [3]. As shown in that work, the evolution equation governing the CDF now takes a very concrete form, namely,
$F_{n+1}-F_{n}=|\nabla_{\mathbb{Z}}F_{n}|\Delta_{\mathbb{Z}}F_{n}\quad\text{in}\,\,\mathbb{Z}$,
where $\nabla_{\mathbb{Z}}$ and $\Delta_{\mathbb{Z}}$ are certain discrete derivative operators. This observation was then used in conjunction with the method of [6] and a regularization effect of the equation to show that, after a suitable rescaling, the CDF converges to the solution of the initial-value problem
$\partial_{t}F=|\partial_{x}F|\partial_{x}^{2}F,\quad F(x,0)=\mathbf{1}_{[0,\infty)}(x).$
After differentiation, this PDE becomes the porous medium equation satisfied by the PDF $\rho=\partial_{x}F$ with Dirac initial data, the solution of which is known to be a self-similarly growing parabolic cap (the so-called Barenblatt solution). In probabilistic terms, this implies the convergence of $N^{-1/3}X^{(N)}$ to a shifted and dilated Beta(2,2) random variable.
In this work, this general strategy — derivation of a discrete-time evolution equation and application of a convergence result in the spirit of [6] — is shown to still apply in the setting of the RDE’s defined above, although its execution is complicated somewhat by the fact that the operator $\mathscr{L}$ is generally nonlocal.
1.3.1. Assumptions
Some assumptions on the law $\mathbf{P}$ of the forcing $(f_{+},f_{-})$ are in order. First, define the class of functions $\mathcal{S}$ so that $f\in\mathcal{S}$ if and only if
(12) $f:\mathbb{R}\to\mathbb{R}\quad\text{is continuous and nonincreasing, and}$
(13) $u\mapsto u+f(u)\,\,\text{is a nondecreasing function in}\,\,[0,+\infty).$
Note that the function $f(u)=\log(1+e^{-u})$ relevant to the series-parallel graph belongs to $\mathcal{S}$ since $1\leq f^{\prime}\leq 0$.
Throughout the paper, $f_{+}$ and $f_{-}$ are assumed to lie in $\mathcal{S}$. For $\rho\in\{+,-\}$, let $g_{\rho}:[0,+\infty)\to[0,+\infty)$ be the function
(14) $g_{\rho}(s)=\sup\left\{u\geq 0\,\mid\,u+f_{\rho}(u)\leq s\right\}\quad\text{if}\,\,s\geq f_{\rho}(0),\quad g_{\rho}(s)=0,\quad\text{otherwise.}$
Notice that the restriction of $g_{\rho}$ to the set $f(0),+\infty)$ is the right-continuous right-inverse of $u\mapsto u+f_{\rho}(u)$.
By definition, if $f_{\rho}(+\infty)=0$, then $g_{\rho}(s)-s\to 0$ as $s\to+\infty$. The next assumption asks that $g_{\rho}(s)-s$ vanishes in a quantitative manner:
(15) $\int_{0}^{\infty}(1+s)\mathbf{E}[|g_{+}(s)-s|+|g_{-}(s)-s|]\,ds<\infty.$
1.3.2. Main Result
Under the assumption (15), it is possible to define effective coefficients $\sigma\in\mathbb{R}$ and $a\geq 0$ according to the formulae
$\sigma$ $=-2p\int_{0}^{\infty}\mathbf{E}[|g_{+}(s)-s|]\,du+2(1-p)\int_{0}^{\infty}\mathbf{E}[|g_{-}(s)-s|]\,ds,\quad$
$a$ $=p\int_{0}^{\infty}\mathbf{E}[(3s-g_{+}(s))(s-g_{+}(s))]\,ds+(1-p)\int_{0}^{+\infty}\mathbf{E}[(3s-g_{-}(s))(s-g_{-}(s))]\,ds.$
The nonnegativity of $a$ follows from the fact that $g_{+}(s),g_{-}(s)\leq s$ for any $s\geq 0$.
Define the exponent $\alpha\in\mathbb{N}$ as follows:
$\alpha=2\quad\text{if}\,\,\sigma\neq 0,\quad\alpha=3,\quad\text{otherwise.}$
The next theorem characterizes the scaling limit of the CDF in the critical case $p=\frac{1}{2}$. The relevant background from PDE theory, particularly the theory of viscosity solutions, is reviewed in Appendix A below.
Theorem 1.
Assume that $p=\frac{1}{2}$, $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, and (15) holds. If $\{F_{n}\}$ is the sequence of CDF’s associated to some solution $\{X^{(n)}\}$ of (9), then, independent of the initial CDF $F_{0}$, if $F_{\epsilon}$ is the rescaled CDF defined by
$F_{\epsilon}(x,t)=F_{[\epsilon^{-\alpha}t]}(\epsilon^{-1}x),$
then $F_{\epsilon}\to F$ locally uniformly in $\mathbb{R}\times(0,\infty)$, where $F$ is determined as follows:
  • (i)
    If $\sigma\neq 0$, then $F$ is the unique bounded discontinuous viscosity solution of the initial value problem
    (18) $\left\{\begin{array}[]{r l}\partial_{t}F-\sigma|\partial_{x}F|^{2}=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=\mathbf{1}_{[0,\infty)}(x).\end{array}\right$.
  • (ii)
    If $\sigma=0$, then $F$ is the unique bounded discontinuous viscosity solution of the initial-value problem
    (21) $\left\{\begin{array}[]{r l}\partial_{t}F-a|\partial_{x}F|\partial_{x}^{2}F=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=\mathbf{1}_{[0,\infty)}(x).\end{array}\right$.
These PDE’s become more familiar when passing to PDF’s: The spatial derivative $\rho=\partial_{x}F$ is a solution of Burger’s equation $\partial_{t}\rho=\sigma\partial_{x}(\rho^{2})$ in case (i) and a solution of the porous medium equation $\partial_{t}\rho=\frac{1}{2}a\partial_{x}^{2}(\rho^{2})$ in case (ii).
As observed in [3, 2], the solutions of the PDE’s above are explicitly known and closely related to the $\text{Beta}(2,1)$ and $\text{Beta}(2,2)$ distributions. The next corollary makes this precise.
Corollary 1.
Assume that $p=\frac{1}{2}$, $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, and (15) holds, and let $\{X^{(n)}\}$ be a solution of the RDE (9) with an arbitrary initial distribution.
  • (i)
    If $\sigma\neq 0$, then
    $-\frac{\text{sgn}(\sigma)}{2\sqrt{|\sigma|}\sqrt{N}}X^{(N)}\overset{d}{\to}\text{Beta}(2,1).$
  • (ii)
    If $\sigma=0$ and $a>0$, then
    $\frac{1}{(36a)^{\frac{1}{3}}N^{\frac{1}{3}}}X^{(N)}+\frac{1}{2}\overset{d}{\to}\text{Beta}(2,2)$
  • (iii)
    If $\sigma=a=0$, then $\mathbf{P}\{f_{+}=f_{-}=0\}=1$ and $X^{(N)}\overset{d}{=}X^{(0)}$ for each $N$.
It is not hard to show, through manipulations of the definitions above, that $\sigma>0$ holds, for instance, if $f_{-}>f_{+}$ $\mathbf{P}$-almost surely, while $\sigma=0$ if $f_{+}=f_{-}$. In the case of the resistance $R^{(n)}$ of the critical series-parallel graph, $\sigma=0$ and $a=2\zeta(3)$. For more on this and further discussion of examples, see Section 4.
In addition to characterizing the asymptotics of the resistance of the critical series-parallel graph, the proof of Theorem 1 also leads to a characterization of the limiting behavior of Pemantle’s min-plus binary tree (or first passage percolation on the critical series-parallel graph). This is discussed in Section 4.2 below.
1.3.3. Idea of Proof
The PDE scaling limit of the CDF is obtained using an extension of the approach of [6], which is broadly applicable to elliptic and parabolic PDE’s studied through the lens of Crandall and Lions’ theory of viscosity solutions. There are two key ingredients in that approach, namely, monotonicity and consistency.
In the present context, monotonicity means the following: If $\{F_{n}\}$ and $\{G_{n}\}$ are two sequences of CDF’s arising from solutions of the RDE and if the inequality $F_{0}\leq G_{0}$ holds pointwise, then $F_{n}\leq G_{n}$ holds for any $n$. For the purposes of this paper, this is the bare minimum structural condition necessary to treat the evolution (11) as a parabolic equation.
Consistency, on the other hand, refers to the behavior of $T$ under rescaling. Specifically, Proposition 8 below shows that if $\Phi$ is a smooth CDF, rescaled according to $\Phi_{\delta}(x)=\Phi(\delta x)$, then
(22) $(\mathscr{L}\Phi_{\delta})(\delta^{-1}x)=\delta^{2}\sigma|\partial_{x}\Phi(x)|^{2}+\delta^{3}a|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)+\cdots.$
Naïvely inserting this expansion in the equation (11) and recalling the definition of the rescaled CDF $F_{\epsilon}$, one is led to guess that the PDE’s (18) and (21) ought to appear in the limit. The approach of [6] provides a framework for making such a formal argument rigorous.
1.3.4. Variants of the Fisher-KPP Equation when $p\neq\frac{1}{2}$
The expansion (22) leads to interesting new questions away from the critical point $p=\frac{1}{2}$. Passing to the ballistically rescaled CDF
$F_{\epsilon}(x,t)=F_{[\epsilon^{-1}t]}(\epsilon^{-1}x),$
one finds that formally
$\partial_{t}F_{\epsilon}$ $=\epsilon^{2}a|\partial_{x}F_{\epsilon}|\partial_{x}^{2}F_{\epsilon}+\epsilon\sigma|\partial_{x}F_{\epsilon}|^{2}+\frac{1}{\epsilon}(1-2p)F_{\epsilon}(1-F_{\epsilon})+\cdots.$
After discarding the error terms, this is the PDE obtained after applying a ballistic rescaling to the following quasilinear, fractional variant of the Fisher-KPP equation
(23) $\partial_{t}G=a|\partial_{x}G|\partial_{x}^{2}G+\sigma|\partial_{x}G|^{2}+(1-2p)G(1-G).$
The asymptotics of this equation, particularly the propagation of “front-like” solutions (such as CDF’s), does not seem to have been treated thus far in the literature. Both the cases where $a=0$ or $\sigma=0$ appear to be relevant to the problem at hand; see Section 4.3 for justification of this.
1.4. Related Literature
There is considerable interest in RDE’s in the probability literature. The reader is referred to the survey article of Aldous and Bandyopadhyay [4], which discusses applications in the study of branching processes, percolation, and mean-field combinatorial optimization problems.
There has been a recent burst of activity concerning RDE’s in which the relevant asymptotic behavior is not convergence to a fixed point as in the examples in [4], but instead a distributional scaling limit in the same spirit as the central limit theorem.
Auffinger and Cable [5] analyzed Pemantle’s Min-Plus Binary Tree (or, equivalently, first passage percolation on the critical series-parallel graph, see [18, Section 3]), proving that solutions of the associated RDE converge to a $\text{Beta}(2,1)$ distribution after a diffusive rescaling provided the initial distribution is concentrated at one. This is improved to the case of an arbitrary initial distribution in Corollary 2 below.
Addario-Berry and coauthors recently introduced two classes of RDE’s, called hipster random walks and cooperative motion, in [1, 2, 3]. In [1], $\text{Beta}(2,1)$ and $\text{Beta}(2,2)$ limit theorems for some hipster random walks were proved by treating the evolution of the PMF as a numerical approximation of the solution of a divergence-form PDE. As discussed therein, some of the motivation to study these RDE’s came from interest in the resistance of the series-parallel graph and Pemantle’s Min-Plus Binary Tree.
Cooperative motion was analyzed in [2, 3], this time through the analysis of the CDF and using techniques developed for PDE’s in nondivergence form, specifically the theory of viscosity solutions. In the present paper, the idea to focus attention on the CDF came from these last two works.
The approach used here is similar in spirit to that in [2, 3]. A few differences are worth highlighting. Many of the RDE’s treated in [2, 3] are not monotone, hence the method of [6] does not apply out of the box. Instead, those works showed that, after a finite time, the law of $X^{(n)}$ is pulled into a region in which the RDE is monotone; at the level of the CDF, the key observation is that the CDF obtains a uniform Lipschitz bound in a (universal) finite time, which mimics the properties of the PDE’s that emerge in the scaling limit. After this finite waiting period, monotonicity can be used as in [6] (or, in the case of [2], following the earlier work [12]).
By contrast, here the RDE’s of interest are all monotone. There is a technical issue, namely, that the result of [6] does not apply directly as the evolution (11) only makes sense as an equation posed in the space of CDF’s, whereas [6] works with evolution equations posed in the space of bounded functions. In [3], this is arguably less of an issue as the equations all involve finite differences (although the relevant details are not discussed there). Here $\mathscr{L}$ behaves like a nonlocal operator with a kernel with infinite range so at first blush one would like to know that [6] can be carried out with globally nondecreasing test functions, which is not obvious. It is, in fact, possible to adapt the method to monotone semigroups in the space of CDF’s. In the hope that it may be useful in other contexts, the result is formulated in an abstract way in Appendix B below.
When $p\neq\frac{1}{2}$, the evolution equation obtained here suggests a strategy for tackling the asymptotics of the RDE. This is of interest both in the analysis of the resistance and the distance on the series-parallel graph. The latter was recently considered by Chen, Derrida, Duquesne, and Shi in [11]. They proved that when $p>\frac{1}{2}$, the logarithm of the expected value of the distance grows at a linear rate with slope $\alpha(p)$, and they proved that $\alpha(1/2+\epsilon)\approx\sqrt{\zeta(2)\epsilon}$ as $\epsilon\to 0$.
As mentioned above, convergence of the rescaled CDF’s to the solution of a PDE is proved here by adapting the classical approach of Barles and Souganidis from [6]. At a high level, as soon as the monotonicity and scaling behavior (“consistency”) of $\mathscr{L}$ is understood, convergence follows. This is not the only setting where some extra tailoring of the method is needed to furnish a proof. For instance, in [9], the same authors adapted their approach to the setting of geometric flows and interface motions.
1.5. Organization of the Paper
Section 2 discusses some preliminaries used throughout the paper. The key monotonicity result is proved there. The equation describing the evolution of CDF’s is derived in Section 3. Section 4 discusses examples that fit into the class described by (9) and (8). The main results are proved in Section 5 conditional on an extension of [6], which is presented in the appendix.
There are three appendices. Necessary terminology and results from the theory of viscosity solutions are presented in Appendix A. A variant of the convergence result of [6], tailored to monotone semigroups in the space of CDF’s, is stated and proved in Appendix B. Finally, Appendix C includes computations of the coefficients $\sigma$ and $a$ arising respectively in Pemantle’s min-plus binary tree and the resistance of the critical series-parallel graph
Acknowledgements
This work was made possible in part by the hospitality and support of the Rényi Institute in Budapest, Hungary. The initial inspiration came from the talk [17] delivered during the Workshop on Disordered Media hosted by the Erdős Center during the Simons Semester on Probability and Statistical Physics in the spring of 2025.
2. Preliminaries
2.1. Extended-Real-Valued Random Variables
Note that the definition (9) still makes sense if the input random variables are permitted to take the values $\infty$ and -$\infty$. It will be convenient to extend the definition to include such extended-real-valued random variables. To that end, for an extended-real-valued random variable $X$, define the CDF $F:\mathbb{R}\to[0,1]$ via the rule
$\mathbb{P}\{X=-\infty\}=\lim_{y\to-\infty}F(y),\quad\mathbb{P}\{X\leq x\}=F(x),\quad\mathbb{P}\{X=+\infty\}=1-\lim_{y\to+\infty}F(y).$
Notice that the space $CDF(\overline{\mathbb{R}})$ of all such CDF’s is simply
$CDF(\overline{\mathbb{R}})=\{F:\mathbb{R}\to[0,1]\,\,\text{increasing, right-continuous}\}.$
Using standard extended-real-number arithmetic, from now on, the formula (10) defining $T$ will be understood to hold in the domain $CDF(\overline{\mathbb{R}})$.
2.2. Symmetry
In order to reduce redundancy in the proofs that follow, it is worth observing that if $Y=\Phi(X_{1},X_{2})$, where $\Phi$ is given by (8), then
$-Y=\Psi(-X_{1},-X_{2}),$
provided $\Psi$ is defined by
$\Psi(x,y)=\left\{\begin{array}[]{r l}\max\{x,y\}+f_{-}(|x-y|),&\text{if}\,\,1-\Theta=1,\\ \min\{x,y\}-f_{+}(|x-y|),&\text{if}\,\,1-\Theta=0.\end{array}\right$.
Thus, in proofs where the events $\{\Theta=1\}$ and $\{\Theta=0\}$ would in principle need to be treated independently, by symmetry, there is no loss of generality in assuming $\Theta=1$.
The symmetry property above has a counterpart at the level of the CDF. If $\{F_{n}\}$ is the sequence of CDF’s associated with a solution $\{X^{(n)}\}$ of the RDE $X^{(n)}=\Phi(X^{(n-1)}_{1},X^{(n-1)}_{2})$, then the functions $\{G_{n}\}$ defined by
$G_{n}(x)=1-\lim_{\delta\downarrow 0}F_{n}(-x-\delta)$
are the CDF’s associated with the solution $\{Y^{(n)}\}$ of the RDE $Y^{(n)}=\Psi(Y^{(n-1)}_{1},Y^{(n-1)}_{2})$ obtained by setting $Y^{(n)}=-X^{(n)}$.
2.3. Monotonicity
In this section, the monotonicity of $T$ is proved. Here monotone means that
$\text{if}\,\,F_{1}\leq F_{2}\,\,\text{pointwise in}\,\,\mathbb{R},\quad\text{then}\,\,TF_{1}\leq TF_{2}\,\,\text{pointwise in}\,\,\mathbb{R}.$
To this end, it is convenient to exploit the quadratic structure inherent in the definition. Specifically, it will be useful to consider a certain natural operator
$S:CDF(\overline{\mathbb{R}})\times CDF(\overline{\mathbb{R}})\to CDF(\overline{\mathbb{R}})$
with the property that $S(F,F)=TF$ for any CDF $F$.
$S$ is defined analogously to $T$: Given two CDF’s $F,G\in CDF(\overline{\mathbb{R}})$, let $X_{F}$ and $X_{G}$ be independent extended-real-valued random variables distributed according to $F$ and $G$, respectively, and sample the data $(\Theta,f_{+},f_{-})$ independently of $(X_{F},X_{G})$. Let $S(F,G)$ denote the CDF of the random variable $Y$ defined as follows:
(26) $Y=\Phi(X_{F},X_{G})=\left\{\begin{array}[]{r l}\max\{X_{F},X_{G}\}+f_{+}(|X_{F}-X_{G}|),&\text{if}\,\,\Theta=1,\\ \min\{X_{F},X_{G}\}-f_{-}(|X_{F}-X_{G}|),&\text{if}\,\,\Theta=0.\end{array}\right$.
Clearly $S(F,F)=TF$ as desired. Further, $S(F,G)=S(G,F)$ by symmetry.
The next result asserts that $S$ is monotone provided $f_{+},f_{-}\in\mathcal{S}$.
Proposition 1.
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$. Given $F_{+},F_{-},G\in CDF(\overline{\mathbb{R}})$, if $F_{+}\leq F_{-}$ holds pointwise in $\mathbb{R}$, then $S(F_{+},G)\leq S(F_{-},G)$ also holds. Therefore, $T$ is monotone.
Proof.
Since $S(F,F)=TF$ and $S$ is symmetric, the desired monotonicity of $T$ readily follows from that of $S$.
It only remains to prove the monotonicity of $S$. Fix $F_{+},F_{-},G\in CDF(\overline{\mathbb{R}})$ such that $F_{+}\leq F_{-}$ pointwise in $\mathbb{R}$. Recall that, by Strassen’s Theorem, this is equivalent to the existence of a Borel probability measure $\Pi$ on $\overline{\mathbb{R}}\times\overline{\mathbb{R}}$ such that $\Pi\{(x,y)\,\mid\,x\leq y\}=1$ and the first and second marginals of $\Pi$ are distributed according to $F_{-}$ and $F_{+}$, respectively.
Let $(X_{1}^{-},X_{1}^{+},X_{2})$ be a random vector such that $(X_{1}^{-},X_{1}^{+})$ is distributed according to $\Pi$, independent of $X_{2}$, and $X_{2}$ is distributed according to $G$. Note that $X_{1}^{-}\leq X_{2}^{+}$ almost surely. Define $Y^{+}=\Phi(X_{1}^{+},X_{2})$ and $Y^{-}=\Phi(X_{1}^{-},X_{2})$, where $\Phi$ is independent of $(X_{1}^{-},X_{1}^{+},X_{2})$. By definition, $S(F_{\rho},G)$ is the CDF of $Y^{\rho}$ for $\rho\in\{+,-\}$. Thus, to conclude, it only remains to show that $Y^{-}\leq Y^{+}$ holds almost surely.
This can be checked via a case analysis. The details for the event $\{\Theta=1\}$ are provided below; the case when $\{\Theta=0\}$ follows by symmetry as in Section 2.2. Assume henceforth that $\Theta=1$.
If $X_{2}\leq X_{1}^{-}\leq X_{1}^{+}$, then the assumption that the map $u\mapsto u+f(u)$ is nondecreasing (see (13)) implies
$Y^{-}=X^{-}_{1}+f_{+}(X^{-}_{1}-X_{2})\leq X_{1}^{+}+f_{+}(X^{+}_{1}-X_{2})=Y^{+}.$
On the other hand, if $X^{-}_{1}\leq X_{2}\leq X^{+}_{1}$, then invoking first (13) and then the fact that $f_{+}$ is nonincreasing (see (12)), one finds
$Y^{-}=X_{2}+f_{+}(X_{2}-X^{-}_{1})$ $\leq X_{2}+f_{+}(0)$
$=X_{2}+f_{+}(X_{2}-X_{2})\leq X^{+}_{1}+f_{+}(X^{+}_{1}-X_{2})=Y^{+}.$
Finally, if $X^{-}_{1}\leq X^{+}_{1}\leq X_{2}$, then, again, the fact that $f_{+}$ is nonincreasing implies
$Y^{-}=X_{2}+f_{+}(X_{2}-X^{-}_{1})\leq X_{2}+f_{+}(X_{2}-X^{+}_{1})=Y^{+}.$
2.4. Continuity
For the sake of developing intuition, it may be useful to note that $T$ maps the space of continuous CDF’s to itself.
Proposition 2.
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$. If $F\in CDF(\overline{\mathbb{R}})$ is continuous, then $TF$ is also continuous.
Proof.
Let $(X_{1},X_{2})$ be i.i.d. random variables distributed according to $F$ and let $Y=\Phi(X_{1},X_{2})$, where $\Phi$ is sampled independently of $(X_{1},X_{2})$. Since $TF$ is the CDF of $Y$, to see that it is continuous, it suffices to prove that $\mathbb{P}\{Y=x_{0}\}=0$ for each $x_{0}\in\mathbb{R}$.
Consider the event $\Theta=1$: Since the law of $(X_{1},X_{2})$ is invariant under permutation of the coordinates,
$\mathbb{P}\{Y=x_{0}\,\mid\,\Theta=1\}$ $=2\int_{-\infty}^{\infty}\mathbb{P}\{X_{2}-x_{1}+f(X_{2}-x_{1})=x_{0}-x_{1},\,\,X_{2}>x_{1}\}\mu(dx_{1}),$
where $\mu$ is the law of $X_{1}$. Since $u\mapsto u+f(u)$ is nondecreasing by (14), there are two possibilities: Either the set $\{u\geq 0\,\mid\,u+f(u)=x_{0}-x_{1}\}$ is an interval, or it is a point. If it is a point, then the corresponding probability in the integral above vanishes since $X_{2}$ is a continuous random variable. On the other hand, there are at most countably many disjoint intervals on which $u+f(u)$ is constant, hence countably many values of $x_{0}-x_{1}$. Since $\mu$ has no atoms, these values of $x_{0}-x_{1}$ do not contribute to the integral either. Therefore, the probability above is zero.
A similar analysis shows that $\mathbb{P}\{Y=x_{0}\,\mid\,\Theta=0\}=0$. ∎
3. Evolution of the CDF
This section concerns the evolution equation (11) satisfied by CDF’s. To this end, it will be useful to decompose the random variable $Y$ in the definition of $T$ (see (10)) in a specific way. Given $F\in CDF(\overline{\mathbb{R}})$, let $X_{1},X_{2}$ be i.i.d. extended-real-valued random variables distributed according to $F$, and sample the data $(\Theta,f_{+},f_{-})$ independently of $(X_{1},X_{2})$. Define the extended-real-valued random variable $Z$ by
(29) $Z$ $=\left\{\begin{array}[]{r l}\max\{X_{1},X_{2}\},&\text{if}\,\,\Theta=1,\\ \min\{X_{1},X_{2}\},&\text{if}\,\,\Theta=0.\end{array}\right$.
so that the random variable $Y$ of (10) can be written in the form
(30) $Y$ $=Z+\Theta f_{+}(|X_{2}-X_{1}|)+(1-\Theta)f_{-}(|X_{2}-X_{1}|).$
The next result describes the difference TFFTF-F in terms of $(Y,Z,\Theta)$. Specifically, let $\mathscr{L}$ be the operator defined on $CDF(\overline{\mathbb{R}})$ by
(31) $\mathscr{L}F(x)=-p\mathbb{P}\{Z\leq x<Y\mid\Theta=1\}+(1-p)\mathbb{P}\{Y\leq x<Z\mid\Theta=0\}.$
Proposition 3.
For any $F\in CDF(\overline{\mathbb{R}})$.
$TF-F=\mathscr{L}F+(1-2p)F(1-F).$
Further, if $p=\frac{1}{2}$, then $TF(\pm\infty)=F(\pm\infty)$ for any $F\in CDF(\overline{\mathbb{R}})$.
Using the explicit formulas (30) and (29), one obtains very concrete expressions for the action of $\mathscr{L}$ on smooth CDF’s, as stated in the next result.
Proposition 4.
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$. If $F\in CDF(\overline{\mathbb{R}})$ is smooth, then
(32) $\mathscr{L}F(x)$ $=2p\int_{-\infty}^{x}\mathbf{E}[F(x_{1}+g_{+}(x-x_{1}))-F(x)]F^{\prime}(x_{1})\,dx_{1}$
$\quad+2(1-p)\int_{x}^{\infty}\mathbf{E}[F(x_{1}-g_{-}(x_{1}-x))-F(x)]F^{\prime}(x_{1})\,dx_{1}.$
Proposition 4 has two immediate consequences that will be used in the sequel. First, it makes it relatively easy to see that $T$ commutes with the addition of constants whenever such addition makes sense.
Proposition 5.
Under the assumptions of Proposition 4, if $F\in CDF(\overline{\mathbb{R}})$, then $T(F+c)=TF+c$ for any constant $c\in[-F(-\infty),1-F(+\infty)]$.
The next consequence of Proposition 4, which follows from Proposition 5 and is stated precisely in Appendix B.3, concerns the contractive property of $T$, namely,
$\|TF-TG\|_{\sup}\leq\|F-G\|_{\sup}.$
This is classical for operators with the same properties as $T$ defined instead on the larger space $B(X)$ of bounded functions on some set $X$, which is precisely the setting considered in [6]. The fact that $T$ is defined only on $CDF(\overline{\mathbb{R}})$ makes the proof slightly less obvious, but contractivity holds nonetheless.
The remainder of this section is organized as follows: In Section 3.1, Proposition 3 is proved. The formula for $\mathscr{L}$ on smooth CDF’s is then proved in Section 3.2, along with Proposition 5.
3.1. Proof of Proposition 3
Fix a CDF $F$ and define $(X_{1},X_{2},Z,Y,\Theta,f_{+},f_{-})$ as in the discussion preceding the statement of the proposition. Start by decomposing the difference TFFTF-F as follows
$TF(x)-F(x)$ $=\mathbb{P}\{Y\leq x\}-F(x)$
(33) $=\Big(\mathbb{P}\{Y\leq x\}-\mathbb{P}\{Z\leq x\}\Big)+\Big(\mathbb{P}\{Z\leq x\}-F(x)\Big).$
Observe that, since $f_{+},f_{-}\geq 0$,
$Y-Z\geq 0\quad\text{if}\,\,\Theta=1,\quad\text{and}\quad Y-Z\leq 0\quad\text{if}\,\,\Theta=0.$
Thus, since $p=\mathbb{P}\{\Theta=1\}$ by definition,
$\mathbb{P}\{Y\leq x\}-\mathbb{P}\{Z\leq x\}$ $=-p\mathbb{P}\{Z\leq x<Y\mid\Theta=1\}+(1-p)\mathbb{P}\{Y\leq x<Z\mid\Theta=0\}$
$=\mathscr{L}F(x).$
It remains to treat the remaining summand, $\mathbb{P}\{Z\leq x\}-F(x)$. Notice that
$\{Z\leq x\}=\{\Theta=1,X_{1}\leq x\,\,\text{and}\,\,X_{2}\leq x\}\cup\{\Theta=0,X_{1}\leq x\,\,\text{or}\,\,X_{2}\leq x\}.$
Thus, since $X_{1}$ and $X_{2}$ are i.i.d. with CDF $F$, one obtains, exactly as in [18, Lemma 1.1],
$\mathbb{P}\{Z\leq x\}=pF(x)^{2}+(1-p)(1-(1-F(x))^{2}).$
Upon subtracting $F(x)$ and simplifying, this becomes
$\mathbb{P}\{Z\leq x\}-F(x)=(1-2p)F(x)(1-F(x)).$
Remark 1.
In [18, Lemma 1.1], it is observed that the fixed points of the map $q\mapsto pq^{2}+(1-p)(1-(1-q)^{2})$ are fundamental in the study of the asymptotics of the series-parallel graph. Yet, as in the derivation above, a fixed point of this map is nothing but a zero of the polynomial (12p)q(1q)(1-2p)q(1-q), the “reaction term” appearing in the equation.
3.1.1. Stationarity at infinity if $p=\frac{1}{2}$
If $p=\frac{1}{2}$, then
$TF(\pm\infty)-F(\pm\infty)=\mathscr{L}F(\pm\infty)$
and inspection of the formula (31) shows that $\mathscr{L}F(\pm\infty)=0$ independently of $F$. ∎
3.2. Proof of Proposition 4
Suppose that $F\in CDF(\overline{\mathbb{R}})$ is smooth. To obtain the formula (32) for $\mathscr{L}F$ in this case, it suffices to establish that
$\mathbb{P}\{Z\leq x<Y\mid\Theta=1\}=-2\int_{-\infty}^{x}\mathbf{E}[F(x_{1}+g_{+}(x-x_{1}))-F(x)]F^{\prime}(x_{1})\,dx_{1},$
$\mathbb{P}\{Y\leq x<Z\mid\Theta=0\}=2\int_{x}^{\infty}\mathbf{E}[F(x_{1}-g_{-}(x_{1}-x))-F(x)]F^{\prime}(x_{1})\,dx_{1}.$
Since the law of $(X_{1},X_{2})$ is invariant under permutation of the coordinates, this is equivalent to
$\mathbb{P}\{Z\leq x<Y,X_{1}\leq X_{2}\mid\Theta=1\}=-\int_{-\infty}^{x}\mathbf{E}[F(x_{1}+g_{+}(x-x_{1}))-F(x)]F^{\prime}(x_{1})\,dx_{1},$
$\mathbb{P}\{Y\leq x<Z,X_{2}\leq X_{1}\mid\Theta=0\}=\int_{x}^{\infty}\mathbf{E}[F(x_{1}-g_{-}(x_{1}-x))-F(x)]F^{\prime}(x_{1})\,dx_{1}.$
In particular, in the analysis that follows, only the case where $Z=X_{2}$ needs to be treated. Further, there is no loss of generality assuming that $x\in\mathbb{R}$ as the integral vanishes for $x\in\{-\infty,+\infty\}$ because $F(x_{1}+g_{+}(+\infty-x_{1}))=F(+\infty)$.
3.2.1. Case: $\Theta=1$
Condition first on the event $\{\Theta=1\}$. Before going further, it is worth noting that the assumption that $f_{+}(+\infty)=0$ implies that $X_{1}>$-$\infty$ almost surely on the event $\{Z\leq x<Y,\,\,X_{1}\leq X_{2},\,\,\Theta=1\}$ for $x\in\mathbb{R}$. This is immediate from the fact that $Y=X_{2}=Z$ on $\{X_{1}=-\infty,\,\,\Theta=1,\,\,f_{+}(+\infty)=0\}$.
Subtracting $X_{1}$ from both sides of the equation defining $Y$ while assuming that $X_{2}\geq X_{1}$ leads to
$Y-X_{1}=(X_{2}-X_{1})+f_{+}(X_{2}-X_{1}).$
At the same time, notice that $g_{+}$ is defined precisely so that, for any $u>0$ and $s\geq 0$,
$g_{+}(s)<u\quad\text{if and only if}\quad s<u+f_{+}(u).$
Thus,
$\{Z\leq x<Y,\,\,X_{1}<X_{2},\,\,\Theta=1\}$ $=\{-\infty<X_{1}<X_{2}\leq x,\,\,g_{+}(x-X_{1})<X_{2}-X_{1}\}$
$=\{-\infty<X_{1}<x,\,\,X_{1}+g_{+}(x-X_{1})<X_{2}\leq x\}.$
Therefore, since $f_{+}$ is independent of $(X_{1},X_{2})$,
$\mathbb{P}\{Z\leq x<Y,X_{1}\leq X_{2}\mid\Theta=1\}$ $=-\int_{-\infty}^{x}\mathbf{E}[F(x_{1}+g_{+}(x-x_{1}))-F(x)]F^{\prime}(x_{1})\,dx_{1}.$
3.2.2. Case: $\Theta=0$
The analysis of $\mathbb{P}\{Y\leq x<Z,X_{2}\leq X_{1}\mid\Theta=0\}$ follows from arguments very similar to those used when $\Theta=1$. In fact, the desired formula can be quickly derived from the fact that the transformation $(X_{1},X_{2},Z,Y)\mapsto(-X_{1},-X_{2},-Z,-Y)$ preserves the structure of the problem (see Section 2.2). ∎
3.3. Proof of Proposition 5
Suppose that $p=\frac{1}{2}$ and $F\in CDF(\overline{\mathbb{R}})$ satisfies $c_{-}\leq F\leq c_{+}$ pointwise for some $c_{-},c_{+}\in[0,1]$. Fix $c\in[-c_{-},1-c_{+}]$. The goal at this stage is to show that $T(F+c)=TF+c$.
Let $X$ be an extended-real-valued random variable with CDF $F$. Consider the CDF $F_{n}$ of the random variable $X+n^{-1}Y$, where $Y$ is a mean-zero, variance-one Gaussian, independent of $X$. Then $F_{n}$ is a smooth function and a standard coupling argument shows that both $F_{n}\to F$ and $TF_{n}\to TF$ pointwise in $\overline{\mathbb{R}}$, except for points of discontinuity, as $n\to\infty$. Therefore, since the same reasoning applies to $F+c$, it suffices to consider the case when $F$ is smooth. Yet direct inspection of formula (32) shows that $\mathscr{L}(F+c)=\mathscr{L}F$ for smooth $F$.
4. Examples
This section describes a few relevant examples, namely, those that have appeared earlier in the literature, and describes some of the useful, concrete forms of $\mathscr{L}$ that can be derived.
4.1. Resistance of the series-parallel graph
As mentioned already in the introduction, if $X^{(n)}=\log R^{(n)}$ is the logarithm of the resistance of the series-parallel graph, then $(X^{(n)})_{n\in\mathbb{N}}$ is a solution of the RDE (9) provided $f_{+}=f_{-}=f$ $\mathbf{P}$-almost surely, where $f$ is given by the formula
$f(u)=\log(1+e^{-u}).$
In this case, the right-continuous, right-inverse $g$ of the function $u\mapsto u+f(u)$ can be computed to be
$g(s)=\left\{\begin{array}[]{r l}\log(e^{s}-1),&\text{if}\,\,s\geq\log 2,\\ 0,&\text{otherwise.}\end{array}\right$.
It is immediate to check that $f\in\mathcal{S}$ and sg(s)=f(g(s))s-g(s)=f(g(s)) for $s\geq\log 2$ so assumption (15) of Theorem 1 is satisfied.
In this case, the constant $a$ is given by $a=2\zeta(3)$. See Appendix C for the proof.
4.2. Distance on the series-parallel graph
Let $D^{(n)}$ be the distance between the two terminal nodes in the series-parallel graph at stage $n$. Due to the recursive structure inherent in the graph, $D^{(n)}$ is a solution of the RDE
$D^{(n)}=\left\{\begin{array}[]{r l}D^{(n-1)}_{1}+D^{(n-1)}_{2},&\text{with probability}\,\,p,\\ \min\{D^{(n-1)}_{1},D^{(n-1)}_{2}\},&\text{with probability}\,\,1-p.\end{array}\right$.
This RDE also arises in Pemantle’s min-plus binary tree, and its asymptotics were analyzed in [5]. Analogous to the case of the resistance, the logarithm $X^{(n)}=\log D^{(n)}$ gives a solution of the RDE (9) with $\Phi$ given as in (8) provided $f_{+}(u)=\log(1+e^{-u})$ and $f_{-}=0$.
A complete analysis of the asymptotics of this RDE in the subcritical case ($p<\frac{1}{2}$) is given in [18]. There is a random variable $K$ such that if $\lambda$ denotes the essential infimum of $D^{(0)}$, then $D^{(n)}\to\lambda K$ as $n\to\infty$. This implies that $K$ is the unique (up to multiplication by a constant) fixed point of the RDE in this case.
In the critical case ($p=\frac{1}{2}$), the main result of [5] obtains a $\text{Beta}(2,1)$ limit for $N^{-1/2}\log D^{(N)}$ provided $D^{(0)}=1$ almost surely. The next corollary of Theorem 1 describes the asymptotics for an arbitrary initial distribution. Notice the contrast with the subcritical case: In the subcritical case, if $D^{(0)}$ can be arbitrarily small with positive probability, then $D^{(n)}\to 0$ as $n\to\infty$, whereas here, conditionally on the event $\{D^{(n)}>0\}$, mass at zero only slows the rate of growth.
Corollary 2.
Assume that $p=\frac{1}{2}$ and $\mathbb{P}\{0\leq D^{(0)}<\infty\}=1$ and let $q=\mathbb{P}\{D^{(0)}>0\}$. For any $n$, $\mathbb{P}\{D^{(n)}=0\}=q$. Further,
$\lim_{N\to\infty}\mathbb{P}\{-\frac{\text{sgn}(\sigma)}{2\sqrt{|\sigma|}\sqrt{N}}\log D^{(N)}\leq x\,\mid\,D^{(N)}>0\}=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ \frac{x^{2}}{q},&\text{if}\,\,0\leq x\leq\sqrt{q},\\ 1,&\text{otherwise.}\end{array}\right$.
Here the constant $\sigma=-\frac{1}{2}\zeta(2)=-\frac{\pi^{2}}{12}$.
Proof.
Let $X^{(n)}=\log D^{(n)}$ so that $\{D^{(n)}>0\}=\{X^{(n)}>-\infty\}$. By Proposition 3, $\mathbb{P}\{X^{(n)}>-\infty\}$ is constant in $n$.
Let $F_{n}$ be the CDF of $X^{(n)}$ and let $G_{n}$ be the CDF conditional on $\{X^{(n)}>-\infty\}$, that is,
$G_{n}(x)=\mathbb{P}\{X^{(n)}\leq x\,\mid\,X^{(n)}>-\infty\}.$
Notice that $G_{n}=q^{-1}(F_{n}-(1-q))$. At the same time, the strengthened version of Theorem 1 from Section 5 (Theorem 2) applies: If $F_{\epsilon}(x,t)=F_{[\epsilon^{-2}t]}(\epsilon^{-1}x)$, then the theorem implies that
$F_{\epsilon}\to F\quad\text{locally uniformly in}\,\,\mathbb{R}\times(0,\infty),$
where $F$ is the unique upper semicontinuous viscosity solution of the PDE $\partial_{t}F-\sigma|\partial_{x}F|^{2}=0$ with initial datum $F(x,0)=(1-q)+q\mathbf{1}_{[0,\infty)}(x)$. After writing $G_{\epsilon}(x,t)=G_{[\epsilon^{-2}t]}(\epsilon^{-1}x)$, it follows from this that $G_{\epsilon}\to G$, where $G$ is the unique discontinuous viscosity solution of $\partial_{t}G-q\sigma|\partial_{x}G|^{2}=0$ with $G(x,0)=\mathbf{1}_{[0,\infty)}$. The proof of Corollary 1 in Section 5 gives an explicit formula for this function, from which the claimed asymptotics of $\log D^{(N)}$ follow.
The claimed value of $\sigma$ is computed in Appendix C and is consistent with [5]. ∎
It is worth emphasizing that $D^{(n)}$ is integer-valued, hence its logarithm takes values in the set $\log\mathbb{N}$ for all $n$. For the present choice of $f_{+}$ and $f_{-}$, it is not hard to show that this remains true for any solution that starts initially in $\log\mathbb{N}$. This is worth mentioning because the equation for the CDF of such a solution takes a particularly simple form, namely, if the backward-difference operator $\nabla^{-}_{\log\mathbb{N}}$ is defined for functions on $\log\mathbb{N}$ by $\nabla^{-}_{\log\mathbb{N}}F(x)=F(x)-F(x_{-})$, where $x_{-}$ denotes the largest element of $\log\mathbb{N}$ smaller than $x$, then
$F_{n+1}-F_{n}=-p(\nabla^{-}_{\log\mathbb{N}}F)^{2}+(1-2p)F_{n}(1-F_{n})\quad\text{in}\,\,\log\mathbb{N}.$
Notice that this equation looks like a discrete-time, discrete-in-space version of (23) with $a=0$.
4.3. $\mathbb{Z}$-valued Solutions and Finite-Difference Equations
This section discusses the subclass of RDE’s described by (9) and (8) in the case when $f_{+}$ and $f_{-}$ take values in the set $\{0,f_{\mathbb{Z}}\}$, where $f_{\mathbb{Z}}(u)=(1-u)_{+}$. For instance, if the law of the data is chosen so that $f_{+}=f_{-}=f_{\mathbb{Z}}$ surely, then $X^{(n)}$ takes a jump only if $X^{(n-1)}_{1}$ and $X^{(n-1)}_{2}$ are at most distance one apart, and the size of the jump decays linearly with the distance.
What is nice about this particular case of the RDE is a $\mathbb{Z}$-valued RDE is embedded within it: If the initial data $X^{(0)}$ takes values in $\mathbb{Z}$, then $X^{(n)}$ remains integer-valued for all $n$ since $f_{\mathbb{Z}}$ maps integers to integers. For this $\mathbb{Z}$-valued RDE, the one-step evolution is described by
(36) $Y\overset{d}{=}\left\{\begin{array}[]{r l}\max\{X_{1},X_{2}\}+\mathbf{1}_{\{0\}}(|X_{2}-X_{1}|),&\text{if}\,\,\Theta=1,\\ \min\{X_{1},X_{2}\}-\mathbf{1}_{\{0\}}(|X_{2}-X_{1}|),&\text{if}\,\,\Theta=0.\end{array}\right$.
The RDE associated with this equation provides a very natural distribution-dependent random walk on $\mathbb{Z}$, and it has the nice feature that the equations describing the CDF are nothing other than finite-difference discretizations of the quasilinear Fisher-KPP equations mentioned in the introduction.
For the rest of this subsection, assume that the joint law $\mathbf{P}$ of $(f_{+},f_{-})$ is concentrated on the set $\{0,f_{\mathbb{Z}}\}\times\{0,f_{\mathbb{Z}}\}$. Notice that
$f_{\mathbb{Z}}\in\mathcal{S}\quad\text{and}\quad g_{\mathbb{Z}}(s)=s\quad\text{for each}\,\,s\geq 1,$
where $g_{\mathbb{Z}}$ denotes the left-inverse of the function $u\mapsto u+f_{\mathbb{Z}}(u)$. For such a law $\mathbf{P}$, all the assumptions of Theorem 1 are satisfied. In particular, as long as $(f_{+},f_{-})$ is not identically zero, either $n^{-\frac{1}{2}}X^{(n)}$ or $n^{-\frac{1}{3}}X^{(n)}$ converges to a nondegenerate limit.
Let $q_{+}=\mathbf{P}\{f_{+}=f_{\mathbb{Z}}\}$ and $q_{-}=\mathbf{P}\{f_{-}=f_{\mathbb{Z}}\}$ in the discussion that follows.
For $\mathbb{Z}$-valued solutions of the RDE, there is no loss restricting the CDF to the domain $\mathbb{Z}$. By invoking the formula (31), one obtains, in this case,
(37) $F_{n+1}-F_{n}$ $=-pq_{+}(\nabla_{\mathbb{Z}}^{-}F_{n})^{2}+(1-p)q_{-}(\nabla_{\mathbb{Z}}^{+}F_{n})^{2}+(1-2p)F_{n}(1-F_{n}),$
where $\nabla_{\mathbb{Z}}^{+}$, $\nabla_{\mathbb{Z}}^{-}$ are the forward- and backward-difference operators
$\nabla_{\mathbb{Z}}^{+}F(x)=F(x+1)-F(x),\quad\nabla_{\mathbb{Z}}^{-}F(x)=F(x)-F(x-1).$
Using the identity $\nabla_{\mathbb{Z}}^{+}-\nabla_{\mathbb{Z}}^{-}=\Delta_{\mathbb{Z}}$ for $\Delta_{\mathbb{Z}}F(x)=(F(x+1)+F(x-1))-F(x)$ and simplifying the difference of the two squares, this becomes
(38) $F_{n+1}-F_{n}$ $=a\nabla^{\text{sym}}_{\mathbb{Z}}F_{n}\Delta_{\mathbb{Z}}F_{n}+\sigma(\nabla_{\mathbb{Z}}^{+}F_{n})^{2}+(1-2p)F_{n}(1-F_{n}).$
where $\nabla^{\text{sym}}_{\mathbb{Z}}=2^{-1}(\nabla_{\mathbb{Z}}^{+}+\nabla_{\mathbb{Z}}^{-})$, $\sigma=(1-p)q_{-}-pq_{+}$, and $a=2pq_{+}$.
Equation (38) looks like a discrete analogue of the quasilinear Fisher-KPP equation (23) derived above. Notice that here $a$ or $\sigma$ may vanish, irrespective of whether or not $p=\frac{1}{2}$. Since replacing $\mathbb{Z}$ by $\mathbb{R}$ and finite-differences by derivatives does not seem likely to fundamentally change the large-scale asymptotic behavior, equation (38) motivates the general study of (23) for arbitrary $\sigma\in\mathbb{R}$ and $a\geq 0$.
It is worth emphasizing that there is no contradiction between the formulas (38) and (32) since the former involves the CDF of a $\mathbb{Z}$-valued random variable while the latter describes smooth CDF’s.
4.4. Hipster Random Walks and Cooperative Motion
Inspection of (38) also reveals a connection with the symmetric hipster random walk and two-player cooperative motion, examples considered in in [1] and [3], respectively.
To understand the connection, consider again the random variable $Z$ defined by $Z=\Theta\max\{X_{1},X_{2}\}+(1-\Theta)\min\{X_{1},X_{2}\}$. Recall that $Z$ and $X_{1}$ have the same distribution if $p=\frac{1}{2}$. At the same time, since the two coincide when $X_{1}=X_{2}$, a stronger statement is true: $Z$ and $X_{1}$ have the same conditional laws on the event $\{X_{1}\neq X_{2}\}$. The same reasoning applies if $X_{1}$ is replaced by the more symmetric $\Theta X_{1}+(1-\Theta)X_{2}$. From this, it follows that the law of the output $Y$ is unchanged if the equation is instead changed to
(41) $Y\overset{d}{=}\left\{\begin{array}[]{r l}X_{1}+\mathbf{1}_{\{0\}}(|X_{2}-X_{1}|),&\text{if}\,\,\Theta=1,\\ X_{2}-\mathbf{1}_{\{0\}}(|X_{2}-X_{1}|),&\text{if}\,\,\Theta=0,\end{array}\right$.
or
(44) $Y\overset{d}{=}\left\{\begin{array}[]{r l}X_{1}+\mathbf{1}_{\{0\}}(|X_{2}-X_{1}|),&\text{if}\,\,\Theta=1,\\ X_{1}-\mathbf{1}_{\{0\}}(|X_{2}-X_{1}|),&\text{if}\,\,\Theta=0.\end{array}\right$.
These are precisely the one-step updates used to define the symmetric hipster random walk and two-player symmetric cooperative motion, respectively. Thus, although it is not immediately obvious from inspection of (41) or (44), those two RDE’s belong to the class considered here.
In fact, by this reasoning, it is also possible to draw a connection with the other RDE considered in [1], called the totally asymmetric hipster random walk. This is the RDE in $\mathbb{Z}$ with the one-step update determined by two independent Bernoulli random variables $\Theta,\Xi$ via the rule
$Y\overset{d}{=}\left\{\begin{array}[]{r l}X_{1}+\Xi\mathbf{1}_{\{0\}}(|X_{2}-X_{1}|),&\text{if}\,\,\Theta=1,\\ X_{2}+\Xi\mathbf{1}_{\{0\}}(|X_{2}-X_{1}|),&\text{if}\,\,\Theta=0.\end{array}\right$.
where $\mathbb{P}\{\Theta=1\}=\frac{1}{2}$. This is a class of RDE’s parametrized by $q:=\mathbb{P}\{\Xi=1\}$. By the same reasoning employed above, as long as $q\leq\frac{1}{2}$, the totally asymmetric hipster random walk belongs to the subclass of RDE’s considered in the previous subsection. This is made precise in the next result.
Proposition 6.
Assume that $p=\frac{1}{2}$ and $\mathbf{P}\{f_{+},f_{-}\in\{0,f_{\mathbb{Z}}\}\}=1$, and define $(q_{+},q_{-})$ by $q_{\rho}=\mathbf{P}\{f_{\rho}=f_{\mathbb{Z}}\}$. If $\{X^{(n)}\}$ is the solution of the RDE (9) with forcing $\Phi$ given by (8) for some integer-valued initial condition $X^{(0)}$, then, for each $n$, $X^{(n)}$ has the same law as (i) the symmetric hipster random walk with the same initial datum if $q_{+}=q_{-}=1$ and (ii) the totally asymmetric hipster random walk with the same initial datum if $q_{-}=0$ provided the parameter $q$ is set to $q=\frac{1}{2}q_{+}$.
Note that the equation $q=\frac{1}{2}q_{+}$ introduces the constraint $q\leq\frac{1}{2}$, so that the previous result only covers the asymmetric hipster random walk for half of the possible values of $q$. In fact, it is possible to show (by manipulation of the evolution equation satisfied by the CDF) that the asymmetric hipster random walk is a monotone RDE only if $q\leq\frac{1}{2}$. Therefore, it cannot be described using (9) outside this regime.
Similar arguments apply to the two-player version of the RDE’s analyzed in [2]. Specifically, two-player totally asymmetric, $q$-lazy cooperative motion is the RDE for $\mathbb{Z}$-valued random variables given by
$X^{(n)}\overset{d}{=}\left\{\begin{array}[]{r l}X_{1}^{(n-1)}+\mathbf{1}_{\{0\}}(|X_{2}^{(n-1)}-X_{1}^{(n-1)}|),&\text{with probability}\,\,rq,\\ X_{1}^{(n-1)}-\mathbf{1}_{\{0\}}(|X_{2}^{(n-1)}-X_{1}^{(n-1)}|),&\text{with probability}\,\,(1-r)q,\\ X_{1}^{(n-1)},&\text{with probability}\,\,1-q.\end{array}\right$.
The RDE is parameterized by the pair $(r,q)$. The next result describes the connection with the RDE’s considered here. As for the asymmetric hipster random walk, only a subset of the parameter values are covered.
Proposition 7.
Assume that $p=\frac{1}{2}$ and $\mathbf{P}\{f_{+},f_{-}\in\{0,f_{\mathbb{Z}}\}\}=1$, and define $(q_{+},q_{-})$ by $q_{\rho}=\mathbb{P}\{f_{\rho}=f_{\mathbb{Z}}\}$. If $\{X^{(n)}\}$ is the solution of the RDE (9) with forcing $\Phi$ given by (8) for some integer-valued initial condition $X^{(0)}$, then, for each $n$, $X^{(n)}$ has the same law as (i) two-player symmetric cooperative motion with that initial condition if $q_{+}=q_{-}=1$ or (ii) two-player asymmetric, $q$-lazy cooperative motion with that initial condition if $q_{+}\neq q_{-}$, provided the parameters $(r,q)$ are set to $q=\frac{1}{2}(q_{+}+q_{-})$ and $r=\frac{q_{+}}{q_{-}+q_{+}}$.
Notice that when $q_{+}\neq q_{-}$, the range of the map $(q_{+},q_{-})\mapsto(r,q)$ is
(45) $\{(r,q)\,\mid\,q\in[0,1],\,\,|r-\frac{1}{2}|\leq\frac{1}{2}(q^{-1}-1)\}.$
In general, this is a proper subset of the unit square. However, for any fixed $q\leq\frac{1}{2}$, it covers all possible values of $r\in[0,1]$.
As in the case of hipster random walks, the restriction in (45) is fundamental. Two-player totally asymmetric, $q$-lazy cooperative motion is monotone only for $(r,q)$ in the set given above; this can readily be proved via differentiation of the evolution equation satisfied by the CDF (see [2, Equation (1.6)]).
In view of the proposition above, the $\text{Beta}(2,2)$ limit theorems for the symmetric RDE’s considered in [1] and [3] follow as a special case of this paper’s main theorem (Theorem 1) and Corollary 1. The $\text{Beta}(2,1)$ limit theorems for asymmetric RDE’s obtained in [1] and [2] also follow, albeit only in the parameter regimes identified above. Interestingly, in the asymmetric case, those works obtain $\text{Beta}(2,1)$ limits, even in the nonmonotone regime.
5. Scaling Limit
This section proves the main result, Theorem 1. In fact, a slightly more general version is presented here, which allows for extended-real-valued solutions and diffuse initial conditions. In order to state the theorem efficiently, as in the introduction, let $\alpha=2$ if $\sigma\neq 0$ and $\alpha=3$, otherwise, and define the nonlinearity $\mathcal{F}$ appearing in the limiting PDE analogously:
(48) $\mathcal{F}(p,X)=\left\{\begin{array}[]{r l}\sigma|p|^{2},&\text{if}\,\,\sigma\neq 0,\\ a|p|X,&\text{otherwise.}\end{array}\right$.
The theorem that follows involves the notion of viscosity solutions of parabolic PDE’s. The relevant definitions and existence and uniqueness results are reviewed in Appendix A below.
Theorem 2.
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds and fix an arbitrary $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$.
For each $\epsilon>0$, let $\{X_{\epsilon}^{(n)}\}$ be an extended-real-valued solution of the RDE (9) and let $\{F^{\epsilon}_{n}\}$ be the associated sequence of CDF’s. Let $F_{\epsilon}:\mathbb{R}\times[0,\infty)\to[0,1]$ be the rescaled CDF given by
$F_{\epsilon}(x,t)=F^{\epsilon}_{[\epsilon^{-\alpha}t]}(\epsilon^{-1}x).$
If convergence holds as $\epsilon\to 0$ at the initial time in the sense that
$F_{\epsilon}(x,0)\to F_{\text{in}}(x)\quad\text{for each point of continuity}\,\,x\,\,\text{of}\,\,F_{\text{in}},$
then
$F_{\epsilon}\to F\quad\text{locally uniformly in}\,\,\mathbb{R}\times(0,\infty),$
where $F$ is the unique bounded discontinuous viscosity solution of the initial-value problem
$\left\{\begin{array}[]{r l}\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
The theorem will be proved by following the approach of [6], suitably extended to monotone semigroups in the space $CDF(\overline{\mathbb{R}})$. Notice that if $T_{\epsilon}$ and $\mathscr{L}_{\epsilon}$ are the operators on $CDF(\overline{\mathbb{R}})$ defined by
(49) $(T_{\epsilon}G)(x)=(TG_{\epsilon})(\epsilon^{-1}x),\quad\text{where}\,\,G_{\epsilon}(x)=G(\epsilon^{-1}x),\quad\mathscr{L}_{\epsilon}=T_{\epsilon}-\text{Id},$
then the rescaled CDF $F_{\epsilon}$ above satisfies the equation
(50) $F_{\epsilon}(x,t+\epsilon^{\alpha})-F_{\epsilon}(x,t)=\mathscr{L}_{\epsilon}F_{\epsilon}(x,t).$
This motivates the consideration of one-parameter families of operators $\{T_{\epsilon}\}$ defined on $CDF(\overline{\mathbb{R}})$.
5.1. Abstract Framework
Let $\{T_{\epsilon}\}$ be a family of operators on $CDF(\overline{\mathbb{R}})$ with the following properties:
  • (i)
    Monotonicity: If $F\leq G$ pointwise in $\mathbb{R}$, then $T_{\epsilon}F\leq T_{\epsilon}G$ also holds.
  • (ii)
    Commutation with addition of constants: If $F\in CDF(\overline{\mathbb{R}})$ satisfies $c_{-}\leq F\leq c_{+}$ pointwise for some constants $c_{-},c_{+}\in[0,1]$, then $T_{\epsilon}(F+c)=T_{\epsilon}F+c$ for any $c\in[-c_{-},1-c_{+}]$.
  • (iii)
    Stationarity at infinity: For any $F\in CDF(\overline{\mathbb{R}})$, $T_{\epsilon}F(+\infty)=F(+\infty)$ and $T_{\epsilon}F(-\infty)=F(-\infty)$.
Motivated by (50), let $\{\mathscr{L}_{\epsilon}\}$ be the operators on $CDF(\overline{\mathbb{R}})$ defined by
$\mathscr{L}_{\epsilon}=T_{\epsilon}-\text{Id}.$
In addition to (i) and (ii), assume that there is an exponent $\alpha>0$ and a continuous function $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, which is increasing in the second variable, for which the following property holds:
  • (iv)
    Consistency: If $\Phi\in CDF(\overline{\mathbb{R}})$ is smooth with bounded, uniformly continuous second derivatives, then
    $\lim_{\epsilon\to 0}\epsilon^{-\alpha}\mathscr{L}_{\epsilon}\Phi(x)=\mathcal{F}(\partial_{x}\Phi(x),\partial_{x}^{2}\Phi(x)),$
    where convergence holds uniformly with respect to $x\in\mathbb{R}$.
Note that (iii) implies $Tc=c$ for any constant (function) $c\in[0,1]$, and, thus, by (iv),
(51) $\mathcal{F}(0,0)=0.$
Under these assumptions, the following convergence result holds for rescaled solutions $\{F^{\epsilon}_{n}\}$ of the recursion $F^{\epsilon}_{n+1}=T_{\epsilon}F^{\epsilon}_{n}$.
Theorem 3.
Assume that $\{T_{\epsilon}\}$ is a family of operators on $CDF(\overline{\mathbb{R}})$ satisfying assumptions (i)-(iv) above. For each $\epsilon$, let $\{F^{\epsilon}_{n}\}$ be a sequence in $CDF(\overline{\mathbb{R}})$ generated by the recursion $F_{n+1}^{\epsilon}=T_{\epsilon}F_{n}^{\epsilon}$. If $F_{\epsilon}:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is the rescaled CDF defined by
$F_{\epsilon}(x,t)=F^{\epsilon}_{[\epsilon^{-\alpha}t]}(x),$
and if there is a continuous $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ such that $F_{\epsilon}(x,0)\to F_{\text{in}}(x)$ pointwise as $\epsilon\to 0$, then
$F_{\epsilon}\to F\quad\text{locally uniformly in}\,\,\mathbb{R}\times(0,\infty),$
where $F$ is the unique bounded continuous viscosity solution of the equation
(52) $\left\{\begin{array}[]{r l}\partial_{t}F=\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
Notice that the one-parameter family $\{T_{\epsilon}\}$ of interest in this paper (see (49)) satisfies assumptions (i)-(iii) according to Propositions 1 and 5. It only remains to verify condition (iv). That is the subject of the next proposition.
Proposition 8.
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds, and consider the operator $\mathscr{L}$ given by (31). If $\Phi:\mathbb{R}\to[0,1]$ is a nondecreasing function with bounded, uniformly continuous second derivative, rescaled according to the rule $\Phi_{\delta}(x)=\Phi(\delta x)$, and if $R_{\Phi}$ is defined by the identity
$(\mathscr{L}\Phi_{\delta})(\delta^{-1}x)=\delta^{2}\sigma|\partial_{x}\Phi(x)|^{2}+\delta^{3}a|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)+\delta^{3}R_{\Phi}(\delta;x),$
then
$\lim_{\delta\to 0}\sup\left\{|R_{\Phi}(\delta;x)|\,\mid\,x\in\mathbb{R}\right\}=0.$
The proof of the proposition is relegated to the end of this section.
There is one last technicality to address, namely, whereas Theorem 3 asks for the continuity of the initial datum $F_{\text{in}}$, the main theorem of this section allows for discontinuities. This gap is bridged using the fact that the equations of interest here are regularizing. This is made precise in the next condition:
  • (v)
    Asymptotically regularizing: If $\underline{F}$ and $\overline{F}$ denote the minimal and maximal viscosity sub- and supersolutions of the initial-value problem (52) for a given $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ (see Appendix A.1), then
    $\underline{F}=\overline{F}\quad\text{in}\,\,\mathbb{R}\times(0,\infty).$
The next corollary asserts that if the family $\{T_{\epsilon}\}$ satisfies assumption (v) in addition to (i)-(iv), then Theorem 3 remains true even if $F_{\text{in}}$ has discontinuities.
Corollary 3.
Assume that $\{T_{\epsilon}\}$ is a family of operators on $CDF(\overline{\mathbb{R}})$ satisfying assumptions (i)-(v) above. Then the statement of Theorem 3 remains true for an arbitrary $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ provided $F$ is taken to be the unique discontinuous viscosity solution of the initial-value problem (52).
See Appendix B for the proof.
In Theorem 2, the nonlinearity $\mathcal{F}$ takes the form (48), which satisfies assumption (v) by Propositions 10 and 11 below. Therefore, the proof of the theorem is complete.
Remark 2.
Theorem 3 is stated for semigroups defined in $CDF(\overline{\mathbb{R}})$, the space of nondecreasing, right-continuous functions taking values in $[0,1]$. There is nothing special about right-continuity, however. In fact, let $F_{-}$ and $F^{+}$ denote the left- and right-increasing versions of a nondecreasing function $F$. If operators $\{T_{\epsilon,-}\}$ are defined on the space $CDF_{\text{left}}(\overline{\mathbb{R}})$ of nondecreasing, left-continuous functions taking values in $[0,1]$ by the formula $T_{\epsilon,-}F=(T_{\epsilon}(F^{+}))_{-}$, then $\{T_{\epsilon,-}\}$ satisfies assumptions (i)-(iv) if and only if $\{T_{\epsilon}\}$ does. Further, it is not hard to see that a PDE scaling limit of $\{T_{\epsilon}\}$ holds as in Theorem 3 if and only if the same conclusion applies to $\{T_{\epsilon,-}\}$.
5.2. Proof of Corollary 1
It only remains to reinterpret Theorem 2 in terms of the $\text{Beta}(2,1)$ and $\text{Beta}(2,2)$ distributions.
Let $\{X^{(n)}\}$ be the solution of the RDE (9) started from an arbitrary initial datum. According to the theorem,
(53) $\lim_{N\to\infty}\mathbb{P}\{N^{-1/\alpha}X^{(N)}\leq x\}=F(x,1),$
where $F$ is the unique bounded discontinuous viscosity solution of (52) with initial datum $F_{\text{in}}=\mathbf{1}_{[0,+\infty)}$.
If $\sigma<0$, then $\mathcal{F}(p,X)=\sigma|p|^{2}$ and
$F(x,t)=\frac{x^{2}}{4|\sigma|t}\quad\text{for each}\,\,x\in[0,2\sqrt{|\sigma|}t^{1/2}],$
as can be proved directly or by using the Hopf-Lax formula (see Proposition 10 below). If $\sigma>0$, then one can instead study the asymptotics of $X^{(n)}$ reasoning as in Section 2.2, which has the effect of flipping the sign of $\sigma$. In any event, (53) implies
$\lim_{N\to\infty}\mathbb{P}\{-\frac{\text{sgn}(\sigma)}{2\sqrt{|\sigma|}N^{1/2}}X^{(N)}\leq x\}=F(2\sqrt{|\sigma|}x,1)=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ x^{2},&\text{if}\,\,0\leq x\leq 1,\\ 1,&\text{otherwise.}\end{array}\right$.
The right-hand side is precisely the CDF of the $\text{Beta}(2,1)$ distribution. Thus, case (i) is complete.
If $\sigma=0$ and $a>0$, then $\mathcal{F}(p,X)=a|p|X$ and Proposition 11 asserts that $F(x,t)=\int_{-\infty}^{x}\rho(y,t)\,dy$, where $\rho$ is the unique continuous distributional solution of the porous medium equation $\partial_{t}\rho=\frac{1}{2}a\partial_{x}^{2}(\rho^{2})$ with initial datum $\rho(0)=\delta_{0}$. This solution, called the Barenblatt (or ZKB) solution, is given by
$\rho(x,t)=\frac{1}{(6at)^{\frac{1}{3}}}\left(\left(\frac{3}{4}\right)^{\frac{2}{3}}-\frac{x^{2}}{(6at)^{\frac{2}{3}}}\right)\mathbf{1}_{[-(9at/2)^{\frac{1}{3}},(9at/2)^{\frac{1}{3}}]}(x).$
By (53), for any bounded continuous function $h$, if $Y^{(N)}=(36a)^{-\frac{1}{3}}N^{-1/3}X^{(N)}+\frac{1}{2}$, then
$\lim_{N\to\infty}\mathbb{E}[h(Y^{(N)})]$ $=\int_{-(9a/2)^{\frac{1}{3}}}^{(9a/2)^{\frac{1}{3}}}h((36a)^{-\frac{1}{3}}x+2^{-1})\rho(x,1)\,dx=6\int_{0}^{1}h(y)y(1-y)\,dy,$
and the integrand $6y(1-y)\mathbf{1}_{[0,1]}(y)$ is nothing but the PDF of the $\text{Beta}(2,2)$ distribution. This proves case (ii).
Finally, if $\sigma=a=0$, then the formula in Section 1.3.2 implies that $(3s-g_{\rho}(s))(s-g_{\rho}(s))=0$ for all $s\geq 0$ and $\rho\in\{+,-\}$ almost surely. Since $g_{+}(s),g_{-}(s)\leq s$ hold automatically by definition, this implies that
$\mathbf{P}\{f_{+}=f_{-}=0\}=\mathbf{P}\{g_{+}(s)=g_{-}(s)=s\,\,\text{for all}\,\,s\geq 0\}=1,$
and thus $X^{(n)}\overset{d}{=}X^{(0)}$ for all $n$. This is case (iii).
5.3. Proof of Proposition 8
Let $\omega_{\Phi}$ denote the modulus of continuity of the second derivative $\partial_{x}^{2}\Phi$:
$\omega_{\Phi}(\epsilon)=\sup\left\{|\partial_{x}^{2}\Phi(x)-\partial_{x}^{2}\Phi(y)|\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\epsilon\right\}.$
In view of the representation of $\mathscr{L}$ from Proposition 4 and symmetry (see Section 2.2), it suffices to prove that
(54) $\int_{-\infty}^{\delta^{-1}x}\mathbf{E}[\Phi_{\delta}(y_{1}+g_{+}(\delta^{-1}x-y_{1}))-\Phi_{\delta}(\delta^{-1}x)]\partial_{x}\Phi_{\delta}(y_{1})\,dy_{1}$
$\quad=-\delta^{2}(\partial_{x}\Phi(x))^{2}\int_{0}^{\infty}\mathbf{E}[|g_{+}(u)-u|]\,du$
$\qquad+\delta^{3}\Phi^{\prime}(x)\Phi^{\prime\prime}(x)\left(\frac{1}{2}\int_{0}^{\infty}\mathbf{E}[(g_{+}(u)-u)^{2}]\,du-\int_{0}^{\infty}u\mathbf{E}[g_{+}(u)-u]\,du\right)$
$\qquad+C_{0}\delta^{3}\|\Phi\|_{C^{2}(\mathbb{R})}\left(\int_{0}^{\infty}\mathbf{E}[\omega_{\Phi}(\delta|g(u)-u|)\{g_{+}(u)-u\}^{2}]\,du\right$.
$\qquad\left.+\int_{0}^{\infty}u\omega_{\Phi}(\delta u)\mathbf{E}[|g(u)-u|]\,du+\delta\|\Phi\|_{C^{2}(\mathbb{R})}\int_{a}^{\infty}\mathbf{E}[\{g_{+}(u)-u\}^{2}]u\,du\right),$
where $\|\Phi\|_{C^{2}(\mathbb{R})}=\|\Phi\|_{\sup}+\|\partial_{x}\Phi\|_{\sup}+\|\partial^{2}_{x}\Phi\|_{\sup}$. To this end, first, notice that
$\int_{-\infty}^{\delta^{-1}x}\mathbf{E}[\Phi_{\delta}(y_{1}+g_{+}(\delta^{-1}x-y_{1}))-\Phi_{\delta}(\delta^{-1}x)]\partial_{x}\Phi_{\delta}(y_{1})\,dy_{1}$
$\quad=\int_{-\infty}^{x}\mathbf{E}[\Phi(x_{1}+\delta g_{+}(\delta^{-1}(x-x_{1})))-\Phi(x)]\partial_{x}\Phi(x_{1})\,dx_{1}.$
Next, it is worth noting that, by definition, the difference $s\mapsto g_{+}(s)-s$ is bounded above and below as follows
$-\sup\left\{|f(u)|\,\mid\,u\geq 0\right\}\leq g_{+}(s)-s\leq 0.$
Therefore, the assumption (15) implies that the integrals above are all finite.
Let $L$ and $Q$ be the Taylor polynomials
$L(x_{1})=\partial_{x}\Phi(x)+\partial_{x}^{2}\Phi(x)(x_{1}-x),\quad Q(x_{1})=\partial_{x}\Phi(x)(x_{1}-x)+\frac{1}{2}\partial_{x}^{2}\Phi(x)(x_{1}-x)^{2}.$
For notational ease, define $x_{1,g}$ according to the formula
$x_{1,g}=x_{1}+\delta g(\delta^{-1}[x-x_{1}]).$
The term of interest then becomes
$\int_{-\infty}^{x}\{\Phi(x_{1,g})-\Phi(x)\}\partial_{x}\Phi(x_{1})\,dx_{1}$ $=\int_{-\infty}^{x}L(x_{1})Q(x_{1,g})\,dx_{1}$
$\quad+\int_{-\infty}^{x}L(x_{1})\{\Phi(x_{1,g})-\Phi(x)-Q(x_{1,g})\}\,dx_{1}$
$\quad+\int_{-\infty}^{x}\{\partial_{x}\Phi(x_{1})-L(x_{1})\}\{\Phi(x_{1,g})-\Phi(x)\}\,dx_{1}.$
Expanding the main term and using the change-of-variables $x_{1}=x+\delta u$ leads to
$\int_{-\infty}^{x}L(x_{1})\mathbf{E}[Q(x_{1,g})]\,dx_{1}$ $=\delta^{2}(\partial_{x}\Phi(x))^{2}\int_{0}^{\infty}\mathbf{E}[g_{+}(u)-u]\,du$
$\quad+\frac{1}{2}\delta^{3}\partial_{x}\Phi(x)\partial_{x}^{2}\Phi(x)\int_{-\infty}^{x}\mathbf{E}[(g_{+}(u)-u)^{2}]\,dx_{1}$
$\quad-\delta^{3}\partial_{x}\Phi(x)\partial_{x}^{2}\Phi(x)\int_{-\infty}^{x}u\mathbf{E}[g_{+}(u)-u]\,dx_{1}$
$\quad-\frac{1}{2}\delta^{4}(\partial_{x}^{2}\Phi(x))^{2}\int_{-\infty}^{x}u\mathbf{E}[(g(u)-u)^{2}]\,dx_{1}.$
Notice that the order $\delta^{2}$ term has the desired form: In general, $g_{+}(s)\leq s$ always holds so $g_{+}(s)-s=-|g_{+}(s)-s|$.
It only remains to estimate the error terms:
$\left|\int_{-\infty}^{x}L(x_{1})\{\Phi(x_{1,g})-\Phi(x)-Q(x_{1,g})\}\,dx_{1}\right|$
$\quad\leq\|\Phi\|_{C^{2}(\mathbb{R})}\int_{-\infty}^{x}\omega_{\Phi}(|x_{1,g}-x|)\left(1+|x_{1}-x|\right)\{x_{1,g}-x\}^{2}\,dx_{1}$
$\quad=\delta^{3}\|\Phi\|_{C^{2}(\mathbb{R})}\int_{0}^{\infty}\omega_{\Phi}(\delta|g(u)-u|)(1+\delta|u|)\{g(u)-u\}^{2}\,du,$
$\left|\int_{-\infty}^{x}\{\partial_{x}\Phi(x_{1})-L(x_{1})\}\{\Phi(x_{1,g})-\Phi(x)\}\,dx_{1}\right|$
$\quad\leq\|\Phi\|_{C^{2}(\mathbb{R})}\int_{-\infty}^{x}\omega_{\Phi}(|x_{1}-x|)|x_{1}-x||x_{1,g}-x|\,dx_{1}$
$\quad=\delta^{3}\|\Phi\|_{C^{2}(\mathbb{R})}\int_{0}^{\infty}\omega_{\Phi}(\delta u)u|g(u)-u|\,du.$
Appendix A Viscosity Solutions
This section reviews relevant definitions and results from the theory of viscosity solutions. General references for this material are [13, 7]. The main interest here is the initial-value problem
(57) $\left\{\begin{array}[]{r l}\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
Since the problem at hand naturally leads to consideration of possibly discontinuous initial data, some care will be needed in defining precisely what is meant by a solution, subsolution, or supersolution.
The next definition recalls what it means for a function to satisfy a partial differential equation in the viscosity sense.
Definition 1.
An upper semicontinuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)\leq 0$ in the viscosity sense if it has the following property: If $\varphi$ is a smooth function defined in some open set $U$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from above (that is, $F(x,t)\leq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))\leq 0.$
A lower semi-continuous function $G:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial_{x}^{2}G)\geq 0$ in the viscosity sense if it has the following property: If $\varphi$ is a smooth function defined in some open set $U$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from below (that is, $F(x,t)\geq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))\geq 0.$
A continuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential equation $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)=0$ in the viscosity sense if it satisfies both $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)\leq 0$ and $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)\geq 0$ in the viscosity sense.
The next definition specifies what is meant by a solution, subsolution, or supersolution of the initial-value problem (57).
Definition 2.
An upper semicontinuous function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ said to be a viscosity subsolution of (57) if it satisfies $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
(58) $\lim_{\delta\downarrow 0}\sup\left\{F(y,0)-F_{\text{in}}(x)\,\mid\,x,y\in\mathbb{R}\,\,\text{such that}\,\,|y-x|\leq\delta\right\}\leq 0.$
A lower semicontinuous function $G:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a viscosity supersolution of (57) if it satisfies $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
$\lim_{\delta\downarrow 0}\inf\left\{G(y,0)-F_{\text{in}}(x)\,\mid\,x,y\in\mathbb{R}\,\,\text{such that}\,\,|y-x|\leq\delta\right\}\geq 0.$
A continuous function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a (continuous) viscosity solution of (57) if it is both a viscosity subsolution and a viscosity supersolution.
Since the main theorem of this work, Theorem 1, involves problems with a discontinuous initial datum, a different definition of solution is needed. In order to state it, it is necessary to recall the notions of upper and lower semicontinuous envelopes. Specifically, for a function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$, the upper and lower semicontinuous envelopes $F^{*}$ and $F_{*}$ are the functions defined in $\mathbb{R}\times[0,\infty)$ via the formulae
$F^{*}(x,t)=\lim_{\delta\downarrow 0}\sup\left\{F(y,s)\,\mid\,y,s\in\mathbb{R}\times[0,\infty)\,\,\text{such that}\,\,|x-y|+|t-s|\leq\delta\right\},$
$F_{*}(x,t)=\lim_{\delta\downarrow 0}\inf\left\{F(y,s)\,\mid\,y,s\in\mathbb{R}\times[0,\infty)\,\,\text{such that}\,\,|x-y|+|t-s|\leq\delta\right\}.$
Notice that $F^{*}$ and $F_{*}$ are always upper and lower semicontinuous, respectively.
Definition 3.
A function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a discontinuous viscosity solution of (57) if $\partial_{t}F^{*}-\mathcal{F}(\partial_{x}F^{*},\partial_{x}^{2}F^{*})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$; $\partial_{t}F_{*}-\mathcal{F}(\partial_{x}F_{*},\partial_{x}^{2}F_{*})\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$; and $F^{*}(\cdot,0)=(F_{\text{in}})^{*}$ and $F_{*}(\cdot,0)=(F_{\text{in}})_{*}$ in $\mathbb{R}$.
A.1. Comparison Principle
This subsection begins by recalling the comparison principle for bounded sub- and supersolutions of the initial-value problem (57), which is standard. An existence and uniqueness result is then stated in terms of the maximal subsolution and minimal supersolution.
Theorem 4 (Comparison Principle).
Assume that the nonlinearity $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ is jointly continuous and nondecreasing in the second argument. If $F,G:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ are respectively a bounded upper and a bounded lower semicontinuous function such that $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)\leq 0$ and $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
$\lim_{\delta\downarrow 0}\sup\left\{F(x,0)-G(y,0)\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\delta\right\}\leq 0,$
then $F\leq G$ pointwise in $\mathbb{R}\times[0,\infty)$.
In particular, if $F$ and $G$ are respectively bounded viscosity sub- and supersolutions of (57) for some initial datum $F_{\text{in}}$, then $F\leq G$ pointwise in $\mathbb{R}\times[0,\infty)$.
The reader is referred to [16, Theorem 2.1], for instance, for the proof. The comparison principle implies that there is at most one continuous viscosity solutions of (57). Existence of the solution is also known in the present setting. However, there is no need to state an existence result here as it actually follows from the main convergence results (see the proof of Theorem 3).
Let $\overline{F}$ and $\underline{F}$ be the maximal sub- and minimal supersolution of (57):
[…]
Notice that the comparison principle implies that $\overline{F}\leq\underline{F}$ and any viscosity solution (continuous or discontinuous) lies in between the two. This immediately implies a uniqueness criterion.
Proposition 9.
Let $\overline{F}$ and $\underline{F}$ be the maximal subsolution and minimal supersolution of (57), respectively. If $\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,\infty)$, then any discontinuous viscosity solution $F$ of (57) must also satisfy $F=\overline{F}$ in $\mathbb{R}\times(0,\infty)$. In particular, in this case, modulo the freedom to choose the value at discontinuities of $F_{\text{in}}$ at $t=0$, there is a unique discontinuous viscosity solution of (57).
Proof.
As in the discussion preceding the statement, if $F$ is any discontinuous viscosity solution, then $\overline{F}\leq F\leq\underline{F}$ in $\mathbb{R}\times[0,\infty)$ in general. Therefore, in the present setting, $F=\overline{F}=\underline{F}$ for $t>0$, implying that there is at most one discontinuous viscosity solution. The fact that at least one exists is similarly immediate: The equality $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,\infty)$ actually implies that both $\overline{F}$ and $\underline{F}$ are discontinuous viscosity solutions. ∎
A.2. Uniqueness for Nondecreasing Data
In this section, the PDE’s of interest here are shown to satisfy the uniqueness criterion of Proposition 9. For the Hamilton-Jacobi equations arising in the $\text{Beta}(2,1)$ case, the fact that almost-everywhere-continuous data leads to a unique solution at positive times is known from [10]. In the case of the second-order equation in the $\text{Beta}(2,2)$ case, the uniqueness of solutions follows from the characterization of nonnegative solutions of the porous medium equation. Short, simplified proofs are presented here for the reader’s convenience.
First, consider the case when $\mathcal{F}$ is a first-order operator of the form
$\mathcal{F}(p,X)=-H(p)$
for some convex or concave function $H:\mathbb{R}\to\mathbb{R}$ exhibiting superlinear growth:
(59) $\lim_{|p|\to\infty}|p|^{-1}H(p)=+\infty.$
At this stage, it is worth noting that there is no loss of generality in assuming that $H$ is convex. Indeed, in general, if $\partial_{t}F+H(\partial_{x}F)=0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$, then the function G(x,t)=1F(x,t)G(x,t)=1-F(-x,t) satisfies
$\partial_{t}G-H(-\partial_{x}G)=0\quad\text{in the viscosity sense in}\,\,\mathbb{R}\times(0,\infty)$
and $H$ is convex if and only if the function $p\mapsto-H(-p)$ is concave.
For superlinear convex/concave $H$, the Hopf-Lax formula furnishes the unique discontinuous viscosity solution of (57).
Proposition 10.
Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. If $\mathcal{F}(p,X)=-H(p)$ for a convex $H:\mathbb{R}\to\mathbb{R}$ satisfying (i) and (ii) above, and if $\overline{F}$ and $\underline{F}$ are the maximal subsolution and minimal supersolution, respectively, of (57), then, for any $(x,t)\in\mathbb{R}\times(0,\infty)$,
(60) $\overline{F}(x,t)=\underline{F}(x,t)=\inf\left\{F_{\text{in}}(y)+tL\left(\frac{x-y}{t}\right)\,\mid\,y\in\mathbb{R}\right\}.$
where $L$ is the Legendre-Fenchel transform of $H$.
Some regularity of $F_{\text{in}}$ is necessary to ensure that $\overline{F}$ and $\underline{F}$ coincide at positive times. For instance, $F_{\text{in}}=\mathbf{1}_{\mathbb{Q}}$ provides a counterexample. An optimal version of the result can be found in [10].
Proof.
When $F_{\text{in}}$ is uniformly Lipschitz continuous, the fact that the right-hand side of (60) determines the (continuous) viscosity solution is classical (see, e.g., [15, Section 10.3.4]). Otherwise, let $\{F^{(N)}_{\text{in}}\}\subseteq CDF(\overline{\mathbb{R}})$ be a sequence of uniformly Lipschitz continuous functions such that
$F^{(N)}_{\text{in}}\leq F^{(N+1)}_{\text{in}}<F_{\text{in}}\quad\text{pointwise for each}\,\,N\in\mathbb{N},$
$\lim_{N\to\infty}F^{(N)}_{\text{in}}(x)=F_{\text{in}}(x)\quad\text{for each point of continuity}\,\,x\,\,\text{of}\,\,F_{\text{in}}.$
If $\{F_{N}\}$ are the corresponding solutions of (57), then the comparison principle implies that $F_{N}\leq F_{N+1}\leq\overline{F}$ for each $N$. By the stability of viscosity solutions, the limit $F=\lim_{N\to\infty}F_{N}$ satisfies $\partial_{t}F-H(\partial_{x}F)=0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$.
At the same time, recall that the superlinearity of $H$ implies that $L$ is also superlinear. Therefore, for any $(x,t)$, it is possible to write $F_{N}(x,t)=F^{(N)}_{\text{in}}(y_{N})+tL((x-y_{N})/t)$ for some bounded sequence $\{y_{N}\}$. If $y_{N}\to y$ as $N\to\infty$ (say, along a subsequence), then
$F_{\text{in}}(y-\delta)=\lim_{N\to\infty}F_{N}(y-\delta)\leq\liminf_{N\to\infty}F^{(N)}_{\text{in}}(y_{N})\quad\text{for almost every}\,\,\delta>0,$
from which it follows that
$\inf\left\{F_{\text{in}}(y)+tL\left(\frac{x-y}{t}\right)\,\mid\,y\in\mathbb{R}\right\}\leq\lim_{N\to\infty}F_{N}(x,t)=F(x,t)\leq\overline{F}(x,t).$
$\underline{F}$ can be treated similarly: Let $\{F^{(N)}_{\text{in}}\}\subseteq CDF(\overline{\mathbb{R}}$ be uniformly Lipschitz functions such that $F_{\text{in}}<F^{(N+1)}_{\text{in}}\leq F^{(N)}_{\text{in}}$ for each $N$ and $F_{\text{in}}(x)=\lim_{N\to\infty}F^{(N)}_{\text{in}}(x)$ at points of continuity $x$ of $F_{\text{in}}$. This time, if $F_{N}$ is the corresponding solution of (57), then, for any point of continuity $y$ of $F_{\text{in}}$ and any $(x,t)\in\mathbb{R}\times(0,\infty)$,
$\underline{F}(x,t)\leq\lim_{N\to\infty}F_{N}(x,t)\leq\lim_{N\to\infty}F_{\text{in}}^{(N)}(y)+tL(\frac{x-y}{t})=F_{\text{in}}(y)+tL(\frac{x-y}{t}).$
Since $F$ is right-continuous and points of continuity are dense, this proves $\underline{F}$ is bounded above by the right-hand side of (60). ∎
Next, consider the case of the nonlinearity
(61) $\mathcal{F}(p,X)=a|p|^{m-1}X$
Notice that if $F$ were a smooth solution of $\partial_{t}F-a|\partial_{x}F|^{m-1}\partial_{x}^{2}F=0$, then the derivative $\rho=\partial_{x}F$ would solve the porous medium equation (PME) $\partial_{t}\rho=\frac{a}{m}\partial_{x}^{2}(\rho^{m})$. To prove the uniqueness of solutions with possibly discontinuous, nondecreasing data, this reasoning will be inverted.
Proposition 11.
Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. If $\mathcal{F}$ has the form (61) for some $m>1$, then the maximal subsolution $\overline{F}$ and minimal supersolution $\underline{F}$ of (57) are given for $(x,t)\in\mathbb{R}\times(0,\infty)$ by the formula
$\overline{F}(x,t)=\underline{F}(x,t)=\int_{-\infty}^{x}\rho(y,t)\,dy,$
where $\rho\geq 0$ is the unique continuous distributional solution of the initial-value problem
(64) $\left\{\begin{array}[]{r l}\partial_{t}\rho-\frac{a}{m}\partial_{x}^{2}(\rho^{m})=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ \rho(0)=\partial_{x}F_{\text{in}}.\end{array}\right$.
When $F_{\text{in}}$ is uniformly Lipschitz, this result is a special case of [3, Theorem 3.1].
Unless otherwise noted, basic properties of the PME used in the proof that follows can be found in the monograph [21].
Proof.
First, assume that $F_{\text{in}}$ is smooth and $\partial_{x}F_{\text{in}}$ is bounded and nowhere vanishing. In this case, classical results on the PME imply that there is a unique smooth solution $\rho$ of (64), which is positive and satisfies $\int_{-\infty}^{\infty}\rho(y,t)\,dy=F_{\text{in}}(+\infty)-F_{\text{in}}(-\infty)$ for all $t>0$. Since $\rho$ is smooth, the function $F(x,t)=\int_{-\infty}^{x}\rho(y,t)\,dy$ is also, and, thus, it satisfies $\partial_{t}F-a|\partial_{x}F|^{m-1}\partial_{x}^{2}F=\partial_{t}F-\frac{a}{m}\partial_{x}(\partial_{x}F)^{m}=0$ classically. As is well-known, being a classical solution, it is also a viscosity solution.
If $F_{\text{in}}$ is merely absolutely continuous with $\partial_{x}F_{\text{in}}\in L^{1}(\mathbb{R})$, nonetheless $\partial_{x}F_{\text{in}}$ can be approximated in $L^{1}(\mathbb{R})$ by smooth, nowhere-vanishing functions. The $L^{1}$ contraction property of the PME then implies that the corresponding solutions converge in $L^{\infty}([0,\infty);L^{1}(\mathbb{R}))$, which implies uniform convergence of the antiderivatives. Thus, since the viscosity solution property is stable under uniform convergence, even if $\partial_{x}F_{\text{in}}$ is merely in $L^{1}(\mathbb{R})$, the unique viscosity solution of (57) in this case is still given as the antiderivative of $\rho$.
Finally, it remains to consider the case when $\partial_{x}F_{\text{in}}$ is only a finite positive measure (so that $F_{\text{in}}$ may have discontinuities). Let $\{F^{(N)}_{\text{in}}\}\subseteq CDF(\overline{\mathbb{R}})$ be smooth functions such that $F^{(N)}_{\text{in}}\leq F^{(N+1)}_{\text{in}}<F_{\text{in}}$ for each $N$, $F^{(N)}_{\text{in}}\to F_{\text{in}}$ at points of continuity as $N\to\infty$. If $\rho_{N}$ is the corresponding solution of the PME (64) and $F_{N}$, its antiderivative, then $F_{N}\leq\overline{F}$ for each $N$ since $F_{N}$ is a viscosity subsolution of (57).
At the same time, some subsequence of $\{\rho_{N}\}$ converges locally uniformly in $\mathbb{R}\times(0,\infty)$ as $N\to\infty$ to a continuous, nonnegative distributional solution $\rho$ of (64) with initial trace equal to the measure $\partial_{x}F_{\text{in}}$; this follows, for instance, from the fact that, in this one-dimensional setting, the pressure $\{\rho_{N}^{m-1}\}$ is both uniformly bounded and uniformly Lipschitz in $\mathbb{R}\times[\tau,+\infty)$ for any $\tau>0$, independently of $N$ (see [21, Chapter 14]). A classical result of Dahlberg and Kenig [14] implies that $\rho$ is unique, hence the entire sequence $\{\rho_{N}\}$ converges to $\rho$. At the level of antiderivatives, this implies $F_{N}\to F$ as $N\to\infty$, where $F$ is given by (64) and is a discontinuous viscosity solution of (57). By comparison, $\overline{F}\leq F$, hence actually $F=\overline{F}$.
Approximating $F_{\text{in}}$ from above, one similarly finds $F=\underline{F}$. ∎
A.3. Viscosity Solutions in $CDF(\overline{\mathbb{R}})$
Recall that throughout this work, the functions of interest belong to the space $CDF(\overline{\mathbb{R}})$ of CDF’s of extended-real-valued random variables. (Recall that this is nothing other than the space of nondecreasing, right-continuous functions taking values in $[0,1]$.) This section covers some specific simplifications of viscosity solutions theory that are applicable as long as the sub- or supersolutions in question are nondecreasing in the spatial variable. Specifically, it shows that, in this case, there is no loss of generality assuming that the test functions involved are themselves nondecreasing in space and defined globally in space.
The first lemma concerns the case when a subsolution is touched above by a test function that is increasing close to the touching point.
Lemma 1.
Fix $T,A>0$ and let $F:\mathbb{R}\times(0,T)\to[0,A)$ be an upper semicontinuous function that is nondecreasing in the spatial variable. Fix $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ and let $\varphi$ be a smooth function defined in some neighborhood of $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ such that $\varphi(x_{0},t_{0})<A$ and $\partial_{x}\varphi(x_{0},t_{0})>0$. If $\varphi$ touches $F$ from above at $(x_{0},t_{0})$, then there is an $r>0$, a nondecreasing function $\psi_{1}:\mathbb{R}\to[0,1]$ with $\psi_{1}(x)=\psi_{1}(x_{0}-r)$ for each $x\leq x_{0}-r$ and $\psi(x)=A$ for each $x\geq x_{0}+r$, and a $C^{2}$ function $\psi_{2}:\mathbb{R}\to\mathbb{R}$ with $\psi_{2}(t_{0})=0$ such that if $\psi$ is the function
$\psi(x,t)=\psi_{1}(x)+\psi_{2}(t),$
then
$F(x,t)\leq\psi(x,t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r]$
and equality holds if and only if $(x,t)=(x_{0},t_{0})$.
Proof.
By Taylor expansion, there is an $r>0$ and degree two polynomials $R$ and $Q$ such that $Q(0)=0$ and
$F(x,t)\leq\varphi(x,t)\leq R(x)+Q(t-t_{0})\quad\text{for each}\,\,(x,t)\in[x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r]$
Adding a term of the form $(x-x_{0})^{2}+(t-t_{0})^{2}$ to $R+Q$, there is no loss of generality in assuming in addition that $\varphi(x,t)=R(x)+Q(t-t_{0})$ if and only if $(x,t)=(x_{0},t_{0})$. Since $0<\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}R(x_{0})$, after possibly making $r$ smaller, it can similarly be assumed that $\partial_{x}R(x)>0$ for each $x\in[x_{0}-2r,x_{0}+2r]$.
Given $a>0$, let $\rho:\mathbb{R}\to\mathbb{R}$ be a smooth function such that $\rho(x)=0$ if $|x-x_{0}|\leq\frac{a}{2}$ and $\rho(x)=1$ if $|x-x_{0}|\geq a$. Define $\psi_{1}$ by
$\psi_{1}(x)=(1-\rho(x))R(x)+A\rho(x)\mathbf{1}_{[x_{0},+\infty)}(x)+R(x_{0}-\frac{a}{2})\rho(x)\mathbf{1}_{(-\infty,x_{0}]}(x).$
Since $R(x_{0})=\varphi(x_{0},t_{0})<A$, $\psi_{1}$ is a nondecreasing function provided $a$ is small enough. Fix such a small $a$ henceforth and let $\psi_{2}(t)=Q(t-t_{0})$. Since $F$ is nondecreasing in the spatial variable and $F(x,t)\leq\varphi(x,t)$ for each $(x,t)\in[x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r]$, the function $\psi(x,t)=\psi_{1}(x)+\psi_{2}(t)$ has the desired properties. ∎
The next lemma concerns the case when a subsolution is touched above by a test function that is flat close to the touching point.
Lemma 2.
Fix $T,A>0$ and let $F:\mathbb{R}\times(0,T)\to[0,A)$ be an upper semicontinuous function that is nondecreasing in the spatial variable. Assume that $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ and $\varphi$ is a smooth function defined in some neighborhood of $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ such that $\varphi(x_{0},t_{0})<A$ and $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$. If $\varphi$ touches $F$ from above at $(x_{0},t_{0})$, then there is an $r>0$, a nondecreasing $C^{2}$ function $\psi_{1}:\mathbb{R}\to[0,1]$ with $\psi(x)=\psi(x_{0})$ for each $x\leq x_{0}$ and $\psi(x)=A$ for each $x\geq x_{0}+r$, and a $C^{2}$ function $\psi_{2}:[t_{0}-r,t_{0}+r]\to\mathbb{R}$ with $\psi_{2}(t_{0})=0$ such that if $\psi$ is the function
$\psi(x,t)=\psi_{1}(x)+\psi_{2}(t),$
then
$F(x,t)\leq\psi(x,t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r],\quad\text{and}$
$(x_{0},t_{0})\in\{(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r]\,\mid\,F(x,t)=\psi(x,t)\}\subseteq(-\infty,x_{0}]\times\{t_{0}\}.$
Proof.
For concreteness, choose $r>0$ such that $F(x,t)\leq\varphi(x,t)$ for each $(x,t)\in[x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r]$ with equality when $(x,t)=(x_{0},t_{0})$. Expanding $\varphi$ in a fourth-order Taylor expansion around $(x_{0},t_{0})$, one finds
$\varphi(x,t)$ $=\varphi(x_{0},t_{0})+Q(t-t_{0})+\frac{1}{3!}\partial^{3}_{x}\varphi(x_{0},t_{0})(x-x_{0})^{3}$
$\quad+\frac{1}{4!}\partial^{4}_{x}\varphi(x_{0},t_{0})(x-x_{0})^{4}+o(|x-x_{0}|^{4}+|t-t_{0}|^{4}),$
where $Q$ is some polynomial of degree four satisfying $Q(0)=0$. Since $F$ is nondecreasing in the $x$ variable; $F(x_{0},t_{0})=\varphi(x_{0},t_{0})$; and $F(x,t)\leq\varphi(x,t)$ for $(x,t)$ close enough to $(x_{0},t_{0})$, it follows that
$\partial^{3}\varphi(x_{0},t_{0})\geq 0.$
Let $\alpha=\max\{1+\partial_{x}^{4}\varphi(x_{0},t_{0}),0\}$. Making $r$ smaller if necessary, there is no loss of generality in assuming that, for each $(x,t)\in[x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r]$,
$\varphi(x,t)$ $\leq\varphi(x_{0},t_{0})+Q(t-t_{0})+\frac{(t-t_{0})^{4}}{4!}+\frac{1}{3!}\partial_{x}^{3}\varphi(x_{0},t_{0})(x-x_{0})^{3}+\frac{1}{4!}\alpha(x-x_{0})^{4}$
$=:P(x,t).$
with equality if and only if $(x,t)=(x_{0},t_{0})$. $P$ has the desired form, but it is not nondecreasing in $x$. This will be rectified in the remainder of the proof.
Define the nondecreasing function $\eta$ by
$\eta(x)=\left\{\begin{array}[]{r l}\varphi(x_{0},t_{0})+\frac{1}{3!}\partial_{x}^{3}\varphi(x_{0},t_{0})(x-x_{0})^{3}+\frac{1}{4!}\alpha(x-x_{0})^{4},&\text{if}\,\,x_{0}\leq x\\ \varphi(x_{0},t_{0}),&\text{otherwise.}\end{array}\right$.
Notice that $\eta$ is $C^{2}$. Given $a\leq r$, let $\rho:\mathbb{R}\to[0,1]$ be a smooth, nondecreasing function such that $\rho(x)=0$ if $x\leq x_{0}+\frac{a}{2}$ and $\rho(x)=1$ if $x\geq x_{0}+a$. Define $h$ by $h(x)=(1-\rho(x))\eta(x)+A\rho(x)$. Since $\eta(x_{0})=\varphi(x_{0},t_{0})<A$, $h$ is a nondecreasing function provided $a$ is small enough. Fix such a small $a$ henceforth and let $\psi$ be the modified version of $P$ given by
$\psi(x,t)=h(x)+Q(t-t_{0})+\frac{(t-t_{0})^{4}}{4!}.$
Since $F\leq\varphi\leq P$ in $x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r]$, $F$ is nondecreasing in the $x$ variable, and $F<A$ globally, it follows that $F\leq\psi$ in $\mathbb{R}\times[t_{0}-r,t_{0}+r]$ with equality only if $x\leq x_{0}$ and $t=t_{0}$. Therefore, the desired conclusion follows with $\psi_{1}(x)=h(x)$ and $\psi_{2}(t)=Q(t-t_{0})$. ∎
Appendix B Scaling Limits of Monotone Semigroups in $CDF(\overline{\mathbb{R}})$
This appendix proves Theorem 3, the extension of the convergence framework of [6] to approximations of PDE’s posed in $CDF(\overline{\mathbb{R}})$. As described already briefly in the introduction, technical issues arise stemming from the fact that $CDF(\overline{\mathbb{R}})$ is not a vector space, and the smooth test functions used in the definition of viscosity solution are not necessarily nondecreasing. These issues are not fatal and, indeed, similar ones have been overcome in other settings, such as geometric flows (see [9]).
In the proof of the theorem, there is one technicality that needs to be overcome. When one checks that, for instance, the limit $F$ satisfies $\partial_{t}F\leq\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)$, it is desirable to invoke the asymptotics of $\mathscr{L}_{\epsilon}\varphi$ for an arbitrary smooth test function $\varphi$ that touches $F$ from above. If $\partial_{x}\varphi>0$ at the contact point, then Lemma 1 above shows that $\varphi$ can be assumed to be nondecreasing in $\mathbb{R}$ and then $\mathscr{L}_{\epsilon}\varphi$ is well-defined. Otherwise, if $\partial_{x}\varphi=0$ at the contact point, of course, $\varphi$ generally will not be monotone at all, and some work is required to proceed.
At the same time, smooth nondecreasing functions have a particular property that is suggestive here: If $\Phi$ is a smooth nondecreasing function in $\mathbb{R}$, and if $\partial_{x}\Phi(x_{0})=0$ at some point $x_{0}$, then $\partial_{x}^{2}\Phi(x_{0})=0$. This motivates the following lemma (which applies even if the subsolution is not monotone in the spatial variable).
Lemma 3.
Let $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be a continuous function that is nondecreasing in the second argument. An upper semicontinuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ satisfies $\partial_{t}F\leq\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)$ in the viscosity sense if and only if it has the following property: If $\varphi$ is a smooth test function that touches $F$ from above at some point $(x_{0},t_{0})$ and
$\text{either (i)}\quad\partial_{x}\varphi(x_{0},t_{0})\neq 0,\quad\text{or (ii)}\quad\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0,$
then
$\partial_{t}\varphi(x_{0},t_{0})\leq\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0})).$
Proof.
The fact that this weaker definition of viscosity subsolution is equivalent to the usual one can be proved exactly as in [8, Proposition 2.2] (see also [20, Proof of Theorem 1.2]). ∎
Remark 3.
A corresponding version of Lemma 3 holds for supersolutions, which can be deduced from the transformation G(x,t)=1F(x,t)G(x,t)=1-F(-x,t).
Now Lemma 3 can be leveraged in conjunction with Lemma 2 above to resolve the issue when $\partial_{x}\varphi=0$ at the contact point: There is no loss of generality assuming that $\partial_{x}^{2}\varphi=0$ also holds there, and then the lemma shows that $\varphi$ can be taken to be globally nondecreasing in the spatial variable.
B.1. Proof of Theorem 3
Since $F_{\epsilon}$ is discontinuous in time, it will clarify the picture somewhat to observe that the statement that $F_{\epsilon}\to F$ locally uniformly in $\mathbb{R}\times[0,\infty)$ is equivalent to the following: For each $(x,t)\in\mathbb{R}\times[0,\infty)$,
(65) $\lim_{\delta\downarrow 0}\sup\left\{|F^{\epsilon}_{n}(y)-F(x,t)|\,\mid\,|y-x|+|\epsilon^{\alpha}n-t|\leq\delta\right\}=0.$
This is the precise statement that will be proved below.
Since the operators $(T_{\epsilon})_{\epsilon>0}$ are monotone and nondecreasing continuous functions can be approximated locally uniformly from above and below by smooth functions, it suffices to consider the case when $F^{\epsilon}_{0}=F_{\text{in}}$ for each $\epsilon>0$ and $F_{\text{in}}$ has a bounded, uniformly continuous second derivative. This will be assumed for the remainder of the proof.
Throughout what follows, define the constants $c_{-},c_{+}\in[0,1]$ by
(66) $c_{-}=F_{\text{in}}(-\infty)=F_{\epsilon}(+\infty,t),\quad c_{+}=F_{\text{in}}(+\infty)=F_{\epsilon}(-\infty,t),$
where property (iii) implies that the identities above hold independently of $t>0$.
Step 1: Lipschitz Estimate in Time. Since $F_{\text{in}}$ is currently assumed to have a bounded, uniformly continuous second derivative, by the consistency condition (iv),
$\epsilon^{-\alpha}\mathscr{L}_{\epsilon}F_{0}^{\epsilon}=\epsilon^{-\alpha}\mathscr{L}_{\epsilon}F_{\text{in}}=\mathcal{F}(\partial_{x}F_{\text{in}},\partial^{2}_{x}F_{\text{in}})+R_{\epsilon}$
where $R_{\epsilon}(x)\to 0$ uniformly with respect to $x$ as $\epsilon\to 0$. Thus, since $\mathcal{F}$ is bounded on compact sets by continuity, there is a constant $C_{0}>0$ such that, for sufficiently small $\epsilon$,
$\|T_{\epsilon}F_{0}^{\epsilon}-F_{0}^{\epsilon}\|_{\text{sup}}\leq C_{0}\epsilon^{\alpha}\quad\text{in}\,\,\mathbb{R}.$
In Proposition 12 below, the assumptions (i)-(iii) are shown to imply the following contractivity property of $T_{\epsilon}$:
$\|T_{\epsilon}F-T_{\epsilon}G\|_{\text{sup}}\leq\|F-G\|_{\text{sup}}.$
Thus, the estimate of $T_{\epsilon}F_{0}^{\epsilon}-F_{0}^{\epsilon}$ may be iterated to find
$\|F^{\epsilon}_{n+1}-F_{n}^{\epsilon}\|_{\text{sup}}=\|T_{\epsilon}F^{\epsilon}_{n}-F_{n}^{\epsilon}\|_{\text{sup}}\leq C_{0}\epsilon^{\alpha}\quad\text{for each}\,\,n\in\mathbb{N}_{0}.$
Note, in particular, that $|F^{\epsilon}_{n}(x)-F_{\text{in}}(x)|\leq C_{0}\epsilon^{\alpha}n$ uniformly in $(x,n)$ as $\epsilon\to 0$.
Step 3: Convergence. The proof uses the classical half-relaxed limit method. Let $F_{\star}$ and $F^{\star}$ be the half-relaxed limits of $\{F^{\epsilon}_{n}\}$ defined by
$F_{\star}(x,t)$ $=\lim_{\delta\downarrow 0}\inf\left\{F_{n}^{\epsilon}(y)\,\mid\,|x-y|+|\epsilon^{\alpha}n-t|+\epsilon\leq\delta\right\},$
$F^{\star}(x,t)$ $=\lim_{\delta\downarrow 0}\sup\left\{F_{n}^{\epsilon}(y)\,\mid\,|x-y|+|\epsilon^{\alpha}n-t|+\epsilon\leq\delta\right\},$
To prove the local convergence statement (65), it suffices to show that $F^{\star}=F_{\star}=F$ (see [13, Remark 6.4] or [7, Lemma 6.2]). To do this, since $F_{\star}\leq F^{\star}$ is immediate, it is only necessary to show that $F^{\star}\leq F\leq F_{\star}$. This will follow from the comparison principle, Theorem 4, as soon as it is established that $F^{\star}$ and $F_{\star}$ define respectively a viscosity sub- and a viscosity supersolution of the initial value problem (57).
Step 4: $F^{\star}$ is a Subsolution. In view of what was proved in Step 1, it is clear that $F^{\star}(\cdot,0)=F_{\text{in}}$ in the sense of (58). Therefore, it only remains to establish that $\partial_{t}F^{\star}-\mathcal{F}(\partial_{x}F^{\star},\partial_{x}^{2}F^{\star})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$.
In order to avoid boundary issues that can arise if $F^{\star}$ is somewhere equal to one, it will be convenient to pull $F^{\star}$ down: To see that $F^{\star}$ is a subsolution, it suffices to prove that, for any $\delta\in(0,c_{+}-c_{-})$, the function $G^{\star}:\mathbb{R}\times[0,\infty)\to[c_{-},c_{+}-\delta]$ defined by $G^{\star}=\max\{c_{-},F^{\star}-\delta\}$ satisfies
$\partial_{t}G^{\star}\leq\mathcal{F}(\partial_{x}G^{\star},\partial_{x}^{2}G^{\star}).$
For the rest of this step, fix $\delta$ and consider $G^{\star}$. Note also that if $\{G^{\epsilon}_{n}\}$ is the sequence defined for $\epsilon>0$ by $G^{\epsilon}_{n}=\max\{c_{-},F^{\epsilon}_{n}-\delta\}$, then Lemma 4 below implies that
(67) $G^{\epsilon}_{n+1}\leq T_{\epsilon}G^{\epsilon}_{n}\quad\text{pointwise for each}\,\,n\in\mathbb{N}_{0}.$
Further, by definition of $G^{\star}$,
(68) $G^{\star}(x,t)=\lim_{\delta\downarrow 0}\sup\left\{G^{\epsilon}_{n}(y)\,\mid\,|x-y|+|\epsilon^{\alpha}n-t|+\epsilon\leq\delta\right\}.$
Thus, in passing from $F$’s to $G$’s, nothing has changed except the tighter upper bound $G^{\epsilon}_{n}\leq c_{+}-\delta<1$ holds for each $n$ and the evolution equation has been replaced by an inequality in (67).
Suppose that $\varphi$ is a smooth function defined in some open subset of space-time that touches $G^{\star}$ from above at some point $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ in its domain. The goal is to show that $\partial_{t}\varphi(x_{0},t_{0})\leq\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))$. Notice that one technicality can be ruled out immediately: Since $G^{\star}\geq c_{-}$ pointwise, if $\varphi(x_{0},t_{0})=c_{-}$, then
$\partial_{t}\varphi(x_{0},t_{0})=\partial_{x}\varphi(x_{0},t_{0})=0,\quad\partial_{x}^{2}\varphi(x_{0},t_{0})\geq 0,$
which implies the desired inequality by (51). Therefore, from now on, assume that $c_{-}<\varphi(x_{0},t_{0})$.
According to Lemma 3, there is no loss of generality in assuming that either (a) $|\partial_{x}\varphi(x_{0},t_{0})|>0$ or (b) $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$. In case (a), since $G^{\star}$ is nondecreasing, it follows that $\partial_{x}\varphi(x_{0},t_{0})>0$, and Lemma 1 is applicable. In case (b), Lemma 2 applies. To avoid repetition, only case (b) will be considered in what follows.
Therefore, assume henceforth that (b) holds. According to Lemma 2 with $A=c_{+}-\frac{\delta}{2}$, there is no loss of generality in fixing an $r>0$ and assuming that $\varphi(x,t)=\varphi_{1}(x)+\varphi_{2}(t)$ for $(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r]$, where $\varphi_{1}$ is a nondecreasing $C^{2}$ function with $\varphi_{1}(x)=\varphi_{1}(x_{0})$ for $x\leq x_{0}$ and $\varphi_{1}(x)=c_{+}-\frac{\delta}{2}$ for $x\geq x_{0}+r$; $\varphi_{2}$ is $C^{2}$ in $t_{0}-r,t_{0}+r]$ with $\varphi_{2}(t_{0})=0$; and
$G^{\star}(x,t)\leq\varphi(x,t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r],$
$(x_{0},t_{0})\in\{(x,t)\,\mid\,G^{\star}(x,t)=\varphi(x,t)\}\subseteq(-\infty,x_{0}]\times\{t_{0}\}.$
Since $\varphi_{2}$ is continuous in $t_{0}-r,t_{0}+r]$ and $\varphi_{1}(-\infty)=\varphi(x_{0},t_{0})>c_{-}$, up to shrinking $r$ if necessary, there is no loss of generality in assuming that there are constants $c_{*},c^{*}\in[0,1]$ such that, for each $t\in[t_{0}-r,t_{0}+r]$,
(69) $c_{-}<c_{*}\leq\varphi(-\infty,t)\quad\text{and}\quad c_{+}-\delta<c_{*}\leq\varphi(+\infty,t)\leq c_{+}-\frac{\delta}{4}.$
Finally, for each $\epsilon>0$, let $c(\epsilon)$ be the constant
$c(\epsilon)=\sup\left\{G^{\epsilon}_{n}(x)-\varphi(x,\epsilon^{\alpha}n)\,\mid\,x\in\mathbb{R},\,\,n\in\mathbb{N}\cap\epsilon^{-\alpha}[t_{0}-r,t_{0}+r]\right\}.$
Notice that (69) and (66) imply that there is an $R>0$ such that, for any $\epsilon>0$,
$\sup\left\{G^{\epsilon}_{n}(x)-\varphi(x,\epsilon^{\alpha}n)\,\mid\,|x|\geq R,\,\,n\in\mathbb{N}\cap\epsilon^{-\alpha}[t_{0}-r,t_{0}+r]\right\}<0.$
Therefore, by the equation (68) relating $G^{\epsilon}_{n}$ to $G^{\star}$, it follows that there is a sequence $\{\epsilon_{k},x_{k},n_{k}\}$ such that $(\epsilon_{k},c(\epsilon_{k}))\to(0,0)$ as $k\to\infty$; $|x_{k}|\leq R$ and $|\epsilon_{k}^{\alpha}n_{k}-t|\leq r$ for each $k$; and
(70) $G^{\epsilon_{k}}_{n_{k}}(x_{k})-\varphi(x_{k},\epsilon_{k}^{\alpha}n_{k})-c(\epsilon_{k})=0.$
Further, since $\{(x,t)\,\mid\,G^{\star}(x,t)=\varphi(x,t)\}$ is contained in $(-\infty,x_{0}]\times\{t_{0}\}$, there is no loss of generality in assuming that there is a $y_{0}\leq x_{0}$ such that $(x_{k},\epsilon^{\alpha}_{k}n_{k})\to(y_{0},t_{0})$ and $G^{\star}(y_{0},t_{0})=\varphi(y_{0},t_{0})$. Note, in particular, that the inequality $y_{0}\leq x_{0}$ implies
(71) $\lim_{k\to\infty}(\partial_{x}\varphi_{1}(x_{k}),\partial_{x}^{2}\varphi_{1}(x_{k}))=(0,0).$
Finally, for notational ease, let $\Phi_{k}:\mathbb{R}\times[t_{0}-r,t_{0}+r]\to\mathbb{R}$ be the function $\Phi_{k}(x,t)=\varphi(x,t)+c(\epsilon_{k})$. Since $c(\epsilon_{k})\to 0$, the bounds (69) implies that $0\leq\Phi_{k}\leq 1$ pointwise for all $k$ large enough. That is, $\Phi_{k}\in CDF(\overline{\mathbb{R}})$ for large enough $k$. By the choice of $c(\epsilon_{k})$,
$G^{\epsilon_{k}}_{n}\leq\Phi_{k}(\cdot,\epsilon^{\alpha}_{k}n)\quad\text{pointwise for each}\,\,n\in\mathbb{N}\cap\epsilon_{k}^{-\alpha}[t_{0}-r,t_{0}+r].$
Therefore, by the monotonicity of $T_{\epsilon}$,
$T_{\epsilon_{k}}G^{\epsilon_{k}}_{n_{k}-1}\leq T_{\epsilon_{k}}\Phi_{\epsilon_{k}}(\cdot,\epsilon^{\alpha}_{k}n_{k}-\epsilon^{\alpha}_{k}).$
By (67), (70), and the definition of $\Phi_{k}$, this implies
$\Phi_{k}(x_{k},\epsilon^{\alpha}_{k}n_{k})=G^{\epsilon_{k}}_{n_{k}}(x_{k})\leq T_{\epsilon_{k}}G^{\epsilon_{k}}_{n_{k}-1}(x_{k})\leq T_{\epsilon_{k}}\Phi_{k}(x_{k},\epsilon^{\alpha}_{k}n_{k}-\epsilon_{k}^{\alpha}).$
Upon rearranging, this becomes
$\epsilon_{k}^{-\alpha}\Big(\Phi_{\epsilon_{k}}(x_{k},\epsilon_{k}^{\alpha}n_{k})-\Phi_{k}(x_{k},\epsilon^{\alpha}_{k}n_{k}-\epsilon^{\alpha}_{k})\Big)-\epsilon_{k}^{-\alpha}(\mathscr{L}_{\epsilon_{k}}\Phi_{k})(x_{k},\epsilon^{\alpha}_{k}n_{k}-\epsilon_{k}^{\alpha})\leq 0.$
By assumption (ii),
$\mathscr{L}_{\epsilon_{k}}\Phi_{k}(x_{k},\epsilon^{\alpha}_{k}n_{k}-\epsilon_{k}^{\alpha})=\mathscr{L}_{\epsilon_{k}}\varphi_{1}(x_{k}),$
hence, by assumption (iv), (71), and the form of $\varphi$, sending $k\to\infty$ yields
$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(0,0)=\partial_{t}\varphi_{2}(t_{0})-\lim_{k\to\infty}\mathcal{F}(\partial_{x}\varphi_{1}(x_{k}),\partial_{x}^{2}\varphi_{1}(x_{k}))\leq 0.$
Since $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$ by assumption, this is the desired inequality.
Step 5: $F_{\star}$ is a Supersolution. The proof that $F_{\star}$ is a supersolution of (57) can be reduced to the subsolution property of $F^{\star}$ after a suitable transformation. As in Remark 2 above, let $G_{-}$ denote the left-continuous version of a nondecreasing function $G$, and define an involution $G\mapsto G_{\text{rev}}$ on $CDF(\overline{\mathbb{R}})$ via the formula
$G_{\text{rev}}(x)=1-G_{-}(-x).$
Let $\{\tilde{T}_{\epsilon}\}$ be the operators on $CDF(\overline{\mathbb{R}})$ obtained from $\{T_{\epsilon}\}$ by
$\tilde{T}_{\epsilon}F=(T_{\epsilon}(F_{\text{rev}}))_{\text{rev}}.$
It is straightforward to check that the family $\{\tilde{T}_{\epsilon}\}$ satisfies assumptions (i)-(iv), the only difference being $\mathcal{F}$ has to be replaced by the function $\tilde{\mathcal{F}}(p,X)=-\mathcal{F}(p,-X)$.
Define $\{\tilde{F}^{\epsilon}_{n}\}$ by the rule $\tilde{F}^{\epsilon}_{n}=(F^{\epsilon}_{n})_{\text{rev}}$. Notice that the recursion $F^{\epsilon}_{n+1}=T_{\epsilon}F^{\epsilon}_{n}$ implies that $\tilde{F}^{\epsilon}_{n+1}=\tilde{T}_{\epsilon}\tilde{F}^{\epsilon}_{n}$. Therefore, if $\tilde{F}^{\star}$ is defined by
$\tilde{F}^{\star}(x,t)=\lim_{\delta\downarrow 0}\sup\left\{\tilde{F}^{\epsilon}_{n}(y)\,\mid\,|y-x|+|\epsilon^{\alpha}n-t|+\epsilon\leq\delta\right\},$
then $\tilde{F}^{\star}$ is a subsolution of the initial value problem (57) with $\mathcal{F}$ replaced by $\tilde{\mathcal{F}}$ and $F_{\text{in}}$ by $(F_{\text{in}})_{\text{rev}}$. Further, by definition, $\tilde{F}^{\star}(x,t)=1-F_{\star}(-x,t)$. From these last two observations, one readily concludes that $F_{\star}$ is a viscosity supersolution of (57).∎
B.2. Proof of Corollary 3
In the proof that follows, it it convenient to work with the half-relaxed limits $\liminf_{*}F_{\epsilon}$ and $\limsup_{*}F_{\epsilon}$ of the rescaled CDF $F_{\epsilon}$ defined by
$(\liminf\nolimits_{*}F_{\epsilon})(x,t)$ $=\lim_{\delta\downarrow 0}\inf\left\{F_{\epsilon}(y,s)\,\mid\,|y-x|+|s-t|+\epsilon\leq\delta\right\},$
$(\limsup\nolimits_{*}F_{\epsilon})(x,t)$ $=\lim_{\delta\downarrow 0}\sup\left\{F_{\epsilon}(y,s)\,\mid\,|y-x|+|s-t|+\epsilon\leq\delta\right\}.$
Proof of Corollary 3.
Fix $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. Suppose that $F_{\epsilon}(x,0)\to F_{\text{in}}(x)$ at each point of continuity $x$ of $F_{\text{in}}$. Let $\overline{F}$ and $\underline{F}$ be the maximal subsolution and the minimal supersolution of the problem (57). In view of Proposition 9 and well-known properties of half-relaxed limits (see [13, Remark 6.4] or [7, Lemma 6.2]), to see that $F_{\epsilon}$ converges locally uniformly to the discontinuous viscosity solution, it suffices to establish that
$\overline{F}\leq\liminf\nolimits_{*}F_{\epsilon}\leq\limsup\nolimits^{*}F_{\epsilon}\leq\underline{F}\quad\text{in}\,\,\mathbb{R}\times(0,\infty).$
The details will be provided only for the lower bound as the upper bound follows similarly.
For each $N\in\mathbb{N}$, define $F_{\epsilon,N}$ like $F_{\epsilon}$, albeit instead with the initial datum
$F_{\epsilon,N}(x,0)=2^{N}\int_{0}^{\infty}F_{\epsilon}(x-y,0)\rho(2^{N}y)\,dy,$
where $\rho:(0,\infty)\to(0,\infty)$ is a smooth function supported in $[1,2]$ with $\int_{0}^{\infty}\rho(y)\,dy=1$. From the fact that $\rho$ is supported in $[1,2]$ and $F_{\epsilon}$ is nondecreasing in $x$, it follows that $F_{\epsilon,N}(x,0)\leq F_{\epsilon,N+1}(x,0)\leq F_{\epsilon}(x,0)$ for every $x\in\mathbb{R}$. Due to the monotonicity assumption (i), this implies that $F_{\epsilon,N}\leq F_{\epsilon,N+1}\leq F_{\epsilon}$ in $\mathbb{R}\times[0,\infty)$ for each $\epsilon>0$ and $N\in\mathbb{N}$. By Theorem 3, $F_{\epsilon,N}\to F_{N}$ as $\epsilon\to 0$, where $F_{N}$ is the solution of (57) with initial datum $F_{\text{in}}^{(N)}$ given by $F_{\text{in}}^{(N)}(x)=2^{N}\int_{0}^{\infty}F_{\text{in}}(x-y)\rho(2^{N}y)\,dy$. Note that $F_{N}\leq\liminf\nolimits_{*}F_{\epsilon}$ pointwise in $\mathbb{R}\times[0,\infty)$ for any fixed $N$.
Standard arguments (see [7, Theorem 6.2]) show that the function $F_{\infty}$ defined by
$F_{\infty}(x,t)=\lim_{\delta\downarrow 0}\inf\left\{F_{N}(y,s)\,\mid\,|y-x|+|t-s|+N^{-1}\leq\delta\right\}$
satisfies $\partial_{t}F_{\infty}-\mathcal{F}(\partial_{x}F_{\infty},\partial_{x}^{2}F_{\infty})\geq 0$ in $\mathbb{R}\times(0,\infty)$ and $F_{\infty}$ equals the left-continuous version of $F_{\text{in}}$ at $t=0$. Therefore, by the comparison principle, $\overline{F}\leq F_{\infty}$ in $\mathbb{R}\times(0,\infty)$, and this implies $\overline{F}\leq F_{\infty}\leq\liminf\nolimits_{*}F_{\epsilon}$ as desired. ∎
B.3. Contractivity
This subsection establishes that operators such as the family $(T_{\epsilon})_{\epsilon>0}$ are contractive on $CDF(\overline{\mathbb{R}})$ with respect to the supremum norm.
Proposition 12.
Suppose that $T$ is an operator on $CDF(\overline{\mathbb{R}})$ satisfying the following two properties:
  • (i)
    Monotonicity: If $F\leq G$ pointwise in $\mathbb{R}$, then $TF\leq TG$ also holds.
  • (ii)
    Commutation with constants: If $F\in CDF(\overline{\mathbb{R}})$ satisfies $c_{-}\leq F\leq c_{+}$ for some constants $c_{-},c_{+}\in[0,1]$, then $T(F+c)=TF$ for any $c\in[-c_{-},1-c_{+}]$.
  • (iii)
    Stationarity at infinity: For any $F\in CDF(\overline{\mathbb{R}})$, $TF(+\infty)=F(+\infty)$ and $TF(-\infty)=F(-\infty)$.
Then $T$ is contractive, that is,
$\|TF-TG\|_{\sup}\leq\|F-G\|_{\sup}\quad\text{for any}\,\,F,G\in CDF(\overline{\mathbb{R}}).$
The proof is based on the following lemma.
Lemma 4.
Let $T$ be an operator on $CDF(\overline{\mathbb{R}})$ satisfying assumptions (i)-(iii) of Proposition 12. If $F_{1},F_{2}\in CDF(\overline{\mathbb{R}})$ are chosen so that
$F_{2}\leq TF_{1}\quad\text{pointwise in}\,\,\mathbb{R}$,
then, for any $\delta\in[0,1]$ and $c\in[0,1-\delta]$, the functions $G_{1},G_{2}$ defined by
$G_{i}=\max\{F_{i}-\delta,c\}$
satisfy $G_{2}\leq TG_{1}$ pointwise in $\mathbb{R}$.
Proof.
Notice that $G_{1}+\delta=\max\{F_{1},c+\delta\}$, which is a function in $CDF(\overline{\mathbb{R}})$ that is pointwise larger than $F_{1}$. Therefore, by monotonicity,
$T(G_{1}+\delta)\geq TF_{1}\geq F_{2}.$
On the other hand, since property (iii) implies $T(G_{1}+\delta)(-\infty)=G_{1}(-\infty)+\delta\geq c+\delta$, one deduces that $T(G_{1}+\delta)\geq c+\delta$ pointwise. Therefore, $T(G_{1}+\delta)\geq\max\{F_{2},c+\delta\}=G_{2}+\delta$ follows. By property (ii), this implies $TG_{1}\geq G_{2}$. ∎
Here is the proof of contractivity:
Proof of Proposition 12.
Let $F,G\in\text{CDF}(\overline{\mathbb{R}})$. Note that
$G-\|F-G\|_{\text{sup}}\leq F\leq G+\|F-G\|_{\text{sup}}$
and the goal is to show that the same inequality holds if $(F,G)$ is replaced by $(TF,TG)$. By symmetry, it suffices to prove that $TF\leq TG+\|F-G\|_{\text{sup}}$.
Since the desired inequality is trivial if $\|F-G\|_{\text{sup}}\geq 1$, assume that $\|F-G\|_{\text{sup}}<1$ in what follows.
Let $G_{1}=\max\{F-\|F-G\|_{\text{sup}},0\}$ and $G_{2}=\max\{TF-\|F-G\|_{\text{sup}},0\}$. By Lemma 4,
$G_{2}\leq TG_{1}.$
Yet $G_{1}\leq G$ holds so $G_{2}\leq TG_{1}\leq TG$ by (i). This implies $TF-\|F-G\|_{\text{sup}}\leq TG$.∎
Appendix C Effective Coefficients in the Series-Parallel Graph
This appendix computes the constants $\sigma_{D}$ and $a_{R}$ for the logarithms of the distance $D^{(n)}$ and the resistance $R^{(n)}$ of the critical series-parallel graph. In either case, the relevant forcing $f$ takes the form
$f(u)=\log(1+e^{-u}).$
The left-inverse $g$ is given by
$g(s)=\log(e^{s}-1)\quad\text{if}\,\,s\geq\log 2,\quad g(s)=0,\quad\text{otherwise.}$
Since $g(s)+f(g(s))=s$ for $s\geq\log 2$, the formulas from Section 1.3.2 become
$\sigma_{D}$ $=-\int_{0}^{\log 2}s\,ds-\int_{\log 2}^{\infty}\log\left(\frac{1}{1-e^{-s}}\right)\,ds,$
$a_{R}$ $=\int_{0}^{\log 2}s^{2}\,ds+\int_{\log 2}^{\infty}\log^{2}\left(\frac{1}{1-e^{-s}}\right)\,ds$
$\quad+2\int_{0}^{\log 2}s^{2}\,ds+2\int_{\log 2}^{\infty}s\log\left(\frac{1}{1-e^{-s}}\right)\,ds$
The integrals above can be computed using polylogarithms, a family of functions related to the Riemann zeta function $\zeta$. The dilogarithm and trilogarithm $Li_{2}$ and $Li_{3}$ are defined for $t\in[0,1]$ by the formulas
$Li_{2}(t)=-\int_{0}^{t}\frac{1}{x}\log(1-x)\,dt,\quad Li_{3}(x)=\int_{0}^{t}\frac{1}{x}Li_{2}(x)\,dx.$
These functions have the power series representation $Li_{k}(t)=\sum_{n=1}^{\infty}n^{-k}t^{n}$, hence, in particular, $Li_{k}(1)=\zeta(k)$. The computations that follow seem to be related to Euler’s formula for $Li_{2}(\frac{1}{2})$ and Landen’s formula for $Li_{3}(\frac{1}{2})$ (see [19, Section 1.1]).
To compute the integrals, use the change-of-variables $x=e^{-s}$, $ds=\frac{1}{x}dx$. In the case of $\sigma_{D}$, this leads to
$\sigma_{D}$ $=\int_{\frac{1}{2}}^{1}\log(x)\frac{1}{x}\,dx+\int_{0}^{\frac{1}{2}}\log(1-x)\frac{1}{x}\,dx$
$=-\frac{1}{2}\log^{2}(\frac{1}{2})+\frac{1}{2}\int_{\frac{1}{2}}^{1}\log(x)\frac{1}{1-x}\,dx+\frac{1}{2}\int_{0}^{\frac{1}{2}}\log(1-x)\frac{1}{x}\,dx$
$=-\frac{1}{2}\log^{2}(\frac{1}{2})+\frac{1}{2}\int_{\frac{1}{2}}^{1}\log(1-x)\frac{1}{x}\,dx+\frac{1}{2}\log^{2}(\frac{1}{2})+\frac{1}{2}\int_{0}^{\frac{1}{2}}\log(1-x)\frac{1}{x}\,dx$
$=-\frac{1}{2}Li_{2}(1)=-\frac{1}{2}\zeta(2).$
In the case of $a_{R}$, this change-of-variables yields
$a_{R}$ $=3\int_{\frac{1}{2}}^{1}\log^{2}(x)\frac{1}{x}\,dx+\int_{0}^{\frac{1}{2}}\log^{2}(1-x)\frac{1}{x}\,dx+2\int_{0}^{\frac{1}{2}}\log(x)\log(1-x)\frac{1}{x}\,dx$
After integrating-by-parts and changing variables in the second integral, one has
$\int_{0}^{\frac{1}{2}}\log^{2}(1-x)\frac{1}{x}\,dx$ $=2\int_{0}^{\frac{1}{2}}\log(x)\log(1-x)\frac{1}{1-x}\,dx+\log^{3}(\frac{1}{2})$
$=2\int_{\frac{1}{2}}^{1}\log(x)\log(1-x)\frac{1}{x}\,dx+\log^{3}(\frac{1}{2}).$
This leads to
$a_{R}$ $=-\log^{3}(\frac{1}{2})+2\int_{\frac{1}{2}}^{1}\log(x)\log(1-x)\frac{1}{x}\,dx+\log^{3}(\frac{1}{2})+2\int_{0}^{\frac{1}{2}}\log(x)\log(1-x)\frac{1}{x}\,dx$
$=2\int_{0}^{1}\log(x)\log(1-x)\frac{1}{x}\,dx.$
Finally, integrating-by-parts again and invoking the definitions of $Li_{2}$ and $Li_{3}$ yields
$a_{R}$ $=2\int_{0}^{1}Li_{2}(x)\frac{1}{x}\,dx-Li_{2}(1)\log(1)=2\int_{0}^{1}Li_{2}(x)\frac{1}{x}\,dx=2Li_{3}(1)=2\zeta(3).$

### v2

@@
 This paper analyzes a class of recursive distributional equations (RDE’s) proposed by Gurel-Gurevich [20] and involving a bias parameter $p$, which includes the logarithm of the resistance of the series-parallel graph.
-A discrete-time evolution equation resembling a nonlinear, fractional Fisher-KPP equation is derived to describe the CDF’s of solutions.
+A discrete-time evolution equation resembling a quasilinear Fisher-KPP equation is derived to describe the CDF’s of solutions.
 When the bias parameter $p=\frac{1}{2}$, this equation is shown to have a PDE scaling limit, from which distributional limit theorems for the RDE are derived.
@@
 Interpreting $G^{(n)}$ as a resistor network where each edge has resistance one, one can study the effective resistance $R^{(n)}$ between $A$ and $B$.
-The asymptotic analysis of $R^{(n)}$ as $n\to\infty$ was initiated in [18].
+The asymptotic analysis of $R^{(n)}$ as $n\to+\infty$ was initiated in [21].
 Since $G^{(n)}$ can be realized by combining two independent copies $G^{(n-1)}_{1}$ and $G^{(n-1)}_{2}$ of $G^{(n-1)}$, say, in series with probability $p$ and in parallel with probability 1p1-p, the law of the resistance $R^{(n)}$ is a solution of the recursive distributional equation (RDE)
 $R^{(n)}\overset{d}{=}\left\{\begin{array}[]{r l}R^{(n-1)}_{1}+R^{(n-1)}_{2},&\text{with probability}\,\,p,\\ \left(\frac{1}{R_{1}^{(n-1)}}+\frac{1}{R_{2}^{(n-1)}}\right)^{-1},&\text{with probability}\,\,1-p.\end{array}\right$.
-Using this fact, a number of results were proved in [18]: When $p>\frac{1}{2}$, the logarithm $\log R^{(n)}$ grows linearly in $n$. More precisely, there is a constant $\lambda(p)>0$ such that
-(1) $\mathbb{P}\left\{(2p-1)\log 2\leq\frac{1}{N}\log R^{(N)}\leq\lambda(p)\right\}\to 1\quad\text{as}\,\,N\to\infty.$
-Due to symmetry, this implies that $\log R^{(n)}\to$-$\infty$ linearly when $p<\frac{1}{2}$.
-At the critical point $p=\frac{1}{2}$, Hambly and Jordan instead proved that $\log R^{(n)}$ exhibits sublinear growth.
-The main results of this paper (see Theorem 1 and Corollary 1 below) show that, in fact, $\log R^{(N)}$ grows like $N^{\frac{1}{3}}$.
+Using this fact, a number of results were proved in [21]: When $p>1/2$, the logarithm $\log R^{(n)}$ grows linearly in $n$. More precisely, there is a constant $\lambda(p)>0$ such that
+(1) $\mathbb{P}\left\{(2p-1)\log 2\leq\frac{1}{N}\log R^{(N)}\leq\lambda(p)\right\}\to 1\quad\text{as}\,\,N\to+\infty.$
+Due to symmetry, this implies that $\log R^{(n)}\to$-$\infty$ linearly when $p<1/2$.
+At the critical point $p=1/2$, Hambly and Jordan instead proved that $\log R^{(n)}$ exhibits sublinear growth.
+The main results of this paper (see Theorem 2 and Corollary 1 below) show that, in fact, $\log R^{(N)}$ grows like $N^{1/3}$.
 Specifially,
-$\frac{1}{(72\zeta(3))^{\frac{1}{3}}N^{\frac{1}{3}}}\log R^{(N)}+\frac{1}{2}\overset{d}{\to}\text{Beta}(2,2)\quad\text{as}\,\,N\to\infty,$
+$\frac{1}{(72\zeta(3))^{\frac{1}{3}}N^{\frac{1}{3}}}\log R^{(N)}+\frac{1}{2}\overset{d}{\to}\text{Beta}(2,2)\quad\text{as}\,\,N\to+\infty,$
 where $\zeta$ is the Riemann zeta function. This confirms a conjecture of Addario-Berry, Cairns, Devroye, Kerriou, and Mitchell [1].
@@
 What can be said about the long-time behavior of the law of solutions?
-This paper endeavors to address this problem in some generality when $p=\frac{1}{2}$, focusing on the distributional limit of $n^{-1/3}X^{(n)}$.
+This paper endeavors to address this problem in some generality when $p=1/2$, focusing on the distributional limit of $n^{-1/3}X^{(n)}$.
 The general framework developed here applies not only to the resistance of the critical series-parallel graph, but also to some other RDE’s considered recently in the literature.
 For instance, as is shown in Section 4 below, the class of RDE’s determined by (2) and (5) includes the symmetric hipster random walk from [1] and two-player symmetric cooperative motion, one of the RDE’s analyzed very recently by Addario-Berry, Beckman, and Lin in [3].
-While the analysis of the asymptotic behavior in the case $p\neq\frac{1}{2}$ is not touched on here, the framework developed in this paper suggests that it is related to the singular limit of certain fractional, quasilinear variants of the Fisher-KPP equation.
-This is made precise in Section 1.3.4 below.
+While the analysis of the asymptotic behavior in the case $p\neq 1/2$ is not touched on here, the framework developed in this paper suggests that it may be amenable to techniques developed in the study of reaction-diffusion equations.
+This is discussed further in Section 1.5 below.
+A few days after this paper first appeared as a preprint, a work of Chen, Duquesne, and Shi [13] appeared, which considers the same class of RDE’s and independently obtains the limit theorem for $n^{-1/3}X^{(n)}$.
 1.2. General Setting
@@
 The joint law of the functions $(f_{+},f_{-})$ will be denoted by $\mathbf{P}$ in what follows.
-$\mathbf{P}$ is assumed to be a Borel probability measure on the space $BC([0,\infty))\times BC([0,\infty))$ of pairs of bounded continuous functions in $0,\infty)$ with the uniform norm topology.
+$\mathbf{P}$ is assumed to be a Borel probability measure on the space $BC([0,+\infty))\times BC([0,+\infty))$ of pairs of bounded continuous functions in $0,+\infty)$ with the uniform norm topology.
 In most of the examples of interest here, however, $f_{+}$ and $f_{-}$ are deterministic; see the discussion in Section 4.
@@
 With this notation, it follows that the random variables $\{X^{(n)}\}$ solve the RDE (9) if and only if the corresponding CDF’s $\{F_{n}\}$ solve the recursive equation
-$F_{n+1}=TF_{n}.$
-In the first step of this work, a discrete-time evolution equation is derived to describe the evolution of the CDF of $X^{(n)}$.
-Namely, there is a nonlinear operator $\mathscr{L}$ defined on the space of CDF’s such that the recursive equation $F_{n+1}=TF_{n}$ can be rewritten in the following suggestive form
-(11) $F_{n+1}-F_{n}=\mathscr{L}F_{n}+(1-2p)F_{n}(1-F_{n}).$
+$F_{n}=TF_{n-1}.$
+In the first step of this work, this recursion is reformulated as a discrete-time evolution equation resembling a parabolic PDE.
+Namely, there is a nonlinear operator $\mathscr{L}$ defined on the space of CDF’s such that
+(11) $F_{n}-F_{n-1}=\mathscr{L}F_{n-1}+(1-2p)F_{n-1}(1-F_{n-1}).$
 As will be argued below, $\mathscr{L}$ has the character of an advection-diffusion operator, albeit a nonlinear one, while the “reaction term” F(1F)F(1-F) is familiar from the Fisher-KPP equation.
 The equation thus resembles a reaction-diffusion equation.
+A key property of the equation (11) is its monotonicity: If $F\leq G$ pointwise, then $TF\leq TG$ also holds.
+This means that (11) is amenable to techniques developed in the theory of parabolic PDE’s.
+The monotonicity of $T$ also plays a fundamental role in the aforementioned work of Chen, Duquesne, and Shi [13].
+Their proof involves controlling CDF’s through the delicate construction of explicit sub- and supersolutions of (11).
+By contrast, the proof here is more conceptual: It starts by establishing that the evolution equation (11) “looks” like a PDE under a certain rescaling of the variables $(x,n)$, then invokes abstract convergence results for monotone semigroups and parabolic PDE’s.
+In addition to the clear intuitive interpretation, the method has the advantage that it applies very generally, including to cases in which explicit formulae are not available.
+1.4. Scaling Limits at the Critical Point $p=1/2$
 As far as convergence in distribution is concerned, deriving a limit for $N^{-1/\alpha}X^{(N)}$ for some exponent $\alpha$ is equivalent to proving a scaling limit for the equation (11).
-In the main result of this paper, which treats the critical point $p=\frac{1}{2}$, this is done by adapting the classical approach to numerical approximations of parabolic PDE’s developed by Barles and Souganidis [6].
+In the main result of this paper, which treats the critical point $p=1/2$, this is done by adapting the classical approach to numerical approximations of parabolic PDE’s developed by Barles and Souganidis [7].
 To have a sense of why it is reasonable to expect a PDE scaling limit, consider what happens in (8) when $f_{+}=f_{-}=f_{\mathbb{Z}}$, where $f_{\mathbb{Z}}(u)=(1-u)_{+}$.
@@
 As shown in that work, the evolution equation governing the CDF now takes a very concrete form, namely,
-$F_{n+1}-F_{n}=|\nabla_{\mathbb{Z}}F_{n}|\Delta_{\mathbb{Z}}F_{n}\quad\text{in}\,\,\mathbb{Z}$,
+$F_{n}-F_{n-1}=|\nabla_{\mathbb{Z}}F_{n-1}|\Delta_{\mathbb{Z}}F_{n-1}\quad\text{in}\,\,\mathbb{Z}$,
 where $\nabla_{\mathbb{Z}}$ and $\Delta_{\mathbb{Z}}$ are certain discrete derivative operators.
-This observation was then used in conjunction with the method of [6] and a regularization effect of the equation to show that, after a suitable rescaling, the CDF converges to the solution of the initial-value problem
+This observation was then used in [3] in conjunction with the method of [7] and a regularization effect of the equation to show that, after a suitable rescaling, the CDF converges to the solution of the initial-value problem
 $\partial_{t}F=|\partial_{x}F|\partial_{x}^{2}F,\quad F(x,0)=\mathbf{1}_{[0,\infty)}(x).$
@@
 In this work, this general strategy — derivation of a discrete-time evolution equation and application of a convergence result in the spirit of [7] — is shown to still apply in the setting of the RDE’s defined above, although its execution is complicated somewhat by the fact that the operator $\mathscr{L}$ is generally nonlocal.
-1.3.1. Assumptions
-Some assumptions on the law $\mathbf{P}$ of the forcing $(f_{+},f_{-})$ are in order. First, define the class of functions $\mathcal{S}$ so that $f\in\mathcal{S}$ if and only if
+1.4.1. Main Results
+The main result holds under some assumptions on the law $\mathbf{P}$ of $(f_{+},f_{-})$. First, define the class of functions $\mathcal{S}$ so that $f\in\mathcal{S}$ if and only if
 (12) $f:\mathbb{R}\to\mathbb{R}\quad\text{is continuous and nonincreasing, and}$
@@
 (14) $g_{\rho}(s)=\sup\left\{u\geq 0\,\mid\,u+f_{\rho}(u)\leq s\right\}\quad\text{if}\,\,s\geq f_{\rho}(0),\quad g_{\rho}(s)=0,\quad\text{otherwise.}$
-Notice that the restriction of $g_{\rho}$ to the set $f(0),+\infty)$ is the right-continuous right-inverse of $u\mapsto u+f_{\rho}(u)$.
-By definition, if $f_{\rho}(+\infty)=0$, then $g_{\rho}(s)-s\to 0$ as $s\to+\infty$. The next assumption asks that $g_{\rho}(s)-s$ vanishes in a quantitative manner:
-(15) $\int_{0}^{\infty}(1+s)\mathbf{E}[|g_{+}(s)-s|+|g_{-}(s)-s|]\,ds<\infty.$
-1.3.2. Main Result
-Under the assumption (15), it is possible to define effective coefficients $\sigma\in\mathbb{R}$ and $a\geq 0$ according to the formulae
-$\sigma$ $=-2p\int_{0}^{\infty}\mathbf{E}[|g_{+}(s)-s|]\,du+2(1-p)\int_{0}^{\infty}\mathbf{E}[|g_{-}(s)-s|]\,ds,\quad$
-$a$ $=p\int_{0}^{\infty}\mathbf{E}[(3s-g_{+}(s))(s-g_{+}(s))]\,ds+(1-p)\int_{0}^{+\infty}\mathbf{E}[(3s-g_{-}(s))(s-g_{-}(s))]\,ds.$
-The nonnegativity of $a$ follows from the fact that $g_{+}(s),g_{-}(s)\leq s$ for any $s\geq 0$.
-Define the exponent $\alpha\in\mathbb{N}$ as follows:
-$\alpha=2\quad\text{if}\,\,\sigma\neq 0,\quad\alpha=3,\quad\text{otherwise.}$
-The next theorem characterizes the scaling limit of the CDF in the critical case $p=\frac{1}{2}$.
+Notice that the restriction of $g_{\rho}$ to the set $f_{\rho}(0),+\infty)$ is the right-continuous right-inverse of the function $u\mapsto u+f_{\rho}(u)$.
+By definition, if $f_{\pm}(+\infty)=0$, then $g_{\pm}(s)-s\to 0$ as $s\to+\infty$. The next assumptions ask that $g_{\pm}(s)-s$ vanishes in a quantitative manner: First, under the assumption that
+(15) $\int_{0}^{+\infty}\mathbf{E}[|g_{+}(s)-s|+|g_{-}(s)-s|]\,ds<\infty,$
+define the constant $\sigma$ by
+$\sigma=-\int_{0}^{+\infty}\mathbf{E}[|g_{+}(s)-s|]\,du+\int_{0}^{+\infty}\mathbf{E}[|g_{-}(s)-s|]\,ds.$
+If in addition the following stronger assumption holds
+(16) $\int_{0}^{+\infty}(1+s)\mathbf{E}[|g_{+}(s)-s|+|g_{-}(s)-s|]\,ds,$
+define also the constant $a\geq 0$ by
+$a=\frac{1}{2}\int_{0}^{+\infty}\mathbf{E}[(3s-g_{+}(s))(s-g_{+}(s))]\,ds+\frac{1}{2}\int_{0}^{+\infty}\mathbf{E}[(3s-g_{-}(s))(s-g_{-}(s))]\,ds.$
+The nonnegativity of $a$ follows from the fact that $g_{\pm}(s)\leq s$ for any $s\geq 0$.
+The next two theorems characterize the scaling limit of the CDF at the critical point $p=1/2$ in terms of a family of parabolic PDE’s parametrized by $\sigma$ and $a$.
 The relevant background from PDE theory, particularly the theory of viscosity solutions, is reviewed in Appendix A below.
+Motivated by recent work of Chen, Derrida, Duquesne, and Shi [12], the theorem allows for the possibility that $p=p^{(N)}$ converges to $1/2$ as $N\to+\infty$.
 Theorem 1.
-Assume that $p=\frac{1}{2}$, $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, and (15) holds.
-If $\{F_{n}\}$ is the sequence of CDF’s associated to some solution $\{X^{(n)}\}$ of (9), then, independent of the initial CDF $F_{0}$, if $F_{\epsilon}$ is the rescaled CDF defined by
-$F_{\epsilon}(x,t)=F_{[\epsilon^{-\alpha}t]}(\epsilon^{-1}x),$
-then $F_{\epsilon}\to F$ locally uniformly in $\mathbb{R}\times(0,\infty)$, where $F$ is determined as follows:
+Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, (15) holds, and $\sigma\neq 0$.
+Fix a $\theta\in\mathbb{R}$ and let $p^{(N)}=1/2+\theta N^{-1}$.
+For each $N\in\mathbb{N}$, let $\{X^{(N,n)}\}$ be a solution of the RDE
+$X^{(N,n)}\overset{d}{=}\Phi^{(N)}(X^{(N,n-1)}_{1},X^{(N,n-1)}_{2})$,
+where $(X^{(N,n-1)}_{1},X^{(N,n-1)}_{2})$ are i.i.d. copies of $X^{(N,n-1)}$ and $\Phi^{(N)}$ is given by (8) with bias parameter $p=p^{(N)}$.
+Assume that there is a function $F_{\text{in}}$ such that, at the initial time $n=0$, the following limit holds:
+$\lim_{N\to+\infty}\mathbb{P}\{N^{-1/2}X^{(N,0)}\leq x\}=F_{\text{in}}(x)\quad\text{for almost every}\,\,x\in\mathbb{R}.$
+If $\{F_{N}\}$ is the sequence of rescaled CDF’s defined in $\mathbb{R}\times[0,+\infty)$ by
+$F_{N}(x,t)=\mathbb{P}\{N^{-1/2}X^{(N,[Nt])}\leq x\},$
+then $F_{N}\to F$ locally uniformly in $\mathbb{R}\times(0,+\infty)$ as $N\to+\infty$, where $F$ is the unique bounded discontinuous viscosity solution of the initial value problem
+(19) $\left\{\begin{array}[]{r l}\partial_{t}F-\sigma|\partial_{x}F|^{2}+2\theta F(1-F)=0&\text{in}\,\,\mathbb{R}\times(0,+\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
+See Section 5.5 for an example in which (15) fails to hold and $N^{-1/2}X^{(N)}\overset{d}{\to}+\infty$.
+If $\theta=\sigma=0$ above, then the proof of the theorem instead implies that $F_{N}$ converges to the initial datum $F_{\text{in}}$, meaning that $\{X^{(N,n)}\}$ exhibits no nontrivial motion in the diffusive scaling limit.
+Under the stronger assumption (16), one finds that instead the appropriate scaling is subdiffusive.
+Theorem 2.
+Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, (16) holds, $\sigma=0$, and $a>0$. For each $N\in\mathbb{N}$, let $\{X^{(N,n)}\}$ be a solution of the RDE
+$X^{(N,n)}\overset{d}{=}\Phi^{(N)}(X^{(N,n-1)}_{1},X^{(N,n-1)}_{2})$,
+where $(X^{(N,n-1)}_{1},X^{(N,n-1)}_{2})$ are i.i.d.
+copies of $X^{(N,n-1)}$ and $\Phi^{(N)}$ is given by (8) with bias parameter $p=p^{(N)}$ satisfying $p^{(N)}=1/2+\theta N^{-1}$ for some $\theta\in\mathbb{R}$.
+Assume that there is a function $F_{\text{in}}$ such that, at the initial time $n=0$, the following limit holds:
+$\lim_{N\to+\infty}\mathbb{P}\{N^{-1/3}X^{(N,0)}\leq x\}=F_{\text{in}}(x)\quad\text{for almost every}\,\,x\in\mathbb{R}.$
+If $\{F_{N}\}$ is the sequence of rescaled CDF’s defined in $\mathbb{R}\times[0,+\infty)$ by
+$F_{N}(x,t)=\mathbb{P}\{N^{-1/3}X^{(N,[Nt])}\leq x\},$
+then $F_{N}\to F$ locally uniformly in $\mathbb{R}\times(0,+\infty)$ as $N\to+\infty$, where $F$ is the unique bounded discontinuous viscosity solution of the initial-value problem
+(22) $\left\{\begin{array}[]{r l}\partial_{t}F-a|\partial_{x}F|\partial_{x}^{2}F+2\theta F(1-F)=0&\text{in}\,\,\mathbb{R}\times(0,+\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
+The PDE’s derived above become more familiar when passing to PDF’s and setting $\theta=0$: The spatial derivative $\rho=\partial_{x}F$ is a solution of Burger’s equation $\partial_{t}\rho=\sigma\partial_{x}(\rho^{2})$ in Theorem 1 and a solution of the porous medium equation $\partial_{t}\rho=\frac{1}{2}a\partial_{x}^{2}(\rho^{2})$ in Theorem 2.
+As observed in [3, 2], when $\theta=0$ and $F_{\text{in}}=\mathbf{1}_{[0,\infty)}$, the solutions of the PDE’s above are explicitly known and closely related to the $\text{Beta}(2,1)$ and $\text{Beta}(2,2)$ distributions.
+This leads to limit theorems for the RDE (9) started from an arbitrary (fixed) initial distribution.
+Corollary 1.
+Assume that $p=1/2$, $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, and (15) holds, and let $\{X^{(n)}\}$ be a solution of the RDE (9) with an arbitrary initial distribution.
   • (i)
-    If $\sigma\neq 0$, then $F$ is the unique bounded discontinuous viscosity solution of the initial value problem
-    (18) $\left\{\begin{array}[]{r l}\partial_{t}F-\sigma|\partial_{x}F|^{2}=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=\mathbf{1}_{[0,\infty)}(x).\end{array}\right$.
+    The sequence $\{N^{-1/2}X^{(N)}\}$ has the following distributional limit:
+    $-\frac{\text{sgn}(\sigma)}{2\sqrt{N}}X^{(N)}\overset{d}{\to}\sqrt{|\sigma|}\text{Beta}(2,1),$
+    where $\sqrt{|\sigma|}\text{Beta}(2,1)$ denotes a $\text{Beta}(2,1)$ random variable dilated by a factor $\sqrt{|\sigma|}$. In particular, if $\sigma=0$, then the limit is zero.
   • (ii)
-    If $\sigma=0$, then $F$ is the unique bounded discontinuous viscosity solution of the initial-value problem
-    (21) $\left\{\begin{array}[]{r l}\partial_{t}F-a|\partial_{x}F|\partial_{x}^{2}F=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=\mathbf{1}_{[0,\infty)}(x).\end{array}\right$.
-These PDE’s become more familiar when passing to PDF’s: The spatial derivative $\rho=\partial_{x}F$ is a solution of Burger’s equation $\partial_{t}\rho=\sigma\partial_{x}(\rho^{2})$ in case (i) and a solution of the porous medium equation $\partial_{t}\rho=\frac{1}{2}a\partial_{x}^{2}(\rho^{2})$ in case (ii).
-As observed in [3, 2], the solutions of the PDE’s above are explicitly known and closely related to the $\text{Beta}(2,1)$ and $\text{Beta}(2,2)$ distributions. The next corollary makes this precise.
-Corollary 1.
-Assume that $p=\frac{1}{2}$, $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, and (15) holds, and let $\{X^{(n)}\}$ be a solution of the RDE (9) with an arbitrary initial distribution.
-  • (i)
-    If $\sigma\neq 0$, then
-    $-\frac{\text{sgn}(\sigma)}{2\sqrt{|\sigma|}\sqrt{N}}X^{(N)}\overset{d}{\to}\text{Beta}(2,1).$
-  • (ii)
-    If $\sigma=0$ and $a>0$, then
+    Under the stronger assumption (16), if $\sigma=0$ and $a>0$, then $\{N^{-1/3}X^{(N)}\}$ instead has a nontrivial distributional limit:
     $\frac{1}{(36a)^{\frac{1}{3}}N^{\frac{1}{3}}}X^{(N)}+\frac{1}{2}\overset{d}{\to}\text{Beta}(2,2)$
   • (iii)
-    If $\sigma=a=0$, then $\mathbf{P}\{f_{+}=f_{-}=0\}=1$ and $X^{(N)}\overset{d}{=}X^{(0)}$ for each $N$.
+    Under (16), if $\sigma=a=0$, then $\mathbf{P}\{f_{+}=f_{-}=0\}=1$ and $\{X^{(n)}\}$ is the constant sequence $X^{(n)}\overset{d}{=}X^{(0)}$.
 It is not hard to show, through manipulations of the definitions above, that $\sigma>0$ holds, for instance, if $f_{-}>f_{+}$ $\mathbf{P}$-almost surely, while $\sigma=0$ if $f_{+}=f_{-}$.
@@
 For more on this and further discussion of examples, see Section 4.
-In addition to characterizing the asymptotics of the resistance of the critical series-parallel graph, the proof of Theorem 1 also leads to a characterization of the limiting behavior of Pemantle’s min-plus binary tree (or first passage percolation on the critical series-parallel graph).
-This is discussed in Section 4.2 below.
-1.3.3. Idea of Proof
+In addition to Corollary 1, Theorems 1 and 2 have two applications that are worth highlighting now.
+In Section 4.1 below, Theorem 2 is used to characterize the asymptotics of the resistance of the critical series-parallel graph in the general case when the edges have i.i.d.
+resistances, which could equal zero or $\infty$.
+Similarly, Section 4.2 gives a characterization of the limiting behavior of Pemantle’s min-plus binary tree (or first passage percolation on the critical series-parallel graph) when the weights on the leaves are i.i.d.
+with an arbitrary distribution, answering a question posed in [6].
+1.4.2. Idea of the Proof
 The PDE scaling limit of the CDF is obtained using an extension of the approach of [7], which is broadly applicable to elliptic and parabolic PDE’s studied through the lens of Crandall and Lions’ theory of viscosity solutions.
 There are two key ingredients in that approach, namely, monotonicity and consistency.
-In the present context, monotonicity means the following: If $\{F_{n}\}$ and $\{G_{n}\}$ are two sequences of CDF’s arising from solutions of the RDE and if the inequality $F_{0}\leq G_{0}$ holds pointwise, then $F_{n}\leq G_{n}$ holds for any $n$.
-For the purposes of this paper, this is the bare minimum structural condition necessary to treat the evolution (11) as a parabolic equation.
-Consistency, on the other hand, refers to the behavior of $T$ under rescaling.
+As mentioned already above, monotonicity refers to the fact that the equation (11) is order-preserving: If $\{F_{n}\}$ and $\{G_{n}\}$ are two solutions, and if $F_{0}\leq G_{0}$ pointwise, then the inequality $F_{n}\leq G_{n}$ remains true for any subsequent $n$.
+Consistency, on the other hand, refers to the behavior of the equation under rescaling.
 Specifically, Proposition 10 below shows that if $\Phi$ is a smooth CDF, rescaled according to $\Phi_{\delta}(x)=\Phi(\delta x)$, then
-(22) $(\mathscr{L}\Phi_{\delta})(\delta^{-1}x)=\delta^{2}\sigma|\partial_{x}\Phi(x)|^{2}+\delta^{3}a|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)+\cdots.$
-Naïvely inserting this expansion in the equation (11) and recalling the definition of the rescaled CDF $F_{\epsilon}$, one is led to guess that the PDE’s (18) and (21) ought to appear in the limit.
-The approach of [6] provides a framework for making such a formal argument rigorous.
-1.3.4. Variants of the Fisher-KPP Equation when $p\neq\frac{1}{2}$
-The expansion (22) leads to interesting new questions away from the critical point $p=\frac{1}{2}$. Passing to the ballistically rescaled CDF
-$F_{\epsilon}(x,t)=F_{[\epsilon^{-1}t]}(\epsilon^{-1}x),$
-one finds that formally
-$\partial_{t}F_{\epsilon}$ $=\epsilon^{2}a|\partial_{x}F_{\epsilon}|\partial_{x}^{2}F_{\epsilon}+\epsilon\sigma|\partial_{x}F_{\epsilon}|^{2}+\frac{1}{\epsilon}(1-2p)F_{\epsilon}(1-F_{\epsilon})+\cdots.$
-After discarding the error terms, this is the PDE obtained after applying a ballistic rescaling to the following quasilinear, fractional variant of the Fisher-KPP equation
-(23) $\partial_{t}G=a|\partial_{x}G|\partial_{x}^{2}G+\sigma|\partial_{x}G|^{2}+(1-2p)G(1-G).$
-The asymptotics of this equation, particularly the propagation of “front-like” solutions (such as CDF’s), does not seem to have been treated thus far in the literature.
-Both the cases where $a=0$ or $\sigma=0$ appear to be relevant to the problem at hand; see Section 4.3 for justification of this.
+(23) $(\mathscr{L}\Phi_{\delta})(\delta^{-1}x)=\delta^{2}\sigma|\partial_{x}\Phi(x)|^{2}+\delta^{3}a|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)+\cdots\quad\text{as}\,\,\delta\downarrow 0.$
+Rescaling the sequence of CDF’s $\{F_{n}\}$ according to $F_{N}(x,t)=F_{[Nt]}(N^{1/\alpha}x)$, naïvely inserting the above expansion into the equation (11), and setting $p=1/2+\theta N^{-1}$, one obtains the formal asymptotic expansion
+$\partial_{t}F_{N}=N^{1-\frac{2}{\alpha}}\sigma|\partial_{x}F_{N}|^{2}+N^{1-\frac{3}{\alpha}}a|\partial_{x}F_{N}|^{2}-2\theta F_{N}(1-F_{N})+\cdots\quad\text{as}\,\,N\to+\infty,$
+which motivates the choices $\alpha=2$ and $\alpha=3$ used above and the appearance of the PDE’s (19) and (22). The approach of [7] provides a framework for making such a formal argument rigorous.
+1.5. Variants of the Fisher-KPP Equation when $p\neq\frac{1}{2}$
+The results of this work suggest interesting new questions away from the critical point $p=\frac{1}{2}$.
+In view of the asymptotic expansion (23) and the PDE’s obtained in Theorems 1 and 2, the equation (11) shares some similarities with the reaction-diffusion equation
+(24) $\partial_{t}G=a|\partial_{x}G|\partial_{x}^{2}G+\sigma|\partial_{x}G|^{2}-2\theta G(1-G).$
+In fact, (11) becomes the discrete-in-space-time, finite-difference version of (24) for suitable choices of the forcing $(f_{+},f_{-})$; see Section 4.3 for more details.
+A number of works in the PDE literature considered the equation (24) in the case when $\sigma=0$ and $a>0$.
+Using an auxiliary ODE, Enguiça, Gavioli, and Sanchez [17] established the existence of traveling waves propagating at any large-enough speed, as in the classical Fisher-KPP equation.
+The work of Audrito and Vázquez [5] establishes existence, uniqueness, and positivity properties of traveling waves and also analyzes the asymptotic behavior of solutions, proving that sufficiently rapidly decaying perturbations of the unstable state (say, 0 if $\theta<0$) converge to the stable state ($1$ if $\theta<0$) inside a ballistically growing region, with the rate of growth determined by the minimal wave speed.
+It would be interesting to extend such results to the present context.
 1.6. Related Literature
@@
 There has been a recent burst of activity concerning RDE’s in which the relevant asymptotic behavior is not convergence to a fixed point as in the examples in [4], but instead a distributional scaling limit in the same spirit as the central limit theorem.
-Auffinger and Cable [5] analyzed Pemantle’s Min-Plus Binary Tree (or, equivalently, first passage percolation on the critical series-parallel graph, see [18, Section 3]), proving that solutions of the associated RDE converge to a $\text{Beta}(2,1)$ distribution after a diffusive rescaling provided the initial distribution is concentrated at one.
-This is improved to the case of an arbitrary initial distribution in Corollary 2 below.
+Auffinger and Cable [6] analyzed Pemantle’s Min-Plus Binary Tree (or, equivalently, first passage percolation on the critical series-parallel graph, see [21, Section 3]), proving that solutions of the associated RDE converge to a $\text{Beta}(2,1)$ distribution after a diffusive rescaling provided the initial distribution is concentrated at one.
+This is improved to the case of an arbitrary initial distribution in Corollary 3 below, answering one of the open problems posed therein.
+The proof in [6] is based on the explicit construction of sub- and supersolutions of (11), which, by monotonicity, can be used to control solutions, or what is referred to as a barrier argument in the PDE literature.
+The same strategy was employed by Chen, Duquesne, and Shi in [13], who derived the same $\text{Beta}(2,2)$ limit theorem as in Corollary 1 above, albeit under slightly more restrictive assumptions on $(f_{+},f_{-})$.
 Addario-Berry and coauthors recently introduced two classes of RDE’s, called hipster random walks and cooperative motion, in [1, 2, 3].
@@
 By contrast, here the RDE’s of interest are all monotone.
-There is a technical issue, namely, that the result of [6] does not apply directly as the evolution (11) only makes sense as an equation posed in the space of CDF’s, whereas [6] works with evolution equations posed in the space of bounded functions.
-In [3], this is arguably less of an issue as the equations all involve finite differences (although the relevant details are not discussed there).
+There is a technical issue, namely, that the result of [7] does not apply directly as the evolution (11) only makes sense as an equation posed in the space of CDF’s, whereas [7] works with evolution equations posed in the (vector) space of bounded functions.
+In [3], this was arguably less of an issue as the equations all involve finite differences (although the relevant details were not discussed there).
 Here $\mathscr{L}$ behaves like a nonlocal operator with a kernel with infinite range so at first blush one would like to know that [7] can be carried out with globally nondecreasing test functions, which is not obvious.
@@
 In the hope that it may be useful in other contexts, the result is formulated in an abstract way in Appendix B below.
-When $p\neq\frac{1}{2}$, the evolution equation obtained here suggests a strategy for tackling the asymptotics of the RDE.
+When $p\neq\frac{1}{2}$, the evolution equation obtained here and the analogy with reaction-diffusion equations suggests a strategy for tackling the asymptotics of the RDE.
 This is of interest both in the analysis of the resistance and the distance on the series-parallel graph.
 The latter was recently considered by Chen, Derrida, Duquesne, and Shi in [12].
-They proved that when $p>\frac{1}{2}$, the logarithm of the expected value of the distance grows at a linear rate with slope $\alpha(p)$, and they proved that $\alpha(1/2+\epsilon)\approx\sqrt{\zeta(2)\epsilon}$ as $\epsilon\to 0$.
+They proved that when $p>\frac{1}{2}$, the logarithm of the expected value of the distance grows at a linear rate with slope $\alpha(p)$, and, again using barrier arguments, they proved that $\alpha(1/2+\epsilon)\approx\sqrt{\zeta(2)\epsilon}$ as $\epsilon\to 0$.
+Their proof is motivated through the heuristic derivation of a PDE that can be shown to be closely related to (19).
 As mentioned above, convergence of the rescaled CDF’s to the solution of a PDE is proved here by adapting the classical approach of Barles and Souganidis from [7].
@@
 Section 4 discusses examples that fit into the class described by (9) and (8).
-The main results are proved in Section 5 conditional on an extension of [6], which is presented in the appendix.
+The main results are proved in Section 5 conditional on an extension of [7] and some uniqueness results for discontinuous viscosity solutions, which are presented in the appendix.
 There are three appendices.
@@
 Using standard extended-real-number arithmetic, from now on, the formula (10) defining $T$ will be understood to hold in the domain $CDF(\overline{\mathbb{R}})$.
+2.2. Continuity under Vague Convergence.
+It is useful to note that $T$ is continuous with respect to vague convergence. Recall that $\{F_{n}\}$ converges vaguely to $F$ if $F_{n}(x)\to F(x)$ at each point of continuity $x$ of $F$.
+Proposition 1.
+If $\{F_{n}\}$ is a sequence in $CDF(\overline{\mathbb{R}})$ converging vaguely to some $F$, then $\{TF_{n}\}$ converges vaguely to $TF$.
+Proof.
+It suffices to recall that this mode of convergence for $(F_{n})_{n\in\mathbb{N}}$ is equivalent to the existence of a probability space $\mathbb{P}$ supporting a sequence $\{(X_{1,n},X_{2,n})\}$ and a vector $(X_{1},X_{2})$ such that, for each $n$, $X_{1,n}$ and $X_{2,n}$ are i.i.d.
+with CDF $F_{n}$; $(X_{1},X_{2})$ are i.i.d.
+with CDF $F$; and $(X_{1,n},X_{2,n})\to(X_{1},X_{2})$ $\mathbb{P}$-almost surely.
+(Note that the possibility that $X_{1}$ or $X_{2}$ are infinite poses no difficulties here.) Letting $Y_{n}=\Phi(X_{1,n},X_{2,n})$ and $Y=\Phi(X_{1},X_{2})$, where $\Phi$ is sampled independently of $\{(X_{1,n},X_{2,n})\}$ and $(X_{1},X_{2})$, the boundedness and continuity of $f_{+}$ and $f_{-}$ are enough to deduce that $Y_{n}\to Y$ almost surely as $n\to\infty$.
+By definition of $T$, this implies $TF_{n}\to TF$.
+∎
 2.3. Symmetry
@@
 In this section, the monotonicity of $T$ is proved. Here monotone means that
-$\text{if}\,\,F_{1}\leq F_{2}\,\,\text{pointwise in}\,\,\mathbb{R},\quad\text{then}\,\,TF_{1}\leq TF_{2}\,\,\text{pointwise in}\,\,\mathbb{R}.$
+$\text{if}\,\,F\leq G\,\,\text{pointwise in}\,\,\mathbb{R},\quad\text{then}\,\,TF\leq TG\,\,\text{pointwise in}\,\,\mathbb{R}.$
 To this end, it is convenient to exploit the quadratic structure inherent in the definition. Specifically, it will be useful to consider a certain natural operator
@@
 $Y^{-}=X_{2}+f_{+}(X_{2}-X^{-}_{1})\leq X_{2}+f_{+}(X_{2}-X^{+}_{1})=Y^{+}.$
-2.4. Continuity
+Using coupling arguments very similar to the one employed above, one readily deduces that the map $T=T(p,\mathbf{P})$ is nondecreasing with respect to $p$ and $f_{+}$ and nonincreasing with respect to $f_{-}$.
+The result is stated next for the sake of precision.
+Proposition 3.
+(i) For any $p_{1}\leq p_{2}$, $T(p_{1},\mathbf{P})F\leq T(p_{2},\mathbf{P})F$ for all $F\in CDF(\overline{\mathbb{R}})$.
+(ii) If $\mathbf{P}$ and $\mathbf{\tilde{P}}$ are measures on $BC([0,\infty)\times BC([0,\infty)$ for which it is possible to find a coupling of the corresponding random variables $(f_{+},f_{-})$ and (f~+,f~)($\tilde$$_{+},\tilde{f}_{-}$) such that $f_{+}\leq\tilde{f}_{+}$ and $f_{-}\geq\tilde{f}_{-}$, then $T(p,\mathbf{P})F\leq T(p,\mathbf{\tilde{P}})F$ for each $F\in CDF(\overline{\mathbb{R}})$.
+2.5. Absolute Continuity
 For the sake of developing intuition, it may be useful to note that $T$ maps the space of continuous CDF’s to itself.
@@
 (31) $Y$ $=Z+\Theta f_{+}(|X_{2}-X_{1}|)+(1-\Theta)f_{-}(|X_{2}-X_{1}|).$
-The next result describes the difference TFFTF-F in terms of $(Y,Z,\Theta)$. Specifically, let $\mathscr{L}$ be the operator defined on $CDF(\overline{\mathbb{R}})$ by
+The next result describes the difference TFFTF-F in terms of $(Y,Z,\Theta)$. To that end, let $\mathscr{L}F$ be the function defined by
 (32) $\mathscr{L}F(x)=-p\mathbb{P}\{Z\leq x<Y\mid\Theta=1\}+(1-p)\mathbb{P}\{Y\leq x<Z\mid\Theta=0\}.$
@@
 $TF-F=\mathscr{L}F+(1-2p)F(1-F).$
-Further, if $p=\frac{1}{2}$, then $TF(\pm\infty)=F(\pm\infty)$ for any $F\in CDF(\overline{\mathbb{R}})$.
 Using the explicit formulas (31) and (30), one obtains very concrete expressions for the action of $\mathscr{L}$ on smooth CDF’s, as stated in the next result.
 Proposition 6.
-Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$. If $F\in CDF(\overline{\mathbb{R}})$ is smooth, then
+Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$. If $F\in CDF(\overline{\mathbb{R}})$ is differentiable, then
 (33) $\mathscr{L}F(x)$ $=2p\int_{-\infty}^{x}\mathbf{E}[F(x_{1}+g_{+}(x-x_{1}))-F(x)]F^{\prime}(x_{1})\,dx_{1}$
-$\quad+2(1-p)\int_{x}^{\infty}\mathbf{E}[F(x_{1}-g_{-}(x_{1}-x))-F(x)]F^{\prime}(x_{1})\,dx_{1}.$
-Proposition 4 has two immediate consequences that will be used in the sequel.
-First, it makes it relatively easy to see that $T$ commutes with the addition of constants whenever such addition makes sense.
+$\quad+2(1-p)\int_{x}^{+\infty}\mathbf{E}[F(x_{1}-g_{-}(x_{1}-x))-F(x)]F^{\prime}(x_{1})\,dx_{1}.$
+Propositions 5 and 6 immediately imply two properties of $\mathscr{L}$ that will be useful in the sequel and are stated in the next proposition.
 Proposition 7.
-Under the assumptions of Proposition 4, if $F\in CDF(\overline{\mathbb{R}})$, then $T(F+c)=TF+c$ for any constant $c\in[-F(-\infty),1-F(+\infty)]$.
-The next consequence of Proposition 4, which follows from Proposition 5 and is stated precisely in Appendix B.3, concerns the contractive property of $T$, namely,
-$\|TF-TG\|_{\sup}\leq\|F-G\|_{\sup}.$
-This is classical for operators with the same properties as $T$ defined instead on the larger space $B(X)$ of bounded functions on some set $X$, which is precisely the setting considered in [6].
-The fact that $T$ is defined only on $CDF(\overline{\mathbb{R}})$ makes the proof slightly less obvious, but contractivity holds nonetheless.
-The remainder of this section is organized as follows: In Section 3.1, Proposition 3 is proved. The formula for $\mathscr{L}$ on smooth CDF’s is then proved in Section 3.2, along with Proposition 5.
+Under the assumption that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$, the function $\mathscr{L}$ has the following two properties:
+  • (i)
+    $(\mathscr{L}F)(+\infty)=(\mathscr{L}F)(-\infty)=0$ for each $F\in CDF(\overline{\mathbb{R}})$.
+  • (ii)
+    Given $F,G\in CDF(\overline{\mathbb{R}})$, if the difference FGF-G is a constant function, then $\mathscr{L}F=\mathscr{L}G$.
+In particular, $T$ maps constant functions to constant functions: Its action on constants is determined by the dynamical system
+$Tq-q=(1-2p)q(1-q)\quad\text{for each}\,\,q\in[0,1].$
+As a consequence of property (ii) and the monotonicity of $T$, the following continuity property of $T$ follows, namely,
+$\|TF-TG\|_{\sup}\leq 2(1-p)\|F-G\|_{\sup}$
+In particular, $T$ is a contraction when $p=1/2$. See Appendix B.3 for the proof.
 3.1. Proof of Proposition 5
@@
 $\{Z\leq x\}=\{\Theta=1,X_{1}\leq x\,\,\text{and}\,\,X_{2}\leq x\}\cup\{\Theta=0,X_{1}\leq x\,\,\text{or}\,\,X_{2}\leq x\}.$
-Thus, since $X_{1}$ and $X_{2}$ are i.i.d. with CDF $F$, one obtains, exactly as in [18, Lemma 1.1],
+Thus, since $X_{1}$ and $X_{2}$ are i.i.d. with CDF $F$, one obtains, exactly as in [21, Lemma 1.1],
 $\mathbb{P}\{Z\leq x\}=pF(x)^{2}+(1-p)(1-(1-F(x))^{2}).$
@@
 Remark 1.
-In [18, Lemma 1.1], it is observed that the fixed points of the map $q\mapsto pq^{2}+(1-p)(1-(1-q)^{2})$ are fundamental in the study of the asymptotics of the series-parallel graph.
+In [21, Lemma 1.1], it is observed that the fixed points of the map $q\mapsto pq^{2}+(1-p)(1-(1-q)^{2})$ are fundamental in the study of the asymptotics of the series-parallel graph.
 Yet, as in the derivation above, a fixed point of this map is nothing but a zero of the polynomial (12p)q(1q)(1-2p)q(1-q), the “reaction term” appearing in the equation.
-3.1.1. Stationarity at infinity if $p=\frac{1}{2}$
-If $p=\frac{1}{2}$, then
-$TF(\pm\infty)-F(\pm\infty)=\mathscr{L}F(\pm\infty)$
-and inspection of the formula (31) shows that $\mathscr{L}F(\pm\infty)=0$ independently of $F$. ∎
 3.2. Proof of Proposition 6
-Suppose that $F\in CDF(\overline{\mathbb{R}})$ is smooth. To obtain the formula (32) for $\mathscr{L}F$ in this case, it suffices to establish that
+Suppose that $F\in CDF(\overline{\mathbb{R}})$ is continuously differentiable. To obtain the formula (33) for $\mathscr{L}F$ in this case, it suffices to establish that
 $\mathbb{P}\{Z\leq x<Y\mid\Theta=1\}=-2\int_{-\infty}^{x}\mathbf{E}[F(x_{1}+g_{+}(x-x_{1}))-F(x)]F^{\prime}(x_{1})\,dx_{1},$
@@
 Before going further, it is worth noting that the assumption that $f_{+}(+\infty)=0$ implies that $X_{1}>$-$\infty$ almost surely on the event $\{Z\leq x<Y,\,\,X_{1}\leq X_{2},\,\,\Theta=1\}$ for $x\in\mathbb{R}$.
-This is immediate from the fact that $Y=X_{2}=Z$ on $\{X_{1}=-\infty,\,\,\Theta=1,\,\,f_{+}(+\infty)=0\}$.
+This is immediate from the fact that $Y=X_{2}=Z$ on the event $\{X_{1}=-\infty,\,\,\Theta=1,\,\,f_{+}(+\infty)=0\}$.
 Subtracting $X_{1}$ from both sides of the equation defining $Y$ while assuming that $X_{2}\geq X_{1}$ leads to
@@
 $=\{-\infty<X_{1}<x,\,\,X_{1}+g_{+}(x-X_{1})<X_{2}\leq x\}.$
-Therefore, since $f_{+}$ is independent of $(X_{1},X_{2})$,
+Therefore, since $f_{+}$ is independent of $(X_{1},X_{2})$ and $\mathbb{P}\{X_{1}\leq X_{2}\}=\mathbb{P}\{X_{1}<X_{2}\}$ by the continuity of $F$,
 $\mathbb{P}\{Z\leq x<Y,X_{1}\leq X_{2}\mid\Theta=1\}$ $=-\int_{-\infty}^{x}\mathbf{E}[F(x_{1}+g_{+}(x-x_{1}))-F(x)]F^{\prime}(x_{1})\,dx_{1}.$
@@
 3.3. Proof of Proposition 7
-Suppose that $p=\frac{1}{2}$ and $F\in CDF(\overline{\mathbb{R}})$ satisfies $c_{-}\leq F\leq c_{+}$ pointwise for some $c_{-},c_{+}\in[0,1]$.
-Fix $c\in[-c_{-},1-c_{+}]$.
-The goal at this stage is to show that $T(F+c)=TF+c$.
-Let $X$ be an extended-real-valued random variable with CDF $F$.
-Consider the CDF $F_{n}$ of the random variable $X+n^{-1}Y$, where $Y$ is a mean-zero, variance-one Gaussian, independent of $X$.
-Then $F_{n}$ is a smooth function and a standard coupling argument shows that both $F_{n}\to F$ and $TF_{n}\to TF$ pointwise in $\overline{\mathbb{R}}$, except for points of discontinuity, as $n\to\infty$.
-Therefore, since the same reasoning applies to $F+c$, it suffices to consider the case when $F$ is smooth.
-Yet direct inspection of formula (32) shows that $\mathscr{L}(F+c)=\mathscr{L}F$ for smooth $F$.
+Here is the proof of (i): Fix $F\in CDF(\overline{\mathbb{R}})$. Observe that, by continuity of measure,
+$\lim_{x\to+\infty}\mathbb{P}\{Z\leq x<Y\,\mid\,\Theta=1\}$ $=\lim_{x\to+\infty}\left(\mathbb{P}\{Z\leq x\,\mid\,\Theta=1\}-\mathbb{P}\{Y\leq x\,\mid\,\Theta=1\}\right)$
+$=\mathbb{P}\{Z=+\infty\,\mid\,\Theta=1\}-\mathbb{P}\{Y=+\infty\,\mid\,\Theta=1\}=0$
+since $\{Y=+\infty\}=\{Z=+\infty\}$ by the boundedness of $f_{+}$ and $f_{-}$. By symmetry,
+$\lim_{x\to-\infty}\mathbb{P}\{Y\leq x<Z\,\mid\,\Theta=0\}=0.$
+Therefore, by (32), $(\mathscr{L}F)(+\infty)=(\mathscr{L}F)(-\infty)=0$ as claimed.
+Next, the proof of (ii): Suppose that $F,G\in CDF(\overline{\mathbb{R}})$ differ by a constant, hence $G=F+c$ for some $c\in\mathbb{R}$.
+First, if $F$ is smooth, then equation (33) applied to $F$ and $F+c$ implies $\mathscr{L}(F+c)=\mathscr{L}F$.
+In general, there is a sequence $\{F_{n}\}$ of smooth CDF’s that converges vaguely to $F$.
+By Proposition 1, $\{TF_{n}\}$ and $\{T(F_{n}+c)\}$ converge vaguely to respective limits $TF$ and $T(F+c)$.
+At the same time, since the polynomial P(q)=q+(12p)q(1q)P(q)=q+(1-2p)q(1-q) is smooth, the functions $\{P(F_{n})\}$ converge vaguely to $\{P(F)\}$.
+Therefore, for a dense set of points $x\in\mathbb{R}$,
+$T(F+c)(x)-P(F+c)(x)=\lim_{n\to\infty}\mathscr{L}(F_{n}+c)(x)=\lim_{n\to\infty}\mathscr{L}F_{n}(x)=TF(x)-PF(x).$
+Rearranging, this implies $T(F+c)+PF=TF+P(F+c)$ on a dense set.
+Since right- and lefthand sides are right-continuous (being sums of such functions), this actually implies they are equal everywhere, and, thus,
+$\mathscr{L}(F+c)=T(F+c)-P(F+c)=TF-PF=\mathscr{L}F.$
+Finally, consider a constant CDF $F\equiv q$ for some $q\in[0,1]$. Writing $TF=\mathcal{L}F+P(F)$ as in the first part of the proof, observe that
+$(TF)(+\infty)=(\mathcal{L}F)(+\infty)+P(q)=P(q),$
+and, similarly, $TF(-\infty)=P(q)$.
+Therefore, being nondecreasing, $TF$ must be a constant function, equal to $P(q)$.
+Identifying $F$ with $q$, this proves that $T$ maps $q$ to the constant $Tq=P(q)$, as claimed.
+∎
 4. Examples
-This section describes a few relevant examples, namely, those that have appeared earlier in the literature, and describes some of the useful, concrete forms of $\mathscr{L}$ that can be derived.
+This section describes a few relevant examples and consequences of the main theorems, Theorems 1 and 2.
 4.1. Resistance of the series-parallel graph
@@
 $g(s)=\left\{\begin{array}[]{r l}\log(e^{s}-1),&\text{if}\,\,s\geq\log 2,\\ 0,&\text{otherwise.}\end{array}\right$.
-It is immediate to check that $f\in\mathcal{S}$ and sg(s)=f(g(s))s-g(s)=f(g(s)) for $s\geq\log 2$ so assumption (15) of Theorem 1 is satisfied.
+It is immediate to check that $f\in\mathcal{S}$ and sg(s)=f(g(s))s-g(s)=f(g(s)) for $s\geq\log 2$ so the stronger assumption (16) of Theorems 1 and 2 is satisfied.
 In this case, the constant $a$ is given by $a=2\zeta(3)$. See Appendix C for the proof.
+The approach of this paper readily implies the following characterization of the asymptotics of the resistance $R^{(n)}$ of the critical series-parallel graph, in which resistors have i.i.d.
+resistances with any given law on $0,+\infty]$.
+The result allows for the possibility that $\mathbb{P}\{R^{(0)}=0\}>0$ or $\mathbb{P}\{R^{(0)}=+\infty\}>0$, that is, any given resistor of the graph could be open or shorted.
+Corollary 2.
+Assume that $p=1/2$ and let $R^{(0)}$ be any random variable taking values in $0,+\infty]$.
+Let $R^{(n)}$ be the resistance of the series-parallel graph at stage $n$, in which each edge has i.i.d.
+resistances with the same law as $R^{(0)}$, and let $q=\mathbb{P}\{0<R^{(0)}<+\infty\}$.
+Then $\mathbb{P}\{0<R^{(n)}<+\infty\}=q$ for each $n$ and the law of $R^{(n)}$ conditional on this event has the following limit:
+$\lim_{N\to\infty}\mathbb{P}\{(36a)^{-\frac{1}{3}}N^{-\frac{1}{3}}\log R^{(N)}\leq x\,\mid\,0<R^{(N)}<+\infty\}=F_{\text{Beta}(2,2)}(q^{-\frac{1}{3}}x+\frac{1}{2}),$
+where $F_{\text{Beta}(2,2)}$ is the CDF of the $\text{Beta}(2,2)$ distribution and $a=2\zeta(3)$.
+Recall that the $\text{Beta}(2,2)$ distribution is determined by the PDF $6y(1-y)\mathbf{1}_{[0,1]}(y)$.
+Proof.
+Let $X^{(n)}=\log R^{(n)}$.
+As in the discussion in the introduction, $\{X^{(n)}\}$ is a solution of the RDE (9) with $f_{+}(u)=f_{-}(u)=\log(1+e^{-u})$.
+Let $q_{-}=\mathbb{P}\{X^{(0)}=-\infty\}$ and $q_{+}=\mathbb{P}\{X^{(0)}=+\infty\}$ so that $1-q=q_{+}+q_{-}$.
+Notice that
+$\lim_{N\to\infty}\mathbb{P}\{N^{-\frac{1}{3}}X^{(0)}\leq x\}=q_{-}+q\mathbf{1}_{[0,+\infty)}(x)\quad\text{for each}\,\,x\in\mathbb{R}\setminus\{0\}.$
+Let $F_{N}:\mathbb{R}\times[0,+\infty)\to[0,1]$ be the rescaled CDF of $\{X^{(n)}\}$ defined as in Theorem 2.
+By that theorem, $F_{N}\to F$ locally uniformly in $\mathbb{R}\times(0,+\infty)$ as $N\to+\infty$, where $F$ is the bounded discontinuous viscosity solution of the PDE
+$\partial_{t}F-a|\partial_{x}F|\partial_{x}^{2}F=0\quad\text{in}\,\,\mathbb{R}\times(0,+\infty),\quad F(x,0)=q_{-}+q\mathbf{1}_{[0,+\infty)}(x).$
+$F$ can be readily related to the solution of the same PDE, but with initial datum $\mathbf{1}_{[0,+\infty)}$ and a different coefficient $a$: Notice that if $G$ is defined by $G(x,t)=q^{-1}(F(x,t)-q_{-})$, then $G(x,0)=\mathbf{1}_{[0,+\infty)}(x)$ and
+$\partial_{t}G-q^{2}a|\partial_{x}G|\partial_{x}^{2}G=0\quad\text{in}\,\,\mathbb{R}\times(0,+\infty),\quad G(x,0)=\mathbf{1}_{[0,+\infty)}(x).$
+As in the proof of Corollary 1 (see Section 5.3) below, $G$ can be explicitly computed to be
+$G(x,t)=F_{\text{Beta}(2,2)}(U(x,t)),\quad\text{where}\quad U(x,t)=\frac{x}{(36qa)^{\frac{1}{3}}}+\frac{1}{2}.$
+At the same time, since $\mathbb{P}\{0<R^{(N)}<+\infty\}=q$ and $\mathbb{P}\{R^{(N)}>0\}=q_{-}$ for any $N$ by Proposition 7,
+$\lim_{N\to+\infty}\mathbb{P}\{N^{-\frac{1}{3}}\log R^{(N)}\leq x\,\mid\,0<R^{(N)}<+\infty\}=\lim_{N\to+\infty}\frac{1}{q}(F_{N}(x,t)-q_{-})=G(x,t).$
 4.2. Distance on the series-parallel graph
@@
 Corollary 3.
-Assume that $p=\frac{1}{2}$ and $\mathbb{P}\{0\leq D^{(0)}<\infty\}=1$ and let $q=\mathbb{P}\{D^{(0)}>0\}$. For any $n$, $\mathbb{P}\{D^{(n)}=0\}=q$. Further,
-$\lim_{N\to\infty}\mathbb{P}\{-\frac{\text{sgn}(\sigma)}{2\sqrt{|\sigma|}\sqrt{N}}\log D^{(N)}\leq x\,\mid\,D^{(N)}>0\}=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ \frac{x^{2}}{q},&\text{if}\,\,0\leq x\leq\sqrt{q},\\ 1,&\text{otherwise.}\end{array}\right$.
+Assume that $p=\frac{1}{2}$ and $\mathbb{P}\{0\leq D^{(0)}<+\infty\}=1$ and let $q=\mathbb{P}\{D^{(0)}>0\}$.
+Then $\mathbb{P}\{D^{(n)}>0\}=q$ for any $n$ and the law of $D^{(n)}$ conditional on this event has the following limit:
+$\lim_{N\to\infty}\mathbb{P}\{N^{-\frac{1}{2}}\log D^{(N)}\leq 2x\sqrt{|\sigma|}\,\mid\,D^{(N)}>0\}=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ \frac{x^{2}}{q},&\text{if}\,\,0\leq x\leq\sqrt{q},\\ 1,&\text{otherwise.}\end{array}\right$.
 Here the constant $\sigma=-\frac{1}{2}\zeta(2)=-\frac{\pi^{2}}{12}$.
 Proof.
-Let $X^{(n)}=\log D^{(n)}$ so that $\{D^{(n)}>0\}=\{X^{(n)}>-\infty\}$. By Proposition 3, $\mathbb{P}\{X^{(n)}>-\infty\}$ is constant in $n$.
-Let $F_{n}$ be the CDF of $X^{(n)}$ and let $G_{n}$ be the CDF conditional on $\{X^{(n)}>-\infty\}$, that is,
-$G_{n}(x)=\mathbb{P}\{X^{(n)}\leq x\,\mid\,X^{(n)}>-\infty\}.$
-Notice that $G_{n}=q^{-1}(F_{n}-(1-q))$.
-At the same time, the strengthened version of Theorem 1 from Section 5 (Theorem 2) applies: If $F_{\epsilon}(x,t)=F_{[\epsilon^{-2}t]}(\epsilon^{-1}x)$, then the theorem implies that
-$F_{\epsilon}\to F\quad\text{locally uniformly in}\,\,\mathbb{R}\times(0,\infty),$
-where $F$ is the unique upper semicontinuous viscosity solution of the PDE $\partial_{t}F-\sigma|\partial_{x}F|^{2}=0$ with initial datum $F(x,0)=(1-q)+q\mathbf{1}_{[0,\infty)}(x)$.
-After writing $G_{\epsilon}(x,t)=G_{[\epsilon^{-2}t]}(\epsilon^{-1}x)$, it follows from this that $G_{\epsilon}\to G$, where $G$ is the unique discontinuous viscosity solution of $\partial_{t}G-q\sigma|\partial_{x}G|^{2}=0$ with $G(x,0)=\mathbf{1}_{[0,\infty)}$.
-The proof of Corollary 1 in Section 5 gives an explicit formula for this function, from which the claimed asymptotics of $\log D^{(N)}$ follow.
-The claimed value of $\sigma$ is computed in Appendix C and is consistent with [5]. ∎
-It is worth emphasizing that $D^{(n)}$ is integer-valued, hence its logarithm takes values in the set $\log\mathbb{N}$ for all $n$.
-For the present choice of $f_{+}$ and $f_{-}$, it is not hard to show that this remains true for any solution that starts initially in $\log\mathbb{N}$.
-This is worth mentioning because the equation for the CDF of such a solution takes a particularly simple form, namely, if the backward-difference operator $\nabla^{-}_{\log\mathbb{N}}$ is defined for functions on $\log\mathbb{N}$ by $\nabla^{-}_{\log\mathbb{N}}F(x)=F(x)-F(x_{-})$, where $x_{-}$ denotes the largest element of $\log\mathbb{N}$ smaller than $x$, then
-$F_{n+1}-F_{n}=-p(\nabla^{-}_{\log\mathbb{N}}F)^{2}+(1-2p)F_{n}(1-F_{n})\quad\text{in}\,\,\log\mathbb{N}.$
-Notice that this equation looks like a discrete-time, discrete-in-space version of (23) with $a=0$.
+The claimed value of $\sigma$ is computed in Appendix C and is consistent with [6].
+The limiting behavior of $D^{(N)}$ can be derived following the same strategy as in Corollary 2 above, once again using the fact that $\mathbb{P}\{D^{(n)}>0\}=q$ for any $n$.
+∎
 4.3. $\mathbb{Z}$-valued Solutions and Finite-Difference Equations
@@
 where $g_{\mathbb{Z}}$ denotes the left-inverse of the function $u\mapsto u+f_{\mathbb{Z}}(u)$.
-For such a law $\mathbf{P}$, all the assumptions of Theorem 1 are satisfied.
-In particular, as long as $(f_{+},f_{-})$ is not identically zero, either $n^{-\frac{1}{2}}X^{(n)}$ or $n^{-\frac{1}{3}}X^{(n)}$ converges to a nondegenerate limit.
+Thus, such a law $\mathbf{P}$ satisfies the assumptions of Theorems 1 and 2.
+In particular, as long as $(f_{+},f_{-})$ are not identically zero, either $n^{-\frac{1}{2}}X^{(n)}$ or $n^{-\frac{1}{3}}X^{(n)}$ converges to a nondegenerate limit.
 Let $q_{+}=\mathbf{P}\{f_{+}=f_{\mathbb{Z}}\}$ and $q_{-}=\mathbf{P}\{f_{-}=f_{\mathbb{Z}}\}$ in the discussion that follows.
@@
 In fact, by this reasoning, it is also possible to draw a connection with the other RDE considered in [1], called the totally asymmetric hipster random walk.
-This is the RDE in $\mathbb{Z}$ with the one-step update determined by two independent Bernoulli random variables $\Theta,\Xi$ via the rule
+This is the RDE in $\mathbb{Z}$ with the one-step update determined by two independent Bernoulli random variables $(\Theta,\Xi)$ via the rule
 $Y\overset{d}{=}\left\{\begin{array}[]{r l}X_{1}+\Xi\mathbf{1}_{\{0\}}(|X_{2}-X_{1}|),&\text{if}\,\,\Theta=1,\\ X_{2}+\Xi\mathbf{1}_{\{0\}}(|X_{2}-X_{1}|),&\text{if}\,\,\Theta=0.\end{array}\right$.
@@
 Two-player totally asymmetric, $q$-lazy cooperative motion is monotone only for $(r,q)$ in the set given above; this can readily be proved via differentiation of the evolution equation satisfied by the CDF (see [2, Equation (1.6)]).
-In view of the proposition above, the $\text{Beta}(2,2)$ limit theorems for the symmetric RDE’s considered in [1] and [3] follow as a special case of this paper’s main theorem (Theorem 1) and Corollary 1.
+In view of the proposition above, the $\text{Beta}(2,2)$ limit theorems for the symmetric RDE’s considered in [1] and [3] follow as a special case of this paper’s main results.
 The $\text{Beta}(2,1)$ limit theorems for asymmetric RDE’s obtained in [1] and [2] also follow, albeit only in the parameter regimes identified above.
-Interestingly, in the asymmetric case, those works obtain $\text{Beta}(2,1)$ limits, even in the nonmonotone regime.
+Interestingly, in the asymmetric case, those works were able to obtain limit theorems even in the nonmonotone regime.
 5. Scaling Limit
-This section proves the main result, Theorem 1.
-In fact, a slightly more general version is presented here, which allows for extended-real-valued solutions and diffuse initial conditions.
-In order to state the theorem efficiently, as in the introduction, let $\alpha=2$ if $\sigma\neq 0$ and $\alpha=3$, otherwise, and define the nonlinearity $\mathcal{F}$ appearing in the limiting PDE analogously:
-(48) $\mathcal{F}(p,X)=\left\{\begin{array}[]{r l}\sigma|p|^{2},&\text{if}\,\,\sigma\neq 0,\\ a|p|X,&\text{otherwise.}\end{array}\right$.
-The theorem that follows involves the notion of viscosity solutions of parabolic PDE’s. The relevant definitions and existence and uniqueness results are reviewed in Appendix A below.
+This section proves the main results, Theorems 1 and 2, as well as Corollary 1.
+Fix data $(f_{+},f_{-})$ satisfying the assumptions of Theorems 1 or 2. Given $\theta\in\mathbb{R}$, define a sequence $\{p^{(N)}\}$ by
+$p^{(N)}=1/2+\theta N^{-1}.$
+Let $T^{(N)}$ be the operator associated with the RDE (9) with $\Phi$ given by (8) and bias parameter $p=p^{(N)}$.
+In view of the results of Section 3, the equation $F_{n}=T^{(N)}F_{n-1}$ can be written in the form
+(47) $F_{n}-F_{n-1}=\mathcal{L}^{(N)}F_{n-1}-2\theta N^{-1}F_{n-1}(1-F_{n-1})$
+for some function $\mathcal{L}^{(N)}$ defined on $CDF(\overline{\mathbb{R}})$.
+This equation, together with the properties of $\mathcal{L}^{(N)}$ obtained in Section 3, motivates the study of scaling limits of certain sequences of monotone operators defined on $CDF(\overline{\mathbb{R}})$ in an abstract setting.
+5.1. Abstract Framework
+Let $\{T^{(N)}\}$ be a family of operators on $CDF(\overline{\mathbb{R}})$ with the following properties:
+  • (i)
+    Monotonicity: If $F\leq G$ pointwise in $\mathbb{R}$, then $T^{(N)}F\leq T^{(N)}G$ also holds.
+  • (ii)
+    Discrete-Time Reaction-Diffusion Equation: There are functions $\{\mathcal{L}^{(N)}\}$ defined on $CDF(\overline{\mathbb{R}})$ and functions $\{Q^{(N)}\}$ defined on $[0,1]$ such that
+    $T^{(N)}F-F=\mathcal{L}^{(N)}F-Q^{(N)}(F).$
+  • (iii)
+    Commutation with Constants: Given $F,G\in CDF(\overline{\mathbb{R}})$, if the difference FGF-G is a constant function, then $\mathcal{L}^{(N)}F=\mathcal{L}^{(N)}G$.
+  • (iv)
+    Action at Infinity: If $F\in CDF(\overline{\mathbb{R}})$, then, for each $\bar{x}\in\{+\infty,-\infty\}$,
+    $(T^{(N)}F)(\bar{x})=F(\bar{x})-Q^{(N)}(F(\bar{x})).$
+Notice that, as in the proof of Proposition 7 above, assumption (iv) implies that $\mathcal{L}^{(N)}$ vanishes on constant functions and, thus, $T^{(N)}$ maps constant functions to constant functions via the dynamical system
+$T^{(N)}q=q-Q^{(N)}(q)\quad\text{for each}\,\,q\in[0,1].$
+Since $T^{(N)}$ is monotone and maps $CDF(\overline{\mathbb{R}})$ to itself, it follows that the function $q\mapsto q-Q^{(N)}(q)$ is a nondecreasing map that sends $[0,1]$ into itself.
+In addition to (i)-(iv) above, assume that there is a sequence of positive numbers $\{\delta^{(N)}\}$ such that $\delta^{(N)}\to 0$ as $N\to+\infty$ and a continuous function $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, which is nondecreasing in the second variable, for which the following property holds:
+  • (v)
+    Consistency of $\mathcal{L}^{(N)}$: If $\Phi\in CDF(\overline{\mathbb{R}})$ is smooth with bounded, uniformly continuous second derivatives, rescaled according to the formula $\Phi_{\delta}(y)=\Phi(\delta y)$, then
+    $\lim_{N\to+\infty}N(\mathscr{L}^{(N)}\Phi_{\delta^{(N)}})\left(\frac{x}{\delta^{(N)}}\right)=\mathcal{F}(\partial_{x}\Phi(x),\partial_{x}^{2}\Phi(x)),$
+    where convergence holds uniformly with respect to $x\in\mathbb{R}$.
+Since $\mathcal{L}^{(N)}$ vanishes on constant functions for any $N$, assumption (v) implies
+(48) $\mathcal{F}(0,0)=0.$
+An analogous assumption will be made on the reaction term $Q^{(N)}$, as follows:
+  • (vi)
+    Consistency of $Q^{(N)}$: There is a function $Q:[0,1]\to[0,1]$ such that
+    $\lim_{N\to\infty}\sup\left\{|NQ^{(N)}(q)-Q(q)|\,\mid\,q\in[0,1]\right\}=0$
+    Further, there is a constant $L>0$ such that, independently of $N$,
+    $|Q^{(N)}(q)-Q^{(N)}(q^{\prime})|\leq LN^{-1}|q-q^{\prime}|\quad\text{for each}\,\,q,q^{\prime}\in[0,1].$
+Under these assumptions, sequences of solutions $\{F^{(N)}_{n}\}$ of the recursion $F^{(N)}_{n}=T^{(N)}F^{(N)}_{n-1}$ undergo a scaling limit, whereby their asymptotic behavior is determined by the PDE $\partial_{t}F=\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)-Q(F)$.
+This is made precise in the next theorem.
 Theorem 3.
-Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds and fix an arbitrary $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$.
-For each $\epsilon>0$, let $\{X_{\epsilon}^{(n)}\}$ be an extended-real-valued solution of the RDE (9) and let $\{F^{\epsilon}_{n}\}$ be the associated sequence of CDF’s.
-Let $F_{\epsilon}:\mathbb{R}\times[0,\infty)\to[0,1]$ be the rescaled CDF given by
-$F_{\epsilon}(x,t)=F^{\epsilon}_{[\epsilon^{-\alpha}t]}(\epsilon^{-1}x).$
-If convergence holds as $\epsilon\to 0$ at the initial time in the sense that
-$F_{\epsilon}(x,0)\to F_{\text{in}}(x)\quad\text{for each point of continuity}\,\,x\,\,\text{of}\,\,F_{\text{in}},$
+Assume that $\{T^{(N)}\}$ is a family of operators on $CDF(\overline{\mathbb{R}})$ satisfying assumptions (i)-(vi) above.
+For each $N$, let $\{F^{(N)}_{n}\}$ be a sequence in $CDF(\overline{\mathbb{R}})$ generated by the recursion $F_{n}^{(N)}=T^{(N)}F^{(N)}_{n-1}$.
+If $F_{N}:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is the rescaled CDF defined by
+$F_{N}(x,t)=F^{(N)}_{[Nt]}\left(\frac{x}{\delta^{(N)}}\right),$
+and if there is a continuous $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ such that, at time $t=0$,
+$F^{(N)}(\cdot,0)\to F_{\text{in}}\quad\text{vaguely as}\,\,N\to+\infty,$
 then
-$F_{\epsilon}\to F\quad\text{locally uniformly in}\,\,\mathbb{R}\times(0,\infty),$
-where $F$ is the unique bounded discontinuous viscosity solution of the initial-value problem
-$\left\{\begin{array}[]{r l}\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
-The theorem will be proved by following the approach of [6], suitably extended to monotone semigroups in the space $CDF(\overline{\mathbb{R}})$.
-Notice that if $T_{\epsilon}$ and $\mathscr{L}_{\epsilon}$ are the operators on $CDF(\overline{\mathbb{R}})$ defined by
-(49) $(T_{\epsilon}G)(x)=(TG_{\epsilon})(\epsilon^{-1}x),\quad\text{where}\,\,G_{\epsilon}(x)=G(\epsilon^{-1}x),\quad\mathscr{L}_{\epsilon}=T_{\epsilon}-\text{Id},$
-then the rescaled CDF $F_{\epsilon}$ above satisfies the equation
-(50) $F_{\epsilon}(x,t+\epsilon^{\alpha})-F_{\epsilon}(x,t)=\mathscr{L}_{\epsilon}F_{\epsilon}(x,t).$
-This motivates the consideration of one-parameter families of operators $\{T_{\epsilon}\}$ defined on $CDF(\overline{\mathbb{R}})$.
-5.1. Abstract Framework
-Let $\{T_{\epsilon}\}$ be a family of operators on $CDF(\overline{\mathbb{R}})$ with the following properties:
-  • (i)
-    Monotonicity: If $F\leq G$ pointwise in $\mathbb{R}$, then $T_{\epsilon}F\leq T_{\epsilon}G$ also holds.
-  • (ii)
-    Commutation with addition of constants: If $F\in CDF(\overline{\mathbb{R}})$ satisfies $c_{-}\leq F\leq c_{+}$ pointwise for some constants $c_{-},c_{+}\in[0,1]$, then $T_{\epsilon}(F+c)=T_{\epsilon}F+c$ for any $c\in[-c_{-},1-c_{+}]$.
-  • (iii)
-    Stationarity at infinity: For any $F\in CDF(\overline{\mathbb{R}})$, $T_{\epsilon}F(+\infty)=F(+\infty)$ and $T_{\epsilon}F(-\infty)=F(-\infty)$.
-Motivated by (50), let $\{\mathscr{L}_{\epsilon}\}$ be the operators on $CDF(\overline{\mathbb{R}})$ defined by
-$\mathscr{L}_{\epsilon}=T_{\epsilon}-\text{Id}.$
-In addition to (i) and (ii), assume that there is an exponent $\alpha>0$ and a continuous function $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, which is increasing in the second variable, for which the following property holds:
-  • (iv)
-    Consistency: If $\Phi\in CDF(\overline{\mathbb{R}})$ is smooth with bounded, uniformly continuous second derivatives, then
-    $\lim_{\epsilon\to 0}\epsilon^{-\alpha}\mathscr{L}_{\epsilon}\Phi(x)=\mathcal{F}(\partial_{x}\Phi(x),\partial_{x}^{2}\Phi(x)),$
-    where convergence holds uniformly with respect to $x\in\mathbb{R}$.
-Note that (iii) implies $Tc=c$ for any constant (function) $c\in[0,1]$, and, thus, by (iv),
-(51) $\mathcal{F}(0,0)=0.$
-Under these assumptions, the following convergence result holds for rescaled solutions $\{F^{\epsilon}_{n}\}$ of the recursion $F^{\epsilon}_{n+1}=T_{\epsilon}F^{\epsilon}_{n}$.
-Theorem 3.
-Assume that $\{T_{\epsilon}\}$ is a family of operators on $CDF(\overline{\mathbb{R}})$ satisfying assumptions (i)-(iv) above.
-For each $\epsilon$, let $\{F^{\epsilon}_{n}\}$ be a sequence in $CDF(\overline{\mathbb{R}})$ generated by the recursion $F_{n+1}^{\epsilon}=T_{\epsilon}F_{n}^{\epsilon}$.
-If $F_{\epsilon}:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is the rescaled CDF defined by
-$F_{\epsilon}(x,t)=F^{\epsilon}_{[\epsilon^{-\alpha}t]}(x),$
-and if there is a continuous $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ such that $F_{\epsilon}(x,0)\to F_{\text{in}}(x)$ pointwise as $\epsilon\to 0$, then
-$F_{\epsilon}\to F\quad\text{locally uniformly in}\,\,\mathbb{R}\times(0,\infty),$
-where $F$ is the unique bounded continuous viscosity solution of the equation
-(52) $\left\{\begin{array}[]{r l}\partial_{t}F=\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
-Notice that the one-parameter family $\{T_{\epsilon}\}$ of interest in this paper (see (49)) satisfies assumptions (i)-(iii) according to Propositions 1 and 5.
-It only remains to verify condition (iv).
+$F_{N}\to F\quad\text{locally uniformly in}\,\,\mathbb{R}\times[0,\infty)\,\,\text{as}\,\,N\to+\infty,$
+where $F$ is the unique bounded viscosity solution of the equation
+(49) $\left\{\begin{array}[]{r l}\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
+It is possible to show that if $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds, as is assumed in Theorems 1 and 2, then the one-parameter family $\{T^{(N)}\}$ defined in the discussion at the beginning of this section satisfies assumptions (i)-(iv); see Section 5.2 for the details.
+In particular, $\{Q^{(N)}\}$ is given by $Q^{(N)}(q)=2\theta N^{-1}q(1-q)$, which trivially satisfies (vi).
+It remains to verify condition (v).
 That is the subject of the next proposition.
+Recall that the constants $\sigma$ and $a$ are defined in Section 1.4.1 above.
 Proposition 10.
-Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds, and consider the operator $\mathscr{L}$ given by (31).
-If $\Phi:\mathbb{R}\to[0,1]$ is a nondecreasing function with bounded, uniformly continuous second derivative, rescaled according to the rule $\Phi_{\delta}(x)=\Phi(\delta x)$, and if $R_{\Phi}$ is defined by the identity
-$(\mathscr{L}\Phi_{\delta})(\delta^{-1}x)=\delta^{2}\sigma|\partial_{x}\Phi(x)|^{2}+\delta^{3}a|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)+\delta^{3}R_{\Phi}(\delta;x),$
-then
-$\lim_{\delta\to 0}\sup\left\{|R_{\Phi}(\delta;x)|\,\mid\,x\in\mathbb{R}\right\}=0.$
-The proof of the proposition is relegated to the end of this section.
-There is one last technicality to address, namely, whereas Theorem 3 asks for the continuity of the initial datum $F_{\text{in}}$, the main theorem of this section allows for discontinuities.
-This gap is bridged using the fact that the equations of interest here are regularizing.
+Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds, and fix $\theta\in\mathbb{R}$.
+Define $p^{(N)}=1/2+\theta N^{-1}$.
+Consider the operator $\mathscr{L}^{(N)}$ given by (32) with $p=p^{(N)}$.
+If $\Phi:\mathbb{R}\to[0,1]$ is a nondecreasing function with bounded, uniformly continuous first and second derivatives, rescaled according to the rule $\Phi_{\delta}(y)=\Phi(\delta y)$, and if $R_{\Phi}^{(N)}:\mathbb{R}\to\mathbb{R}$ is the remainder defined by
+$(\mathscr{L}^{(N)}\Phi_{N^{-\frac{1}{2}}})(N^{\frac{1}{2}}x)=N^{-1}\sigma|\partial_{x}\Phi(x)|^{2}+N^{-1}R_{\Phi}^{(N)}(x),$
+then $R^{(N)}_{\Phi}\to 0$ uniformly in $\mathbb{R}$ as $N\to+\infty$.
+Further, if the stronger assumption (16) holds, and if $R^{(N)}_{\Phi}$ is instead defined by
+$(\mathscr{L}^{(N)}\Phi_{N^{-\frac{1}{3}}})(N^{\frac{1}{3}}x)=N^{-\frac{2}{3}}\sigma|\partial_{x}\Phi(x)|^{2}+N^{-1}a|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)+N^{-1}R_{\Phi}^{(N)}(x),$
+then, as before, $R^{(N)}_{\Phi}\to 0$ uniformly as $N\to+\infty$.
+The proof of the proposition is relegated to the end of this section (Section 5.4).
+There is one last technicality to address, namely, whereas Theorem 3 asks for the continuity of the initial datum $F_{\text{in}}$, Theorems 1 and 2 allow for discontinuities.
+This gap is bridged using the fact that the equations of interest here are well-posed even for discontinuous data.
 This is made precise in the next condition:
-  • (v)
-    Asymptotically regularizing: If $\underline{F}$ and $\overline{F}$ denote the minimal and maximal viscosity sub- and supersolutions of the initial-value problem (52) for a given $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ (see Appendix A.1), then
-    $\underline{F}=\overline{F}\quad\text{in}\,\,\mathbb{R}\times(0,\infty).$
-The next corollary asserts that if the family $\{T_{\epsilon}\}$ satisfies assumption (v) in addition to (i)-(iv), then Theorem 3 remains true even if $F_{\text{in}}$ has discontinuities.
+  • (vii)
+    Asymptotically regularizing: If $\overline{F}$ and $\underline{F}$ denote the maximal and minimal viscosity sub- and supersolutions, respectively, of the initial-value problem (49) for a given $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ (see Appendix A.1), then
+    $\overline{F}=\underline{F}\quad\text{in}\,\,\mathbb{R}\times(0,\infty).$
+As explained in Appendix A, from the fact that $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,+\infty)$, it follows that there is a unique solution of (49) started from any initial datum $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$, which is continuous for positive times.
+The next corollary asserts that if the family $\{T^{(N)}\}$ satisfies assumption (vii) in addition to (i)-(vi), then Theorem 3 remains true even if $F_{\text{in}}$ has discontinuities.
 Corollary 4.
-Assume that $\{T_{\epsilon}\}$ is a family of operators on $CDF(\overline{\mathbb{R}})$ satisfying assumptions (i)-(v) above.
-Then the statement of Theorem 3 remains true for an arbitrary $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ provided $F$ is taken to be the unique discontinuous viscosity solution of the initial-value problem (52).
+Assume that $\{T^{(N)}\}$ is a family of operators on $CDF(\overline{\mathbb{R}})$ satisfying assumptions (i)-(vii) above.
+Then the statement of Theorem 3 remains true for an arbitrary $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ provided $F$ is taken to be the unique bounded discontinuous viscosity solution of the initial-value problem (49).
 See Appendix B for the proof.
-In Theorem 2, the nonlinearity $\mathcal{F}$ takes the form (48), which satisfies assumption (v) by Propositions 10 and 11 below. Therefore, the proof of the theorem is complete.
+It follows from classical results that, in the PDE’s of interest here, assumption (vii) holds as long as either $\sigma\neq 0$ or $a>0$.
+(The degenerate PDE’s $\partial_{t}F-Q(F)=0$ do not have this property and therefore require a slightly different treatment.) This is made precise in the next proposition.
+Proposition 11.
+Let $Q$ be an arbitrary function for which there is an $L>0$ such that $|Q(q)-Q(q^{\prime})|\leq L|q-q^{\prime}|$ for each $q,q^{\prime}\in[0,1]$ and satisfying both $Q(0)\geq 0$ and $Q(1)\leq 0$.
+If $\mathcal{F}(v,w)=\sigma|v|^{2}$ for some $\sigma\neq 0$ or $\mathcal{F}(v,w)=a|v|w$ for some $a>0$, then assumption (vii) is satisfied.
+Proof.
+When $\theta=0$, this is established in Propositions 14 and 15 in Appendix A. The improvement to the case $\theta\neq 0$ is covered by Proposition 16. ∎
 Remark 2.
@@
 In fact, let $F_{-}$ and $F^{+}$ denote the left- and right-increasing versions of a nondecreasing function $F$.
-If operators $\{T_{\epsilon,-}\}$ are defined on the space $CDF_{\text{left}}(\overline{\mathbb{R}})$ of nondecreasing, left-continuous functions taking values in $[0,1]$ by the formula $T_{\epsilon,-}F=(T_{\epsilon}(F^{+}))_{-}$, then $\{T_{\epsilon,-}\}$ satisfies assumptions (i)-(iv) if and only if $\{T_{\epsilon}\}$ does.
-Further, it is not hard to see that a PDE scaling limit of $\{T_{\epsilon}\}$ holds as in Theorem 3 if and only if the same conclusion applies to $\{T_{\epsilon,-}\}$.
+If operators $\{T^{(N)}_{-}\}$ are defined on the space $CDF_{\text{left}}(\overline{\mathbb{R}})$ of nondecreasing, left-continuous functions taking values in $[0,1]$ by the formula $T^{(N)}_{-}F=(T^{(N)}(F^{+}))_{-}$, then $\{T^{(N)}_{-}\}$ satisfies assumptions (i)-(vi) if and only if $\{T^{(N)}\}$ does.
+Further, it is not hard to see that a PDE scaling limit of $\{T^{(N)}\}$ holds as in Theorem 3 if and only if the same conclusion applies to $\{T^{(N)}_{-}\}$.
+5.2. Proof of Theorems 1 and 2
+Both theorems are direct applications of Theorem 3, or rather its corollary, Corollary 4.
+As in the discussion at the beginning of this section, let $T^{(N)}$ be the operator associated with the RDE (9) with data $(f_{+},f_{-})$ satisfying $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) and bias parameter $p^{(N)}=1/2+\theta N^{-1}$ for some fixed constant $\theta\in\mathbb{R}$.
+Assumptions (i) and (ii) follow from Propositions 2 and 5, respectively.
+The assumptions (iii) and (iv) hold by Proposition 7.
+To see that this proposition is applicable, it only remains to check that $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$.
+Fix $\rho\in\{+,-\}$.
+Since $s-g_{\rho}(s)=f_{\rho}(g_{\rho}(s))$ for $s\geq f_{\rho}(0)$ and $g_{\rho}(s)\to+\infty$ as $s\to+\infty$, the integral $\int_{f_{\rho}(0)}^{+\infty}|g_{\rho}(s)-s|\,ds$ is finite only if $f_{\rho}(+\infty)=0$.
+At the same time, (15) implies that this integral is finite almost surely, hence $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$ holds as claimed.
+Thus, the hypotheses of Proposition 7 are satisfied, and it follows that assumptions (iii) and (iv) both apply to the sequence $\{T^{(N)}\}$.
+In the context of Theorem 1 (i.e., assuming both (15) and $\sigma\neq 0$), Proposition 10 implies that assumption (v) holds with
+$\delta^{(N)}=N^{-\frac{1}{2}}\quad\text{and}\quad\mathcal{F}(v,w)=\sigma|v|^{2}.$
+According to Proposition 11 above, from the fact that $\sigma\neq 0$, assumption (vii) also holds.
+Therefore, Corollary 4 applies, establishing convergence of the rescaled CDF $F_{N}$ as desired to the solution of the initial-value problem associated to the PDE $\partial_{t}F-\sigma|\partial_{x}F|^{2}+2\theta F(1-F)=0$, as desired.
+Similarly, in Theorem 2 (that is, assuming that (16) holds, $\sigma=0$, and $a>0$), Proposition 10 instead establishes that assumption (v) holds with
+$\delta^{(N)}=N^{-\frac{1}{3}}\quad\text{and}\quad\mathcal{F}(v,w)=a|v|w.$
+Once again, because $a>0$, Proposition 11 asserts that assumption (vii) is satisfied.
+Therefore, by Corollary 4, the rescaled CDF $F_{N}$ instead converges to the solution of the initial-value problem associated to $\partial_{t}F-a|\partial_{x}F|\partial_{x}^{2}F+2\theta F(1-F)=0$.
+∎
 5.3. Proof of Corollary 1
-It only remains to reinterpret Theorem 2 in terms of the $\text{Beta}(2,1)$ and $\text{Beta}(2,2)$ distributions.
-Let $\{X^{(n)}\}$ be the solution of the RDE (9) started from an arbitrary initial datum. According to the theorem,
-(53) $\lim_{N\to\infty}\mathbb{P}\{N^{-1/\alpha}X^{(N)}\leq x\}=F(x,1),$
+It only remains to reinterpret Theorems 1 and 2 in terms of the $\text{Beta}(2,1)$ and $\text{Beta}(2,2)$ distributions.
+In what follows, consider a solution $\{X^{(n)}\}$ of the RDE (9) with $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, (constant) bias parameter $p=1/2$, and started from an arbitrary real-valued initial datum $X^{(0)}$.
+Recall that assumption (15) is in effect.
+Case (i): Diffusive Asymptotics. Let $\{F_{n}\}$ be the sequence of CDF’s associated with $\{X^{(n)}\}$. Define $F_{N}(x,t)=F_{[Nt]}(N^{1/2}x)$ as in Theorem 1. Since $X^{(0)}$ is real-valued,
+$\lim_{N\to\infty}F_{N}(x,0)=\lim_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(0)}\leq x\}=\mathbf{1}_{[0,+\infty)}(x)\quad\text{for each}\,\,x\in\mathbb{R}\setminus\{0\}.$
+First, consider the case when $\sigma\neq 0$. By Theorem 1,
+(50) $\lim_{N\to\infty}\mathbb{P}\{N^{-\frac{1}{2}}X^{(N)}\leq x\}=F(x,1),$
 where $F$ is the unique bounded discontinuous viscosity solution of (49) with initial datum $F_{\text{in}}=\mathbf{1}_{[0,+\infty)}$.
-If $\sigma<0$, then $\mathcal{F}(p,X)=\sigma|p|^{2}$ and
+If $\sigma<0$, then $F$ is determined by the formula
 $F(x,t)=\frac{x^{2}}{4|\sigma|t}\quad\text{for each}\,\,x\in[0,2\sqrt{|\sigma|}t^{1/2}],$
@@
 In any event, (50) implies
-$\lim_{N\to\infty}\mathbb{P}\{-\frac{\text{sgn}(\sigma)}{2\sqrt{|\sigma|}N^{1/2}}X^{(N)}\leq x\}=F(2\sqrt{|\sigma|}x,1)=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ x^{2},&\text{if}\,\,0\leq x\leq 1,\\ 1,&\text{otherwise.}\end{array}\right$.
-The right-hand side is precisely the CDF of the $\text{Beta}(2,1)$ distribution. Thus, case (i) is complete.
-If $\sigma=0$ and $a>0$, then $\mathcal{F}(p,X)=a|p|X$ and Proposition 11 asserts that $F(x,t)=\int_{-\infty}^{x}\rho(y,t)\,dy$, where $\rho$ is the unique continuous distributional solution of the porous medium equation $\partial_{t}\rho=\frac{1}{2}a\partial_{x}^{2}(\rho^{2})$ with initial datum $\rho(0)=\delta_{0}$.
-This solution, called the Barenblatt (or ZKB) solution, is given by
+$\lim_{N\to\infty}\mathbb{P}\{-\frac{\text{sgn}(\sigma)}{2\sqrt{|\sigma|}N^{1/2}}X^{(N)}\leq x\}=F(2x\sqrt{|\sigma|},1)=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ x^{2},&\text{if}\,\,0\leq x\leq 1,\\ 1,&\text{otherwise.}\end{array}\right$.
+The righthand side is precisely the CDF of the $\text{Beta}(2,1)$ distribution. Therefore, $2^{-1}N^{-1/2}X^{(N)}\overset{d}{\to}\sqrt{|\sigma|}\text{Beta}(2,1)$.
+It only remains to consider the case when $\sigma=0$.
+Fix a continuous function $G_{\text{in}}\in CDF(\overline{\mathbb{R}})$ such that $G_{\text{in}}(x)\leq\mathbf{1}_{[0+\infty)}(x)$ for each $x\in\mathbb{R}$.
+Consider the sequence $\{Y^{(N,n)}\}$ obtained by solving the RDE (9) with an $N$-dependent initial datum $Y^{(N,0)}$ chosen in such a way that, for every $x\in\mathbb{R}$,
+(51) $\mathbb{P}\{Y^{(N,0)}\leq x\}\leq\mathbb{P}\{X^{(0)}\leq x\}\quad\text{for each}\,\,N\in\mathbb{N},$
+$\lim_{N\to\infty}\mathbb{P}\{N^{-1/2}Y^{(N,0)}\leq x\}=G_{\text{in}}(x).$
+This is equivalent to asking that $Y^{(N,0)}=X^{(0)}+N^{1/2}Z^{(N)}$ for some sequence of nonnegative random variables $\{Z^{(N)}\}$ with law converging to $G_{\text{in}}$.
+Since $G_{\text{in}}$ is continuous, $\sigma=0$, and $p=1/2$ independently of $N$, applying Theorem 3 with $T^{(N)}=T$, $\delta^{(N)}=N^{-1/2}$, $\mathcal{F}=0$, and $Q^{(N)}=Q=0$ yields
+$\lim_{N\to\infty}\mathbb{P}\{N^{-1/2}Y^{(N,N)}\leq x\}=G(x,1),$
+where $G$ is the unique bounded viscosity solution of the equation $\partial_{t}G=0$ in $\mathbb{R}\times(0,+\infty)$ with initial condition $G(x,0)=G_{\text{in}}(x)$.
+This is just the time-independent function $G(x,t)=G_{\text{in}}(x)$.
+In particular, by (51) and monotonicity (Proposition 2),
+$G_{\text{in}}(x)=\lim_{N\to\infty}\mathbb{P}\{N^{-1/2}Y^{(N,N)}\leq x\}\leq\liminf_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(N)}\leq x\}.$
+Since $G_{\text{in}}$ was an arbitrary continuous CDF lying below $\mathbf{1}_{[0,+\infty)}$, this implies
+$\mathbf{1}_{[0,+\infty)}(x)\leq\liminf_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(N)}\leq x\}\quad\text{for each}\,\,x\in\mathbb{R}\setminus\{0\}.$
+Approximating the step function $\mathbf{1}_{[0,+\infty)}$ this time from above rather than below, one similarly finds
+$\limsup_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(N)}\leq x\}\leq\mathbf{1}_{[0,+\infty)}(x)\quad\text{for each}\,\,x\in\mathbb{R}\setminus\{0\}.$
+Therefore, $N^{-1/2}X^{(N)}\overset{d}{\to}0$ as $N\to\infty$, and case (i) is complete.
+Case (ii): Subdiffusive Asymptotics.
+Assume the stronger condition (15) and that $\sigma=0$ and $a>0$.
+Since as in the previous step $N^{-1/3}X^{(0)}\to 0$ in distribution as $N\to+\infty$, Theorem 2 implies that
+(52) $\lim_{N\to\infty}\mathbb{P}\{N^{-\frac{1}{3}}X^{(N)}\leq x\}=F(x,1)$
+provided $F$ is the unique bounded discontinuous viscosity solution of (49) with initial datum $F_{\text{in}}=\mathbf{1}_{[0,\infty)}$.
+By Proposition 15 below, $F(x,t)=\int_{-\infty}^{x}\rho(y,t)\,dy$, where $\rho$ is the unique continuous distributional solution of the porous medium equation $\partial_{t}\rho=\frac{1}{2}a\partial_{x}^{2}(\rho^{2})$ with initial datum $\rho(0)=\delta_{0}$.
+According to [24], this solution, called the Barenblatt (or ZKB) solution, is given by
 $\rho(x,t)=\frac{1}{(6at)^{\frac{1}{3}}}\left(\left(\frac{3}{4}\right)^{\frac{2}{3}}-\frac{x^{2}}{(6at)^{\frac{2}{3}}}\right)\mathbf{1}_{[-(9at/2)^{\frac{1}{3}},(9at/2)^{\frac{1}{3}}]}(x).$
@@
 $\lim_{N\to\infty}\mathbb{E}[h(Y^{(N)})]$ $=\int_{-(9a/2)^{\frac{1}{3}}}^{(9a/2)^{\frac{1}{3}}}h((36a)^{-\frac{1}{3}}x+2^{-1})\rho(x,1)\,dx=6\int_{0}^{1}h(y)y(1-y)\,dy,$
-and the integrand $6y(1-y)\mathbf{1}_{[0,1]}(y)$ is nothing but the PDF of the $\text{Beta}(2,2)$ distribution. This proves case (ii).
-Finally, if $\sigma=a=0$, then the formula in Section 1.3.2 implies that $(3s-g_{\rho}(s))(s-g_{\rho}(s))=0$ for all $s\geq 0$ and $\rho\in\{+,-\}$ almost surely.
-Since $g_{+}(s),g_{-}(s)\leq s$ hold automatically by definition, this implies that
+and the integrand $6y(1-y)\mathbf{1}_{[0,1]}(y)$ is nothing but the PDF of the $\text{Beta}(2,2)$ distribution. This completes the proof in case (ii).
+Case (iii): Triviality when $\sigma=a=0$.
+Finally, suppose that (16) holds and $\sigma=a=0$.
+Since $g_{\pm}(s)\leq s$ holds automatically by definition, the identity $a=0$ implies that $(3s-g_{\rho}(s))(s-g_{\rho}(s))=0$ for all $s\geq 0$ and $\rho\in\{+,-\}$ $\mathbf{P}$-almost surely.
+In particular,
 $\mathbf{P}\{f_{+}=f_{-}=0\}=\mathbf{P}\{g_{+}(s)=g_{-}(s)=s\,\,\text{for all}\,\,s\geq 0\}=1,$
-and thus $X^{(n)}\overset{d}{=}X^{(0)}$ for all $n$. This is case (iii).
+and thus $X^{(n)}\overset{d}{=}(1-\Theta)\max\{X^{(n-1)}_{1},X^{(n-1)}_{2}\}+\Theta\min\{X^{(n-1)}_{1},X^{(n-1)}_{2}\}\overset{d}{=}X^{(n-1)}$ for all $n$ (since $\mathbb{P}\{\Theta=1\}=p=1/2$).
+In other words, $\{X^{(n)}\}$ is the constant sequence $X^{(n)}\overset{d}{=}X^{(0)}$, completing the proof of case (iii).
+∎
 5.4. Proof of Proposition 10
-Let $\omega_{\Phi}$ denote the modulus of continuity of the second derivative $\partial_{x}^{2}\Phi$:
-$\omega_{\Phi}(\epsilon)=\sup\left\{|\partial_{x}^{2}\Phi(x)-\partial_{x}^{2}\Phi(y)|\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\epsilon\right\}.$
-In view of the representation of $\mathscr{L}$ from Proposition 4 and symmetry (see Section 2.2), it suffices to prove that
-(54) $\int_{-\infty}^{\delta^{-1}x}\mathbf{E}[\Phi_{\delta}(y_{1}+g_{+}(\delta^{-1}x-y_{1}))-\Phi_{\delta}(\delta^{-1}x)]\partial_{x}\Phi_{\delta}(y_{1})\,dy_{1}$
-$\quad=-\delta^{2}(\partial_{x}\Phi(x))^{2}\int_{0}^{\infty}\mathbf{E}[|g_{+}(u)-u|]\,du$
-$\qquad+\delta^{3}\Phi^{\prime}(x)\Phi^{\prime\prime}(x)\left(\frac{1}{2}\int_{0}^{\infty}\mathbf{E}[(g_{+}(u)-u)^{2}]\,du-\int_{0}^{\infty}u\mathbf{E}[g_{+}(u)-u]\,du\right)$
-$\qquad+C_{0}\delta^{3}\|\Phi\|_{C^{2}(\mathbb{R})}\left(\int_{0}^{\infty}\mathbf{E}[\omega_{\Phi}(\delta|g(u)-u|)\{g_{+}(u)-u\}^{2}]\,du\right$.
-$\qquad\left.+\int_{0}^{\infty}u\omega_{\Phi}(\delta u)\mathbf{E}[|g(u)-u|]\,du+\delta\|\Phi\|_{C^{2}(\mathbb{R})}\int_{a}^{\infty}\mathbf{E}[\{g_{+}(u)-u\}^{2}]u\,du\right),$
-where $\|\Phi\|_{C^{2}(\mathbb{R})}=\|\Phi\|_{\sup}+\|\partial_{x}\Phi\|_{\sup}+\|\partial^{2}_{x}\Phi\|_{\sup}$. To this end, first, notice that
-$\int_{-\infty}^{\delta^{-1}x}\mathbf{E}[\Phi_{\delta}(y_{1}+g_{+}(\delta^{-1}x-y_{1}))-\Phi_{\delta}(\delta^{-1}x)]\partial_{x}\Phi_{\delta}(y_{1})\,dy_{1}$
-$\quad=\int_{-\infty}^{x}\mathbf{E}[\Phi(x_{1}+\delta g_{+}(\delta^{-1}(x-x_{1})))-\Phi(x)]\partial_{x}\Phi(x_{1})\,dx_{1}.$
-Next, it is worth noting that, by definition, the difference $s\mapsto g_{+}(s)-s$ is bounded above and below as follows
-$-\sup\left\{|f(u)|\,\mid\,u\geq 0\right\}\leq g_{+}(s)-s\leq 0.$
-Therefore, the assumption (15) implies that the integrals above are all finite.
-Let $L$ and $Q$ be the Taylor polynomials
-$L(x_{1})=\partial_{x}\Phi(x)+\partial_{x}^{2}\Phi(x)(x_{1}-x),\quad Q(x_{1})=\partial_{x}\Phi(x)(x_{1}-x)+\frac{1}{2}\partial_{x}^{2}\Phi(x)(x_{1}-x)^{2}.$
-For notational ease, define $x_{1,g}$ according to the formula
-$x_{1,g}=x_{1}+\delta g(\delta^{-1}[x-x_{1}]).$
-The term of interest then becomes
-$\int_{-\infty}^{x}\{\Phi(x_{1,g})-\Phi(x)\}\partial_{x}\Phi(x_{1})\,dx_{1}$ $=\int_{-\infty}^{x}L(x_{1})Q(x_{1,g})\,dx_{1}$
-$\quad+\int_{-\infty}^{x}L(x_{1})\{\Phi(x_{1,g})-\Phi(x)-Q(x_{1,g})\}\,dx_{1}$
-$\quad+\int_{-\infty}^{x}\{\partial_{x}\Phi(x_{1})-L(x_{1})\}\{\Phi(x_{1,g})-\Phi(x)\}\,dx_{1}.$
-Expanding the main term and using the change-of-variables $x_{1}=x+\delta u$ leads to
-$\int_{-\infty}^{x}L(x_{1})\mathbf{E}[Q(x_{1,g})]\,dx_{1}$ $=\delta^{2}(\partial_{x}\Phi(x))^{2}\int_{0}^{\infty}\mathbf{E}[g_{+}(u)-u]\,du$
-$\quad+\frac{1}{2}\delta^{3}\partial_{x}\Phi(x)\partial_{x}^{2}\Phi(x)\int_{-\infty}^{x}\mathbf{E}[(g_{+}(u)-u)^{2}]\,dx_{1}$
-$\quad-\delta^{3}\partial_{x}\Phi(x)\partial_{x}^{2}\Phi(x)\int_{-\infty}^{x}u\mathbf{E}[g_{+}(u)-u]\,dx_{1}$
-$\quad-\frac{1}{2}\delta^{4}(\partial_{x}^{2}\Phi(x))^{2}\int_{-\infty}^{x}u\mathbf{E}[(g(u)-u)^{2}]\,dx_{1}.$
-Notice that the order $\delta^{2}$ term has the desired form: In general, $g_{+}(s)\leq s$ always holds so $g_{+}(s)-s=-|g_{+}(s)-s|$.
-It only remains to estimate the error terms:
-$\left|\int_{-\infty}^{x}L(x_{1})\{\Phi(x_{1,g})-\Phi(x)-Q(x_{1,g})\}\,dx_{1}\right|$
-$\quad\leq\|\Phi\|_{C^{2}(\mathbb{R})}\int_{-\infty}^{x}\omega_{\Phi}(|x_{1,g}-x|)\left(1+|x_{1}-x|\right)\{x_{1,g}-x\}^{2}\,dx_{1}$
-$\quad=\delta^{3}\|\Phi\|_{C^{2}(\mathbb{R})}\int_{0}^{\infty}\omega_{\Phi}(\delta|g(u)-u|)(1+\delta|u|)\{g(u)-u\}^{2}\,du,$
-$\left|\int_{-\infty}^{x}\{\partial_{x}\Phi(x_{1})-L(x_{1})\}\{\Phi(x_{1,g})-\Phi(x)\}\,dx_{1}\right|$
-$\quad\leq\|\Phi\|_{C^{2}(\mathbb{R})}\int_{-\infty}^{x}\omega_{\Phi}(|x_{1}-x|)|x_{1}-x||x_{1,g}-x|\,dx_{1}$
-$\quad=\delta^{3}\|\Phi\|_{C^{2}(\mathbb{R})}\int_{0}^{\infty}\omega_{\Phi}(\delta u)u|g(u)-u|\,du.$
+It will be convenient to start by fixing some notation. Throughout this proof, define $h_{+},h_{-}:[0,+\infty)\to[0,+\infty)$ by
+$h_{\pm}(s)=s-g_{\pm}(s).$
+Notice that $0\leq h_{\pm}(s)\leq s$ for any $s\geq 0$ by the definition of $g_{\pm}$ (Section 1.4.1).
+Define coefficients $\sigma_{+}$ and $\sigma_{-}$ by
+(53) $\sigma_{+}=-\int_{0}^{+\infty}\mathbf{E}h_{+}(s)\,ds,\quad\sigma_{-}=\int_{0}^{+\infty}\mathbf{E}h_{-}(s)\,ds,$
+which are finite by (15), so that $\sigma=\sigma_{+}+\sigma_{-}$. In the case of the stronger assumption (16), also define coefficients $a_{+}$ and $a_{-}$ by
+$a_{\pm}=\int_{0}^{+\infty}s\mathbf{E}h_{\pm}(s)\,ds+\frac{1}{2}\int_{0}^{+\infty}\mathbf{E}h_{\pm}(s)^{2}\,ds=\frac{1}{2}\int_{0}^{+\infty}\mathbf{E}(2s+h_{\pm}(s))h_{\pm}(s)\,ds.$
+By definition, $a=a_{+}+a_{-}$.
+Assumption (15) says that $\int_{0}^{+\infty}\mathbf{E}h_{\pm}(s)\,ds<+\infty$, while assumption (16) says that $\int_{0}^{+\infty}(1+s)\mathbf{E}h_{\pm}(s)\,ds<+\infty$.
+Notice that, in the latter case, the bound $0\leq h_{\pm}(s)\leq s$ implies, in addition, that
+(54) $\int_{0}^{+\infty}\mathbf{E}h_{\pm}(s)^{2}\,ds\leq\int_{0}^{+\infty}s\mathbf{E}h_{\pm}(s)\,ds<+\infty.$
+In particular, $a_{+}$ and $a_{-}$ are finite.
+With these preliminaries out of the way, fix a $\Phi\in CDF(\overline{\mathbb{R}})$ that has bounded, uniformly continuous first and second derivatives, and define $\Phi_{\delta}(y)=\Phi(\delta y)$ for $\delta>0$.
+Let $\omega_{\Phi}$ be the modulus of continuity
+$\omega_{\Phi}(\delta)=\sup\left\{|\partial_{x}\Phi(x)-\partial_{x}\Phi(y)|+|\partial_{x}^{2}\Phi(x)-\partial_{x}^{2}\Phi(y)|\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\delta\right\}.$
+Recall from Section 3 that it is possible to write
+(55) $\mathscr{L}^{(N)}=2p^{(N)}\mathscr{L}_{+}+2(1-p^{(N)})\mathscr{L}_{-}$
+for some $N$-independent functions $\mathscr{L}_{+}$ and $\mathscr{L}_{-}$ defined on $CDF(\overline{\mathbb{R}})$.
+Thus, the scaling behavior of $\mathscr{L}^{(N)}$ is determined by that of $\mathscr{L}_{+}$ and $\mathscr{L}_{-}$.
+Further, since the assumption (15) implies that $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=1\}$ (as proved in Section 5.2 above), the formula (33) can be invoked to find
+$(\mathscr{L}_{+}\Phi_{\delta})(\delta^{-1}x)$ $=\delta\int_{0}^{+\infty}\mathbf{E}[\Phi(x-\delta h_{+}(s))-\Phi(x)]\partial_{x}\Phi(x-\delta s)\,ds,$
+$(\mathscr{L}_{-}\Phi_{\delta})(\delta^{-1}x)$ $=\delta\int_{0}^{+\infty}\mathbf{E}[\Phi(x+\delta h_{-}(s))-\Phi(x)]\partial_{x}\Phi(x+\delta s)\,ds.$
+In view of (55), it only remains to prove the following two claims: First, under assumption (15), for any $\rho\in\{+,-\}$,
+(56) $\lim_{\delta\downarrow 0}\sup\left\{\left|\delta^{-2}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}\right|\,\mid\,x\in\mathbb{R}\right\}=0.$
+Further, if assumption (16) holds, then
+(57) $\lim_{\delta\downarrow 0}\sup\left\{\left|\delta^{-3}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\delta^{-1}\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}-a_{\rho}|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)\right|\,\mid\,x\in\mathbb{R}\right\}=0.$
+$\delta^{2}$ Asymptotics. Assume (15) holds and fix a $\rho\in\{+,-\}$. First, observe that, by the definition (53) of $\sigma_{\rho}$,
+(58) $\delta^{-2}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}$
+$=\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}\partial_{x}\Phi(x-\rho\delta s)+\rho|\partial_{x}\Phi(x)|^{2}\right)h_{\rho}(s)\,ds.$
+By the assumptions on $\Phi$ and the bound $h_{\rho}(s)\leq s$,
+$\left|\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}\partial_{x}\Phi(x-\rho\delta s)+\rho|\partial_{x}\Phi(x)|^{2}\right)h_{\rho}(s)\,ds\right|$
+$\qquad\leq\|\partial_{x}\Phi\|_{\text{sup}}\mathbf{E}\int_{0}^{+\infty}\left(\omega_{\Phi}(\delta h_{+}(s))+\omega_{\Phi}(\delta s)\right)h_{\rho}(s)\,ds$
+$\qquad\leq 2\|\partial_{x}\Phi\|_{\text{sup}}\int_{0}^{+\infty}\omega_{\Phi}(\delta s)\mathbf{E}h_{\rho}(s)\,ds$
+Since assumption (15) implies $\mathbf{E}h_{\rho}$ is integrable, Lebesgue’s dominated convergence theorem implies that the righthand side vanishes as $\delta\downarrow 0$. Therefore, (56) follows.
+$\delta^{3}$ Asymptotics. Finally, assume that assumption (16) holds. In this case, notice that the main error term from the previous step can be rewritten in the form
+$\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}\partial_{x}\Phi(x-\rho\delta s)+\rho|\partial_{x}\Phi(x)|^{2}\right)h_{\rho}(s)\,ds$
+$\quad=\rho\partial_{x}\Phi(x)\mathbf{E}\int_{0}^{+\infty}(\partial_{x}\Phi(x)-\partial_{x}\Phi(x-\rho\delta s))h_{\rho}(s)\,ds$
+$\quad\quad+\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}+\rho\partial_{x}\Phi(x)\right)\partial_{x}\Phi(x-\rho\delta s)h_{\rho}(s)\,ds.$
+In the first term, one finds
+$\left|\frac{\rho}{\delta}\mathbf{E}\int_{0}^{+\infty}\partial_{x}\Phi(x)(\partial_{x}\Phi(x)-\partial_{x}\Phi(x-\rho\delta s))h_{\rho}(s)\,ds-\partial_{x}\Phi(x)\partial_{x}^{2}\Phi(x)\int_{0}^{+\infty}s\mathbf{E}h_{\rho}(s)\,ds\right|$
+$\qquad\leq\|\partial_{x}\Phi\|_{\text{sup}}\int_{0}^{+\infty}\omega_{\Phi}(\delta s)s\mathbf{E}h_{\rho}(s)\,ds.$
+Similarly, in the second term, again using $\omega_{\Phi}(\delta h_{+}(s))\leq\omega_{\Phi}(\delta s)$,
+$\left|\frac{1}{\delta}\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}+\rho\partial_{x}\Phi(x)\right)\partial_{x}\Phi(x-\rho\delta s)h_{\rho}(s)\,ds\right$.
+$\qquad\left.-\frac{1}{2}\partial_{x}\Phi(x)\partial_{x}^{2}\Phi(x)\int_{0}^{+\infty}\mathbf{E}h_{\rho}(s)^{2}\,ds\right|$
+$\qquad\leq\frac{1}{2}\left(\|\partial_{x}\Phi\|_{\text{sup}}+\|\partial_{x}^{2}\Phi\|_{\text{sup}}\right)\int_{0}^{+\infty}\omega_{\Phi}(\delta s)\mathbf{E}h_{\rho}(s)^{2}\,ds.$
+Combining these estimates with the identity (58) and estimate (54), one finds
+$\sup\left\{\left|\delta^{-3}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\delta^{-1}\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}-a_{\rho}|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)\right|\,\mid\,x\in\mathbb{R}\right\}$
+$\qquad\leq\left(1+\|\partial_{x}\Phi\|_{\text{sup}}+\|\partial_{x}^{2}\Phi\|_{\text{sup}}\right)\int_{0}^{+\infty}\omega_{\Phi}(\delta s)s\mathbf{E}h_{\rho}(s)\,ds.$
+Since the righthand side vanishes as $\delta\downarrow 0$ by dominated convergence, (57) follows. ∎
+5.5. Counterexample
+Here is a counterexample, showing that $N^{-1/2}X^{(N)}$ can blow up if (15) fails to hold. Consider the law $\mathbf{P}$ under which $(f_{+},f_{-})$ are given by
+$f_{+}(u)=(Z-u)_{+},\quad f_{-}(u)=0,$
+where $Z$ is a nonnegative random variable such that $\mathbb{E}Z^{2}=+\infty$.
+With this choice of forcing, the RDE (9) with $p=1/2$ is superdiffusive in the sense that if $\{X^{(n)}\}$ is a solution started from any initial datum and $p=1/2$, then
+$N^{-\frac{1}{2}}X^{(N)}\overset{d}{\to}+\infty\quad\text{as}\,\,N\to+\infty.$
+To see this, it is convenient to define $f^{(m)}_{+}$ for $m\in\mathbb{N}$ by
+$f^{(m)}_{+}(u)=(\max\{Z,m\}-u)_{+}.$
+Let $\{X^{(m,n)}\}$ be a solution of the RDE (9) started from the same initial distribution as $\{X^{(n)}\}$, but with the forcing $(f_{+},0)$ replaced by $(f_{+}^{(m)},0)$.
+Note that the inverse $g_{+}^{(m)}$ associated with $f_{+}^{(m)}$ is given by
+$g_{+}^{(m)}(s)=0\quad\text{if}\,\,s<\max\{Z,m\},\quad g_{+}^{(m)}(s)=s,\quad\text{otherwise,}$
+and the inverse $g_{-}$ associated with $f_{-}=0$ is $g_{-}(s)=s$. Therefore, the RDE with $(f^{(m)}_{+},0)$ satisfies (15) and the constant $\sigma^{(m)}$ determining the asymptotics is given by
+$\sigma^{(m)}=-\mathbf{E}\int_{0}^{\max\{Z,m\}}s\,ds=-\frac{1}{2}\mathbf{E}\max\{Z^{2},m^{2}\}.$
+Note, in particular, that $|\sigma^{(m)}|\to+\infty$ as $m\to+\infty$ by the choice of $Z$.
+Since $f_{+}\geq f^{(m)}_{+}$, monotonicity with respect to the data $f_{+}$ (Proposition 3) implies that, for any $x\in\mathbb{R}$,
+$\mathbb{P}\{N^{-1/2}X^{(N)}\leq x\}\leq\mathbb{P}\{N^{-1/2}X^{(m,N)}\leq x\}.$
+Therefore, by Corollary 1,
+$\limsup_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(m,N)}\leq x\}\leq F(2^{-1}|\sigma^{(m)}|^{-\frac{1}{2}}x).$
+where $F$ is the CDF of the $\text{Beta}(2,1)$ random variable Sending $m\to+\infty$, this implies
+$\limsup_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(N)}\leq x\}\leq F(0)=0\quad\text{for each}\,\,x\in\mathbb{R}.$
 Appendix A Viscosity Solutions
-This section reviews relevant definitions and results from the theory of viscosity solutions. General references for this material are [13, 7]. The main interest here is the initial-value problem
-(57) $\left\{\begin{array}[]{r l}\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
+This section reviews relevant definitions and results from the theory of viscosity solutions. General references for this material are [15, 10]. The interest here is in the initial-value problem
+(61) $\left\{\begin{array}[]{r l}\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=F_{\text{in}}(x),\end{array}\right$.
+where $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ and the diffusion term $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ and the reaction term $Q:\mathbb{R}\to\mathbb{R}$ are continuous functions, which satisfy the assumptions, for some $L\geq 0$,
+(62) $\mathcal{F}(v,w_{1})\leq\mathcal{F}(v,w_{2})\quad\text{for each}\,\,v,w_{1},w_{2}\in\mathbb{R}\,\,\text{with}\,\,w_{1}<w_{2},$
+(63) $|Q(q)-Q(q^{\prime})|\leq L|q-q^{\prime}|\,\,\text{for each}\,\,q,q^{\prime}\in\mathbb{R},$
+(64) $\mathcal{F}(0,0)=0,\quad Q(0)\geq 0,\quad\text{and}\quad Q(1)\leq 0.$
 Since the problem at hand naturally leads to consideration of possibly discontinuous initial data, some care will be needed in defining precisely what is meant by a solution, subsolution, or supersolution.
@@
 Definition 1.
-An upper semicontinuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)\leq 0$ in the viscosity sense if it has the following property: If $\varphi$ is a smooth function defined in some open set $U$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from above (that is, $F(x,t)\leq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
-$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))\leq 0.$
-A lower semi-continuous function $G:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial_{x}^{2}G)\geq 0$ in the viscosity sense if it has the following property: If $\varphi$ is a smooth function defined in some open set $U$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from below (that is, $F(x,t)\geq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
-$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))\geq 0.$
-A continuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential equation $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)=0$ in the viscosity sense if it satisfies both $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)\leq 0$ and $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)\geq 0$ in the viscosity sense.
+An upper semicontinuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ in the viscosity sense if it has the following property: If $\varphi$ is a smooth function defined in some open set $U$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from above (that is, $F(x,t)\leq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
+$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))+Q(\varphi(x_{0},t_{0}))\leq 0.$
+A lower semi-continuous function $G:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial_{x}^{2}G)+Q(G)\geq 0$ in the viscosity sense if it has the following property: If $\varphi$ is a smooth function defined in some open set $U$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from below (that is, $F(x,t)\geq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
+$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))+Q(\varphi(x_{0},t_{0}))\geq 0.$
+A continuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential equation $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)=0$ in the viscosity sense if it satisfies both $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ and $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)\geq 0$ in the viscosity sense.
 The next definition specifies what is meant by a solution, subsolution, or supersolution of the initial-value problem (61).
 Definition 2.
-An upper semicontinuous function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ said to be a viscosity subsolution of (57) if it satisfies $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
+An upper semicontinuous function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ said to be a viscosity subsolution of (61) if it satisfies $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
 (65) $\lim_{\delta\downarrow 0}\sup\left\{F(y,0)-F_{\text{in}}(x)\,\mid\,x,y\in\mathbb{R}\,\,\text{such that}\,\,|y-x|\leq\delta\right\}\leq 0.$
-A lower semicontinuous function $G:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a viscosity supersolution of (57) if it satisfies $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
+A lower semicontinuous function $G:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a viscosity supersolution of (61) if it satisfies $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)+Q(G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
 $\lim_{\delta\downarrow 0}\inf\left\{G(y,0)-F_{\text{in}}(x)\,\mid\,x,y\in\mathbb{R}\,\,\text{such that}\,\,|y-x|\leq\delta\right\}\geq 0.$
 A continuous function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a (continuous) viscosity solution of (61) if it is both a viscosity subsolution and a viscosity supersolution.
-Since the main theorem of this work, Theorem 1, involves problems with a discontinuous initial datum, a different definition of solution is needed.
+Since the main results of this work, Theorems 1 and 2, involve problems with a discontinuous initial datum, a different definition of solution is needed.
 In order to state it, it is necessary to recall the notions of upper and lower semicontinuous envelopes.
@@
 Definition 3.
-A function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a discontinuous viscosity solution of (57) if $\partial_{t}F^{*}-\mathcal{F}(\partial_{x}F^{*},\partial_{x}^{2}F^{*})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$; $\partial_{t}F_{*}-\mathcal{F}(\partial_{x}F_{*},\partial_{x}^{2}F_{*})\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$; and $F^{*}(\cdot,0)=(F_{\text{in}})^{*}$ and $F_{*}(\cdot,0)=(F_{\text{in}})_{*}$ in $\mathbb{R}$.
+A function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a discontinuous viscosity solution of (61) if $\partial_{t}F^{*}-\mathcal{F}(\partial_{x}F^{*},\partial_{x}^{2}F^{*})+Q(F^{*})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$; $\partial_{t}F_{*}-\mathcal{F}(\partial_{x}F_{*},\partial_{x}^{2}F_{*})+Q(F_{*})\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$; and $F^{*}(\cdot,0)\leq(F_{\text{in}})^{*}$ and $F_{*}(\cdot,0)\geq(F_{\text{in}})_{*}$ in $\mathbb{R}$.
 A.1. Comparison Principle
@@
 Theorem 4 (Comparison Principle).
-Assume that the nonlinearity $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ is jointly continuous and nondecreasing in the second argument.
-If $F,G:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ are respectively a bounded upper and a bounded lower semicontinuous function such that $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)\leq 0$ and $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
-$\lim_{\delta\downarrow 0}\sup\left\{F(x,0)-G(y,0)\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\delta\right\}\leq 0,$
-then $F\leq G$ pointwise in $\mathbb{R}\times[0,\infty)$.
-In particular, if $F$ and $G$ are respectively bounded viscosity sub- and supersolutions of (57) for some initial datum $F_{\text{in}}$, then $F\leq G$ pointwise in $\mathbb{R}\times[0,\infty)$.
-The reader is referred to [16, Theorem 2.1], for instance, for the proof.
-The comparison principle implies that there is at most one continuous viscosity solutions of (57).
-Existence of the solution is also known in the present setting.
-However, there is no need to state an existence result here as it actually follows from the main convergence results (see the proof of Theorem 3).
-Let $\overline{F}$ and $\underline{F}$ be the maximal sub- and minimal supersolution of (57):
+Assume that $\mathcal{F}$ and $Q$ satisfy assumptions (62) and (63).
+If $F,G:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ are respectively a bounded upper and a bounded lower semicontinuous function such that $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ and $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)+Q(G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and if $M$ is the constant defined by
+$M=\lim_{\delta\downarrow 0}\sup\left\{(F(x,0)-G(y,0))_{+}\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\delta\right\},$
+then, for any $t>0$,
+(66) $\sup\left\{F(x,t)-G(x,t)\,\mid\,x\in\mathbb{R}\right\}\leq Me^{Lt}$
+In particular, if $F$ and $G$ are respectively bounded viscosity sub- and supersolutions of (61) for some fixed initial datum $F_{\text{in}}$, then $F\leq G$ pointwise in $\mathbb{R}\times[0,\infty)$.
+Proof.
+The reader can find a proof of this version of the theorem with $L=0$, $M=0$ in [19].
+The extension to the case $L>0$, $M=0$ is routine (see also [10, Section 5.2]).
+The case $L>0$, $M>0$ can be reduced to the case $L>0$, $M=0$ using an analogue of the argument appearing in the proof of Lemma 5 below, namely, using that the function $\tilde{F}=F-Me^{Lt}$ satisfies $\partial_{t}\tilde{F}-\mathcal{F}(\partial_{x}\tilde{F},\partial_{x}^{2}\tilde{F})+Q(\tilde{F})\leq 0$ in $\mathbb{R}\times(0,+\infty)$.
+∎
+The comparison principle implies, in particular, that there is at most one viscosity solution of (61) if $F_{\text{in}}$ is continuous.
+It is known that, indeed, a solution exists in this case, and the assumption (64) ensures that, for each $t>0$, $F(\cdot,t)\in CDF(\overline{\mathbb{R}})$.
+Proposition 12.
+Assume that $\mathcal{F}$ and $Q$ satisfy (62), (63), and (64).
+For any continuous $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$, there is a unique viscosity solution $F$ of (61).
+Further, for any $t>0$, the function $F(\cdot,t)$ is in $CDF(\overline{\mathbb{R}})$.
+Proof.
+Existence of a solution taking values in $[0,1]$ follows from Perron’s Method, which is applicable since the constant functions 0 and $1$ are, respectively, viscosity sub- and supersolutions by (64).
+By the comparison principle, the solution is unique and it takes values in $[0,1]$.
+Finally, for any $y\in\mathbb{R}$, the function $(x,t)\mapsto F(x+y,t)$ is a viscosity solution of (61) with the initial datum $x\mapsto F_{\text{in}}(x+y)$.
+Thus, since $F_{\text{in}}$ is nondecreasing, by the comparison principle, $F(x+y,t)\leq F(x,t)$ for all $t\geq 0$ if $y\geq 0$.
+Since $F$ is continuous, this proves $F(\cdot,t)\in CDF(\overline{\mathbb{R}})$ for each $t>0$.
+∎
+In general, let $\underline{F}$ and $\overline{F}$ be the maximal sub- and minimal supersolution of (61):
 […]
 […]
-Notice that the comparison principle implies that $\overline{F}\leq\underline{F}$ and any viscosity solution (continuous or discontinuous) lies in between the two.
+Notice that the comparison principle implies that $\underline{F}\leq\overline{F}$ and any viscosity solution (continuous or discontinuous) lies in between the two.
 This immediately implies a uniqueness criterion.
 Proposition 13.
-Let $\overline{F}$ and $\underline{F}$ be the maximal subsolution and minimal supersolution of (57), respectively.
-If $\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,\infty)$, then any discontinuous viscosity solution $F$ of (57) must also satisfy $F=\overline{F}$ in $\mathbb{R}\times(0,\infty)$.
-In particular, in this case, modulo the freedom to choose the value at discontinuities of $F_{\text{in}}$ at $t=0$, there is a unique discontinuous viscosity solution of (57).
+Assume that $\mathcal{F}$ and $Q$ satisfy assumptions (62) and (63).
+Let $\underline{F}$ and $\overline{F}$ be the maximal subsolution and minimal supersolution of (61), respectively, for some given $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$.
+If $\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,\infty)$, then any discontinuous viscosity solution $F$ of (61) must also satisfy $F=\underline{F}$ in $\mathbb{R}\times(0,\infty)$.
+In particular, in this case, modulo the freedom to choose the value at discontinuities of $F_{\text{in}}$ at $t=0$, there is a unique discontinuous viscosity solution of (61), which is continuous in $\mathbb{R}\times(0,+\infty)$.
 Proof.
-As in the discussion preceding the statement, if $F$ is any discontinuous viscosity solution, then $\overline{F}\leq F\leq\underline{F}$ in $\mathbb{R}\times[0,\infty)$ in general.
-Therefore, in the present setting, $F=\overline{F}=\underline{F}$ for $t>0$, implying that there is at most one discontinuous viscosity solution.
-The fact that at least one exists is similarly immediate: The equality $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,\infty)$ actually implies that both $\overline{F}$ and $\underline{F}$ are discontinuous viscosity solutions.
+As in the discussion preceding the statement, if $F$ is any discontinuous viscosity solution, then $\underline{F}\leq F\leq\overline{F}$ in $\mathbb{R}\times[0,\infty)$ in general.
+Therefore, in the present setting, $F=\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,+\infty)$, implying that there is at most one discontinuous viscosity solution (up to arbitrary choices at $t=0$ at points where $F_{\text{in}}$ is discontinuous).
+The fact that at least one exists is almost immediate: Since $\underline{F}$ is a supremum of subsolutions, the upper semicontinuous envelope $\underline{F}^{*}$ satisfies $\partial_{t}\underline{F}^{*}-\mathcal{F}(\partial_{x}\underline{F}^{*},\partial_{x}^{2}\underline{F}^{*})+Q(\underline{F}^{*})\leq 0$ in $\mathbb{R}\times(0,+\infty)$, and solutions of (61) with initial data larger than $F_{\text{in}}$ can be used to prove $\underline{F}^{*}(\cdot,0)=F_{\text{in}}$ at $t=0$.
+Therefore, by comparison, $\underline{F}^{*}\leq\overline{F}$ in $\mathbb{R}\times[0,+\infty)$.
+A similar argument shows that $\partial_{t}\overline{F}_{*}-\mathcal{F}(\partial_{x}\overline{F}_{*},\partial^{2}_{x}\overline{F}_{*})+Q(\overline{F}_{*})\geq 0$ in $\mathbb{R}\times(0,+\infty)$ and $\overline{F}_{*}(\cdot,0)$ equals the lower semicontinuous envelope of $F_{\text{in}}$, so $\overline{F}_{*}\geq\underline{F}$.
+Since $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,+\infty)$ and the upper (resp.
+lower) semicontinuous envelope of a function is always larger (resp.
+smaller) than the function itself, this proves $\underline{F}$ is continuous in $\mathbb{R}\times(0,+\infty)$ and a discontinuous viscosity solution of (61).
 ∎
+The following lemma will be used below to verify the identity $\underline{F}=\overline{F}$.
+Lemma 1.
+Fix $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$.
+Suppose that $\{F^{(N),+}_{\text{in}}\}$ and $\{F^{(N),-}_{\text{in}}\}$ are two sequences of continuous functions in $CDF(\overline{\mathbb{R}})$ such that
+(67) $F^{(N-1),-}_{\text{in}}\leq F^{(N),-}_{\text{in}}\leq F_{\text{in}}\leq F^{(N),+}_{\text{in}}\leq F^{(N-1),+}_{\text{in}}\quad\text{pointwise for each}\,\,N,$
+(68) $\lim_{N\to+\infty}F^{(N),+}_{\text{in}}=\lim_{N\to+\infty}F^{(N),-}_{\text{in}}=F_{\text{in}}\quad\text{vaguely.}$
+Let $\{F^{(N),+}\}$ and $\{F^{(N),-}\}$ be the (continuous) viscosity solutions of (61) with respective initial data $\{F^{(N),+}_{\text{in}}\}$ and $\{F^{(N),-}_{\text{in}}\}$.
+If the limit of the two sequences coincides, that is, if
+$F:=\lim_{N\to+\infty}F^{(N),+}=\lim_{N\to+\infty}F^{(N),-}\quad\text{pointwise in}\,\,\mathbb{R}\times(0,+\infty),$
+then the maximal subsolution and minimal supersolution of (61) with initial datum $F_{\text{in}}$ are equal in $\mathbb{R}\times(0,+\infty)$ and, in particular, they both equal the limit $F$.
+Proof.
+This is immediate from the fact that $F^{(N),-}\leq\underline{F}\leq\overline{F}\leq F^{(N),+}$ for any $N$ by comparison. ∎
 A.2. Uniqueness for Nondecreasing Data
@@
 For the Hamilton-Jacobi equations arising in the $\text{Beta}(2,1)$ case, the fact that almost-everywhere-continuous data leads to a unique solution at positive times is known from [11].
-In the case of the second-order equation in the $\text{Beta}(2,2)$ case, the uniqueness of solutions follows from the characterization of nonnegative solutions of the porous medium equation.
+In the case of the second-order equation in the $\text{Beta}(2,2)$ case, the uniqueness of solutions follows, for instance, from the characterization of nonnegative solutions of the porous medium equation.
 Short, simplified proofs are presented here for the reader’s convenience.
+Case: $Q=0$. The discussion begins with the classical case when the reaction term $Q=0$.
 First, consider the case when $\mathcal{F}$ is a first-order operator of the form
-$\mathcal{F}(p,X)=-H(p)$
+$\mathcal{F}(v,w)=-H(v)$
 for some convex or concave function $H:\mathbb{R}\to\mathbb{R}$ exhibiting superlinear growth:
@@
 Proposition 14.
-Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$.
-If $\mathcal{F}(p,X)=-H(p)$ for a convex $H:\mathbb{R}\to\mathbb{R}$ satisfying (i) and (ii) above, and if $\overline{F}$ and $\underline{F}$ are the maximal subsolution and minimal supersolution, respectively, of (57), then, for any $(x,t)\in\mathbb{R}\times(0,\infty)$,
-(60) $\overline{F}(x,t)=\underline{F}(x,t)=\inf\left\{F_{\text{in}}(y)+tL\left(\frac{x-y}{t}\right)\,\mid\,y\in\mathbb{R}\right\}.$
+Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ and assume $Q=0$.
+If $\mathcal{F}(v,w)=-H(v)$ for a convex $H:\mathbb{R}\to\mathbb{R}$ satisfying $H(0)=0$ and (69), and if $\underline{F}$ and $\overline{F}$ are the maximal subsolution and minimal supersolution, respectively, of (61), then, for any $(x,t)\in\mathbb{R}\times(0,\infty)$,
+(70) $\underline{F}(x,t)=\overline{F}(x,t)=\inf\left\{F_{\text{in}}(y)+tL\left(\frac{x-y}{t}\right)\,\mid\,y\in\mathbb{R}\right\}.$
 where $L$ is the Legendre-Fenchel transform of $H$.
+Furthermore, the following statements hold:
+  • (i)
+    For any $t>0$, the functions $\underline{F}(\cdot,t)$ and $F_{\text{in}}$ coincide on the set $\{-\infty,+\infty\}$.
+  • (ii)
+    If $\{F^{(n)}_{\text{in}}\}$ is a sequence in $CDF(\overline{\mathbb{R}})$ converging vaguely to $F_{\text{in}}$, and if $\{\underline{F}^{(n)}\}$ are the corresponding maximal subsolutions of (61), then, for any compact set $K\subseteq\mathbb{R}\times(0,+\infty)$,
+    $\lim_{n\to\infty}\sup\left\{|\underline{F}^{(n)}(x,t)-\underline{F}^{(n)}(x,t)|\,\mid\,(x,t)\in K\right\}=0.$
 Some regularity of $F_{\text{in}}$ is necessary to ensure that $\overline{F}$ and $\underline{F}$ coincide at positive times.
@@
 Proof.
-When $F_{\text{in}}$ is uniformly Lipschitz continuous, the fact that the right-hand side of (60) determines the (continuous) viscosity solution is classical (see, e.g., [15, Section 10.3.4]).
-Otherwise, let $\{F^{(N)}_{\text{in}}\}\subseteq CDF(\overline{\mathbb{R}})$ be a sequence of uniformly Lipschitz continuous functions such that
-$F^{(N)}_{\text{in}}\leq F^{(N+1)}_{\text{in}}<F_{\text{in}}\quad\text{pointwise for each}\,\,N\in\mathbb{N},$
-$\lim_{N\to\infty}F^{(N)}_{\text{in}}(x)=F_{\text{in}}(x)\quad\text{for each point of continuity}\,\,x\,\,\text{of}\,\,F_{\text{in}}.$
-If $\{F_{N}\}$ are the corresponding solutions of (57), then the comparison principle implies that $F_{N}\leq F_{N+1}\leq\overline{F}$ for each $N$.
-By the stability of viscosity solutions, the limit $F=\lim_{N\to\infty}F_{N}$ satisfies $\partial_{t}F-H(\partial_{x}F)=0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$.
-At the same time, recall that the superlinearity of $H$ implies that $L$ is also superlinear.
-Therefore, for any $(x,t)$, it is possible to write $F_{N}(x,t)=F^{(N)}_{\text{in}}(y_{N})+tL((x-y_{N})/t)$ for some bounded sequence $\{y_{N}\}$.
+Let $F$ be the function defined in the righthand side of (70).
+When $F_{\text{in}}$ is continuous, the fact that $F$ is the (continuous) viscosity solution is classical (see, e.g., [18, Section 10.3.4]).
+Otherwise, let $\{F^{(N),+}_{\text{in}}\}$ and $\{F_{\text{in}}^{(N),-}\}$ be two sequences converging to $F_{\text{in}}$ as in Lemma 1, and let $\{F^{(N),+}\}$ and $\{F^{(N),-}\}$ be the corresponding solutions given by (70).
+For any point of continuity $y$ of $F_{\text{in}}$, any $x\in\mathbb{R}$, and any $t>0$,
+$F_{\text{in}}(y)+tL\left(\frac{x-y}{t}\right)=\lim_{N\to\infty}F_{\text{in}}^{(N),-}(y)+tL\left(\frac{x-y}{t}\right)\geq\lim_{N\to\infty}F^{(N),+}(x,t).$
+Therefore, since $F_{\text{in}}$ is right-continuous, this proves $\lim_{N\to\infty}F^{(N),+}\leq F$ pointwise in $\mathbb{R}\times(0,+\infty)$.
+Next, by superlinearity of $L$, for any $(x,t)$, it is possible to write $F^{(N),-}(x,t)=F^{(N),-}_{\text{in}}(y_{N})+tL((x-y_{N})/t)$ for some bounded sequence $\{y_{N}\}$.
 If $y_{N}\to y$ as $N\to\infty$ (say, along a subsequence), then
-$F_{\text{in}}(y-\delta)=\lim_{N\to\infty}F_{N}(y-\delta)\leq\liminf_{N\to\infty}F^{(N)}_{\text{in}}(y_{N})\quad\text{for almost every}\,\,\delta>0,$
-from which it follows that
-$\inf\left\{F_{\text{in}}(y)+tL\left(\frac{x-y}{t}\right)\,\mid\,y\in\mathbb{R}\right\}\leq\lim_{N\to\infty}F_{N}(x,t)=F(x,t)\leq\overline{F}(x,t).$
-$\underline{F}$ can be treated similarly: Let $\{F^{(N)}_{\text{in}}\}\subseteq CDF(\overline{\mathbb{R}}$ be uniformly Lipschitz functions such that $F_{\text{in}}<F^{(N+1)}_{\text{in}}\leq F^{(N)}_{\text{in}}$ for each $N$ and $F_{\text{in}}(x)=\lim_{N\to\infty}F^{(N)}_{\text{in}}(x)$ at points of continuity $x$ of $F_{\text{in}}$.
-This time, if $F_{N}$ is the corresponding solution of (57), then, for any point of continuity $y$ of $F_{\text{in}}$ and any $(x,t)\in\mathbb{R}\times(0,\infty)$,
-$\underline{F}(x,t)\leq\lim_{N\to\infty}F_{N}(x,t)\leq\lim_{N\to\infty}F_{\text{in}}^{(N)}(y)+tL(\frac{x-y}{t})=F_{\text{in}}(y)+tL(\frac{x-y}{t}).$
-Since $F$ is right-continuous and points of continuity are dense, this proves $\underline{F}$ is bounded above by the right-hand side of (60). ∎
-Next, consider the case of the nonlinearity
-(61) $\mathcal{F}(p,X)=a|p|^{m-1}X$
+$F_{\text{in}}(y-\delta)=\lim_{N\to\infty}F^{(N),-}(y-\delta)\leq\liminf_{N\to\infty}F^{(N),-}_{\text{in}}(y_{N})\quad\text{for almost every}\,\,\delta>0,$
+from which it follows that $F(x,t)\leq\lim_{N\to+\infty}F^{(N),-}(x,t)$. Therefore, by Lemma 1, $F=\underline{F}=\overline{F}$.
+Item (i) follows directly from the representation (70), the superlinearity of $L$, and the fact that $0\leq F_{\text{in}}\leq 1$.
+Item (ii) can be proved using the comparison principle, Theorem 4, in conjunction with the half-relaxed limit method; the interested reader can compare with the proof of Corollary 4 below.
+∎
+Next, let $m>1$ and consider the case of the nonlinearity
+(71) $\mathcal{F}(v,w)=a|v|^{m-1}w$
 Notice that if $F$ were a smooth solution of $\partial_{t}F-a|\partial_{x}F|^{m-1}\partial_{x}^{2}F=0$, then the derivative $\rho=\partial_{x}F$ would solve the porous medium equation (PME) $\partial_{t}\rho=\frac{a}{m}\partial_{x}^{2}(\rho^{m})$.
@@
 Proposition 15.
-Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$.
-If $\mathcal{F}$ has the form (61) for some $m>1$, then the maximal subsolution $\overline{F}$ and minimal supersolution $\underline{F}$ of (57) are given for $(x,t)\in\mathbb{R}\times(0,\infty)$ by the formula
-$\overline{F}(x,t)=\underline{F}(x,t)=\int_{-\infty}^{x}\rho(y,t)\,dy,$
+Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ and assume $Q=0$.
+If $\mathcal{F}$ has the form (71) for some $m>1$, then the maximal subsolution $\underline{F}$ and minimal supersolution $\overline{F}$ of (61) are given for $(x,t)\in\mathbb{R}\times(0,\infty)$ by the formula
+(72) $\underline{F}(x,t)=\overline{F}(x,t)=F_{\text{in}}(-\infty)+\int_{-\infty}^{x}\rho(y,t)\,dy,$
 where $\rho\geq 0$ is the unique continuous distributional solution of the initial-value problem
 (75) $\left\{\begin{array}[]{r l}\partial_{t}\rho-\frac{a}{m}\partial_{x}^{2}(\rho^{m})=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ \rho(0)=\partial_{x}F_{\text{in}}.\end{array}\right$.
-When $F_{\text{in}}$ is uniformly Lipschitz, this result is a special case of [3, Theorem 3.1].
-Unless otherwise noted, basic properties of the PME used in the proof that follows can be found in the monograph [21].
+Furthermore, the following statements hold:
+  • (i)
+    For any $t>0$, the functions $\underline{F}(\cdot,t)$ and $F_{\text{in}}$ coincide on the set $\{-\infty,+\infty\}$.
+  • (ii)
+    If $\{F_{\text{in}}^{(n)}\}$ is some sequence in $CDF(\overline{\mathbb{R}})$ converging vaguely to $F_{\text{in}}$, and if $\{\underline{F}^{(n)}\}$ are the corresponding maximal subsolutions of (61), then, for any compact set $K\subseteq\mathbb{R}\times(0,+\infty)$,
+    $\lim_{n\to+\infty}\sup\left\{|\underline{F}^{(n)}(x,t)-\underline{F}(x,t)|\,\mid\,(x,t)\in K\right\}=0.$
+When $F_{\text{in}}$ is uniformly Lipschitz, the representation formula obtained above is a special case of [3, Theorem 3.1].
 Proof.
-First, assume that $F_{\text{in}}$ is smooth and $\partial_{x}F_{\text{in}}$ is bounded and nowhere vanishing.
-In this case, classical results on the PME imply that there is a unique smooth solution $\rho$ of (64), which is positive and satisfies $\int_{-\infty}^{\infty}\rho(y,t)\,dy=F_{\text{in}}(+\infty)-F_{\text{in}}(-\infty)$ for all $t>0$.
-Since $\rho$ is smooth, the function $F(x,t)=\int_{-\infty}^{x}\rho(y,t)\,dy$ is also, and, thus, it satisfies $\partial_{t}F-a|\partial_{x}F|^{m-1}\partial_{x}^{2}F=\partial_{t}F-\frac{a}{m}\partial_{x}(\partial_{x}F)^{m}=0$ classically.
-As is well-known, being a classical solution, it is also a viscosity solution.
-If $F_{\text{in}}$ is merely absolutely continuous with $\partial_{x}F_{\text{in}}\in L^{1}(\mathbb{R})$, nonetheless $\partial_{x}F_{\text{in}}$ can be approximated in $L^{1}(\mathbb{R})$ by smooth, nowhere-vanishing functions.
-The $L^{1}$ contraction property of the PME then implies that the corresponding solutions converge in $L^{\infty}([0,\infty);L^{1}(\mathbb{R}))$, which implies uniform convergence of the antiderivatives.
-Thus, since the viscosity solution property is stable under uniform convergence, even if $\partial_{x}F_{\text{in}}$ is merely in $L^{1}(\mathbb{R})$, the unique viscosity solution of (57) in this case is still given as the antiderivative of $\rho$.
-Finally, it remains to consider the case when $\partial_{x}F_{\text{in}}$ is only a finite positive measure (so that $F_{\text{in}}$ may have discontinuities).
-Let $\{F^{(N)}_{\text{in}}\}\subseteq CDF(\overline{\mathbb{R}})$ be smooth functions such that $F^{(N)}_{\text{in}}\leq F^{(N+1)}_{\text{in}}<F_{\text{in}}$ for each $N$, $F^{(N)}_{\text{in}}\to F_{\text{in}}$ at points of continuity as $N\to\infty$.
-If $\rho_{N}$ is the corresponding solution of the PME (64) and $F_{N}$, its antiderivative, then $F_{N}\leq\overline{F}$ for each $N$ since $F_{N}$ is a viscosity subsolution of (57).
-At the same time, some subsequence of $\{\rho_{N}\}$ converges locally uniformly in $\mathbb{R}\times(0,\infty)$ as $N\to\infty$ to a continuous, nonnegative distributional solution $\rho$ of (64) with initial trace equal to the measure $\partial_{x}F_{\text{in}}$; this follows, for instance, from the fact that, in this one-dimensional setting, the pressure $\{\rho_{N}^{m-1}\}$ is both uniformly bounded and uniformly Lipschitz in $\mathbb{R}\times[\tau,+\infty)$ for any $\tau>0$, independently of $N$ (see [21, Chapter 14]).
-A classical result of Dahlberg and Kenig [14] implies that $\rho$ is unique, hence the entire sequence $\{\rho_{N}\}$ converges to $\rho$.
-At the level of antiderivatives, this implies $F_{N}\to F$ as $N\to\infty$, where $F$ is given by (64) and is a discontinuous viscosity solution of (57).
-By comparison, $\overline{F}\leq F$, hence actually $F=\overline{F}$.
-Approximating $F_{\text{in}}$ from above, one similarly finds $F=\underline{F}$. ∎
+The uniqueness and claimed representation formula for the solution follows from the classical uniqueness theorem for the PME by Dahlberg and Kenig [16].
+To see this, first, note that if $\partial_{x}F_{\text{in}}$ is smooth and positive, then it is well-known that the solution $\rho$ of the PME is smooth (up to time $t=0$), hence the antiderivative given by (72) is a smooth solution of (61).
+It is classical that smooth solutions are viscosity solutions.
+It is also classical in this setting that $\int_{-\infty}^{\infty}\rho(y,t)\,dy=F_{\text{in}}(+\infty)-F_{\text{in}}(-\infty)$ for each $t>0$, as in item (i).
+It will be convenient later in the proof to note that, in this one-dimensional setting, since the initial data is a subprobability measure, the pressure $\rho^{m-1}$ associated with any of the smooth solutions above satisfies the following estimates
+(76) $\tau^{\frac{m-1}{m+1}}|\rho^{m-1}|+\tau^{\frac{m}{m+1}}|\partial_{x}(\rho^{m-1})|+\tau^{\frac{2m}{m+1}}|\partial_{t}(\rho^{m-1})|\leq C\quad\text{in}\,\,\mathbb{R}\times(\tau,+\infty),$
+where $C>0$ is a universal constant and $\tau>0$ is arbitrary (see [24, Chapter 15]).
+If $F_{\text{in}}$ is absolutely continuous with $\partial_{x}F_{\text{in}}\in L^{1}(\mathbb{R})$, then it can be approximated by smooth, strictly increasing functions $\{F_{\text{in}}^{(N)}\}$, with $\{\partial_{x}F_{\text{in}}^{(N)}\}$ converging in $L^{1}$, and the corresponding solutions $\{\rho^{(N)}\}$ of the PME then converge in $L^{\infty}([0,+\infty);L^{1}(\mathbb{R}))$ by the $L^{1}$ contractivity property of the equation (see [24, Chapter 9]) to the solution $\rho$ started from $\partial_{x}F_{\text{in}}$.
+This convergence implies uniform convergence of the antiderivatives: Since the uniform limit of viscosity solutions is a viscosity solution, once again (72) provides the viscosity solution in this case.
+Note also that item (i) remains true and the bounds (76) are preserved in the limit.
+Finally, if $\partial_{x}F_{\text{in}}$ is merely a finite measure, consider sequences of smooth functions $\{F^{(N),+}_{\text{in}}\}$ and $\{F^{(N),-}_{\text{in}}\}$ as in Lemma 1, chosen in such a way that $F^{(N),\pm}_{\text{in}}(\bar{x})\to F_{\text{in}}(\bar{x})$ as $N\to+\infty$ for each $\bar{x}\in\{-\infty,+\infty\}$.
+By the bounds (76), the corresponding solutions of the PME $\{\rho^{(N),+}\}$ and $\{\rho^{(N),-}\}$ started from the data $\{\partial_{x}F^{(N),+}_{\text{in}}\}$ and $\{\partial_{x}F^{(N),-}\}$ are uniformly bounded and equicontinuous in $\mathbb{R}\times(\tau,+\infty)$ for any $\tau>0$.
+Therefore, the antiderivatives $\{F^{(N),+}\}$ and $\{F^{(N),-}\}$ become equicontinuous after a positive amount of time.
+Since they are, respectively, monotone decreasing and monotone increasing in $N$, the limits $F^{+}=\lim_{N\to\infty}F^{(N),+}$ and $F^{-}=\lim_{N\to\infty}F^{(N),-}$ exist.
+Since $F^{(N),-}\leq F^{-}\leq F^{+}\leq F^{(N),+}$ holds in $\mathbb{R}\times[0,+\infty)$ for any $N$, $F^{-}(\cdot,t)$ and $F^{+}(\cdot,t)$ both converge vaguely to $F_{\text{in}}$ as $t\downarrow 0$.
+By Arzelà-Ascoli, the derivatives $\partial_{x}F^{-}$ and $\partial_{x}F^{+}$ are continuous distributional solutions of the PME in $\mathbb{R}\times(0,+\infty)$, and the vague convergence of $F^{-}$ and $F^{+}$ at time $t=0$ implies that the initial trace of $\partial_{x}F^{-}$ and $\partial_{x}F^{+}$ equals $\partial_{x}F_{\text{in}}$.
+Therefore, by Dahlberg-Kenig, $\partial_{x}F^{-}=\partial_{x}F^{+}$ in $\mathbb{R}\times(0,+\infty)$, which implies the difference $F^{+}-F^{-}$ depends only on time.
+Since item (i) holds for regular solutions, and, in the construction above, $F^{(N),\pm}_{\text{in}}(\bar{x})\to F_{\text{in}}(\bar{x})$ for each $\bar{x}\in\{-\infty,+\infty\}$, it follows that $F^{+}(\bar{x},t)=F^{-}(\bar{x},t)=F_{\text{in}}(\bar{x})$ also holds for any $t$.
+Therefore, $F^{+}=F^{-}$, and so (72) follows by Lemma 1 and item (i) holds.
+Item (ii) follows as in the previous proposition.∎
+Case $Q\neq 0$. Next, using only the comparison principle, uniqueness in the case $Q=0$ is upgraded to the case when $Q$ is nonzero.
+Proposition 16.
+The identity $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,\infty)$ still holds for the initial-value problem (61) with arbitrary data in $CDF(\overline{\mathbb{R}})$ if $\mathcal{F}$ satisfies the assumptions of Propositions 14 or 15 and $Q$ satisfies (63) and (64).
+Proof.
+Fix $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$.
+If $F_{\text{in}}$ is continuous, then Proposition 12 implies there is a unique viscosity solution $F$ of (61) and $\overline{F}=\underline{F}=F$.
+Otherwise, if it has jumps, fix sequences $\{F_{\text{in}}^{(N),+}\}$ and $\{F_{\text{in}}^{(N),-}\}$ of continuous functions in $CDF(\overline{\mathbb{R}})$ satisfying (67) and (68) as in Lemma 1 and, in addition,
+(77) $\lim_{N\to\infty}F_{\text{in}}^{(N),-}(\pm\infty)=\lim_{N\to\infty}F_{\text{in}}^{(N),+}(\pm\infty)=F_{\text{in}}(\pm\infty).$
+For each $N$, let $F^{(N),+}$ and $F^{(N),-}$ denote the unique viscosity solutions of (61) with initial data $F_{\text{in}}^{(N),+}$ and $F_{\text{in}}^{(N),-}$, respectively.
+By construction and the comparison principle, the bounds $F^{(N),-}\leq\underline{F}\leq\overline{F}\leq F^{(N),+}$ hold pointwise in $\mathbb{R}\times[0,+\infty)$.
+Therefore, it only remains to show that, for any $t>0$,
+(78) $\lim_{N\to\infty}\sup\left\{F^{(N),+}(x,t)-F^{(N),-}(x,t)\,\mid\,x\in\mathbb{R}\right\}=0.$
+To this end, since $Q$ is bounded in $[0,1]$, notice that the modified functions $\tilde{F}^{(N),+}$ and $\tilde{F}^{(N),-}$ given by
+$\tilde{F}^{(N),\pm}(x,t)=F^{(N),\pm}(x,t)\mp C(Q)t,\quad\text{where}\quad C(Q)=\sup\left\{|Q(q)|\,\mid\,0\leq q\leq 1\right\},$
+are, respectively, sub- and supersolutions of the PDE $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)=0$ with $Q$ removed.
+Therefore, if $G^{(N),+}$ and $G^{(N),-}$ are the viscosity solutions of this PDE corresponding to initial data $F^{(N),+}_{\text{in}}$ and $F^{(N),-}_{\text{in}}$, respectively, then the comparison principle implies
+(79) $F^{(N),+}-C(Q)t=\tilde{F}^{(N),+}\leq G^{(N),+}\quad\text{and}\quad G^{(N),-}\leq\tilde{F}^{(N),-}=F^{(N),-}+C(Q)t,$
+where all the inequalities above hold pointwise in $\mathbb{R}\times[0,+\infty)$.
+On the other hand, by Propositions 14 or 15, if $G$ is the discontinuous viscosity solution of the PDE $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)=0$ with initial datum $F_{\text{in}}$, then, for any $\delta>0$, the functions $G^{(N),+}(\cdot,\delta)$ and $G^{(N),-}(\cdot,\delta)$ converge pointwise in $\mathbb{R}$ to $G(\cdot,\delta)$.
+Since the constants $G^{(N),+}(\pm\infty,\delta)$ and $G^{(N),-}(\pm\infty,\delta)$ also converge to $G(\pm\infty,\delta)$ by (77) and item (i) in the aforementioned propositions, Dini’s Theorem (applied to the compact metric space $\mathbb{R}\cup\{-\infty,+\infty\}$) implies that this convergence is actually uniform:
+$\lim_{N\to\infty}\sup\left\{G^{(N),+}(x,\delta)-G^{(N),-}(x,\delta)\,\mid\,x\in\mathbb{R}\right\}=0.$
+Therefore, by (79),
+(80) $\lim_{N\to\infty}\sup\left\{F^{(N),+}(x,\delta)-F^{(N),-}(x,\delta)\,\mid\,x\in\mathbb{R}\right\}\leq 2C(Q)\delta.$
+Finally, moving back to the original problem (i.e., with the reaction term $Q$), the estimate (66) in Theorem 4 implies that, for any $t>\delta$,
+$\limsup_{N\to\infty}\sup\left\{F^{(N),+}(x,t)-F^{(N),-}(x,t)\,\mid\,x\in\mathbb{R}\right\}\leq 2C(B)\delta e^{L(t-\delta)}.$
+After sending $\delta\downarrow 0$, this becomes (78), as desired. ∎
 A.3. Viscosity Solutions in $CDF(\overline{\mathbb{R}})$
@@
 ∎
-Appendix B Scaling Limits of Monotone Semigroups in $CDF(\overline{\mathbb{R}})$
-This appendix proves Theorem 3, the extension of the convergence framework of [6] to approximations of PDE’s posed in $CDF(\overline{\mathbb{R}})$.
+Appendix B Scaling Limits of Some Monotone Semigroups in $CDF(\overline{\mathbb{R}})$
+This appendix proves Theorem 3, an extension of the convergence framework of [7] to approximations of PDE’s posed in $CDF(\overline{\mathbb{R}})$.
 As described already briefly in the introduction, technical issues arise stemming from the fact that $CDF(\overline{\mathbb{R}})$ is not a vector space, and the smooth test functions used in the definition of viscosity solution are not necessarily nondecreasing.
@@
 Lemma 4.
-Let $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be a continuous function that is nondecreasing in the second argument.
-An upper semicontinuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ satisfies $\partial_{t}F\leq\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)$ in the viscosity sense if and only if it has the following property: If $\varphi$ is a smooth test function that touches $F$ from above at some point $(x_{0},t_{0})$ and
+Let $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be a continuous function that is nondecreasing in the second argument and $Q:\mathbb{R}\to\mathbb{R}$ be continuous.
+An upper semicontinuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ satisfies $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)\leq 0$ in the viscosity sense if and only if it has the following property: If $\varphi$ is a smooth test function that touches $F$ from above at some point $(x_{0},t_{0})$ and
 $\text{either (i)}\quad\partial_{x}\varphi(x_{0},t_{0})\neq 0,\quad\text{or (ii)}\quad\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0,$
@@
 Proof.
-The fact that this weaker definition of viscosity subsolution is equivalent to the usual one can be proved exactly as in [8, Proposition 2.2] (see also [20, Proof of Theorem 1.2]). ∎
+The fact that this weaker definition of viscosity subsolution is equivalent to the usual one can be proved exactly as in [8, Proposition 2.2] (see also [23, Proof of Theorem 1.2]). ∎
 Remark 3.
-A corresponding version of Lemma 3 holds for supersolutions, which can be deduced from the transformation G(x,t)=1F(x,t)G(x,t)=1-F(-x,t).
+A corresponding version of Lemma 4 holds for supersolutions, as can be readily checked using the transformation G(x,t)=1F(x,t)G(x,t)=1-F(-x,t).
 Now Lemma 4 can be leveraged in conjunction with Lemma 3 above to resolve the issue when $\partial_{x}\varphi=0$ at the contact point: There is no loss of generality assuming that $\partial_{x}^{2}\varphi=0$ also holds there, and then the lemma shows that $\varphi$ can be taken to be globally nondecreasing in the spatial variable.
 B.1. Proof of Theorem 3
-Since $F_{\epsilon}$ is discontinuous in time, it will clarify the picture somewhat to observe that the statement that $F_{\epsilon}\to F$ locally uniformly in $\mathbb{R}\times[0,\infty)$ is equivalent to the following: For each $(x,t)\in\mathbb{R}\times[0,\infty)$,
-(65) $\lim_{\delta\downarrow 0}\sup\left\{|F^{\epsilon}_{n}(y)-F(x,t)|\,\mid\,|y-x|+|\epsilon^{\alpha}n-t|\leq\delta\right\}=0.$
-This is the precise statement that will be proved below.
-Since the operators $(T_{\epsilon})_{\epsilon>0}$ are monotone and nondecreasing continuous functions can be approximated locally uniformly from above and below by smooth functions, it suffices to consider the case when $F^{\epsilon}_{0}=F_{\text{in}}$ for each $\epsilon>0$ and $F_{\text{in}}$ has a bounded, uniformly continuous second derivative.
+It is technically convenient to proceed by proving the following statement: For each $(x,t)\in\mathbb{R}\times[0,\infty)$,
+(81) $\lim_{\delta\downarrow 0}\sup\left\{|F^{(N)}_{n}(y)-F(x,t)|\,\mid\,|\frac{y}{\delta^{(N)}}-x|+|N^{-1}n-t|+N^{-1}\leq\delta\right\}=0.$
+Notice that this implies that $F_{N}\to F$ locally uniformly in $\mathbb{R}\times[0,+\infty)$ as $N\to+\infty$.
+Since the operators $\{T^{(N)}\}$ are monotone, and nondecreasing continuous functions can be approximated locally uniformly from above and below by smooth functions, it suffices to consider the case when equality holds $F^{(N)}_{0}=F_{\text{in}}$ for each $N$ and $F_{\text{in}}$ has bounded, uniformly continuous first and second derivatives.
 This will be assumed for the remainder of the proof.
-Throughout what follows, define the constants $c_{-},c_{+}\in[0,1]$ by
-(66) $c_{-}=F_{\text{in}}(-\infty)=F_{\epsilon}(+\infty,t),\quad c_{+}=F_{\text{in}}(+\infty)=F_{\epsilon}(-\infty,t),$
-where property (iii) implies that the identities above hold independently of $t>0$.
-Step 1: Lipschitz Estimate in Time. Since $F_{\text{in}}$ is currently assumed to have a bounded, uniformly continuous second derivative, by the consistency condition (iv),
-$\epsilon^{-\alpha}\mathscr{L}_{\epsilon}F_{0}^{\epsilon}=\epsilon^{-\alpha}\mathscr{L}_{\epsilon}F_{\text{in}}=\mathcal{F}(\partial_{x}F_{\text{in}},\partial^{2}_{x}F_{\text{in}})+R_{\epsilon}$
-where $R_{\epsilon}(x)\to 0$ uniformly with respect to $x$ as $\epsilon\to 0$.
-Thus, since $\mathcal{F}$ is bounded on compact sets by continuity, there is a constant $C_{0}>0$ such that, for sufficiently small $\epsilon$,
-$\|T_{\epsilon}F_{0}^{\epsilon}-F_{0}^{\epsilon}\|_{\text{sup}}\leq C_{0}\epsilon^{\alpha}\quad\text{in}\,\,\mathbb{R}.$
-In Proposition 12 below, the assumptions (i)-(iii) are shown to imply the following contractivity property of $T_{\epsilon}$:
-$\|T_{\epsilon}F-T_{\epsilon}G\|_{\text{sup}}\leq\|F-G\|_{\text{sup}}.$
-Thus, the estimate of $T_{\epsilon}F_{0}^{\epsilon}-F_{0}^{\epsilon}$ may be iterated to find
-$\|F^{\epsilon}_{n+1}-F_{n}^{\epsilon}\|_{\text{sup}}=\|T_{\epsilon}F^{\epsilon}_{n}-F_{n}^{\epsilon}\|_{\text{sup}}\leq C_{0}\epsilon^{\alpha}\quad\text{for each}\,\,n\in\mathbb{N}_{0}.$
-Note, in particular, that $|F^{\epsilon}_{n}(x)-F_{\text{in}}(x)|\leq C_{0}\epsilon^{\alpha}n$ uniformly in $(x,n)$ as $\epsilon\to 0$.
-Step 3: Convergence. The proof uses the classical half-relaxed limit method. Let $F_{\star}$ and $F^{\star}$ be the half-relaxed limits of $\{F^{\epsilon}_{n}\}$ defined by
-$F_{\star}(x,t)$ $=\lim_{\delta\downarrow 0}\inf\left\{F_{n}^{\epsilon}(y)\,\mid\,|x-y|+|\epsilon^{\alpha}n-t|+\epsilon\leq\delta\right\},$
-$F^{\star}(x,t)$ $=\lim_{\delta\downarrow 0}\sup\left\{F_{n}^{\epsilon}(y)\,\mid\,|x-y|+|\epsilon^{\alpha}n-t|+\epsilon\leq\delta\right\},$
-To prove the local convergence statement (65), it suffices to show that $F^{\star}=F_{\star}=F$ (see [13, Remark 6.4] or [7, Lemma 6.2]).
-To do this, since $F_{\star}\leq F^{\star}$ is immediate, it is only necessary to show that $F^{\star}\leq F\leq F_{\star}$.
+Throughout what follows, define functions $c_{-},c_{+}:[0,\infty)\to[0,1]$ by
+(82) $\frac{dc_{\pm}}{dt}=-Q(c_{\pm}),\quad c_{+}(0)=F_{\text{in}}(-\infty),\quad c_{+}(0)=F_{\text{in}}(+\infty).$
+Assumption (vi) ensures that these ODE’s are wellposed. For each $N$, let $\{c_{-}^{(N)}(n)\}$ and $\{c_{+}^{(N)}(n)\}$ be the sequences of numbers in $[0,1]$ defined analogously by
+$c_{\pm}^{(N)}(n)=T^{(N)}c_{\pm}(n-1)=c_{\pm}^{(N)}(n-1)-Q^{(N)}(c_{\pm}^{(N)}(n-1)),$
+$c_{+}^{(N)}(0)=c_{+}(0),\quad c_{-}^{(N)}(0)=c_{-}(0).$
+(Assumptions (i) and (iv) imply that $T^{(N)}$ maps constant functions to constant functions, and $\mathcal{L}^{(N)}$ vanishes on constants, as explained in the discussion in Section 5.1.) By assumption (vi) and standard ODE arguments, for any $T>0$,
+(83) $\lim_{\delta\downarrow 0}\sup\left\{|c_{\pm}^{(N)}(n)-c_{\pm}(t)|\,\mid\,|N^{-1}n-t|+N^{-1}\leq\delta,\,\,t\in[0,T]\right\}=0.$
+Further, by assumption (iv), for any $N$,
+(84) $c_{-}^{(N)}(n)=F^{(N)}_{n}(-\infty)\leq F^{(N)}_{n}\leq F^{(N)}_{n}(+\infty)=c_{+}^{(N)}(n)\quad\text{for each}\,\,n\geq 0.$
+Step 1: Lipschitz Estimate in Time. Since $F_{\text{in}}$ is currently assumed to have bounded, uniformly continuous first and second derivatives, by the consistency condition (v),
+$N\mathscr{L}^{(N)}F_{0}^{(N)}=N\mathscr{L}^{(N)}F_{\text{in}}=\mathcal{F}(\partial_{x}F_{\text{in}},\partial^{2}_{x}F_{\text{in}})+R^{(N)}$
+where $R^{(N)}\to 0$ uniformly as $N\to+\infty$.
+Thus, since $\mathcal{F}$ is bounded on compact sets by continuity and the reaction term $Q^{(N)}$ is of order $N^{-1}$ by assumption (vi), there is a constant $C_{0}>0$ such that, for sufficiently large $N$,
+$\|T^{(N)}F_{0}^{(N)}-F_{0}^{(N)}\|_{\text{sup}}\leq C_{0}N^{-1}.$
+By Proposition 17 below, the assumptions (i)-(iv) and (vi) imply the following contractivity property of $T^{(N)}$:
+$\|T^{(N)}F-T^{(N)}G\|_{\text{sup}}\leq(1+LN^{-1})\|F-G\|_{\text{sup}}.$
+Thus, the estimate of $F^{(N)}_{1}-F_{0}^{(N)}=T^{(N)}F_{0}^{(N)}-F_{0}^{(N)}$ can be iterated to find
+$\|F^{(N)}_{n+1}-F_{n}^{(N)}\|_{\text{sup}}=\|T^{(N)}F^{(N)}_{n}-T^{(N)}F_{n-1}^{(N)}\|_{\text{sup}}\leq C_{0}N^{-1}(1+N^{-1}L)^{n}\quad\text{for each}\,\,n\in\mathbb{N}.$
+Summing this estimate in $n$, one obtains a constant $C_{0}^{\prime\prime}>0$ such that that $|F^{(N)}_{n}(x)-F_{\text{in}}(x)|\leq C_{0}^{\prime\prime}N^{-1}(1+N^{-1}L)^{n}$ uniformly in $(x,n)$ as $N\to+\infty$.
+In view of the definition of $F_{N}$, this implies that, for each $T>0$,
+$\sup\left\{|F_{N}(x,t)-F_{\text{in}}(x)|\,\mid\,(x,t)\in\mathbb{R}\times[0,T]\right\}\leq C_{0}^{\prime\prime}N^{-1}(1+LN^{-1})^{[NT]+1}.$
+Step 2: Convergence. The remainder of the proof uses the classical half-relaxed limit method. Let $F_{\star}$ and $F^{\star}$ be the half-relaxed limits of $\{F^{(N)}_{n}\}$ defined by
+$F_{\star}(x,t)$ $=\lim_{\delta\downarrow 0}\inf\left\{F_{n}^{(N)}(y)\,\mid\,|\delta^{(N)}y-x|+|N^{-1}n-t|+N^{-1}\leq\delta\right\},$
+$F^{\star}(x,t)$ $=\lim_{\delta\downarrow 0}\sup\left\{F_{n}^{(N)}(y)\,\mid\,|\delta^{(N)}y-x|+|N^{-1}n-t|+N^{-1}\leq\delta\right\},$
+As is well-known, to prove the local convergence statement (81), it suffices to show that $F^{\star}=F_{\star}=F$ (see [15, Remark 6.4] or [10, Lemma 6.2]).
+To do this, since the bound $F_{\star}\leq F^{\star}$ follows immediately from the definitions, it is only necessary to show that $F^{\star}\leq F\leq F_{\star}$.
 This will follow from the comparison principle, Theorem 4, as soon as it is established that $F^{\star}$ and $F_{\star}$ define respectively a viscosity sub- and a viscosity supersolution of the initial value problem (61).
-Step 4: $F^{\star}$ is a Subsolution.
-In view of what was proved in Step 1, it is clear that $F^{\star}(\cdot,0)=F_{\text{in}}$ in the sense of (58).
-Therefore, it only remains to establish that $\partial_{t}F^{\star}-\mathcal{F}(\partial_{x}F^{\star},\partial_{x}^{2}F^{\star})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$.
-In order to avoid boundary issues that can arise if $F^{\star}$ is somewhere equal to one, it will be convenient to pull $F^{\star}$ down: To see that $F^{\star}$ is a subsolution, it suffices to prove that, for any $\delta\in(0,c_{+}-c_{-})$, the function $G^{\star}:\mathbb{R}\times[0,\infty)\to[c_{-},c_{+}-\delta]$ defined by $G^{\star}=\max\{c_{-},F^{\star}-\delta\}$ satisfies
-$\partial_{t}G^{\star}\leq\mathcal{F}(\partial_{x}G^{\star},\partial_{x}^{2}G^{\star}).$
-For the rest of this step, fix $\delta$ and consider $G^{\star}$.
-Note also that if $\{G^{\epsilon}_{n}\}$ is the sequence defined for $\epsilon>0$ by $G^{\epsilon}_{n}=\max\{c_{-},F^{\epsilon}_{n}-\delta\}$, then Lemma 4 below implies that
-(67) $G^{\epsilon}_{n+1}\leq T_{\epsilon}G^{\epsilon}_{n}\quad\text{pointwise for each}\,\,n\in\mathbb{N}_{0}.$
+Step 3: $F^{\star}$ is a Subsolution.
+In view of what was proved in Step 1 and the uniform continuity of $F_{\text{in}}$, it is clear that $F^{\star}$ coincides with $F_{\text{in}}$ in the sense of (65).
+Therefore, to see that $F^{\star}$ is a viscosity subsolution of (61), it only remains to establish that $\partial_{t}F^{\star}-\mathcal{F}(\partial_{x}F^{\star},\partial_{x}^{2}F^{\star})+Q(F^{\star})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$.
+In order to avoid boundary issues that can arise if $F^{\star}$ is somewhere equal to one, it will be convenient to pull $F^{\star}$ down: To see that $F^{\star}$ is a subsolution, it suffices to prove that, for any $\delta\in(0,1)$, the function $G^{\star}:\mathbb{R}\times[0,\infty)\to[0,1-\delta]$ defined by $G^{\star}(x,t)=\max\{F^{\star}(x,t)-\delta e^{Lt},0\}$ satisfies
+$\partial_{t}G^{\star}-\mathcal{F}(\partial_{x}G^{\star},\partial_{x}^{2}G^{\star})+Q(G^{\star})\leq 0\quad\text{in the viscosity sense in}\,\,\mathbb{R}\times(0,+\infty).$
+(By standard arguments, the subsolution property of $F^{\star}$ can then be recovered upon sending $\delta\downarrow 0$.) For the rest of this step, fix $\delta$ and consider $G^{\star}$.
+Note also that if $\{G^{(N)}_{n}\}$ is the sequence defined for a given $N$ by $G^{(N)}_{n}=\max\{F^{(N)}_{n}-\delta(1+LN^{-1})^{n},0\}$, then Lemma 5 below implies that
+(85) $G^{(N)}_{n}\leq T^{(N)}G^{(N)}_{n-1}\quad\text{pointwise for each}\,\,n\in\mathbb{N}.$
 Further, by definition of $G^{\star}$,
-(68) $G^{\star}(x,t)=\lim_{\delta\downarrow 0}\sup\left\{G^{\epsilon}_{n}(y)\,\mid\,|x-y|+|\epsilon^{\alpha}n-t|+\epsilon\leq\delta\right\}.$
-Thus, in passing from $F$’s to $G$’s, nothing has changed except the tighter upper bound $G^{\epsilon}_{n}\leq c_{+}-\delta<1$ holds for each $n$ and the evolution equation has been replaced by an inequality in (67).
+(86) $G^{\star}(x,t)=\lim_{\delta\downarrow 0}\sup\left\{G^{(N)}_{n}(y)\,\mid\,|\delta^{(N)}y-x|+|N^{-1}n-t|+N^{-1}\leq\delta\right\}.$
+Thus, in passing from $F$’s to $G$’s, nothing has changed except the stricter upper bound $G^{(N)}_{n}\leq 1-\delta(1+LN^{-1})^{n}$ holds for each $n$ and the evolution equation has been weakened to an inequality in (85).
 Suppose that $\varphi$ is a smooth function defined in some open subset of space-time that touches $G^{\star}$ from above at some point $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ in its domain.
-The goal is to show that $\partial_{t}\varphi(x_{0},t_{0})\leq\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))$.
-Notice that one technicality can be ruled out immediately: Since $G^{\star}\geq c_{-}$ pointwise, if $\varphi(x_{0},t_{0})=c_{-}$, then
-$\partial_{t}\varphi(x_{0},t_{0})=\partial_{x}\varphi(x_{0},t_{0})=0,\quad\partial_{x}^{2}\varphi(x_{0},t_{0})\geq 0,$
-which implies the desired inequality by (51). Therefore, from now on, assume that $c_{-}<\varphi(x_{0},t_{0})$.
-According to Lemma 3, there is no loss of generality in assuming that either (a) $|\partial_{x}\varphi(x_{0},t_{0})|>0$ or (b) $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$.
+The goal is to show that
+(87) $\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))+Q(\varphi(x_{0},t_{0}))\leq 0.$
+Since $G^{\star}(x,t)\geq\max\{c_{-}(t)-\delta e^{Lt},0\}$ pointwise in $\mathbb{R}\times[0,+\infty)$ by (84) and the definition of $G^{\star}$, it is convenient to consider two cases separately:
+$\text{(I)}\quad\varphi(x_{0},t_{0})=\max\{c_{-}(t_{0})-\delta e^{Lt_{0}},0\}\quad\text{and}\quad\text{(II)}\quad\varphi(x_{0},t_{0})>\max\{c_{-}(t_{0})-\delta e^{Lt_{0}},0\}.$
+Step 3, Case (I).
+First, consider case (I).
+Since $\varphi\geq G^{\star}$ holds in a neighborhood of $(x_{0},t_{0})$, $\varphi(x_{0},t_{0})=G^{*}(x_{0},t_{0})$, and $G^{\star}(x,t)\geq\max\{c_{-}(t)-\delta e^{Lt},0\}$ globally, in this case, $(x_{0},t_{0})$ is a local minimum of the function $(x,t)\mapsto\varphi(x,t)-\max\{c_{-}(t)-\delta e^{Lt},0\}$.
+In particular, $x_{0}$ is a local minimum of $x\mapsto\varphi(x,t_{0})$, hence
+$\partial_{x}\varphi(x_{0},t_{0})=0,\quad\partial_{x}^{2}\varphi(x_{0},t_{0})\geq 0.$
+Therefore, since $\mathcal{F}(0,0)=0$ by (48) and $\mathcal{F}$ is nondecreasing in the second argument,
+(88) $\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))=\mathcal{F}(0,\partial_{x}^{2}\varphi(x_{0},t_{0}))\geq 0.$
+Restricting attention this time to the slice $x=x_{0}$, the function $t\mapsto\varphi(x_{0},t)-\max\{c_{-}(t)-\delta e^{Lt},0\}$ has a local minimum at $t=t_{0}$, from which it readily follows that either
+$\partial_{t}\varphi(x_{0},t_{0})=\frac{d}{dt}\{c_{-}(t)-\delta e^{Lt}\}=-Q(c_{-}(t_{0}))-L\delta e^{Lt_{0}}\leq-Q(\varphi(x_{0},t_{0}))$
+if $c_{-}(t_{0})-\delta e^{Lt_{0}}>0$, or
+$\partial_{t}\varphi(x_{0},t_{0})=0\leq-Q(0)=-Q(\varphi(x_{0},t_{0}))$
+if $c_{-}(t_{0})-\delta e^{Lt_{0}}\leq 0$.
+(That $Q(0)\leq 0$ is a consequence of the fact that the function $c\mapsto c-Q(c)$ maps $[0,1]$ into itself by assumptions (i) and (iv); see the discussion in Section 5.1.) Combining the inequality $\partial_{t}\varphi(x_{0},t_{0})\leq-Q(\varphi(x_{0},t_{0}))$ with (88), one obtains the desired inequality (87).
+Step 3, Case (II). It only remains to consider case (II). That is, from now on, assume that $\varphi(x_{0},t_{0})>\max\{c_{-}(t_{0})-\delta e^{Lt_{0}},0\}$.
+According to Lemma 4, there is no loss of generality in assuming that either
+$\text{(a)}\quad\partial_{x}\varphi(x_{0},t_{0})\neq 0\quad\text{or}\quad\text{(b)}\quad\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0.$
 In case (a), since $G^{\star}$ is nondecreasing, it follows that $\partial_{x}\varphi(x_{0},t_{0})>0$, and Lemma 2 is applicable.
@@
 To avoid repetition, only case (b) will be considered in what follows.
-Therefore, assume henceforth that (b) holds.
-According to Lemma 2 with $A=c_{+}-\frac{\delta}{2}$, there is no loss of generality in fixing an $r>0$ and assuming that $\varphi(x,t)=\varphi_{1}(x)+\varphi_{2}(t)$ for $(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r]$, where $\varphi_{1}$ is a nondecreasing $C^{2}$ function with $\varphi_{1}(x)=\varphi_{1}(x_{0})$ for $x\leq x_{0}$ and $\varphi_{1}(x)=c_{+}-\frac{\delta}{2}$ for $x\geq x_{0}+r$; $\varphi_{2}$ is $C^{2}$ in $t_{0}-r,t_{0}+r]$ with $\varphi_{2}(t_{0})=0$; and
+Therefore, assume henceforth that (b) holds. According to Lemma 3 with $A=1-\frac{\delta}{2}$, there is no loss of generality in fixing an $r>0$ and assuming that
+$\varphi(x,t)=\varphi_{1}(x)+\varphi_{2}(t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r],$
+where $\varphi_{1}$ is a nondecreasing $C^{2}$ function with $\varphi_{1}(x)=\varphi_{1}(x_{0})$ for $x\leq x_{0}$ and $\varphi_{1}(x)=1-\frac{\delta}{2}$ for $x\geq x_{0}+r$; $\varphi_{2}$ is $C^{2}$ in $t_{0}-r,t_{0}+r]$ with $\varphi_{2}(t_{0})=0$; and
 $G^{\star}(x,t)\leq\varphi(x,t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r],$
 $(x_{0},t_{0})\in\{(x,t)\,\mid\,G^{\star}(x,t)=\varphi(x,t)\}\subseteq(-\infty,x_{0}]\times\{t_{0}\}.$
-Since $\varphi_{2}$ is continuous in $t_{0}-r,t_{0}+r]$ and $\varphi_{1}(-\infty)=\varphi(x_{0},t_{0})>c_{-}$, up to shrinking $r$ if necessary, there is no loss of generality in assuming that there are constants $c_{*},c^{*}\in[0,1]$ such that, for each $t\in[t_{0}-r,t_{0}+r]$,
-(69) $c_{-}<c_{*}\leq\varphi(-\infty,t)\quad\text{and}\quad c_{+}-\delta<c_{*}\leq\varphi(+\infty,t)\leq c_{+}-\frac{\delta}{4}.$
-Finally, for each $\epsilon>0$, let $c(\epsilon)$ be the constant
-$c(\epsilon)=\sup\left\{G^{\epsilon}_{n}(x)-\varphi(x,\epsilon^{\alpha}n)\,\mid\,x\in\mathbb{R},\,\,n\in\mathbb{N}\cap\epsilon^{-\alpha}[t_{0}-r,t_{0}+r]\right\}.$
-Notice that (69) and (66) imply that there is an $R>0$ such that, for any $\epsilon>0$,
-$\sup\left\{G^{\epsilon}_{n}(x)-\varphi(x,\epsilon^{\alpha}n)\,\mid\,|x|\geq R,\,\,n\in\mathbb{N}\cap\epsilon^{-\alpha}[t_{0}-r,t_{0}+r]\right\}<0.$
-Therefore, by the equation (68) relating $G^{\epsilon}_{n}$ to $G^{\star}$, it follows that there is a sequence $\{\epsilon_{k},x_{k},n_{k}\}$ such that $(\epsilon_{k},c(\epsilon_{k}))\to(0,0)$ as $k\to\infty$; $|x_{k}|\leq R$ and $|\epsilon_{k}^{\alpha}n_{k}-t|\leq r$ for each $k$; and
-(70) $G^{\epsilon_{k}}_{n_{k}}(x_{k})-\varphi(x_{k},\epsilon_{k}^{\alpha}n_{k})-c(\epsilon_{k})=0.$
-Further, since $\{(x,t)\,\mid\,G^{\star}(x,t)=\varphi(x,t)\}$ is contained in $(-\infty,x_{0}]\times\{t_{0}\}$, there is no loss of generality in assuming that there is a $y_{0}\leq x_{0}$ such that $(x_{k},\epsilon^{\alpha}_{k}n_{k})\to(y_{0},t_{0})$ and $G^{\star}(y_{0},t_{0})=\varphi(y_{0},t_{0})$.
-Note, in particular, that the inequality $y_{0}\leq x_{0}$ implies
-(71) $\lim_{k\to\infty}(\partial_{x}\varphi_{1}(x_{k}),\partial_{x}^{2}\varphi_{1}(x_{k}))=(0,0).$
-Finally, for notational ease, let $\Phi_{k}:\mathbb{R}\times[t_{0}-r,t_{0}+r]\to\mathbb{R}$ be the function $\Phi_{k}(x,t)=\varphi(x,t)+c(\epsilon_{k})$.
-Since $c(\epsilon_{k})\to 0$, the bounds (69) implies that $0\leq\Phi_{k}\leq 1$ pointwise for all $k$ large enough.
+Since $\varphi_{2}$ is continuous in $t_{0}-r,t_{0}+r]$ and $\varphi_{1}(-\infty)=\varphi(x_{0},t_{0})>\max\{c_{-}(t_{0})-\delta e^{Lt_{0}},0\}$, up to shrinking $r$ if necessary, there is no loss of generality in assuming that there are constants $c_{*},c^{*}\in[0,1]$ such that, for each $t\in[t_{0}-r,t_{0}+r]$,
+(89) $\max\{c_{-}(t)-\delta e^{Lt},0\}<c_{*}\leq\varphi(-\infty,t)\quad\text{and}\quad 1-\delta<c_{*}\leq\varphi(+\infty,t)\leq 1-\frac{\delta}{4}.$
+Finally, for each $N>0$, let $\nu(N)$ be the constant
+$\nu(N)=\sup\left\{G^{(N)}_{n}(y)-\varphi(\delta^{(N)}y,N^{-1}n)\,\mid\,y\in\mathbb{R},\,\,N^{-1}n\in[t_{0}-r,t_{0}+r]\right\}.$
+Notice that (89), (84), and (83) imply that there is an $R>0$ such that, for any $N$,
+$\sup\left\{G^{(N)}_{n}(y)-\varphi(\delta^{(N)}y,N^{-1}n)\,\mid\,\delta^{(N)}|y|\geq R,\,\,N^{-1}n\in[t_{0}-r,t_{0}+r]\right\}<0.$
+Therefore, by the equation (86) relating $\{G^{(N)}_{n}\}$ to $G^{\star}$, it follows that there is a sequence $\{(N_{k},y_{k},n_{k})\}$ such that $(N^{-1}_{k},\nu(N_{k}))\to(0,0)$ as $k\to\infty$; $\delta^{(N)}|y_{k}|\leq R$ and $|N^{-1}n_{k}-t_{0}|\leq r$ for each $k$; and
+(90) $G^{(N_{k})}_{n_{k}}(y_{k})-\varphi(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k})=\nu(N_{k})\quad\text{for each}\,\,k.$
+Further, since the set $\{(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r]\,\mid\,G^{\star}(x,t)=\varphi(x,t)\}$ is contained in $(-\infty,x_{0}]\times\{t_{0}\}$, there is no loss of generality in assuming that there is an $x^{\prime}_{0}\leq x_{0}$ such that $(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k})\to(x^{\prime}_{0},t_{0})$ and $G^{\star}(x^{\prime}_{0},t_{0})=\varphi(x^{\prime}_{0},t_{0})$.
+Note, in particular, that the inequality $x^{\prime}_{0}\leq x_{0}$ implies, by the constancy of $\varphi_{1}$ in $(-\infty,x_{0}]$,
+(91) $\lim_{k\to\infty}(\varphi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}\varphi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}^{2}\varphi_{1}(\delta^{(N_{k})}y_{k}))=(\varphi_{1}(x_{0}),0,0).$
+Finally, for notational ease, let $\Phi_{k}:\mathbb{R}\times[t_{0}-r,t_{0}+r]\to\mathbb{R}$ be the function $\Phi_{k}(y,t)=\varphi(\delta^{(N_{k})}y,t)+\nu(N_{k})$.
+Since $\nu(N_{k})\to 0$, the bounds (89) imply that $0\leq\Phi_{k}\leq 1$ pointwise for all $k$ large enough.
 That is, $\Phi_{k}\in CDF(\overline{\mathbb{R}})$ for large enough $k$.
-By the choice of $c(\epsilon_{k})$,
-$G^{\epsilon_{k}}_{n}\leq\Phi_{k}(\cdot,\epsilon^{\alpha}_{k}n)\quad\text{pointwise for each}\,\,n\in\mathbb{N}\cap\epsilon_{k}^{-\alpha}[t_{0}-r,t_{0}+r].$
-Therefore, by the monotonicity of $T_{\epsilon}$,
-$T_{\epsilon_{k}}G^{\epsilon_{k}}_{n_{k}-1}\leq T_{\epsilon_{k}}\Phi_{\epsilon_{k}}(\cdot,\epsilon^{\alpha}_{k}n_{k}-\epsilon^{\alpha}_{k}).$
+By the choice of $\nu(N_{k})$,
+$G^{(N_{k})}_{n}\leq\Phi_{k}(\cdot,N_{k}^{-1}n)\quad\text{pointwise for each}\,\,n\in[N(t_{0}-r),N(t_{0}+r)].$
+Therefore, by monotonicity (assumption (i)),
+$T^{(N_{k})}G^{(N_{k})}_{n_{k}-1}\leq T^{(N_{k})}\Phi_{k}(\cdot,N_{k}^{-1}(n_{k}-1)).$
 By (85), (90), and the definition of $\Phi_{k}$, this implies
-$\Phi_{k}(x_{k},\epsilon^{\alpha}_{k}n_{k})=G^{\epsilon_{k}}_{n_{k}}(x_{k})\leq T_{\epsilon_{k}}G^{\epsilon_{k}}_{n_{k}-1}(x_{k})\leq T_{\epsilon_{k}}\Phi_{k}(x_{k},\epsilon^{\alpha}_{k}n_{k}-\epsilon_{k}^{\alpha}).$
+$\Phi_{k}(y_{k},N_{k}^{-1}n_{k})=G^{(N_{k})}_{n_{k}}(y_{k})\leq T^{(N_{k})}G^{(N_{k})}_{n_{k}-1}(y_{k})\leq T^{(N_{k})}\Phi_{k}(y_{k},N_{k}^{-1}n_{k}-N_{k}^{-1}).$
 Upon rearranging, this becomes
-$\epsilon_{k}^{-\alpha}\Big(\Phi_{\epsilon_{k}}(x_{k},\epsilon_{k}^{\alpha}n_{k})-\Phi_{k}(x_{k},\epsilon^{\alpha}_{k}n_{k}-\epsilon^{\alpha}_{k})\Big)-\epsilon_{k}^{-\alpha}(\mathscr{L}_{\epsilon_{k}}\Phi_{k})(x_{k},\epsilon^{\alpha}_{k}n_{k}-\epsilon_{k}^{\alpha})\leq 0.$
-By assumption (ii),
-$\mathscr{L}_{\epsilon_{k}}\Phi_{k}(x_{k},\epsilon^{\alpha}_{k}n_{k}-\epsilon_{k}^{\alpha})=\mathscr{L}_{\epsilon_{k}}\varphi_{1}(x_{k}),$
-hence, by assumption (iv), (71), and the form of $\varphi$, sending $k\to\infty$ yields
-$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(0,0)=\partial_{t}\varphi_{2}(t_{0})-\lim_{k\to\infty}\mathcal{F}(\partial_{x}\varphi_{1}(x_{k}),\partial_{x}^{2}\varphi_{1}(x_{k}))\leq 0.$
-Since $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$ by assumption, this is the desired inequality.
-Step 5: $F_{\star}$ is a Supersolution.
+$N_{k}\Big(\Phi_{k}(y_{k},N_{k}^{-1}n_{k})-\Phi_{k}(y_{k},N_{k}^{-1}n_{k}-N_{k}^{-1})\Big)$
+$\quad\quad\quad\leq N_{k}(\mathscr{L}^{(N_{k})}\Phi_{k})(y_{k},N_{k}^{-1}(n_{k}-1))+N_{k}Q^{(N_{k})}(\Phi_{k}(y_{k},N_{k}^{-1}(n_{k}-1)).$
+By assumption (iii), the definition of $\Phi_{k}$, and the form of $\varphi$,
+$(\mathscr{L}^{(N_{k})}\Phi_{k})(y_{k},N_{k}^{-1}(n_{k}-1))=(\mathscr{L}^{(N_{k})}(\varphi_{1})_{\delta^{(N_{k})}})(y_{k}),$
+hence, by assumptions (v) and (vi), (91), and the form of $\varphi$, sending $k\to\infty$ above yields
+$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(0,0)+Q(\varphi(x_{0},t_{0})))$ $=\partial_{t}\varphi_{2}(t_{0})-\lim_{k\to\infty}\mathcal{F}(\partial_{x}\varphi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}^{2}\varphi_{1}(\delta^{(N_{k})}y_{k}))$
+$\quad+\lim_{k\to\infty}N_{k}Q^{(N_{k})}(\varphi(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k}))$
+$\leq 0.$
+Since $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$ by assumption, this yields (87) as desired.
+Step 4: $F_{\star}$ is a Supersolution.
 The proof that $F_{\star}$ is a supersolution of (61) can be reduced to the subsolution property of $F^{\star}$ after a suitable transformation.
@@
 $G_{\text{rev}}(x)=1-G_{-}(-x).$
-Let $\{\tilde{T}_{\epsilon}\}$ be the operators on $CDF(\overline{\mathbb{R}})$ obtained from $\{T_{\epsilon}\}$ by
-$\tilde{T}_{\epsilon}F=(T_{\epsilon}(F_{\text{rev}}))_{\text{rev}}.$
-It is straightforward to check that the family $\{\tilde{T}_{\epsilon}\}$ satisfies assumptions (i)-(iv), the only difference being $\mathcal{F}$ has to be replaced by the function $\tilde{\mathcal{F}}(p,X)=-\mathcal{F}(p,-X)$.
-Define $\{\tilde{F}^{\epsilon}_{n}\}$ by the rule $\tilde{F}^{\epsilon}_{n}=(F^{\epsilon}_{n})_{\text{rev}}$.
-Notice that the recursion $F^{\epsilon}_{n+1}=T_{\epsilon}F^{\epsilon}_{n}$ implies that $\tilde{F}^{\epsilon}_{n+1}=\tilde{T}_{\epsilon}\tilde{F}^{\epsilon}_{n}$.
-Therefore, if $\tilde{F}^{\star}$ is defined by
-$\tilde{F}^{\star}(x,t)=\lim_{\delta\downarrow 0}\sup\left\{\tilde{F}^{\epsilon}_{n}(y)\,\mid\,|y-x|+|\epsilon^{\alpha}n-t|+\epsilon\leq\delta\right\},$
-then $\tilde{F}^{\star}$ is a subsolution of the initial value problem (57) with $\mathcal{F}$ replaced by $\tilde{\mathcal{F}}$ and $F_{\text{in}}$ by $(F_{\text{in}})_{\text{rev}}$.
+Let $\{\tilde{T}^{(N)}\}$ be the operators on $CDF(\overline{\mathbb{R}})$ obtained from $\{T^{(N)}\}$ by
+$\tilde{T}^{(N)}F=(T^{(N)}(F_{\text{rev}}))_{\text{rev}}.$
+It is straightforward to check that the family $\{\tilde{T}^{(N)}\}$ satisfies assumptions (i)-(vi), the only difference being $\mathcal{F}$ has to be replaced by the function $\tilde{\mathcal{F}}(v,w)=-\mathcal{F}(v,-w)$ and $Q$ by $\tilde{Q}(q)=-Q(1-q)$.
+Define $\{\tilde{F}^{(N)}_{n}\}$ by the rule $\tilde{F}^{(N)}_{n}=(F^{(N)}_{n})_{\text{rev}}$.
+Notice that the recursion $F^{(N)}_{n}=T^{(N)}F^{(N)}_{n-1}$ implies that $\tilde{F}^{(N)}_{n}=\tilde{T}^{(N)}\tilde{F}^{(N)}_{n-1}$.
+Therefore, by Step 3, if $\tilde{F}^{\star}$ is defined by
+$\tilde{F}^{\star}(x,t)=\lim_{\delta\downarrow 0}\sup\left\{\tilde{F}^{(N)}_{n}(y)\,\mid\,|\delta^{(N)}y-x|+|N^{-1}n-t|+N^{-1}\leq\delta\right\},$
+then $\tilde{F}^{\star}$ is a subsolution of the initial value problem (61) with $\mathcal{F}$ replaced by $\tilde{\mathcal{F}}$, $Q$ by $\tilde{Q}$, and $F_{\text{in}}$ by $(F_{\text{in}})_{\text{rev}}$.
 Further, by definition, $\tilde{F}^{\star}(x,t)=1-F_{\star}(-x,t)$.
@@
 B.2. Proof of Corollary 4
-In the proof that follows, it it convenient to work with the half-relaxed limits $\liminf_{*}F_{\epsilon}$ and $\limsup_{*}F_{\epsilon}$ of the rescaled CDF $F_{\epsilon}$ defined by
-$(\liminf\nolimits_{*}F_{\epsilon})(x,t)$ $=\lim_{\delta\downarrow 0}\inf\left\{F_{\epsilon}(y,s)\,\mid\,|y-x|+|s-t|+\epsilon\leq\delta\right\},$
-$(\limsup\nolimits_{*}F_{\epsilon})(x,t)$ $=\lim_{\delta\downarrow 0}\sup\left\{F_{\epsilon}(y,s)\,\mid\,|y-x|+|s-t|+\epsilon\leq\delta\right\}.$
+In the proof that follows, it it convenient to work with the half-relaxed limits $\liminf_{*}F_{N}$ and $\limsup_{*}F_{N}$ of the rescaled CDF $F_{N}$ defined by
+$(\liminf\nolimits_{*}F_{N})(x,t)$ $=\lim_{\delta\downarrow 0}\inf\left\{F_{N}(x^{\prime},s)\,\mid\,|x^{\prime}-x|+|s-t|+N^{-1}\leq\delta\right\},$
+$(\limsup\nolimits_{*}F_{N})(x,t)$ $=\lim_{\delta\downarrow 0}\sup\left\{F_{\epsilon}(x^{\prime},s)\,\mid\,|x^{\prime}-x|+|s-t|+N^{-1}\leq\delta\right\}.$
 Proof of Corollary 4.
 Fix $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$.
-Suppose that $F_{\epsilon}(x,0)\to F_{\text{in}}(x)$ at each point of continuity $x$ of $F_{\text{in}}$.
-Let $\overline{F}$ and $\underline{F}$ be the maximal subsolution and the minimal supersolution of the problem (57).
-In view of Proposition 9 and well-known properties of half-relaxed limits (see [13, Remark 6.4] or [7, Lemma 6.2]), to see that $F_{\epsilon}$ converges locally uniformly to the discontinuous viscosity solution, it suffices to establish that
-$\overline{F}\leq\liminf\nolimits_{*}F_{\epsilon}\leq\limsup\nolimits^{*}F_{\epsilon}\leq\underline{F}\quad\text{in}\,\,\mathbb{R}\times(0,\infty).$
+Suppose that, at time $t=0$, $F_{N}(\cdot,0)\to F_{\text{in}}$ vaguely as $N\to+\infty$.
+Let $\underline{F}$ and $\overline{F}$ be the maximal subsolution and the minimal supersolution of the problem (61).
+In view of Proposition 13 and well-known properties of half-relaxed limits (see [15, Remark 6.4] or [10, Lemma 6.2]), to see that $F_{N}$ converges locally uniformly in $\mathbb{R}\times(0,+\infty)$ to the discontinuous viscosity solution as $N\to+\infty$, it suffices to establish that
+$\underline{F}\leq\liminf\nolimits_{*}F_{\epsilon}\leq\limsup\nolimits^{*}F_{\epsilon}\leq\overline{F}\quad\text{in}\,\,\mathbb{R}\times[0,\infty).$
 The details will be provided only for the lower bound as the upper bound follows similarly.
-For each $N\in\mathbb{N}$, define $F_{\epsilon,N}$ like $F_{\epsilon}$, albeit instead with the initial datum
-$F_{\epsilon,N}(x,0)=2^{N}\int_{0}^{\infty}F_{\epsilon}(x-y,0)\rho(2^{N}y)\,dy,$
+For each $M\in\mathbb{N}$, generate a rescaled CDF $F_{N,M}$ like $F_{N}$, albeit instead starting from the initial datum
+$F_{N,M}(x,0)=2^{M}\int_{0}^{\infty}F_{N}(x-y,0)\rho(2^{M}y)\,dy,$
 where $\rho:(0,\infty)\to(0,\infty)$ is a smooth function supported in $[1,2]$ with $\int_{0}^{\infty}\rho(y)\,dy=1$.
-From the fact that $\rho$ is supported in $[1,2]$ and $F_{\epsilon}$ is nondecreasing in $x$, it follows that $F_{\epsilon,N}(x,0)\leq F_{\epsilon,N+1}(x,0)\leq F_{\epsilon}(x,0)$ for every $x\in\mathbb{R}$.
-Due to the monotonicity assumption (i), this implies that $F_{\epsilon,N}\leq F_{\epsilon,N+1}\leq F_{\epsilon}$ in $\mathbb{R}\times[0,\infty)$ for each $\epsilon>0$ and $N\in\mathbb{N}$.
-By Theorem 3, $F_{\epsilon,N}\to F_{N}$ as $\epsilon\to 0$, where $F_{N}$ is the solution of (57) with initial datum $F_{\text{in}}^{(N)}$ given by $F_{\text{in}}^{(N)}(x)=2^{N}\int_{0}^{\infty}F_{\text{in}}(x-y)\rho(2^{N}y)\,dy$.
-Note that $F_{N}\leq\liminf\nolimits_{*}F_{\epsilon}$ pointwise in $\mathbb{R}\times[0,\infty)$ for any fixed $N$.
-Standard arguments (see [7, Theorem 6.2]) show that the function $F_{\infty}$ defined by
-$F_{\infty}(x,t)=\lim_{\delta\downarrow 0}\inf\left\{F_{N}(y,s)\,\mid\,|y-x|+|t-s|+N^{-1}\leq\delta\right\}$
-satisfies $\partial_{t}F_{\infty}-\mathcal{F}(\partial_{x}F_{\infty},\partial_{x}^{2}F_{\infty})\geq 0$ in $\mathbb{R}\times(0,\infty)$ and $F_{\infty}$ equals the left-continuous version of $F_{\text{in}}$ at $t=0$.
-Therefore, by the comparison principle, $\overline{F}\leq F_{\infty}$ in $\mathbb{R}\times(0,\infty)$, and this implies $\overline{F}\leq F_{\infty}\leq\liminf\nolimits_{*}F_{\epsilon}$ as desired.
+From the fact that $\rho$ is supported in $[1,2]$ and $F_{N}$ is nondecreasing in $x$, it follows that $F_{N,M}(x,0)\leq F_{N,M+1}(x,0)\leq F_{N}(x,0)$ for every $x\in\mathbb{R}$.
+Due to the monotonicity assumption (i), this implies that $F_{N,M}\leq F_{N,M+1}\leq F_{N}$ in $\mathbb{R}\times[0,\infty)$ for each $N$ and $M$.
+By Theorem 3, $F_{N,M}\to F_{M}$ locally uniformly in $\mathbb{R}\times[0,+\infty)$ as $N\to+\infty$, where $F_{M}$ is the solution of (61) with initial datum $F_{\text{in},M}$ given by $F_{\text{in},M}(x)=2^{M}\int_{0}^{\infty}F_{\text{in}}(x-y)\rho(2^{M}y)\,dy$.
+Further, by construction, the bound $F_{M}\leq F_{M+1}\leq\liminf\nolimits_{*}F_{N}$ holds pointwise in $\mathbb{R}\times[0,\infty)$ for any fixed $M$.
+Standard arguments (see [10, Theorem 6.2]) show that the function $F_{\infty}$ defined by
+$F_{\infty}(x,t)=\lim_{\delta\downarrow 0}\inf\left\{F_{M}(x^{\prime},s)\,\mid\,|x^{\prime}-x|+|t-s|+M^{-1}\leq\delta\right\}$
+satisfies $\partial_{t}F_{\infty}-\mathcal{F}(\partial_{x}F_{\infty},\partial_{x}^{2}F_{\infty})+Q(F_{\infty})\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and $F_{\infty}(\cdot,0)\geq F_{M}(\cdot,0)$ for any $M$.
+In particular, by the comparison principle, $F_{\infty}$ is at least as large as any viscosity subsolution of the initial-value problem (61).
+Therefore, $F_{\infty}\geq\underline{F}$ in $\mathbb{R}\times[0,\infty)$, and this implies $\liminf\nolimits_{*}F_{N}\geq F_{\infty}\geq\underline{F}$ as desired.
 ∎
-B.3. Contractivity
-This subsection establishes that operators such as the family $(T_{\epsilon})_{\epsilon>0}$ are contractive on $CDF(\overline{\mathbb{R}})$ with respect to the supremum norm.
+B.3. Lipschitz Estimate
+This subsection establishes that operators such as the family $\{T^{(N)}\}$ are uniformly Lipschitz on $CDF(\overline{\mathbb{R}})$ with the supremum norm.
 Proposition 17.
@@
   • (ii)
-    Commutation with constants: If $F\in CDF(\overline{\mathbb{R}})$ satisfies $c_{-}\leq F\leq c_{+}$ for some constants $c_{-},c_{+}\in[0,1]$, then $T(F+c)=TF$ for any $c\in[-c_{-},1-c_{+}]$.
+    Discrete-Time Reaction-Diffusion Equation: There is a function $\mathcal{L}$ defined on $CDF(\overline{\mathbb{R}})$ and a function $Q$ defined on $[0,1]$ such that
+    $TF-F=\mathcal{L}F-Q(F)\quad\text{for each}\,\,F\in CDF(\overline{\mathbb{R}}).$
   • (iii)
-    Stationarity at infinity: For any $F\in CDF(\overline{\mathbb{R}})$, $TF(+\infty)=F(+\infty)$ and $TF(-\infty)=F(-\infty)$.
-Then $T$ is contractive, that is,
-$\|TF-TG\|_{\sup}\leq\|F-G\|_{\sup}\quad\text{for any}\,\,F,G\in CDF(\overline{\mathbb{R}}).$
-The proof is based on the following lemma.
+    Commutation with Constants: Given $F,G\in CDF(\overline{\mathbb{R}})$, if the difference FGF-G is a constant function, then $\mathcal{L}F=\mathcal{L}G$.
+If there is a constant $L>0$ such that $|Q(q)-Q(q^{\prime})|\leq L|q-q^{\prime}|$ for each $q,q^{\prime}\in[0,1]$, then
+$\|T^{n}F-T^{n}G\|_{\sup}\leq(1+L)^{n}\|F-G\|_{\sup}\quad\text{for any}\,\,F,G\in CDF(\overline{\mathbb{R}}),\,\,n\in\mathbb{N}.$
+The proposition follows immediately from the following lemma, which was needed in the proof of Theorem 3 above:
 Lemma 5.
-Let $T$ be an operator on $CDF(\overline{\mathbb{R}})$ satisfying assumptions (i)-(iii) of Proposition 12. If $F_{1},F_{2}\in CDF(\overline{\mathbb{R}})$ are chosen so that
-$F_{2}\leq TF_{1}\quad\text{pointwise in}\,\,\mathbb{R}$,
-then, for any $\delta\in[0,1]$ and $c\in[0,1-\delta]$, the functions $G_{1},G_{2}$ defined by
-$G_{i}=\max\{F_{i}-\delta,c\}$
-satisfy $G_{2}\leq TG_{1}$ pointwise in $\mathbb{R}$.
+Suppose that $T$ is a monotone operator on $CDF(\overline{\mathbb{R}})$ satisfying the assumptions of Proposition 17.
+If $\{F_{n}\}$ is any sequence with $F_{n}\leq TF_{n-1}$ for each $n$, then, for any $\delta>0$, the sequence $\{\tilde{F}_{n}\}$ given by $\tilde{F}_{n}=\max\{F_{n}-\delta(1+L)^{n},0\}$ satisfies
+(92) $\tilde{F}_{n}\leq T\tilde{F}_{n-1}\quad\text{for each}\,\,n.$
+In particular, if $\{F_{n}\}$ and $\{G_{n}\}$ are any two sequences in $CDF(\overline{\mathbb{R}})$ such that
+$F_{n}\leq TF_{n-1}\quad\text{and}\quad G_{n}\geq TG_{n-1}\quad\text{ for each}\,\,n\in\mathbb{N},$
+then
+$F_{n}\leq G_{n}+(1+L)^{n}\sup\left\{(F_{0}(x)-G_{0}(x))_{+}\,\mid\,x\in\mathbb{R}\right\}\quad\text{for each}\,\,n.$
 Proof.
-Notice that $G_{1}+\delta=\max\{F_{1},c+\delta\}$, which is a function in $CDF(\overline{\mathbb{R}})$ that is pointwise larger than $F_{1}$. Therefore, by monotonicity,
-$T(G_{1}+\delta)\geq TF_{1}\geq F_{2}.$
-On the other hand, since property (iii) implies $T(G_{1}+\delta)(-\infty)=G_{1}(-\infty)+\delta\geq c+\delta$, one deduces that $T(G_{1}+\delta)\geq c+\delta$ pointwise.
-Therefore, $T(G_{1}+\delta)\geq\max\{F_{2},c+\delta\}=G_{2}+\delta$ follows.
-By property (ii), this implies $TG_{1}\geq G_{2}$.
-∎
-Here is the proof of contractivity:
-Proof of Proposition 12.
-Let $F,G\in\text{CDF}(\overline{\mathbb{R}})$. Note that
-$G-\|F-G\|_{\text{sup}}\leq F\leq G+\|F-G\|_{\text{sup}}$
-and the goal is to show that the same inequality holds if $(F,G)$ is replaced by $(TF,TG)$. By symmetry, it suffices to prove that $TF\leq TG+\|F-G\|_{\text{sup}}$.
-Since the desired inequality is trivial if $\|F-G\|_{\text{sup}}\geq 1$, assume that $\|F-G\|_{\text{sup}}<1$ in what follows.
-Let $G_{1}=\max\{F-\|F-G\|_{\text{sup}},0\}$ and $G_{2}=\max\{TF-\|F-G\|_{\text{sup}},0\}$. By Lemma 4,
-$G_{2}\leq TG_{1}.$
-Yet $G_{1}\leq G$ holds so $G_{2}\leq TG_{1}\leq TG$ by (i). This implies $TF-\|F-G\|_{\text{sup}}\leq TG$.∎
+First, consider the sequence $\{\tilde{F}_{n}\}$. Let $\{c_{n}\}_{n\in\mathbb{N}}$ be the sequence of nonnegative numbers determined by the recursion
+$c_{n}=(1+L)c_{n-1},\quad c_{0}=\delta,$
+so that $\tilde{F}_{n}=\max\{F_{n}-c_{n},0\}$.
+Fix $n\in\mathbb{N}$.
+If $c_{n-1}\geq 1$, then (92) holds trivially as $\tilde{F}_{n}=\tilde{F}_{n-1}=0$ in this case.
+Thus, assume $c_{n-1}\in(0,1)$.
+By monotonicity and the assumption on $\mathcal{L}$,
+$TF_{n-1}-c_{n}$ $\leq T(\max\{F_{n-1},c_{n-1}\})-c_{n}$
+$=\max\{F_{n-1},c_{n-1}\}+\mathcal{L}(\max\{F_{n-1}-c_{n-1},0\})-Q(\max\{F_{n-1}-c_{n-1},0\})$
+$\qquad-Q(\max\{F_{n-1},c_{n-1}\})+Q(\max\{F_{n-1}-c_{n-1},0\})-c_{n}$
+$\leq T(\max\{F_{n-1}-c_{n-1},0\})+c_{n-1}-c_{n}+Lc_{n-1}.$
+In view of the formula for $\{c_{n}\}$ and the property of $\{F_{n}\}$, this implies
+$F_{n}-c_{n}\leq T(\max\{F_{n-1}-c_{n-1},0\})=T\tilde{F}_{n-1}.$
+In view of the fact that $TG\geq 0$ for any $G\in CDF(\overline{\mathbb{R}})$, this last bound implies (92).
+Finally, suppose that, in addition, $G_{n}\geq TG_{n-1}$ for any $n$, and let $\delta=\sup\{(F_{0}(x)-G_{0}(x))_{+}\,\mid\,x\in\mathbb{R}\}$.
+Notice that $\tilde{F}_{0}=\max\{F_{0}-c_{0},0\}\leq G_{0}$ by definition.
+Thus, by the monotonicity of $T$,
+$\max\{F_{n}-c_{n},0\}=\tilde{F}_{n}\leq G_{n}\quad\text{for each}\,\,n.$
+In particular, $F_{n}\leq G_{n}+c_{n}$ for any $n$, as desired. ∎
 Appendix C Effective Coefficients in the Series-Parallel Graph
@@
 These functions have the power series representation $Li_{k}(t)=\sum_{n=1}^{\infty}n^{-k}t^{n}$, hence, in particular, $Li_{k}(1)=\zeta(k)$.
-The computations that follow seem to be related to Euler’s formula for $Li_{2}(\frac{1}{2})$ and Landen’s formula for $Li_{3}(\frac{1}{2})$ (see [19, Section 1.1]).
+The computations that follow seem to be related to Euler’s formula for $Li_{2}(\frac{1}{2})$ and Landen’s formula for $Li_{3}(\frac{1}{2})$ (see [22, Section 1.1]).
 To compute the integrals, use the change-of-variables $x=e^{-s}$, $ds=\frac{1}{x}dx$. In the case of $\sigma_{D}$, this leads to

### v3

@@
 (8) $\Phi(x,y)=\left\{\begin{array}[]{r l}\max\{x,y\}+f_{+}(|x-y|),\quad\text{if}\,\,\Theta=1,\\ \min\{x,y\}-f_{-}(|x-y|),\quad\text{if}\,\,\Theta=0.\end{array}\right$.
-For simplicity, the pair $(f_{+},f_{-})$ will be assumed to be independent of $\Theta$ henceforth.
+Without loss of generality, the pair $(f_{+},f_{-})$ will be assumed to be independent of $\Theta$ henceforth.
 The object of interest in this paper is the RDE determining the law of random variables $(X^{(n)})_{n\in\mathbb{N}}$ through the recursion
-(9) $X^{(n)}\overset{d}{=}\Phi(X^{(n)}_{1},X^{(n)}_{2})$,
-where $X^{(n)}_{1},X^{(n)}_{2}$ are i.i.d.
+(9) $X^{(n)}\overset{d}{=}\Phi(X^{(n-1)}_{1},X^{(n-1)}_{2})$,
+where $X^{(n-1)}_{1},X^{(n-1)}_{2}$ are i.i.d.
 copies of $X^{(n-1)}$, which are independent of $\Phi$.
@@
 By definition, if $f_{\pm}(+\infty)=0$, then $g_{\pm}(s)-s\to 0$ as $s\to+\infty$. The next assumptions ask that $g_{\pm}(s)-s$ vanishes in a quantitative manner: First, under the assumption that
-(15) $\int_{0}^{+\infty}\mathbf{E}[|g_{+}(s)-s|+|g_{-}(s)-s|]\,ds<\infty,$
+(15) $\int_{0}^{+\infty}\mathbf{E}[|g_{+}(s)-s|+|g_{-}(s)-s|]\,ds<+\infty,$
 define the constant $\sigma$ by
-$\sigma=-\int_{0}^{+\infty}\mathbf{E}[|g_{+}(s)-s|]\,du+\int_{0}^{+\infty}\mathbf{E}[|g_{-}(s)-s|]\,ds.$
+$\sigma=-\int_{0}^{+\infty}\mathbf{E}[|g_{+}(s)-s|]\,ds+\int_{0}^{+\infty}\mathbf{E}[|g_{-}(s)-s|]\,ds.$
 If in addition the following stronger assumption holds
-(16) $\int_{0}^{+\infty}(1+s)\mathbf{E}[|g_{+}(s)-s|+|g_{-}(s)-s|]\,ds,$
-define also the constant $a\geq 0$ by
+(16) $\int_{0}^{+\infty}(1+s)\mathbf{E}[|g_{+}(s)-s|+|g_{-}(s)-s|]\,ds<+\infty,$
+let $a\geq 0$ be the constant defined by
 $a=\frac{1}{2}\int_{0}^{+\infty}\mathbf{E}[(3s-g_{+}(s))(s-g_{+}(s))]\,ds+\frac{1}{2}\int_{0}^{+\infty}\mathbf{E}[(3s-g_{-}(s))(s-g_{-}(s))]\,ds.$
@@
 Theorem 2.
-Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, (16) holds, $\sigma=0$, and $a>0$. For each $N\in\mathbb{N}$, let $\{X^{(N,n)}\}$ be a solution of the RDE
+Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, (16) holds, $\sigma=0$, and $a>0$.
+Fix a $\theta\in\mathbb{R}$ and let $p^{(N)}=1/2+\theta N^{-1}$.
+For each $N\in\mathbb{N}$, let $\{X^{(N,n)}\}$ be a solution of the RDE
 $X^{(N,n)}\overset{d}{=}\Phi^{(N)}(X^{(N,n-1)}_{1},X^{(N,n-1)}_{2})$,
-where $(X^{(N,n-1)}_{1},X^{(N,n-1)}_{2})$ are i.i.d.
-copies of $X^{(N,n-1)}$ and $\Phi^{(N)}$ is given by (8) with bias parameter $p=p^{(N)}$ satisfying $p^{(N)}=1/2+\theta N^{-1}$ for some $\theta\in\mathbb{R}$.
+where $(X^{(N,n-1)}_{1},X^{(N,n-1)}_{2})$ are i.i.d. copies of $X^{(N,n-1)}$ and $\Phi^{(N)}$ is given by (8) with bias parameter $p=p^{(N)}$.
 Assume that there is a function $F_{\text{in}}$ such that, at the initial time $n=0$, the following limit holds:
@@
 Consistency, on the other hand, refers to the behavior of the equation under rescaling.
-Specifically, Proposition 10 below shows that if $\Phi$ is a smooth CDF, rescaled according to $\Phi_{\delta}(x)=\Phi(\delta x)$, then
-(23) $(\mathscr{L}\Phi_{\delta})(\delta^{-1}x)=\delta^{2}\sigma|\partial_{x}\Phi(x)|^{2}+\delta^{3}a|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)+\cdots\quad\text{as}\,\,\delta\downarrow 0.$
+Specifically, Proposition 10 below shows that if $G$ is a smooth CDF, rescaled according to $G_{\delta}(x)=G(\delta x)$, then
+(23) $(\mathscr{L}G_{\delta})(\delta^{-1}x)=\delta^{2}\sigma|\partial_{x}G(x)|^{2}+\delta^{3}a|\partial_{x}G(x)|\partial_{x}^{2}G(x)+\cdots\quad\text{as}\,\,\delta\downarrow 0.$
 Rescaling the sequence of CDF’s $\{F_{n}\}$ according to $F_{N}(x,t)=F_{[Nt]}(N^{1/\alpha}x)$, naïvely inserting the above expansion into the equation (11), and setting $p=1/2+\theta N^{-1}$, one obtains the formal asymptotic expansion
 $\partial_{t}F_{N}=N^{1-\frac{2}{\alpha}}\sigma|\partial_{x}F_{N}|^{2}+N^{1-\frac{3}{\alpha}}a|\partial_{x}F_{N}|^{2}-2\theta F_{N}(1-F_{N})+\cdots\quad\text{as}\,\,N\to+\infty,$
-which motivates the choices $\alpha=2$ and $\alpha=3$ used above and the appearance of the PDE’s (19) and (22). The approach of [7] provides a framework for making such a formal argument rigorous.
-1.5. Variants of the Fisher-KPP Equation when $p\neq\frac{1}{2}$
-The results of this work suggest interesting new questions away from the critical point $p=\frac{1}{2}$.
+which motivates the choices $\alpha=2$ and $\alpha=3$ used above and the appearance of the PDE’s (19) and (22).
+The approach of [7] provides a robust framework for making such a formal argument rigorous.
+1.5. Variants of the Fisher-KPP Equation when $p\neq 1/2$
+The results of this work suggest interesting new questions away from the critical point $p=1/2$.
 In view of the asymptotic expansion (23) and the PDE’s obtained in Theorems 1 and 2, the equation (11) shares some similarities with the reaction-diffusion equation
@@
 The work of Audrito and Vázquez [5] establishes existence, uniqueness, and positivity properties of traveling waves and also analyzes the asymptotic behavior of solutions, proving that sufficiently rapidly decaying perturbations of the unstable state (say, 0 if $\theta<0$) converge to the stable state ($1$ if $\theta<0$) inside a ballistically growing region, with the rate of growth determined by the minimal wave speed.
-It would be interesting to extend such results to the present context.
+It would be interesting to adapt the techniques of [5] to the study of (11) away from the critical point.
 1.6. Related Literature
@@
 The proof in [6] is based on the explicit construction of sub- and supersolutions of (11), which, by monotonicity, can be used to control solutions, or what is referred to as a barrier argument in the PDE literature.
-The same strategy was employed by Chen, Duquesne, and Shi in [13], who derived the same $\text{Beta}(2,2)$ limit theorem as in Corollary 1 above, albeit under slightly more restrictive assumptions on $(f_{+},f_{-})$.
+The same strategy was employed by Chen, Duquesne, and Shi in [13], who independently derived the $\text{Beta}(2,2)$ limit theorem of Corollary 1 above, albeit under slightly more restrictive assumptions on $(f_{+},f_{-})$.
 Addario-Berry and coauthors recently introduced two classes of RDE’s, called hipster random walks and cooperative motion, in [1, 2, 3].
@@
 2. Preliminaries
-2.1. Extended-Real-Valued Random Variables
+2.1. Extended Real-Valued Random Variables
 Note that the definition (9) still makes sense if the input random variables are permitted to take the values $\infty$ and -$\infty$.
-It will be convenient to extend the definition to include such extended-real-valued random variables.
-To that end, for an extended-real-valued random variable $X$, define the CDF $F:\mathbb{R}\to[0,1]$ via the rule
+It will be convenient to extend the definition to include such extended real-valued random variables.
+To that end, for an extended real-valued random variable $X$, define the CDF $F:\mathbb{R}\to[0,1]$ via the rule
 $\mathbb{P}\{X=-\infty\}=\lim_{y\to-\infty}F(y),\quad\mathbb{P}\{X\leq x\}=F(x),\quad\mathbb{P}\{X=+\infty\}=1-\lim_{y\to+\infty}F(y).$
 Notice that the space $CDF(\overline{\mathbb{R}})$ of all such CDF’s is simply
-$CDF(\overline{\mathbb{R}})=\{F:\mathbb{R}\to[0,1]\,\,\text{increasing, right-continuous}\}.$
-Using standard extended-real-number arithmetic, from now on, the formula (10) defining $T$ will be understood to hold in the domain $CDF(\overline{\mathbb{R}})$.
+$CDF(\overline{\mathbb{R}})=\{F:\mathbb{R}\to[0,1]\,\,\text{nondecreasing, right-continuous}\}.$
+Using standard extended real number arithmetic, from now on, the formula (10) defining $T$ will be understood to hold in the domain $CDF(\overline{\mathbb{R}})$.
 2.2. Continuity under Vague Convergence.
-It is useful to note that $T$ is continuous with respect to vague convergence. Recall that $\{F_{n}\}$ converges vaguely to $F$ if $F_{n}(x)\to F(x)$ at each point of continuity $x$ of $F$.
+It is useful to note that $T$ is continuous with respect to vague convergence.
+Recall that $\{F_{n}\}$ converges vaguely to $F$ if $F_{n}(x)\to F(x)$ at each point of continuity $x\in\mathbb{R}$ of $F$.
 Proposition 1.
@@
 Proof.
-It suffices to recall that this mode of convergence for $(F_{n})_{n\in\mathbb{N}}$ is equivalent to the existence of a probability space $\mathbb{P}$ supporting a sequence $\{(X_{1,n},X_{2,n})\}$ and a vector $(X_{1},X_{2})$ such that, for each $n$, $X_{1,n}$ and $X_{2,n}$ are i.i.d.
+It suffices to recall that this mode of convergence for $\{F_{n}\}$ is equivalent to the existence of a probability space $\mathbb{P}$ supporting a sequence $\{(X_{1,n},X_{2,n})\}$ and a vector $(X_{1},X_{2})$ such that, for each $n$, $X_{1,n}$ and $X_{2,n}$ are i.i.d.
 with CDF $F_{n}$; $(X_{1},X_{2})$ are i.i.d.
@@
 with the property that $S(F,F)=TF$ for any CDF $F$.
-$S$ is defined analogously to $T$: Given two CDF’s $F,G\in CDF(\overline{\mathbb{R}})$, let $X_{F}$ and $X_{G}$ be independent extended-real-valued random variables distributed according to $F$ and $G$, respectively, and sample the data $(\Theta,f_{+},f_{-})$ independently of $(X_{F},X_{G})$.
+$S$ is defined analogously to $T$: Given two CDF’s $F,G\in CDF(\overline{\mathbb{R}})$, let $X_{F}$ and $X_{G}$ be independent extended real-valued random variables distributed according to $F$ and $G$, respectively, and sample the data $(\Theta,f_{+},f_{-})$ independently of $(X_{F},X_{G})$.
 Let $S(F,G)$ denote the CDF of the random variable $Y$ defined as follows:
@@
 $Y^{-}=X^{-}_{1}+f_{+}(X^{-}_{1}-X_{2})\leq X_{1}^{+}+f_{+}(X^{+}_{1}-X_{2})=Y^{+}.$
-On the other hand, if $X^{-}_{1}\leq X_{2}\leq X^{+}_{1}$, then invoking first (13) and then the fact that $f_{+}$ is nonincreasing (see (12)), one finds
+On the other hand, if $X^{-}_{1}\leq X_{2}\leq X^{+}_{1}$, then invoking first the fact that $f_{+}$ is nonincreasing (see (12)) and then (13), one finds
 $Y^{-}=X_{2}+f_{+}(X_{2}-X^{-}_{1})$ $\leq X_{2}+f_{+}(0)$
@@
 $Y^{-}=X_{2}+f_{+}(X_{2}-X^{-}_{1})\leq X_{2}+f_{+}(X_{2}-X^{+}_{1})=Y^{+}.$
-Using coupling arguments very similar to the one employed above, one readily deduces that the map $T=T(p,\mathbf{P})$ is nondecreasing with respect to $p$ and $f_{+}$ and nonincreasing with respect to $f_{-}$.
+Using coupling arguments very similar to the one employed above, one readily deduces that the map $T=T(p,\mathbf{P})$ is nonincreasing with respect to $p$ and $f_{+}$ and nondecreasing with respect to $f_{-}$.
 The result is stated next for the sake of precision.
 Proposition 3.
-(i) For any $p_{1}\leq p_{2}$, $T(p_{1},\mathbf{P})F\leq T(p_{2},\mathbf{P})F$ for all $F\in CDF(\overline{\mathbb{R}})$.
-(ii) If $\mathbf{P}$ and $\mathbf{\tilde{P}}$ are measures on $BC([0,\infty)\times BC([0,\infty)$ for which it is possible to find a coupling of the corresponding random variables $(f_{+},f_{-})$ and (f~+,f~)($\tilde$$_{+},\tilde{f}_{-}$) such that $f_{+}\leq\tilde{f}_{+}$ and $f_{-}\geq\tilde{f}_{-}$, then $T(p,\mathbf{P})F\leq T(p,\mathbf{\tilde{P}})F$ for each $F\in CDF(\overline{\mathbb{R}})$.
-2.5. Absolute Continuity
+(i) For any $p_{1}\leq p_{2}$, $T(p_{1},\mathbf{P})F\geq T(p_{2},\mathbf{P})F$ for all $F\in CDF(\overline{\mathbb{R}})$.
+(ii) If $\mathbf{P}$ and $\mathbf{\tilde{P}}$ are measures on $BC([0,\infty)\times BC([0,\infty)$ for which it is possible to find a coupling of the corresponding random variables $(f_{+},f_{-})$ and (f~+,f~)($\tilde$$_{+},\tilde{f}_{-}$) such that $f_{+}\leq\tilde{f}_{+}$ and $f_{-}\geq\tilde{f}_{-}$, then $T(p,\mathbf{P})F\geq T(p,\mathbf{\tilde{P}})F$ for each $F\in CDF(\overline{\mathbb{R}})$.
+2.5. Continuity
 For the sake of developing intuition, it may be useful to note that $T$ maps the space of continuous CDF’s to itself.
@@
 Given $F\in CDF(\overline{\mathbb{R}})$, let $X_{1},X_{2}$ be i.i.d.
-extended-real-valued random variables distributed according to $F$, and sample the data $(\Theta,f_{+},f_{-})$ independently of $(X_{1},X_{2})$.
-Define the extended-real-valued random variable $Z$ by
+extended real-valued random variables distributed according to $F$, and sample the data $(\Theta,f_{+},f_{-})$ independently of $(X_{1},X_{2})$.
+Define the extended real-valued random variable $Z$ by
 (30) $Z$ $=\left\{\begin{array}[]{r l}\max\{X_{1},X_{2}\},&\text{if}\,\,\Theta=1,\\ \min\{X_{1},X_{2}\},&\text{if}\,\,\Theta=0.\end{array}\right$.
@@
 $TF-F=\mathscr{L}F+(1-2p)F(1-F).$
-Using the explicit formulas (31) and (30), one obtains very concrete expressions for the action of $\mathscr{L}$ on smooth CDF’s, as stated in the next result.
+Using the explicit expressions (30) and (31), one obtains very concrete expressions for the action of $\mathscr{L}$ on smooth CDF’s, as stated in the next result.
 Proposition 6.
-Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$. If $F\in CDF(\overline{\mathbb{R}})$ is differentiable, then
+Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$. If $F\in CDF(\overline{\mathbb{R}})$ is absolutely continuous, then
 (33) $\mathscr{L}F(x)$ $=2p\int_{-\infty}^{x}\mathbf{E}[F(x_{1}+g_{+}(x-x_{1}))-F(x)]F^{\prime}(x_{1})\,dx_{1}$
@@
 In [21, Lemma 1.1], it is observed that the fixed points of the map $q\mapsto pq^{2}+(1-p)(1-(1-q)^{2})$ are fundamental in the study of the asymptotics of the series-parallel graph.
-Yet, as in the derivation above, a fixed point of this map is nothing but a zero of the polynomial (12p)q(1q)(1-2p)q(1-q), the “reaction term” appearing in the equation.
+Yet, as in the derivation above, a fixed point of this map is nothing but a zero of the polynomial (12p)q(1q)(1-2p)q(1-q), the “reaction term” appearing in the evolution equation (11).
 3.2. Proof of Proposition 6
-Suppose that $F\in CDF(\overline{\mathbb{R}})$ is continuously differentiable. To obtain the formula (33) for $\mathscr{L}F$ in this case, it suffices to establish that
+Suppose that $F\in CDF(\overline{\mathbb{R}})$ is absolutely continuous and fix $x\in\mathbb{R}$. To obtain the formula (33) for $\mathscr{L}F(x)$ in this case, it suffices to establish that
 $\mathbb{P}\{Z\leq x<Y\mid\Theta=1\}=-2\int_{-\infty}^{x}\mathbf{E}[F(x_{1}+g_{+}(x-x_{1}))-F(x)]F^{\prime}(x_{1})\,dx_{1},$
@@
 In particular, in the analysis that follows, only the case where $Z=X_{2}$ needs to be treated.
-Further, there is no loss of generality assuming that $x\in\mathbb{R}$ as the integral vanishes for $x\in\{-\infty,+\infty\}$ because $F(x_{1}+g_{+}(+\infty-x_{1}))=F(+\infty)$.
 3.2.1. Case: $\Theta=1$
@@
 Thus,
-$\{Z\leq x<Y,\,\,X_{1}<X_{2},\,\,\Theta=1\}$ $=\{-\infty<X_{1}<X_{2}\leq x,\,\,g_{+}(x-X_{1})<X_{2}-X_{1}\}$
-$=\{-\infty<X_{1}<x,\,\,X_{1}+g_{+}(x-X_{1})<X_{2}\leq x\}.$
-Therefore, since $f_{+}$ is independent of $(X_{1},X_{2})$ and $\mathbb{P}\{X_{1}\leq X_{2}\}=\mathbb{P}\{X_{1}<X_{2}\}$ by the continuity of $F$,
+$\{Z\leq x<Y,\,\,X_{1}<X_{2},\,\,\Theta=1\}$
+$\qquad=\{-\infty<X_{1}<X_{2}\leq x,\,\,g_{+}(x-X_{1})<X_{2}-X_{1},\,\,\Theta=1\}$
+$\qquad=\{-\infty<X_{1}<x,\,\,X_{1}+g_{+}(x-X_{1})<X_{2}\leq x,\,\,\Theta=1\}.$
+Therefore, since $(\Theta,f_{+},X_{1},X_{2})$ are mutually independent and $\mathbb{P}\{X_{1}\leq X_{2}\}=\mathbb{P}\{X_{1}<X_{2}\}$ by the continuity of $F$,
 $\mathbb{P}\{Z\leq x<Y,X_{1}\leq X_{2}\mid\Theta=1\}$ $=-\int_{-\infty}^{x}\mathbf{E}[F(x_{1}+g_{+}(x-x_{1}))-F(x)]F^{\prime}(x_{1})\,dx_{1}.$
@@
 Here is the proof of (i): Fix $F\in CDF(\overline{\mathbb{R}})$. Observe that, by continuity of measure,
-$\lim_{x\to+\infty}\mathbb{P}\{Z\leq x<Y\,\mid\,\Theta=1\}$ $=\lim_{x\to+\infty}\left(\mathbb{P}\{Z\leq x\,\mid\,\Theta=1\}-\mathbb{P}\{Y\leq x\,\mid\,\Theta=1\}\right)$
-$=\mathbb{P}\{Z=+\infty\,\mid\,\Theta=1\}-\mathbb{P}\{Y=+\infty\,\mid\,\Theta=1\}=0$
-since $\{Y=+\infty\}=\{Z=+\infty\}$ by the boundedness of $f_{+}$ and $f_{-}$. By symmetry,
-$\lim_{x\to-\infty}\mathbb{P}\{Y\leq x<Z\,\mid\,\Theta=0\}=0.$
+$\lim_{x\to+\infty}\mathbb{P}\{Z\leq x<Y\,\mid\,\Theta=1\}$ $=\lim_{x\to+\infty}\left(\mathbb{P}\{Y>x\,\mid\,\Theta=1\}-\mathbb{P}\{Z>x\,\mid\,\Theta=1\}\right)$
+$=\mathbb{P}\{Y=+\infty\,\mid\,\Theta=1\}-\mathbb{P}\{Z=+\infty\,\mid\,\Theta=1\}=0$
+since $\{Y=+\infty\}=\{Z=+\infty\}$ by the boundedness of $f_{+}$ and $f_{-}$. Similarly,
+$\lim_{x\to-\infty}\mathbb{P}\{Z\leq x<Y\,\mid\,\Theta=1\}=\mathbb{P}\{Z=-\infty\,\mid\,\Theta=1\}-\mathbb{P}\{Y=-\infty\,\mid\,\Theta=1\}=0$
+as $\{Y=-\infty\}=\{Z=-\infty\}$. By symmetry,
+$\lim_{x\to+\infty}\mathbb{P}\{Y\leq x<Z\,\mid\,\Theta=0\}=\lim_{x\to-\infty}\mathbb{P}\{Y\leq x<Z\,\mid\,\Theta=0\}=0.$
 Therefore, by (32), $(\mathscr{L}F)(+\infty)=(\mathscr{L}F)(-\infty)=0$ as claimed.
@@
 First, if $F$ is smooth, then equation (33) applied to $F$ and $F+c$ implies $\mathscr{L}(F+c)=\mathscr{L}F$.
-In general, there is a sequence $\{F_{n}\}$ of smooth CDF’s that converges vaguely to $F$.
+In general, there is a sequence $\{F_{n}\}$ of smooth CDF’s that converges vaguely to $F$ and such that $F_{n}(\pm\infty)=F(\pm\infty)$ for any $n$.
 By Proposition 1, $\{TF_{n}\}$ and $\{T(F_{n}+c)\}$ converge vaguely to respective limits $TF$ and $T(F+c)$.
@@
 Therefore, being nondecreasing, $TF$ must be a constant function, equal to $P(q)$.
-Identifying $F$ with $q$, this proves that $T$ maps $q$ to the constant $Tq=P(q)$, as claimed.
+Identifying $F$ with $q$, this proves that $T$ maps $q$ to the constant $P(q)$, as claimed.
 ∎
@@
 In this case, the right-continuous, right-inverse $g$ of the function $u\mapsto u+f(u)$ can be computed to be
-$g(s)=\left\{\begin{array}[]{r l}\log(e^{s}-1),&\text{if}\,\,s\geq\log 2,\\ 0,&\text{otherwise.}\end{array}\right$.
+$g(s)=\log(e^{s}-1)\quad\text{for each}\,\,s\geq\log 2.$
 It is immediate to check that $f\in\mathcal{S}$ and sg(s)=f(g(s))s-g(s)=f(g(s)) for $s\geq\log 2$ so the stronger assumption (16) of Theorems 1 and 2 is satisfied.
-In this case, the constant $a$ is given by $a=2\zeta(3)$. See Appendix C for the proof.
+By symmery, $\sigma=0$ in this setting and $a$ can be explicitly computed to be $a=2\zeta(3)$. See Appendix C for the proof.
 The approach of this paper readily implies the following characterization of the asymptotics of the resistance $R^{(n)}$ of the critical series-parallel graph, in which resistors have i.i.d.
@@
 Then $\mathbb{P}\{0<R^{(n)}<+\infty\}=q$ for each $n$ and the law of $R^{(n)}$ conditional on this event has the following limit:
-$\lim_{N\to\infty}\mathbb{P}\{(36a)^{-\frac{1}{3}}N^{-\frac{1}{3}}\log R^{(N)}\leq x\,\mid\,0<R^{(N)}<+\infty\}=F_{\text{Beta}(2,2)}(q^{-\frac{1}{3}}x+\frac{1}{2}),$
-where $F_{\text{Beta}(2,2)}$ is the CDF of the $\text{Beta}(2,2)$ distribution and $a=2\zeta(3)$.
+$\lim_{N\to\infty}\mathbb{P}\{(36a)^{-\frac{1}{3}}N^{-\frac{1}{3}}\log R^{(N)}\leq x\,\mid\,0<R^{(N)}<+\infty\}=F_{\text{Beta}(2,2)}(q^{-\frac{1}{3}}x+\frac{1}{2}).$
+Here $F_{\text{Beta}(2,2)}$ is the CDF of the $\text{Beta}(2,2)$ distribution and $a=2\zeta(3)$.
 Recall that the $\text{Beta}(2,2)$ distribution is determined by the PDF $6y(1-y)\mathbf{1}_{[0,1]}(y)$.
@@
 Let $X^{(n)}=\log R^{(n)}$.
-As in the discussion in the introduction, $\{X^{(n)}\}$ is a solution of the RDE (9) with $f_{+}(u)=f_{-}(u)=\log(1+e^{-u})$.
+As in the discussion above, $\{X^{(n)}\}$ is a solution of the RDE (9) with $f_{+}(u)=f_{-}(u)=\log(1+e^{-u})$.
 Let $q_{-}=\mathbb{P}\{X^{(0)}=-\infty\}$ and $q_{+}=\mathbb{P}\{X^{(0)}=+\infty\}$ so that $1-q=q_{+}+q_{-}$.
@@
 Let $F_{N}:\mathbb{R}\times[0,+\infty)\to[0,1]$ be the rescaled CDF of $\{X^{(n)}\}$ defined as in Theorem 2.
-By that theorem, $F_{N}\to F$ locally uniformly in $\mathbb{R}\times(0,+\infty)$ as $N\to+\infty$, where $F$ is the bounded discontinuous viscosity solution of the PDE
+By that theorem (which applies even though $\{X^{(n)}\}$ are extended real-valued, see Remark 3 below), $F_{N}\to F$ locally uniformly in $\mathbb{R}\times(0,+\infty)$ as $N\to+\infty$, where $F$ is the bounded discontinuous viscosity solution of the PDE
 $\partial_{t}F-a|\partial_{x}F|\partial_{x}^{2}F=0\quad\text{in}\,\,\mathbb{R}\times(0,+\infty),\quad F(x,0)=q_{-}+q\mathbf{1}_{[0,+\infty)}(x).$
 $F$ can be readily related to the solution of the same PDE, but with initial datum $\mathbf{1}_{[0,+\infty)}$ and a different coefficient $a$: Notice that if $G$ is defined by $G(x,t)=q^{-1}(F(x,t)-q_{-})$, then $G(x,0)=\mathbf{1}_{[0,+\infty)}(x)$ and
-$\partial_{t}G-q^{2}a|\partial_{x}G|\partial_{x}^{2}G=0\quad\text{in}\,\,\mathbb{R}\times(0,+\infty),\quad G(x,0)=\mathbf{1}_{[0,+\infty)}(x).$
+$\partial_{t}G-qa|\partial_{x}G|\partial_{x}^{2}G=0\quad\text{in}\,\,\mathbb{R}\times(0,+\infty),\quad G(x,0)=\mathbf{1}_{[0,+\infty)}(x).$
 As in the proof of Corollary 1 (see Section 5.3) below, $G$ can be explicitly computed to be
-$G(x,t)=F_{\text{Beta}(2,2)}(U(x,t)),\quad\text{where}\quad U(x,t)=\frac{x}{(36qa)^{\frac{1}{3}}}+\frac{1}{2}.$
+$G(x,t)=F_{\text{Beta}(2,2)}(U(x,t)),\quad\text{where}\quad U(x,t)=\frac{x}{(36qat)^{\frac{1}{3}}}+\frac{1}{2}.$
 At the same time, since $\mathbb{P}\{0<R^{(N)}<+\infty\}=q$ and $\mathbb{P}\{R^{(N)}>0\}=q_{-}$ for any $N$ by Proposition 7,
-$\lim_{N\to+\infty}\mathbb{P}\{N^{-\frac{1}{3}}\log R^{(N)}\leq x\,\mid\,0<R^{(N)}<+\infty\}=\lim_{N\to+\infty}\frac{1}{q}(F_{N}(x,t)-q_{-})=G(x,t).$
+$\lim_{N\to+\infty}\mathbb{P}\{N^{-\frac{1}{3}}\log R^{(N)}\leq x\,\mid\,0<R^{(N)}<+\infty\}=\lim_{N\to+\infty}\frac{1}{q}(F_{N}(x,1)-q_{-})=G(x,1).$
 4.2. Distance on the series-parallel graph
@@
 Analogous to the case of the resistance, the logarithm $X^{(n)}=\log D^{(n)}$ gives a solution of the RDE (9) with $\Phi$ given as in (8) provided $f_{+}(u)=\log(1+e^{-u})$ and $f_{-}=0$.
-A complete analysis of the asymptotics of this RDE in the subcritical case ($p<\frac{1}{2}$) is given in [21].
-There is a random variable $K$ such that if $\lambda$ denotes the essential infimum of $D^{(0)}$, then $D^{(n)}\to\lambda K$ as $n\to\infty$.
+A complete analysis of the asymptotics of this RDE in the subcritical case ($p<1/2$) was given in [21]: There is a random variable $K$ such that if $\lambda$ denotes the essential infimum of $D^{(0)}$, then $D^{(n)}\to\lambda K$ as $n\to\infty$.
 This implies that $K$ is the unique (up to multiplication by a constant) fixed point of the RDE in this case.
-In the critical case ($p=\frac{1}{2}$), the main result of [6] obtains a $\text{Beta}(2,1)$ limit for $N^{-1/2}\log D^{(N)}$ provided $D^{(0)}=1$ almost surely.
-The next corollary of Theorem 1 describes the asymptotics for an arbitrary initial distribution.
-Notice the contrast with the subcritical case: In the subcritical case, if $D^{(0)}$ can be arbitrarily small with positive probability, then $D^{(n)}\to 0$ as $n\to\infty$, whereas here, conditionally on the event $\{D^{(n)}>0\}$, mass at zero only slows the rate of growth.
+At the critical point $p=1/2$, the main result of [6] obtains a $\text{Beta}(2,1)$ limit for $N^{-1/2}\log D^{(N)}$ provided $D^{(0)}=1$ almost surely.
+The next corollary of Theorem 1 describes the asymptotics for an arbitrary initial distribution, allowing for the possibility that the i.i.d.
+edge weights can be zero or $\infty$.
+Notice the contrast with the subcritical case: In the subcritical case, if $D^{(0)}$ can be arbitrarily small with positive probability, then $D^{(n)}\to 0$ as $n\to+\infty$, whereas here, conditionally on the event $\{0<D^{(n)}<+\infty\}$, mass at zero only slows the rate of growth.
 Corollary 3.
-Assume that $p=\frac{1}{2}$ and $\mathbb{P}\{0\leq D^{(0)}<+\infty\}=1$ and let $q=\mathbb{P}\{D^{(0)}>0\}$.
-Then $\mathbb{P}\{D^{(n)}>0\}=q$ for any $n$ and the law of $D^{(n)}$ conditional on this event has the following limit:
-$\lim_{N\to\infty}\mathbb{P}\{N^{-\frac{1}{2}}\log D^{(N)}\leq 2x\sqrt{|\sigma|}\,\mid\,D^{(N)}>0\}=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ \frac{x^{2}}{q},&\text{if}\,\,0\leq x\leq\sqrt{q},\\ 1,&\text{otherwise.}\end{array}\right$.
+Assume that $p=\frac{1}{2}$ and $\mathbb{P}\{0\leq D^{(0)}\leq+\infty\}=1$ and let $q=\mathbb{P}\{0<D^{(0)}<+\infty\}$.
+Then $\mathbb{P}\{0<D^{(n)}<+\infty\}=q$ for any $n$ and the law of $D^{(n)}$ conditional on this event has the following limit:
+$\lim_{N\to\infty}\mathbb{P}\{N^{-\frac{1}{2}}\log D^{(N)}\leq 2x\sqrt{|\sigma|}\,\mid\,0<D^{(N)}<+\infty\}=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ \frac{x^{2}}{q},&\text{if}\,\,0\leq x\leq\sqrt{q},\\ 1,&\text{otherwise.}\end{array}\right$.
 Here the constant $\sigma=-\frac{1}{2}\zeta(2)=-\frac{\pi^{2}}{12}$.
@@
 The claimed value of $\sigma$ is computed in Appendix C and is consistent with [6].
-The limiting behavior of $D^{(N)}$ can be derived following the same strategy as in Corollary 2 above, once again using the fact that $\mathbb{P}\{D^{(n)}>0\}=q$ for any $n$.
+The limiting behavior of $D^{(N)}$ can be derived following the same strategy as in Corollary 2 above, once again using the fact that $\mathbb{P}\{0<D^{(n)}<+\infty\}=q$ for any $n$ by Proposition 7.
 ∎
@@
 (37) $Y\overset{d}{=}\left\{\begin{array}[]{r l}\max\{X_{1},X_{2}\}+\mathbf{1}_{\{0\}}(|X_{2}-X_{1}|),&\text{if}\,\,\Theta=1,\\ \min\{X_{1},X_{2}\}-\mathbf{1}_{\{0\}}(|X_{2}-X_{1}|),&\text{if}\,\,\Theta=0.\end{array}\right$.
-The RDE associated with this equation provides a very natural distribution-dependent random walk on $\mathbb{Z}$, and it has the nice feature that the equations describing the CDF are nothing other than finite-difference discretizations of the quasilinear Fisher-KPP equations mentioned in the introduction.
+The RDE associated with this equation provides a very natural distribution-dependent random walk on $\mathbb{Z}$, and it has the nice feature that the equations describing the CDF are nothing other than finite-difference discretizations of the quasilinear Fisher-KPP equations discussed in the introduction above.
 For the rest of this subsection, assume that the joint law $\mathbf{P}$ of $(f_{+},f_{-})$ is concentrated on the set $\{0,f_{\mathbb{Z}}\}\times\{0,f_{\mathbb{Z}}\}$. Notice that
 $f_{\mathbb{Z}}\in\mathcal{S}\quad\text{and}\quad g_{\mathbb{Z}}(s)=s\quad\text{for each}\,\,s\geq 1,$
-where $g_{\mathbb{Z}}$ denotes the left-inverse of the function $u\mapsto u+f_{\mathbb{Z}}(u)$.
+where, as in Section 1.4.1, $g_{\mathbb{Z}}$ denotes the right-continuous, right-inverse of the function $u\mapsto u+f_{\mathbb{Z}}(u)$.
 Thus, such a law $\mathbf{P}$ satisfies the assumptions of Theorems 1 and 2.
@@
 For $\mathbb{Z}$-valued solutions of the RDE, there is no loss restricting the CDF to the domain $\mathbb{Z}$. By invoking the formula (32), one obtains, in this case,
-(38) $F_{n+1}-F_{n}$ $=-pq_{+}(\nabla_{\mathbb{Z}}^{-}F_{n})^{2}+(1-p)q_{-}(\nabla_{\mathbb{Z}}^{+}F_{n})^{2}+(1-2p)F_{n}(1-F_{n}),$
-where $\nabla_{\mathbb{Z}}^{+}$, $\nabla_{\mathbb{Z}}^{-}$ are the forward- and backward-difference operators
+(38) $F_{n}-F_{n-1}$ $=-pq_{+}(\nabla_{\mathbb{Z}}^{-}F_{n-1})^{2}+(1-p)q_{-}(\nabla_{\mathbb{Z}}^{+}F_{n-1})^{2}+(1-2p)F_{n-1}(1-F_{n-1}),$
+where $\nabla_{\mathbb{Z}}^{+}$ and $\nabla_{\mathbb{Z}}^{-}$ are the forward- and backward-difference operators
 $\nabla_{\mathbb{Z}}^{+}F(x)=F(x+1)-F(x),\quad\nabla_{\mathbb{Z}}^{-}F(x)=F(x)-F(x-1).$
 Using the identity $\nabla_{\mathbb{Z}}^{+}-\nabla_{\mathbb{Z}}^{-}=\Delta_{\mathbb{Z}}$ for $\Delta_{\mathbb{Z}}F(x)=(F(x+1)+F(x-1))-F(x)$ and simplifying the difference of the two squares, this becomes
-(39) $F_{n+1}-F_{n}$ $=a\nabla^{\text{sym}}_{\mathbb{Z}}F_{n}\Delta_{\mathbb{Z}}F_{n}+\sigma(\nabla_{\mathbb{Z}}^{+}F_{n})^{2}+(1-2p)F_{n}(1-F_{n}).$
+(39) $F_{n}-F_{n-1}$ $=a\nabla^{\text{sym}}_{\mathbb{Z}}F_{n-1}\Delta_{\mathbb{Z}}F_{n-1}+\sigma(\nabla_{\mathbb{Z}}^{+}F_{n-1})^{2}+(1-2p)F_{n-1}(1-F_{n-1}).$
 where $\nabla^{\text{sym}}_{\mathbb{Z}}=2^{-1}(\nabla_{\mathbb{Z}}^{+}+\nabla_{\mathbb{Z}}^{-})$, $\sigma=(1-p)q_{-}-pq_{+}$, and $a=2pq_{+}$.
-Equation (39) looks like a discrete analogue of the quasilinear Fisher-KPP equation (24) derived above.
-Notice that here $a$ or $\sigma$ may vanish, irrespective of whether or not $p=\frac{1}{2}$.
+Equation (39) looks like a discrete analogue of the quasilinear Fisher-KPP equation (24) discussed in Section 1.5.
+Notice that here $a$ or $\sigma$ may vanish, irrespective of whether or not $p=1/2$.
 Since replacing $\mathbb{Z}$ by $\mathbb{R}$ and finite-differences by derivatives does not seem likely to fundamentally change the large-scale asymptotic behavior, equation (39) motivates the general study of (24) for arbitrary $\sigma\in\mathbb{R}$ and $a\geq 0$.
-It is worth emphasizing that there is no contradiction between the formulas (39) and (33) since the former involves the CDF of a $\mathbb{Z}$-valued random variable while the latter describes smooth CDF’s.
+It is worth emphasizing that there is no contradiction between the formulae (39) and (33) since the former involves the CDF of a $\mathbb{Z}$-valued random variable while the latter describes smooth CDF’s.
 4.4. Hipster Random Walks and Cooperative Motion
-Inspection of (39) also reveals a connection with the symmetric hipster random walk and two-player cooperative motion, examples considered in in [1] and [3], respectively.
+Inspection of (39) also reveals a connection with the symmetric hipster random walk and two-player cooperative motion, examples considered in [1] and [3], respectively.
+For the rest of this section, let $p=1/2$.
 To understand the connection, consider again the random variable $Z$ defined by $Z=\Theta\max\{X_{1},X_{2}\}+(1-\Theta)\min\{X_{1},X_{2}\}$.
-Recall that $Z$ and $X_{1}$ have the same distribution if $p=\frac{1}{2}$.
-At the same time, since the two coincide when $X_{1}=X_{2}$, a stronger statement is true: $Z$ and $X_{1}$ have the same conditional laws on the event $\{X_{1}\neq X_{2}\}$.
-The same reasoning applies if $X_{1}$ is replaced by the more symmetric $\Theta X_{1}+(1-\Theta)X_{2}$.
-From this, it follows that the law of the output $Y$ is unchanged if the equation is instead changed to
+Recall that $Z$ and $X_{1}$ have the same distribution since $p=1/2$.
+At the same time, in view of the fact that the two coincide when $X_{1}=X_{2}$, a stronger statement is true: $Z$ and $X_{1}$ have the same conditional laws on the event $\{X_{1}\neq X_{2}\}$.
+This reasoning also applies if $X_{1}$ is replaced by the more symmetric $\Theta X_{1}+(1-\Theta)X_{2}$.
+From this, it follows that the law of the output $Y$ from (37) is unchanged if the equation is instead changed to
 (42) $Y\overset{d}{=}\left\{\begin{array}[]{r l}X_{1}+\mathbf{1}_{\{0\}}(|X_{2}-X_{1}|),&\text{if}\,\,\Theta=1,\\ X_{2}-\mathbf{1}_{\{0\}}(|X_{2}-X_{1}|),&\text{if}\,\,\Theta=0,\end{array}\right$.
@@
 Note that the equation $q=\frac{1}{2}q_{+}$ introduces the constraint $q\leq\frac{1}{2}$, so that the previous result only covers the asymmetric hipster random walk for half of the possible values of $q$.
-In fact, it is possible to show (by manipulation of the evolution equation satisfied by the CDF) that the asymmetric hipster random walk is a monotone RDE only if $q\leq\frac{1}{2}$.
-Therefore, it cannot be described using (9) outside this regime.
+In fact, it is possible to show (by manipulation of the evolution equation satisfied by the CDF) that the asymmetric hipster random walk is a monotone RDE only if $q\leq 1/2$.
+Therefore, it falls outside the class of RDE’s considered in this paper when $q>1/2$.
 Similar arguments apply to the two-player version of the RDE’s analyzed in [2].
@@
 The next result describes the connection with the RDE’s considered here.
-As for the asymmetric hipster random walk, only a subset of the parameter values are covered.
+As for the asymmetric hipster random walk, only a subset of the parameter values are covered; the proof is left to the interested reader.
 Proposition 9.
@@
 Notice that when $q_{+}\neq q_{-}$, the range of the map $(q_{+},q_{-})\mapsto(r,q)$ is
-(46) $\{(r,q)\,\mid\,q\in[0,1],\,\,|r-\frac{1}{2}|\leq\frac{1}{2}(q^{-1}-1)\}.$
+(46) $\left\{(r,q)\,\mid\,q\in[0,1],\,\,\left|r-\frac{1}{2}\right|\leq\frac{1}{2}(q^{-1}-1)\right\}.$
 In general, this is a proper subset of the unit square. However, for any fixed $q\leq\frac{1}{2}$, it covers all possible values of $r\in[0,1]$.
@@
 The $\text{Beta}(2,1)$ limit theorems for asymmetric RDE’s obtained in [1] and [2] also follow, albeit only in the parameter regimes identified above.
-Interestingly, in the asymmetric case, those works were able to obtain limit theorems even in the nonmonotone regime.
+Interestingly, in the asymmetric case, those works were able to prove limit theorems even in the nonmonotone regime.
 5. Scaling Limit
 This section proves the main results, Theorems 1 and 2, as well as Corollary 1.
-Fix data $(f_{+},f_{-})$ satisfying the assumptions of Theorems 1 or 2. Given $\theta\in\mathbb{R}$, define a sequence $\{p^{(N)}\}$ by
-$p^{(N)}=1/2+\theta N^{-1}.$
+Fix data $(f_{+},f_{-})$ satisfying the assumptions of Theorems 1 or 2. Given $\theta\in\mathbb{R}$, define the sequence $\{p^{(N)}\}$ by
+$p^{(N)}=\frac{1}{2}+\theta N^{-1}.$
 Let $T^{(N)}$ be the operator associated with the RDE (9) with $\Phi$ given by (8) and bias parameter $p=p^{(N)}$.
-In view of the results of Section 3, the equation $F_{n}=T^{(N)}F_{n-1}$ can be written in the form
-(47) $F_{n}-F_{n-1}=\mathcal{L}^{(N)}F_{n-1}-2\theta N^{-1}F_{n-1}(1-F_{n-1})$
+In view of the results of Section 3, the equation $F^{(N)}_{n}=T^{(N)}F^{(N)}_{n-1}$ can be written in the form
+(47) $F^{(N)}_{n}-F^{(N)}_{n-1}=\mathcal{L}^{(N)}F^{(N)}_{n-1}-2\theta N^{-1}F^{(N)}_{n-1}(1-F^{(N)}_{n-1})$
 for some function $\mathcal{L}^{(N)}$ defined on $CDF(\overline{\mathbb{R}})$.
@@
   • (ii)
-    Discrete-Time Reaction-Diffusion Equation: There are functions $\{\mathcal{L}^{(N)}\}$ defined on $CDF(\overline{\mathbb{R}})$ and functions $\{Q^{(N)}\}$ defined on $[0,1]$ such that
+    Discrete-Time Reaction-Diffusion Equation: There are functions $\{\mathcal{L}^{(N)}\}$ defined on $CDF(\overline{\mathbb{R}})$ and continuous functions $\{Q^{(N)}\}$ defined on $[0,1]$ such that
     $T^{(N)}F-F=\mathcal{L}^{(N)}F-Q^{(N)}(F).$
@@
   • (v)
-    Consistency of $\mathcal{L}^{(N)}$: If $\Phi\in CDF(\overline{\mathbb{R}})$ is smooth with bounded, uniformly continuous second derivatives, rescaled according to the formula $\Phi_{\delta}(y)=\Phi(\delta y)$, then
-    $\lim_{N\to+\infty}N(\mathscr{L}^{(N)}\Phi_{\delta^{(N)}})\left(\frac{x}{\delta^{(N)}}\right)=\mathcal{F}(\partial_{x}\Phi(x),\partial_{x}^{2}\Phi(x)),$
+    Consistency of $\mathcal{L}^{(N)}$: If $G\in CDF(\overline{\mathbb{R}})$ is smooth with bounded, uniformly continuous second derivatives, rescaled according to the formula $G_{\delta}(y)=G(\delta y)$, then
+    $\lim_{N\to+\infty}N(\mathscr{L}^{(N)}G_{\delta^{(N)}})\left(\frac{x}{\delta^{(N)}}\right)=\mathcal{F}(\partial_{x}G(x),\partial_{x}^{2}G(x)),$
     where convergence holds uniformly with respect to $x\in\mathbb{R}$.
@@
   • (vi)
-    Consistency of $Q^{(N)}$: There is a function $Q:[0,1]\to[0,1]$ such that
+    Consistency of $Q^{(N)}$: There is a function $Q:[0,1]\to\mathbb{R}$ such that
     $\lim_{N\to\infty}\sup\left\{|NQ^{(N)}(q)-Q(q)|\,\mid\,q\in[0,1]\right\}=0$
@@
 and if there is a continuous $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ such that, at time $t=0$,
-$F^{(N)}(\cdot,0)\to F_{\text{in}}\quad\text{vaguely as}\,\,N\to+\infty,$
+$F_{N}(\cdot,0)\to F_{\text{in}}\quad\text{vaguely as}\,\,N\to+\infty,$
 then
-$F_{N}\to F\quad\text{locally uniformly in}\,\,\mathbb{R}\times[0,\infty)\,\,\text{as}\,\,N\to+\infty,$
-where $F$ is the unique bounded viscosity solution of the equation
-(49) $\left\{\begin{array}[]{r l}\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
-It is possible to show that if $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds, as is assumed in Theorems 1 and 2, then the one-parameter family $\{T^{(N)}\}$ defined in the discussion at the beginning of this section satisfies assumptions (i)-(iv); see Section 5.2 for the details.
+$F_{N}\to F\quad\text{locally uniformly in}\,\,\mathbb{R}\times[0,+\infty)\,\,\text{as}\,\,N\to+\infty,$
+where $F$ is the unique bounded (continuous) viscosity solution of the equation
+(49) $\left\{\begin{array}[]{r l}\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)=0&\text{in}\,\,\mathbb{R}\times(0,+\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
+The results of the previous sections imply that if $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds, as is assumed in Theorems 1 and 2, then the one-parameter family $\{T^{(N)}\}$ defined in the discussion at the beginning of this section satisfies assumptions (i)-(iv); see Section 5.2 for the details.
 In particular, $\{Q^{(N)}\}$ is given by $Q^{(N)}(q)=2\theta N^{-1}q(1-q)$, which trivially satisfies (vi).
@@
 Consider the operator $\mathscr{L}^{(N)}$ given by (32) with $p=p^{(N)}$.
-If $\Phi:\mathbb{R}\to[0,1]$ is a nondecreasing function with bounded, uniformly continuous first and second derivatives, rescaled according to the rule $\Phi_{\delta}(y)=\Phi(\delta y)$, and if $R_{\Phi}^{(N)}:\mathbb{R}\to\mathbb{R}$ is the remainder defined by
-$(\mathscr{L}^{(N)}\Phi_{N^{-\frac{1}{2}}})(N^{\frac{1}{2}}x)=N^{-1}\sigma|\partial_{x}\Phi(x)|^{2}+N^{-1}R_{\Phi}^{(N)}(x),$
-then $R^{(N)}_{\Phi}\to 0$ uniformly in $\mathbb{R}$ as $N\to+\infty$.
-Further, if the stronger assumption (16) holds, and if $R^{(N)}_{\Phi}$ is instead defined by
-$(\mathscr{L}^{(N)}\Phi_{N^{-\frac{1}{3}}})(N^{\frac{1}{3}}x)=N^{-\frac{2}{3}}\sigma|\partial_{x}\Phi(x)|^{2}+N^{-1}a|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)+N^{-1}R_{\Phi}^{(N)}(x),$
-then, as before, $R^{(N)}_{\Phi}\to 0$ uniformly as $N\to+\infty$.
+If $G:\mathbb{R}\to[0,1]$ is a nondecreasing function with bounded, uniformly continuous first and second derivatives, rescaled according to the rule $G_{\delta}(y)=G(\delta y)$, and if $R_{G}^{(N)}:\mathbb{R}\to\mathbb{R}$ is the remainder defined so that
+$(\mathscr{L}^{(N)}G_{N^{-\frac{1}{2}}})(N^{\frac{1}{2}}x)=N^{-1}\sigma|\partial_{x}G(x)|^{2}+N^{-1}R_{G}^{(N)}(x),$
+then $R^{(N)}_{G}\to 0$ uniformly in $\mathbb{R}$ as $N\to+\infty$.
+Further, if the stronger assumption (16) holds, and if $R^{(N)}_{G}$ is instead defined by
+$(\mathscr{L}^{(N)}G_{N^{-\frac{1}{3}}})(N^{\frac{1}{3}}x)=N^{-\frac{2}{3}}\sigma|\partial_{x}G(x)|^{2}+N^{-1}a|\partial_{x}G(x)|\partial_{x}^{2}G(x)+N^{-1}R_{G}^{(N)}(x),$
+then, as before, $R^{(N)}_{G}\to 0$ uniformly as $N\to+\infty$.
 The proof of the proposition is relegated to the end of this section (Section 5.4).
@@
   • (vii)
-    Asymptotically regularizing: If $\overline{F}$ and $\underline{F}$ denote the maximal and minimal viscosity sub- and supersolutions, respectively, of the initial-value problem (49) for a given $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ (see Appendix A.1), then
-    $\overline{F}=\underline{F}\quad\text{in}\,\,\mathbb{R}\times(0,\infty).$
-As explained in Appendix A, from the fact that $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,+\infty)$, it follows that there is a unique solution of (49) started from any initial datum $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$, which is continuous for positive times.
+    Asymptotically regularizing: If $\underline{F}$ and $\overline{F}$ denote the maximal and minimal viscosity sub- and supersolutions, respectively, of the initial-value problem (49) for a given $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ (see Appendix A.1), then
+    $\underline{F}=\overline{F}\quad\text{in}\,\,\mathbb{R}\times(0,+\infty).$
+As explained in Appendix A, from the fact that $\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,+\infty)$, it follows that there is a unique solution of (49) started from any initial datum $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$, which is continuous for positive times.
 The next corollary asserts that if the family $\{T^{(N)}\}$ satisfies assumption (vii) in addition to (i)-(vi), then Theorem 3 remains true even if $F_{\text{in}}$ has discontinuities.
@@
 It follows from classical results that, in the PDE’s of interest here, assumption (vii) holds as long as either $\sigma\neq 0$ or $a>0$.
-(The degenerate PDE’s $\partial_{t}F-Q(F)=0$ do not have this property and therefore require a slightly different treatment.) This is made precise in the next proposition.
+(The degenerate PDE $\partial_{t}F+Q(F)=0$ does not have this property and therefore requires a slightly different treatment.) This is made precise in the next proposition.
 Proposition 11.
-Let $Q$ be an arbitrary function for which there is an $L>0$ such that $|Q(q)-Q(q^{\prime})|\leq L|q-q^{\prime}|$ for each $q,q^{\prime}\in[0,1]$ and satisfying both $Q(0)\geq 0$ and $Q(1)\leq 0$.
+Let $Q$ be an arbitrary function for which there is an $L>0$ such that $|Q(q)-Q(q^{\prime})|\leq L|q-q^{\prime}|$ for each $q,q^{\prime}\in[0,1]$ and satisfying both $Q(0)\leq 0$ and $Q(1)\geq 0$.
 If $\mathcal{F}(v,w)=\sigma|v|^{2}$ for some $\sigma\neq 0$ or $\mathcal{F}(v,w)=a|v|w$ for some $a>0$, then assumption (vii) is satisfied.
@@
 There is nothing special about right-continuity, however.
-In fact, let $F_{-}$ and $F^{+}$ denote the left- and right-increasing versions of a nondecreasing function $F$.
+In fact, let $F_{-}$ and $F^{+}$ denote the left- and right-continuous versions of a nondecreasing function $F$.
 If operators $\{T^{(N)}_{-}\}$ are defined on the space $CDF_{\text{left}}(\overline{\mathbb{R}})$ of nondecreasing, left-continuous functions taking values in $[0,1]$ by the formula $T^{(N)}_{-}F=(T^{(N)}(F^{+}))_{-}$, then $\{T^{(N)}_{-}\}$ satisfies assumptions (i)-(vi) if and only if $\{T^{(N)}\}$ does.
@@
 Since $s-g_{\rho}(s)=f_{\rho}(g_{\rho}(s))$ for $s\geq f_{\rho}(0)$ and $g_{\rho}(s)\to+\infty$ as $s\to+\infty$, the integral $\int_{f_{\rho}(0)}^{+\infty}|g_{\rho}(s)-s|\,ds$ is finite only if $f_{\rho}(+\infty)=0$.
-At the same time, (15) implies that this integral is finite almost surely, hence $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$ holds as claimed.
+At the same time, (15) implies that this integral is finite almost surely, hence $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$ as claimed.
 Thus, the hypotheses of Proposition 7 are satisfied, and it follows that assumptions (iii) and (iv) both apply to the sequence $\{T^{(N)}\}$.
@@
 According to Proposition 11 above, from the fact that $\sigma\neq 0$, assumption (vii) also holds.
-Therefore, Corollary 4 applies, establishing convergence of the rescaled CDF $F_{N}$ as desired to the solution of the initial-value problem associated to the PDE $\partial_{t}F-\sigma|\partial_{x}F|^{2}+2\theta F(1-F)=0$, as desired.
+Therefore, Corollary 4 applies, establishing convergence of the rescaled CDF $F_{N}$ to the solution of the initial-value problem associated to the PDE $\partial_{t}F-\sigma|\partial_{x}F|^{2}+2\theta F(1-F)=0$, as desired.
 Similarly, in Theorem 2 (that is, assuming that (16) holds, $\sigma=0$, and $a>0$), Proposition 10 instead establishes that assumption (v) holds with
@@
 ∎
+Remark 3.
+Notice that, in the proof of Theorems 1 and 2 above, nowhere was it necessary to assume that $\{X^{(N,n)}\}$ was real-valued.
+Therefore, since Theorem 3 and Corollary 4 in particular do not require that $F^{(N)}_{n}(+\infty)=1$ or $F^{(N)}_{n}(-\infty)=0$, the conclusions of the theorems remain true if $\{X^{(N,n)}\}$ are extended real-valued.
 5.3. Proof of Corollary 1
@@
 In any event, (50) implies
-$\lim_{N\to\infty}\mathbb{P}\{-\frac{\text{sgn}(\sigma)}{2\sqrt{|\sigma|}N^{1/2}}X^{(N)}\leq x\}=F(2x\sqrt{|\sigma|},1)=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ x^{2},&\text{if}\,\,0\leq x\leq 1,\\ 1,&\text{otherwise.}\end{array}\right$.
-The righthand side is precisely the CDF of the $\text{Beta}(2,1)$ distribution. Therefore, $2^{-1}N^{-1/2}X^{(N)}\overset{d}{\to}\sqrt{|\sigma|}\text{Beta}(2,1)$.
+$\lim_{N\to\infty}\mathbb{P}\{-\text{sgn}(\sigma)(2\sqrt{|\sigma|})^{-1}N^{-\frac{1}{2}}X^{(N)}\leq x\}=F(2x\sqrt{|\sigma|},1)=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ x^{2},&\text{if}\,\,0\leq x\leq 1,\\ 1,&\text{otherwise.}\end{array}\right$.
+The righthand side is precisely the CDF of the $\text{Beta}(2,1)$ distribution. Therefore, $2^{-1}N^{-1/2}X^{(N)}\overset{d}{\to}-\text{sgn}(\sigma)\sqrt{|\sigma|}\text{Beta}(2,1)$.
 It only remains to consider the case when $\sigma=0$.
@@
 $\lim_{N\to\infty}\mathbb{P}\{N^{-1/2}Y^{(N,0)}\leq x\}=G_{\text{in}}(x).$
-This is equivalent to asking that $Y^{(N,0)}=X^{(0)}+N^{1/2}Z^{(N)}$ for some sequence of nonnegative random variables $\{Z^{(N)}\}$ with law converging to $G_{\text{in}}$.
+This is equivalent to asking that $Y^{(N,0)}\overset{d}{=}X^{(0)}+N^{1/2}Z^{(N)}$ for some sequence of nonnegative random variables $\{Z^{(N)}\}$ with law converging to $G_{\text{in}}$.
 Since $G_{\text{in}}$ is continuous, $\sigma=0$, and $p=1/2$ independently of $N$, applying Theorem 3 with $T^{(N)}=T$, $\delta^{(N)}=N^{-1/2}$, $\mathcal{F}=0$, and $Q^{(N)}=Q=0$ yields
@@
 and thus $X^{(n)}\overset{d}{=}(1-\Theta)\max\{X^{(n-1)}_{1},X^{(n-1)}_{2}\}+\Theta\min\{X^{(n-1)}_{1},X^{(n-1)}_{2}\}\overset{d}{=}X^{(n-1)}$ for all $n$ (since $\mathbb{P}\{\Theta=1\}=p=1/2$).
-In other words, $\{X^{(n)}\}$ is the constant sequence $X^{(n)}\overset{d}{=}X^{(0)}$, completing the proof of case (iii).
+In other words, $\{X^{(n)}\}$ is the constant sequence $X^{(n)}\overset{d}{=}X^{(0)}$, completing the proof in case (iii).
 ∎
@@
 In particular, $a_{+}$ and $a_{-}$ are finite.
-With these preliminaries out of the way, fix a $\Phi\in CDF(\overline{\mathbb{R}})$ that has bounded, uniformly continuous first and second derivatives, and define $\Phi_{\delta}(y)=\Phi(\delta y)$ for $\delta>0$.
-Let $\omega_{\Phi}$ be the modulus of continuity
-$\omega_{\Phi}(\delta)=\sup\left\{|\partial_{x}\Phi(x)-\partial_{x}\Phi(y)|+|\partial_{x}^{2}\Phi(x)-\partial_{x}^{2}\Phi(y)|\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\delta\right\}.$
+With these preliminaries out of the way, fix a $G\in CDF(\overline{\mathbb{R}})$ that has bounded, uniformly continuous first and second derivatives, and define $G_{\delta}(y)=G(\delta y)$ for $\delta>0$.
+Let $\omega_{G}$ be the modulus of continuity
+$\omega_{G}(\delta)=\sup\left\{|\partial_{x}G(x)-\partial_{x}G(y)|+|\partial_{x}^{2}G(x)-\partial_{x}^{2}G(y)|\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\delta\right\}.$
 Recall from Section 3 that it is possible to write
@@
 Thus, the scaling behavior of $\mathscr{L}^{(N)}$ is determined by that of $\mathscr{L}_{+}$ and $\mathscr{L}_{-}$.
-Further, since the assumption (15) implies that $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=1\}$ (as proved in Section 5.2 above), the formula (33) can be invoked to find
-$(\mathscr{L}_{+}\Phi_{\delta})(\delta^{-1}x)$ $=\delta\int_{0}^{+\infty}\mathbf{E}[\Phi(x-\delta h_{+}(s))-\Phi(x)]\partial_{x}\Phi(x-\delta s)\,ds,$
-$(\mathscr{L}_{-}\Phi_{\delta})(\delta^{-1}x)$ $=\delta\int_{0}^{+\infty}\mathbf{E}[\Phi(x+\delta h_{-}(s))-\Phi(x)]\partial_{x}\Phi(x+\delta s)\,ds.$
+Further, since the assumption (15) implies that $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$ (as proved in Section 5.2 above), the formula (33) can be invoked to find
+$(\mathscr{L}_{+}G_{\delta})(\delta^{-1}x)$ $=\delta\int_{0}^{+\infty}\mathbf{E}[G(x-\delta h_{+}(s))-G(x)]\partial_{x}G(x-\delta s)\,ds,$
+$(\mathscr{L}_{-}G_{\delta})(\delta^{-1}x)$ $=\delta\int_{0}^{+\infty}\mathbf{E}[G(x+\delta h_{-}(s))-G(x)]\partial_{x}G(x+\delta s)\,ds.$
 In view of (55), it only remains to prove the following two claims: First, under assumption (15), for any $\rho\in\{+,-\}$,
-(56) $\lim_{\delta\downarrow 0}\sup\left\{\left|\delta^{-2}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}\right|\,\mid\,x\in\mathbb{R}\right\}=0.$
+(56) $\lim_{\delta\downarrow 0}\sup\left\{\left|\delta^{-2}(\mathscr{L}_{\rho}G_{\delta})(\delta^{-1}x)-\sigma_{\rho}|\partial_{x}G(x)|^{2}\right|\,\mid\,x\in\mathbb{R}\right\}=0.$
 Further, if assumption (16) holds, then
-(57) $\lim_{\delta\downarrow 0}\sup\left\{\left|\delta^{-3}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\delta^{-1}\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}-a_{\rho}|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)\right|\,\mid\,x\in\mathbb{R}\right\}=0.$
+(57) $\lim_{\delta\downarrow 0}\sup\left\{\left|\delta^{-3}(\mathscr{L}_{\rho}G_{\delta})(\delta^{-1}x)-\delta^{-1}\sigma_{\rho}|\partial_{x}G(x)|^{2}-a_{\rho}|\partial_{x}G(x)|\partial_{x}^{2}G(x)\right|\,\mid\,x\in\mathbb{R}\right\}=0.$
 $\delta^{2}$ Asymptotics. Assume (15) holds and fix a $\rho\in\{+,-\}$. First, observe that, by the definition (53) of $\sigma_{\rho}$,
-(58) $\delta^{-2}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}$
-$=\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}\partial_{x}\Phi(x-\rho\delta s)+\rho|\partial_{x}\Phi(x)|^{2}\right)h_{\rho}(s)\,ds.$
-By the assumptions on $\Phi$ and the bound $h_{\rho}(s)\leq s$,
-$\left|\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}\partial_{x}\Phi(x-\rho\delta s)+\rho|\partial_{x}\Phi(x)|^{2}\right)h_{\rho}(s)\,ds\right|$
-$\qquad\leq\|\partial_{x}\Phi\|_{\text{sup}}\mathbf{E}\int_{0}^{+\infty}\left(\omega_{\Phi}(\delta h_{+}(s))+\omega_{\Phi}(\delta s)\right)h_{\rho}(s)\,ds$
-$\qquad\leq 2\|\partial_{x}\Phi\|_{\text{sup}}\int_{0}^{+\infty}\omega_{\Phi}(\delta s)\mathbf{E}h_{\rho}(s)\,ds$
+(58) $\delta^{-2}(\mathscr{L}_{\rho}G_{\delta})(\delta^{-1}x)-\sigma_{\rho}|\partial_{x}G(x)|^{2}$
+$=\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{G(x-\rho\delta h_{\rho}(s))-G(x)\}\partial_{x}G(x-\rho\delta s)+\rho|\partial_{x}G(x)|^{2}\right)h_{\rho}(s)\,ds.$
+By the assumptions on $G$ and the bound $h_{\rho}(s)\leq s$,
+$\left|\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{G(x-\rho\delta h_{\rho}(s))-G(x)\}\partial_{x}G(x-\rho\delta s)+\rho|\partial_{x}G(x)|^{2}\right)h_{\rho}(s)\,ds\right|$
+$\qquad\leq\|\partial_{x}G\|_{\text{sup}}\mathbf{E}\int_{0}^{+\infty}\left(\omega_{G}(\delta h_{\rho}(s))+\omega_{G}(\delta s)\right)h_{\rho}(s)\,ds$
+$\qquad\leq 2\|\partial_{x}G\|_{\text{sup}}\int_{0}^{+\infty}\omega_{G}(\delta s)\mathbf{E}h_{\rho}(s)\,ds$
 Since assumption (15) implies $\mathbf{E}h_{\rho}$ is integrable, Lebesgue’s dominated convergence theorem implies that the righthand side vanishes as $\delta\downarrow 0$. Therefore, (56) follows.
 $\delta^{3}$ Asymptotics. Finally, assume that assumption (16) holds. In this case, notice that the main error term from the previous step can be rewritten in the form
-$\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}\partial_{x}\Phi(x-\rho\delta s)+\rho|\partial_{x}\Phi(x)|^{2}\right)h_{\rho}(s)\,ds$
-$\quad=\rho\partial_{x}\Phi(x)\mathbf{E}\int_{0}^{+\infty}(\partial_{x}\Phi(x)-\partial_{x}\Phi(x-\rho\delta s))h_{\rho}(s)\,ds$
-$\quad\quad+\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}+\rho\partial_{x}\Phi(x)\right)\partial_{x}\Phi(x-\rho\delta s)h_{\rho}(s)\,ds.$
+$\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{G(x-\rho\delta h_{\rho}(s))-G(x)\}\partial_{x}G(x-\rho\delta s)+\rho|\partial_{x}G(x)|^{2}\right)h_{\rho}(s)\,ds$
+$\quad=\rho\partial_{x}G(x)\mathbf{E}\int_{0}^{+\infty}(\partial_{x}G(x)-\partial_{x}G(x-\rho\delta s))h_{\rho}(s)\,ds$
+$\quad\quad+\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{G(x-\rho\delta h_{\rho}(s))-G(x)\}+\rho\partial_{x}G(x)\right)\partial_{x}G(x-\rho\delta s)h_{\rho}(s)\,ds.$
 In the first term, one finds
-$\left|\frac{\rho}{\delta}\mathbf{E}\int_{0}^{+\infty}\partial_{x}\Phi(x)(\partial_{x}\Phi(x)-\partial_{x}\Phi(x-\rho\delta s))h_{\rho}(s)\,ds-\partial_{x}\Phi(x)\partial_{x}^{2}\Phi(x)\int_{0}^{+\infty}s\mathbf{E}h_{\rho}(s)\,ds\right|$
-$\qquad\leq\|\partial_{x}\Phi\|_{\text{sup}}\int_{0}^{+\infty}\omega_{\Phi}(\delta s)s\mathbf{E}h_{\rho}(s)\,ds.$
-Similarly, in the second term, again using $\omega_{\Phi}(\delta h_{+}(s))\leq\omega_{\Phi}(\delta s)$,
-$\left|\frac{1}{\delta}\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}+\rho\partial_{x}\Phi(x)\right)\partial_{x}\Phi(x-\rho\delta s)h_{\rho}(s)\,ds\right$.
-$\qquad\left.-\frac{1}{2}\partial_{x}\Phi(x)\partial_{x}^{2}\Phi(x)\int_{0}^{+\infty}\mathbf{E}h_{\rho}(s)^{2}\,ds\right|$
-$\qquad\leq\frac{1}{2}\left(\|\partial_{x}\Phi\|_{\text{sup}}+\|\partial_{x}^{2}\Phi\|_{\text{sup}}\right)\int_{0}^{+\infty}\omega_{\Phi}(\delta s)\mathbf{E}h_{\rho}(s)^{2}\,ds.$
+$\left|\frac{\rho}{\delta}\mathbf{E}\int_{0}^{+\infty}\partial_{x}G(x)(\partial_{x}G(x)-\partial_{x}G(x-\rho\delta s))h_{\rho}(s)\,ds-\partial_{x}G(x)\partial_{x}^{2}G(x)\int_{0}^{+\infty}s\mathbf{E}h_{\rho}(s)\,ds\right|$
+$\qquad\leq\|\partial_{x}G\|_{\text{sup}}\int_{0}^{+\infty}\omega_{G}(\delta s)s\mathbf{E}h_{\rho}(s)\,ds.$
+Similarly, in the second term, again using $\omega_{G}(\delta h_{\rho}(s))\leq\omega_{G}(\delta s)$,
+$\left|\frac{1}{\delta}\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{G(x-\rho\delta h_{\rho}(s))-G(x)\}+\rho\partial_{x}G(x)\right)\partial_{x}G(x-\rho\delta s)h_{\rho}(s)\,ds\right$.
+$\qquad\left.-\frac{1}{2}\partial_{x}G(x)\partial_{x}^{2}G(x)\int_{0}^{+\infty}\mathbf{E}h_{\rho}(s)^{2}\,ds\right|$
+$\qquad\leq\frac{1}{2}\left(\|\partial_{x}G\|_{\text{sup}}+\|\partial_{x}^{2}G\|_{\text{sup}}\right)\int_{0}^{+\infty}\omega_{G}(\delta s)\mathbf{E}h_{\rho}(s)^{2}\,ds.$
 Combining these estimates with the identity (58) and estimate (54), one finds
-$\sup\left\{\left|\delta^{-3}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\delta^{-1}\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}-a_{\rho}|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)\right|\,\mid\,x\in\mathbb{R}\right\}$
-$\qquad\leq\left(1+\|\partial_{x}\Phi\|_{\text{sup}}+\|\partial_{x}^{2}\Phi\|_{\text{sup}}\right)\int_{0}^{+\infty}\omega_{\Phi}(\delta s)s\mathbf{E}h_{\rho}(s)\,ds.$
+$\sup\left\{\left|\delta^{-3}(\mathscr{L}_{\rho}G_{\delta})(\delta^{-1}x)-\delta^{-1}\sigma_{\rho}|\partial_{x}G(x)|^{2}-a_{\rho}|\partial_{x}G(x)|\partial_{x}^{2}G(x)\right|\,\mid\,x\in\mathbb{R}\right\}$
+$\qquad\leq\left(1+\|\partial_{x}G\|_{\text{sup}}+\|\partial_{x}^{2}G\|_{\text{sup}}\right)\int_{0}^{+\infty}\omega_{G}(\delta s)s\mathbf{E}h_{\rho}(s)\,ds.$
 Since the righthand side vanishes as $\delta\downarrow 0$ by dominated convergence, (57) follows. ∎
@@
 where $Z$ is a nonnegative random variable such that $\mathbb{E}Z^{2}=+\infty$.
-With this choice of forcing, the RDE (9) with $p=1/2$ is superdiffusive in the sense that if $\{X^{(n)}\}$ is a solution started from any initial datum and $p=1/2$, then
+With this choice of forcing, the RDE (9) with $p=1/2$ is superdiffusive in the sense that if $\{X^{(n)}\}$ is a solution started from any real-valued initial datum, then
 $N^{-\frac{1}{2}}X^{(N)}\overset{d}{\to}+\infty\quad\text{as}\,\,N\to+\infty.$
@@
 Let $\{X^{(m,n)}\}$ be a solution of the RDE (9) started from the same initial distribution as $\{X^{(n)}\}$, but with the forcing $(f_{+},0)$ replaced by $(f_{+}^{(m)},0)$.
-Note that the inverse $g_{+}^{(m)}$ associated with $f_{+}^{(m)}$ is given by
+Note that the function $g_{+}^{(m)}$ associated to $f_{+}^{(m)}$ via (14) is given by
 $g_{+}^{(m)}(s)=0\quad\text{if}\,\,s<\max\{Z,m\},\quad g_{+}^{(m)}(s)=s,\quad\text{otherwise,}$
-and the inverse $g_{-}$ associated with $f_{-}=0$ is $g_{-}(s)=s$. Therefore, the RDE with $(f^{(m)}_{+},0)$ satisfies (15) and the constant $\sigma^{(m)}$ determining the asymptotics is given by
+and the function $g_{-}$ associated to $f_{-}=0$ is $g_{-}(s)=s$. Therefore, the RDE with $(f^{(m)}_{+},0)$ satisfies (15) and the constant $\sigma^{(m)}$ determining the asymptotics is given by
 $\sigma^{(m)}=-\mathbf{E}\int_{0}^{\max\{Z,m\}}s\,ds=-\frac{1}{2}\mathbf{E}\max\{Z^{2},m^{2}\}.$
@@
 Therefore, by Corollary 1,
-$\limsup_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(m,N)}\leq x\}\leq F(2^{-1}|\sigma^{(m)}|^{-\frac{1}{2}}x).$
-where $F$ is the CDF of the $\text{Beta}(2,1)$ random variable Sending $m\to+\infty$, this implies
-$\limsup_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(N)}\leq x\}\leq F(0)=0\quad\text{for each}\,\,x\in\mathbb{R}.$
+$\limsup_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(m,N)}\leq x\}\leq F_{\text{Beta}(2,1)}(2^{-1}|\sigma^{(m)}|^{-\frac{1}{2}}x).$
+where $F_{\text{Beta}(2,1)}$ is the CDF of the $\text{Beta}(2,1)$ random variable. Sending $m\to+\infty$, this implies
+$\limsup_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(N)}\leq x\}\leq F_{\text{Beta}(2,1)}(0)=0\quad\text{for each}\,\,x\in\mathbb{R}.$
 Appendix A Viscosity Solutions
@@
 (63) $|Q(q)-Q(q^{\prime})|\leq L|q-q^{\prime}|\,\,\text{for each}\,\,q,q^{\prime}\in\mathbb{R},$
-(64) $\mathcal{F}(0,0)=0,\quad Q(0)\geq 0,\quad\text{and}\quad Q(1)\leq 0.$
+(64) $\mathcal{F}(0,0)=0,\quad Q(0)\leq 0,\quad\text{and}\quad Q(1)\geq 0.$
 Since the problem at hand naturally leads to consideration of possibly discontinuous initial data, some care will be needed in defining precisely what is meant by a solution, subsolution, or supersolution.
@@
 Definition 1.
-An upper semicontinuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ in the viscosity sense if it has the following property: If $\varphi$ is a smooth function defined in some open set $U$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from above (that is, $F(x,t)\leq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
+An upper semicontinuous function $F:\mathbb{R}\times(0,+\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$ if it has the following property: If $\varphi$ is a smooth function defined in some open set $U\subseteq\mathbb{R}\times(0,+\infty)$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from above (that is, $F(x,t)\leq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
 $\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))+Q(\varphi(x_{0},t_{0}))\leq 0.$
-A lower semi-continuous function $G:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial_{x}^{2}G)+Q(G)\geq 0$ in the viscosity sense if it has the following property: If $\varphi$ is a smooth function defined in some open set $U$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from below (that is, $F(x,t)\geq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
+A lower semicontinuous function $G:\mathbb{R}\times(0,+\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial_{x}^{2}G)+Q(G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$ if it has the following property: If $\varphi$ is a smooth function defined in some open set $U\subseteq\mathbb{R}\times(0,+\infty)$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from below (that is, $F(x,t)\geq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
 $\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))+Q(\varphi(x_{0},t_{0}))\geq 0.$
-A continuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential equation $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)=0$ in the viscosity sense if it satisfies both $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ and $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)\geq 0$ in the viscosity sense.
+A continuous function $F:\mathbb{R}\times(0,+\infty)\to\mathbb{R}$ is said to satisfy the differential equation $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)=0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$ if it satisfies both $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ and $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$.
 The next definition specifies what is meant by a solution, subsolution, or supersolution of the initial-value problem (61).
 Definition 2.
-An upper semicontinuous function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ said to be a viscosity subsolution of (61) if it satisfies $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
+An upper semicontinuous function $F:\mathbb{R}\times[0,+\infty)\to\mathbb{R}$ is said to be a viscosity subsolution of (61) if it satisfies $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$ and
 (65) $\lim_{\delta\downarrow 0}\sup\left\{F(y,0)-F_{\text{in}}(x)\,\mid\,x,y\in\mathbb{R}\,\,\text{such that}\,\,|y-x|\leq\delta\right\}\leq 0.$
-A lower semicontinuous function $G:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a viscosity supersolution of (61) if it satisfies $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)+Q(G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
+A lower semicontinuous function $G:\mathbb{R}\times[0,+\infty)\to\mathbb{R}$ is said to be a viscosity supersolution of (61) if it satisfies $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)+Q(G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$ and
 $\lim_{\delta\downarrow 0}\inf\left\{G(y,0)-F_{\text{in}}(x)\,\mid\,x,y\in\mathbb{R}\,\,\text{such that}\,\,|y-x|\leq\delta\right\}\geq 0.$
-A continuous function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a (continuous) viscosity solution of (61) if it is both a viscosity subsolution and a viscosity supersolution.
-Since the main results of this work, Theorems 1 and 2, involve problems with a discontinuous initial datum, a different definition of solution is needed.
+A continuous function $F:\mathbb{R}\times[0,+\infty)\to\mathbb{R}$ is said to be a (continuous) viscosity solution of (61) if it is both a viscosity subsolution and a viscosity supersolution.
+Since the main results of this work, Theorems 1 and 2, consider the problem (61) with initial data that may be discontinuous, a different definition of solution is needed.
 In order to state it, it is necessary to recall the notions of upper and lower semicontinuous envelopes.
-Specifically, for a function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$, the upper and lower semicontinuous envelopes $F^{*}$ and $F_{*}$ are the functions defined in $\mathbb{R}\times[0,\infty)$ via the formulae
+Specifically, for a function $F:\mathbb{R}\times[0,+\infty)\to\mathbb{R}$, the upper and lower semicontinuous envelopes $F^{*}$ and $F_{*}$ are the functions defined in $\mathbb{R}\times[0,+\infty)$ via the formulae
 $F^{*}(x,t)=\lim_{\delta\downarrow 0}\sup\left\{F(y,s)\,\mid\,y,s\in\mathbb{R}\times[0,\infty)\,\,\text{such that}\,\,|x-y|+|t-s|\leq\delta\right\},$
@@
 Definition 3.
-A function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a discontinuous viscosity solution of (61) if $\partial_{t}F^{*}-\mathcal{F}(\partial_{x}F^{*},\partial_{x}^{2}F^{*})+Q(F^{*})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$; $\partial_{t}F_{*}-\mathcal{F}(\partial_{x}F_{*},\partial_{x}^{2}F_{*})+Q(F_{*})\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$; and $F^{*}(\cdot,0)\leq(F_{\text{in}})^{*}$ and $F_{*}(\cdot,0)\geq(F_{\text{in}})_{*}$ in $\mathbb{R}$.
+A function $F:\mathbb{R}\times[0,+\infty)\to\mathbb{R}$ is said to be a discontinuous viscosity solution of (61) if $\partial_{t}F^{*}-\mathcal{F}(\partial_{x}F^{*},\partial_{x}^{2}F^{*})+Q(F^{*})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$; $\partial_{t}F_{*}-\mathcal{F}(\partial_{x}F_{*},\partial_{x}^{2}F_{*})+Q(F_{*})\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$; and $F^{*}(\cdot,0)\leq(F_{\text{in}})^{*}$ and $F_{*}(\cdot,0)\geq(F_{\text{in}})_{*}$ in $\mathbb{R}$.
 A.1. Comparison Principle
@@
 Assume that $\mathcal{F}$ and $Q$ satisfy assumptions (62) and (63).
-If $F,G:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ are respectively a bounded upper and a bounded lower semicontinuous function such that $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ and $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)+Q(G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and if $M$ is the constant defined by
-$M=\lim_{\delta\downarrow 0}\sup\left\{(F(x,0)-G(y,0))_{+}\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\delta\right\},$
+If $F,G:\mathbb{R}\times[0,+\infty)\to\mathbb{R}$ are respectively a bounded upper and a bounded lower semicontinuous function such that $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ and $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)+Q(G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$ and if $M$ is the constant defined by
+$M=\lim_{\delta\downarrow 0}\sup\left\{(F(x,0)-G(y,0))_{+}\,\mid\,x,y\in\mathbb{R}\,\,\text{such that}\,\,|x-y|\leq\delta\right\},$
 then, for any $t>0$,
-(66) $\sup\left\{F(x,t)-G(x,t)\,\mid\,x\in\mathbb{R}\right\}\leq Me^{Lt}$
-In particular, if $F$ and $G$ are respectively bounded viscosity sub- and supersolutions of (61) for some fixed initial datum $F_{\text{in}}$, then $F\leq G$ pointwise in $\mathbb{R}\times[0,\infty)$.
+(66) $\sup\left\{F(x,t)-G(x,t)\,\mid\,x\in\mathbb{R}\right\}\leq Me^{Lt}.$
+In particular, if $F$ and $G$ are respectively bounded viscosity sub- and supersolutions of (61) for some fixed initial datum $F_{\text{in}}$, then $F\leq G$ pointwise in $\mathbb{R}\times[0,+\infty)$.
 Proof.
@@
 ∎
-The comparison principle implies, in particular, that there is at most one viscosity solution of (61) if $F_{\text{in}}$ is continuous.
+The comparison principle implies, in particular, that there is at most one viscosity solution of (61) if $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ is continuous.
 It is known that, indeed, a solution exists in this case, and the assumption (64) ensures that, for each $t>0$, $F(\cdot,t)\in CDF(\overline{\mathbb{R}})$.
@@
 Assume that $\mathcal{F}$ and $Q$ satisfy (62), (63), and (64).
-For any continuous $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$, there is a unique viscosity solution $F$ of (61).
+For any continuous $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$, there is a unique bounded (continuous) viscosity solution $F$ of (61).
 Further, for any $t>0$, the function $F(\cdot,t)$ is in $CDF(\overline{\mathbb{R}})$.
 Proof.
-Existence of a solution taking values in $[0,1]$ follows from Perron’s Method, which is applicable since the constant functions 0 and $1$ are, respectively, viscosity sub- and supersolutions by (64).
-By the comparison principle, the solution is unique and it takes values in $[0,1]$.
+The existence of a bounded solution follows from Perron’s Method, see [15] and the references therein.
+By Theorem 4, this is the unique bounded solution.
+By (64), since $0\leq F_{\text{in}}\leq 1$, the constant functions $G(x,t)\equiv 1$ and $G(x,t)\equiv 0$ are respectively viscosity super- and subsolutions of (61), which implies that $0\leq F(x,t)\leq 1$ for any $(x,t)$ by comparison.
 Finally, for any $y\in\mathbb{R}$, the function $(x,t)\mapsto F(x+y,t)$ is a viscosity solution of (61) with the initial datum $x\mapsto F_{\text{in}}(x+y)$.
-Thus, since $F_{\text{in}}$ is nondecreasing, by the comparison principle, $F(x+y,t)\leq F(x,t)$ for all $t\geq 0$ if $y\geq 0$.
-Since $F$ is continuous, this proves $F(\cdot,t)\in CDF(\overline{\mathbb{R}})$ for each $t>0$.
+Thus, since $F_{\text{in}}$ is nondecreasing, the comparison principle implies that $F(x+y,t)\leq F(x,t)$ for all $t\geq 0$ if $y\geq 0$.
+Altogether, this proves that $F(\cdot,t)\in CDF(\overline{\mathbb{R}})$ for each $t>0$.
 ∎
@@
 Let $\underline{F}$ and $\overline{F}$ be the maximal subsolution and minimal supersolution of (61), respectively, for some given $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$.
-If $\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,\infty)$, then any discontinuous viscosity solution $F$ of (61) must also satisfy $F=\underline{F}$ in $\mathbb{R}\times(0,\infty)$.
+If $\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,+\infty)$, then any discontinuous viscosity solution $F$ of (61) must also satisfy $F=\underline{F}$ in $\mathbb{R}\times(0,+\infty)$.
 In particular, in this case, modulo the freedom to choose the value at discontinuities of $F_{\text{in}}$ at $t=0$, there is a unique discontinuous viscosity solution of (61), which is continuous in $\mathbb{R}\times(0,+\infty)$.
 Proof.
-As in the discussion preceding the statement, if $F$ is any discontinuous viscosity solution, then $\underline{F}\leq F\leq\overline{F}$ in $\mathbb{R}\times[0,\infty)$ in general.
-Therefore, in the present setting, $F=\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,+\infty)$, implying that there is at most one discontinuous viscosity solution (up to arbitrary choices at $t=0$ at points where $F_{\text{in}}$ is discontinuous).
+As in the discussion preceding the statement, if $F$ is any discontinuous viscosity solution, then $\underline{F}\leq F\leq\overline{F}$ in $\mathbb{R}\times[0,+\infty)$ in general.
+Therefore, in the present setting, $F=\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,+\infty)$, implying that there is at most one discontinuous viscosity solution (up to arbitrary choices at $t=0$ at points where $F_{\text{in}}$ jumps).
 The fact that at least one exists is almost immediate: Since $\underline{F}$ is a supremum of subsolutions, the upper semicontinuous envelope $\underline{F}^{*}$ satisfies $\partial_{t}\underline{F}^{*}-\mathcal{F}(\partial_{x}\underline{F}^{*},\partial_{x}^{2}\underline{F}^{*})+Q(\underline{F}^{*})\leq 0$ in $\mathbb{R}\times(0,+\infty)$, and solutions of (61) with initial data larger than $F_{\text{in}}$ can be used to prove $\underline{F}^{*}(\cdot,0)=F_{\text{in}}$ at $t=0$.
@@
 A similar argument shows that $\partial_{t}\overline{F}_{*}-\mathcal{F}(\partial_{x}\overline{F}_{*},\partial^{2}_{x}\overline{F}_{*})+Q(\overline{F}_{*})\geq 0$ in $\mathbb{R}\times(0,+\infty)$ and $\overline{F}_{*}(\cdot,0)$ equals the lower semicontinuous envelope of $F_{\text{in}}$, so $\overline{F}_{*}\geq\underline{F}$.
-Since $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,+\infty)$ and the upper (resp.
+Since $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,+\infty)$ by assumption and the upper (resp.
 lower) semicontinuous envelope of a function is always larger (resp.
@@
 Lemma 1.
-Fix $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$.
+Assume that $\mathcal{F}$ and $Q$ satisfy assumptions (62) and (63) and fix $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$.
 Suppose that $\{F^{(N),+}_{\text{in}}\}$ and $\{F^{(N),-}_{\text{in}}\}$ are two sequences of continuous functions in $CDF(\overline{\mathbb{R}})$ such that
@@
 In the case of the second-order equation in the $\text{Beta}(2,2)$ case, the uniqueness of solutions follows, for instance, from the characterization of nonnegative solutions of the porous medium equation.
-Short, simplified proofs are presented here for the reader’s convenience.
+Short proofs are presented here for the reader’s convenience.
 Case: $Q=0$. The discussion begins with the classical case when the reaction term $Q=0$.
@@
 for some convex or concave function $H:\mathbb{R}\to\mathbb{R}$ exhibiting superlinear growth:
-(69) $\lim_{|p|\to\infty}|p|^{-1}H(p)=+\infty.$
+(69) $\lim_{|v|\to\infty}|v|^{-1}|H(v)|=+\infty.$
 At this stage, it is worth noting that there is no loss of generality in assuming that $H$ is convex.
-Indeed, in general, if $\partial_{t}F+H(\partial_{x}F)=0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$, then the function G(x,t)=1F(x,t)G(x,t)=1-F(-x,t) satisfies
-$\partial_{t}G-H(-\partial_{x}G)=0\quad\text{in the viscosity sense in}\,\,\mathbb{R}\times(0,\infty)$
-and $H$ is convex if and only if the function $p\mapsto-H(-p)$ is concave.
-For superlinear convex/concave $H$, the Hopf-Lax formula furnishes the unique discontinuous viscosity solution of (61).
+Indeed, in general, if $\partial_{t}F+H(\partial_{x}F)=0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$, then the function G(x,t)=1F(x,t)G(x,t)=1-F(-x,t) satisfies
+$\partial_{t}G-H(\partial_{x}G)=0\quad\text{in the viscosity sense in}\,\,\mathbb{R}\times(0,+\infty)$
+and $H$ is convex if and only if the function $v\mapsto-H(v)$ is concave.
+For superlinear convex/concave $H$, the Hopf-Lax formula furnishes the unique discontinuous viscosity solution of (61) when $Q=0$.
 Proposition 14.
 Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ and assume $Q=0$.
-If $\mathcal{F}(v,w)=-H(v)$ for a convex $H:\mathbb{R}\to\mathbb{R}$ satisfying $H(0)=0$ and (69), and if $\underline{F}$ and $\overline{F}$ are the maximal subsolution and minimal supersolution, respectively, of (61), then, for any $(x,t)\in\mathbb{R}\times(0,\infty)$,
+If $\mathcal{F}(v,w)=-H(v)$ for a convex $H:\mathbb{R}\to\mathbb{R}$ satisfying $H(0)=0$ and (69), and if $\underline{F}$ and $\overline{F}$ are the maximal subsolution and minimal supersolution, respectively, of (61), then, for any $(x,t)\in\mathbb{R}\times(0,+\infty)$,
 (70) $\underline{F}(x,t)=\overline{F}(x,t)=\inf\left\{F_{\text{in}}(y)+tL\left(\frac{x-y}{t}\right)\,\mid\,y\in\mathbb{R}\right\}.$
@@
     $\lim_{n\to\infty}\sup\left\{|\underline{F}^{(n)}(x,t)-\underline{F}^{(n)}(x,t)|\,\mid\,(x,t)\in K\right\}=0.$
-Some regularity of $F_{\text{in}}$ is necessary to ensure that $\overline{F}$ and $\underline{F}$ coincide at positive times.
+Some regularity of $F_{\text{in}}$ is necessary to ensure that $\underline{F}$ and $\overline{F}$ coincide at positive times.
 For instance, $F_{\text{in}}=\mathbf{1}_{\mathbb{Q}}$ provides a counterexample.
@@
 Therefore, since $F_{\text{in}}$ is right-continuous, this proves $\lim_{N\to\infty}F^{(N),+}\leq F$ pointwise in $\mathbb{R}\times(0,+\infty)$.
-Next, by superlinearity of $L$, for any $(x,t)$, it is possible to write $F^{(N),-}(x,t)=F^{(N),-}_{\text{in}}(y_{N})+tL((x-y_{N})/t)$ for some bounded sequence $\{y_{N}\}$.
+Next, since the superlinearity of $H$ implies that $L$ is superlinear (see [18, Theorem 3, Section 3.3.2]), for any $(x,t)$, it is possible to write $F^{(N),-}(x,t)=F^{(N),-}_{\text{in}}(y_{N})+tL((x-y_{N})/t)$ for some bounded sequence $\{y_{N}\}$.
 If $y_{N}\to y$ as $N\to\infty$ (say, along a subsequence), then
@@
 (71) $\mathcal{F}(v,w)=a|v|^{m-1}w$
-Notice that if $F$ were a smooth solution of $\partial_{t}F-a|\partial_{x}F|^{m-1}\partial_{x}^{2}F=0$, then the derivative $\rho=\partial_{x}F$ would solve the porous medium equation (PME) $\partial_{t}\rho=\frac{a}{m}\partial_{x}^{2}(\rho^{m})$.
+Notice that if $F$ were a smooth strictly increasing solution of $\partial_{t}F-a|\partial_{x}F|^{m-1}\partial_{x}^{2}F=0$, then the derivative $\rho=\partial_{x}F$ would solve the porous medium equation (PME) $\partial_{t}\rho=\frac{a}{m}\partial_{x}^{2}(\rho^{m})$.
 To prove the uniqueness of solutions with possibly discontinuous, nondecreasing data, this reasoning will be inverted.
@@
 Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ and assume $Q=0$.
-If $\mathcal{F}$ has the form (71) for some $m>1$, then the maximal subsolution $\underline{F}$ and minimal supersolution $\overline{F}$ of (61) are given for $(x,t)\in\mathbb{R}\times(0,\infty)$ by the formula
+If $\mathcal{F}$ has the form (71) for some $m>1$, then the maximal subsolution $\underline{F}$ and minimal supersolution $\overline{F}$ of (61) are given for $(x,t)\in\mathbb{R}\times(0,+\infty)$ by the formula
 (72) $\underline{F}(x,t)=\overline{F}(x,t)=F_{\text{in}}(-\infty)+\int_{-\infty}^{x}\rho(y,t)\,dy,$
 where $\rho\geq 0$ is the unique continuous distributional solution of the initial-value problem
-(75) $\left\{\begin{array}[]{r l}\partial_{t}\rho-\frac{a}{m}\partial_{x}^{2}(\rho^{m})=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ \rho(0)=\partial_{x}F_{\text{in}}.\end{array}\right$.
+(75) $\left\{\begin{array}[]{r l}\partial_{t}\rho-\frac{a}{m}\partial_{x}^{2}(\rho^{m})=0&\text{in}\,\,\mathbb{R}\times(0,+\infty),\\ \rho(0)=\partial_{x}F_{\text{in}}.\end{array}\right$.
 Furthermore, the following statements hold:
@@
 The uniqueness and claimed representation formula for the solution follows from the classical uniqueness theorem for the PME by Dahlberg and Kenig [16].
-To see this, first, note that if $\partial_{x}F_{\text{in}}$ is smooth and positive, then it is well-known that the solution $\rho$ of the PME is smooth (up to time $t=0$), hence the antiderivative given by (72) is a smooth solution of (61).
+To see this, first, note that if $\partial_{x}F_{\text{in}}$ is smooth and positive, then it is well-known that the solution $\rho$ of the PME is smooth, hence the antiderivative given by (72) is a smooth solution of (61).
 It is classical that smooth solutions are viscosity solutions.
@@
 Since $F^{(N),-}\leq F^{-}\leq F^{+}\leq F^{(N),+}$ holds in $\mathbb{R}\times[0,+\infty)$ for any $N$, $F^{-}(\cdot,t)$ and $F^{+}(\cdot,t)$ both converge vaguely to $F_{\text{in}}$ as $t\downarrow 0$.
-By Arzelà-Ascoli, the derivatives $\partial_{x}F^{-}$ and $\partial_{x}F^{+}$ are continuous distributional solutions of the PME in $\mathbb{R}\times(0,+\infty)$, and the vague convergence of $F^{-}$ and $F^{+}$ at time $t=0$ implies that the initial trace of $\partial_{x}F^{-}$ and $\partial_{x}F^{+}$ equals $\partial_{x}F_{\text{in}}$.
+By Arzelà-Ascoli, the derivatives $\partial_{x}F^{-}$ and $\partial_{x}F^{+}$ are continuous distributional solutions of the PME in $\mathbb{R}\times(0,+\infty)$, and the vague convergence of $F^{-}$ and $F^{+}$ at time $t=0$ implies that the initial traces of $\partial_{x}F^{-}$ and $\partial_{x}F^{+}$ both equal $\partial_{x}F_{\text{in}}$.
 Therefore, by Dahlberg-Kenig, $\partial_{x}F^{-}=\partial_{x}F^{+}$ in $\mathbb{R}\times(0,+\infty)$, which implies the difference $F^{+}-F^{-}$ depends only on time.
@@
 Proposition 16.
-The identity $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,\infty)$ still holds for the initial-value problem (61) with arbitrary data in $CDF(\overline{\mathbb{R}})$ if $\mathcal{F}$ satisfies the assumptions of Propositions 14 or 15 and $Q$ satisfies (63) and (64).
+The identity $\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,+\infty)$ still holds for the initial-value problem (61) with arbitrary data in $CDF(\overline{\mathbb{R}})$ if $\mathcal{F}$ satisfies the assumptions of Propositions 14 or 15 and $Q$ satisfies (63) and (64).
 Proof.
 Fix $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$.
-If $F_{\text{in}}$ is continuous, then Proposition 12 implies there is a unique viscosity solution $F$ of (61) and $\overline{F}=\underline{F}=F$.
+If $F_{\text{in}}$ is continuous, then Proposition 12 implies there is a unique viscosity solution $F$ of (61) and $\underline{F}=\overline{F}=F$.
 Otherwise, if it has jumps, fix sequences $\{F_{\text{in}}^{(N),+}\}$ and $\{F_{\text{in}}^{(N),-}\}$ of continuous functions in $CDF(\overline{\mathbb{R}})$ satisfying (67) and (68) as in Lemma 1 and, in addition,
@@
 Finally, moving back to the original problem (i.e., with the reaction term $Q$), the estimate (66) in Theorem 4 implies that, for any $t>\delta$,
-$\limsup_{N\to\infty}\sup\left\{F^{(N),+}(x,t)-F^{(N),-}(x,t)\,\mid\,x\in\mathbb{R}\right\}\leq 2C(B)\delta e^{L(t-\delta)}.$
+$\limsup_{N\to\infty}\sup\left\{F^{(N),+}(x,t)-F^{(N),-}(x,t)\,\mid\,x\in\mathbb{R}\right\}\leq 2C(Q)\delta e^{L(t-\delta)}.$
 After sending $\delta\downarrow 0$, this becomes (78), as desired. ∎
 A.3. Viscosity Solutions in $CDF(\overline{\mathbb{R}})$
-Recall that throughout this work, the functions of interest belong to the space $CDF(\overline{\mathbb{R}})$ of CDF’s of extended-real-valued random variables.
+Recall that throughout this work, the functions of interest belong to the space $CDF(\overline{\mathbb{R}})$ of CDF’s of extended real-valued random variables.
 (Recall that this is nothing other than the space of nondecreasing, right-continuous functions taking values in $[0,1]$.) This section covers some specific simplifications of viscosity solutions theory that are applicable as long as the sub- or supersolutions in question are nondecreasing in the spatial variable.
-Specifically, it shows that, in this case, there is no loss of generality assuming that the test functions involved are themselves nondecreasing in space and defined globally in space.
+Specifically, it shows that, in this case, there is no loss of generality assuming that the test functions involved are both globally defined and nondecreasing in the spatial variable.
 The first lemma concerns the case when a subsolution is touched above by a test function that is increasing close to the touching point.
@@
 Fix $T,A>0$ and let $F:\mathbb{R}\times(0,T)\to[0,A)$ be an upper semicontinuous function that is nondecreasing in the spatial variable.
-Fix $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ and let $\varphi$ be a smooth function defined in some neighborhood of $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ such that $\varphi(x_{0},t_{0})<A$ and $\partial_{x}\varphi(x_{0},t_{0})>0$.
-If $\varphi$ touches $F$ from above at $(x_{0},t_{0})$, then there is an $r>0$, a nondecreasing function $\psi_{1}:\mathbb{R}\to[0,1]$ with $\psi_{1}(x)=\psi_{1}(x_{0}-r)$ for each $x\leq x_{0}-r$ and $\psi(x)=A$ for each $x\geq x_{0}+r$, and a $C^{2}$ function $\psi_{2}:\mathbb{R}\to\mathbb{R}$ with $\psi_{2}(t_{0})=0$ such that if $\psi$ is the function
+Fix $(x_{0},t_{0})\in\mathbb{R}\times(0,+\infty)$ and let $\varphi$ be a smooth function defined in some neighborhood of $(x_{0},t_{0})$ such that $\varphi(x_{0},t_{0})<A$ and $\partial_{x}\varphi(x_{0},t_{0})>0$.
+If $\varphi$ touches $F$ from above at $(x_{0},t_{0})$, then there is an $r>0$, a nondecreasing $C^{2}$ function $\psi_{1}:\mathbb{R}\to[0,1]$ with $\psi_{1}(x)=\psi_{1}(x_{0}-r)$ for each $x\leq x_{0}-r$ and $\psi_{1}(x)=A$ for each $x\geq x_{0}+r$, and a $C^{2}$ function $\psi_{2}:\mathbb{R}\to\mathbb{R}$ with $\psi_{2}(t_{0})=0$ such that if $\psi$ is the function
 $\psi(x,t)=\psi_{1}(x)+\psi_{2}(t),$
@@
 $F(x,t)\leq\psi(x,t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r]$
-and equality holds if and only if $(x,t)=(x_{0},t_{0})$.
+and equality holds if and only if $(x,t)=(x_{0},t_{0})$. Further, for any $\delta>0$, it is possible to construct $\psi$ in such a way that
+$\psi(x_{0},t_{0})=\varphi(x_{0},t_{0}),\quad\partial_{x}\psi(x_{0},t_{0})=\partial_{x}\varphi(x_{0},t_{0}),\quad\partial_{t}\varphi(x_{0},t_{0})=\partial_{t}\varphi(x_{0},t_{0}),$
+$\left|\partial_{x}^{2}\psi(x_{0},t_{0})-\partial_{x}^{2}\varphi(x_{0},t_{0})\right|\leq\delta.$
 Proof.
-By Taylor expansion, there is an $r>0$ and degree two polynomials $R$ and $Q$ such that $Q(0)=0$ and
-$F(x,t)\leq\varphi(x,t)\leq R(x)+Q(t-t_{0})\quad\text{for each}\,\,(x,t)\in[x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r]$
-Adding a term of the form $(x-x_{0})^{2}+(t-t_{0})^{2}$ to $R+Q$, there is no loss of generality in assuming in addition that $\varphi(x,t)=R(x)+Q(t-t_{0})$ if and only if $(x,t)=(x_{0},t_{0})$.
+Fix $\delta>0$.
+By Taylor’s theorem, there is an $r>0$ and second-degree polynomials $R$ and $S$ such that $\partial^{m}_{x}R(x_{0})=\partial_{x}^{m}\varphi(x_{0},t_{0})$ for $m\in\{0,1\}$, $S(0)=0$, $\partial_{t}S(0)=\partial_{t}\varphi(x_{0},t_{0})$, and
+$F(x,t)\leq\varphi(x,t)\leq R(x)+S(t-t_{0})\quad\text{for each}\,\,(x,t)\in[x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r].$
+Further, $R$ can be chosen in such a way that $|\partial_{x}^{2}R(x_{0})-\partial_{x}^{2}\varphi(x_{0},t_{0})|\leq\delta/2$.
+Adding a term of the form $\delta(x-x_{0})^{2}/4+t^{2}$ to $R+S$, there is no loss of generality in assuming in addition that $\varphi(x,t)=R(x)+S(t-t_{0})$ if and only if $(x,t)=(x_{0},t_{0})$.
 Since $0<\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}R(x_{0})$, after possibly making $r$ smaller, it can similarly be assumed that $\partial_{x}R(x)>0$ for each $x\in[x_{0}-2r,x_{0}+2r]$.
-Given $a>0$, let $\rho:\mathbb{R}\to\mathbb{R}$ be a smooth function such that $\rho(x)=0$ if $|x-x_{0}|\leq\frac{a}{2}$ and $\rho(x)=1$ if $|x-x_{0}|\geq a$. Define $\psi_{1}$ by
-$\psi_{1}(x)=(1-\rho(x))R(x)+A\rho(x)\mathbf{1}_{[x_{0},+\infty)}(x)+R(x_{0}-\frac{a}{2})\rho(x)\mathbf{1}_{(-\infty,x_{0}]}(x).$
+Given $a\in(0,r)$, let $\rho:\mathbb{R}\to\mathbb{R}$ be a smooth function such that $\rho(x)=0$ if $|x-x_{0}|\leq\frac{a}{2}$ and $\rho(x)=1$ if $|x-x_{0}|\geq a$. Define $\psi_{1}$ by
+$\psi_{1}(x)=(1-\rho(x))R(x)+A\rho(x)\mathbf{1}_{[x_{0},+\infty)}(x)+R\left(x_{0}-\frac{a}{2}\right)\rho(x)\mathbf{1}_{(-\infty,x_{0}]}(x).$
 Since $R(x_{0})=\varphi(x_{0},t_{0})<A$, $\psi_{1}$ is a nondecreasing function provided $a$ is small enough.
-Fix such a small $a$ henceforth and let $\psi_{2}(t)=Q(t-t_{0})$.
+Fix such a small $a$ henceforth and let $\psi_{2}(t)=S(t-t_{0})$.
 Since $F$ is nondecreasing in the spatial variable and $F(x,t)\leq\varphi(x,t)$ for each $(x,t)\in[x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r]$, the function $\psi(x,t)=\psi_{1}(x)+\psi_{2}(t)$ has the desired properties.
@@
 Fix $T,A>0$ and let $F:\mathbb{R}\times(0,T)\to[0,A)$ be an upper semicontinuous function that is nondecreasing in the spatial variable.
-Assume that $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ and $\varphi$ is a smooth function defined in some neighborhood of $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ such that $\varphi(x_{0},t_{0})<A$ and $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$.
+Assume that $(x_{0},t_{0})\in\mathbb{R}\times(0,+\infty)$ and $\varphi$ is a smooth function defined in some neighborhood of $(x_{0},t_{0})$ such that $\varphi(x_{0},t_{0})<A$ and $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$.
 If $\varphi$ touches $F$ from above at $(x_{0},t_{0})$, then there is an $r>0$, a nondecreasing $C^{2}$ function $\psi_{1}:\mathbb{R}\to[0,1]$ with $\psi(x)=\psi(x_{0})$ for each $x\leq x_{0}$ and $\psi(x)=A$ for each $x\geq x_{0}+r$, and a $C^{2}$ function $\psi_{2}:[t_{0}-r,t_{0}+r]\to\mathbb{R}$ with $\psi_{2}(t_{0})=0$ such that if $\psi$ is the function
 $\psi(x,t)=\psi_{1}(x)+\psi_{2}(t),$
-then
+then $\psi(x_{0},t_{0})=\varphi(x_{0},t_{0})$, $\partial_{t}\psi(x_{0},t_{0})=\partial_{t}\varphi(x_{0},t_{0})$,
 $F(x,t)\leq\psi(x,t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r],\quad\text{and}$
@@
 Expanding $\varphi$ in a fourth-order Taylor expansion around $(x_{0},t_{0})$, one finds
-$\varphi(x,t)$ $=\varphi(x_{0},t_{0})+Q(t-t_{0})+\frac{1}{3!}\partial^{3}_{x}\varphi(x_{0},t_{0})(x-x_{0})^{3}$
-$\quad+\frac{1}{4!}\partial^{4}_{x}\varphi(x_{0},t_{0})(x-x_{0})^{4}+o(|x-x_{0}|^{4}+|t-t_{0}|^{4}),$
-where $Q$ is some polynomial of degree four satisfying $Q(0)=0$.
+$\varphi(x,t)$ $=\varphi(x_{0},t_{0})+S(t-t_{0})+\frac{1}{3!}\partial^{3}_{x}\varphi(x_{0},t_{0})(x-x_{0})^{3}$
+$\qquad+\frac{1}{4!}\partial^{4}_{x}\varphi(x_{0},t_{0})(x-x_{0})^{4}+o(|x-x_{0}|^{4}+|t-t_{0}|^{4}),$
+where $S$ is some polynomial of degree four satisfying $S(0)=0$.
 Since $F$ is nondecreasing in the $x$ variable; $F(x_{0},t_{0})=\varphi(x_{0},t_{0})$; and $F(x,t)\leq\varphi(x,t)$ for $(x,t)$ close enough to $(x_{0},t_{0})$, it follows that
@@
 Making $r$ smaller if necessary, there is no loss of generality in assuming that, for each $(x,t)\in[x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r]$,
-$\varphi(x,t)$ $\leq\varphi(x_{0},t_{0})+Q(t-t_{0})+\frac{(t-t_{0})^{4}}{4!}+\frac{1}{3!}\partial_{x}^{3}\varphi(x_{0},t_{0})(x-x_{0})^{3}+\frac{1}{4!}\alpha(x-x_{0})^{4}$
+$\varphi(x,t)$ $\leq\varphi(x_{0},t_{0})+S(t-t_{0})+\frac{(t-t_{0})^{4}}{4!}+\frac{1}{3!}\partial_{x}^{3}\varphi(x_{0},t_{0})(x-x_{0})^{3}+\frac{1}{4!}\alpha(x-x_{0})^{4}$
 $=:P(x,t).$
@@
 Fix such a small $a$ henceforth and let $\psi$ be the modified version of $P$ given by
-$\psi(x,t)=h(x)+Q(t-t_{0})+\frac{(t-t_{0})^{4}}{4!}.$
+$\psi(x,t)=h(x)+S(t-t_{0})+\frac{(t-t_{0})^{4}}{4!}.$
 Since $F\leq\varphi\leq P$ in $x_{0}-r,x_{0}+r]\times[t_{0}-r,t_{0}+r]$, $F$ is nondecreasing in the $x$ variable, and $F<A$ globally, it follows that $F\leq\psi$ in $\mathbb{R}\times[t_{0}-r,t_{0}+r]$ with equality only if $x\leq x_{0}$ and $t=t_{0}$.
-Therefore, the desired conclusion follows with $\psi_{1}(x)=h(x)$ and $\psi_{2}(t)=Q(t-t_{0})$.
+Therefore, the desired conclusion follows with $\psi_{1}(x)=h(x)$ and $\psi_{2}(t)=S(t-t_{0})+(t-t_{0})^{4}/4!$.
 ∎
@@
 In the proof of the theorem, there is one technicality that needs to be overcome.
-When one checks that, for instance, the limit $F$ satisfies $\partial_{t}F\leq\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)$, it is desirable to invoke the asymptotics of $\mathscr{L}_{\epsilon}\varphi$ for an arbitrary smooth test function $\varphi$ that touches $F$ from above.
+When one checks that, for instance, the limit $F$ satisfies $\partial_{t}F\leq\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)-Q(F)$, it is desirable to invoke the asymptotics of $\mathscr{L}_{\epsilon}\varphi$ for an arbitrary smooth test function $\varphi$ that touches $F$ from above.
 If $\partial_{x}\varphi>0$ at the contact point, then Lemma 2 above shows that $\varphi$ can be assumed to be nondecreasing in $\mathbb{R}$ and then $\mathscr{L}_{\epsilon}\varphi$ is well-defined.
 Otherwise, if $\partial_{x}\varphi=0$ at the contact point, of course, $\varphi$ generally will not be monotone at all, and some work is required to proceed.
-At the same time, smooth nondecreasing functions have a particular property that is suggestive here: If $\Phi$ is a smooth nondecreasing function in $\mathbb{R}$, and if $\partial_{x}\Phi(x_{0})=0$ at some point $x_{0}$, then $\partial_{x}^{2}\Phi(x_{0})=0$.
+At the same time, smooth nondecreasing functions have a particular property that is suggestive here: If $G$ is a smooth nondecreasing function in $\mathbb{R}$, and if $\partial_{x}G(x_{0})=0$ at some point $x_{0}$, then $\partial_{x}^{2}G(x_{0})=0$.
 This motivates the following lemma (which applies even if the subsolution is not monotone in the spatial variable).
@@
 Let $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be a continuous function that is nondecreasing in the second argument and $Q:\mathbb{R}\to\mathbb{R}$ be continuous.
-An upper semicontinuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ satisfies $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)\leq 0$ in the viscosity sense if and only if it has the following property: If $\varphi$ is a smooth test function that touches $F$ from above at some point $(x_{0},t_{0})$ and
+An upper semicontinuous function $F:\mathbb{R}\times(0,+\infty)\to\mathbb{R}$ satisfies $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)\leq 0$ in the viscosity sense if and only if it has the following property: If $\varphi$ is a smooth test function that touches $F$ from above at some point $(x_{0},t_{0})$ and
 $\text{either (i)}\quad\partial_{x}\varphi(x_{0},t_{0})\neq 0,\quad\text{or (ii)}\quad\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0,$
@@
 A corresponding version of Lemma 4 holds for supersolutions, as can be readily checked using the transformation G(x,t)=1F(x,t)G(x,t)=1-F(-x,t).
-Now Lemma 4 can be leveraged in conjunction with Lemma 3 above to resolve the issue when $\partial_{x}\varphi=0$ at the contact point: There is no loss of generality assuming that $\partial_{x}^{2}\varphi=0$ also holds there, and then the lemma shows that $\varphi$ can be taken to be globally nondecreasing in the spatial variable.
+Now Lemma 4 can be leveraged in conjunction with Lemma 3 above to resolve the issue when $\partial_{x}\varphi=0$ at the contact point: There is no loss of generality assuming that $\partial_{x}^{2}\varphi=0$ also holds there, and then Lemma 3 shows that $\varphi$ can be taken to be globally nondecreasing in the spatial variable.
 B.1. Proof of Theorem 3
-It is technically convenient to proceed by proving the following statement: For each $(x,t)\in\mathbb{R}\times[0,\infty)$,
-(81) $\lim_{\delta\downarrow 0}\sup\left\{|F^{(N)}_{n}(y)-F(x,t)|\,\mid\,|\frac{y}{\delta^{(N)}}-x|+|N^{-1}n-t|+N^{-1}\leq\delta\right\}=0.$
+It is technically convenient to proceed by proving the following statement: For each $(x,t)\in\mathbb{R}\times[0,+\infty)$,
+(81) $\lim_{\delta\downarrow 0}\sup\left\{|F^{(N)}_{n}(y)-F(x,t)|\,\mid\,|\delta^{(N)}y-x|+|N^{-1}n-t|+N^{-1}\leq\delta\right\}=0.$
 Notice that this implies that $F_{N}\to F$ locally uniformly in $\mathbb{R}\times[0,+\infty)$ as $N\to+\infty$.
@@
 This will be assumed for the remainder of the proof.
-Throughout what follows, define functions $c_{-},c_{+}:[0,\infty)\to[0,1]$ by
-(82) $\frac{dc_{\pm}}{dt}=-Q(c_{\pm}),\quad c_{+}(0)=F_{\text{in}}(-\infty),\quad c_{+}(0)=F_{\text{in}}(+\infty).$
-Assumption (vi) ensures that these ODE’s are wellposed. For each $N$, let $\{c_{-}^{(N)}(n)\}$ and $\{c_{+}^{(N)}(n)\}$ be the sequences of numbers in $[0,1]$ defined analogously by
+Throughout what follows, define functions $c_{-},c_{+}:[0,+\infty)\to\mathbb{R}$ by
+(82) $\frac{dc_{\pm}}{dt}=-Q(c_{\pm}),\quad c_{-}(0)=F_{\text{in}}(-\infty),\quad c_{+}(0)=F_{\text{in}}(+\infty).$
+Assumption (vi) ensures that these ODE’s are wellposed.
+As discussed in Section 5.1, for any $N$, the function $q\mapsto q-Q^{(N)}(q)$ maps $[0,1]$ into itself.
+Taken together with assumption (vi), this implies that $Q(0)\leq 0$ and $Q(1)\geq 0$ for each $N$, and, thus,
+$0\leq c_{+}(t),c_{-}(t)\leq 1\quad\text{for each}\,\,t\geq 0.$
+For each $N$, let $\{c_{-}^{(N)}(n)\}$ and $\{c_{+}^{(N)}(n)\}$ be the sequences of numbers in $[0,1]$ defined analogously by
 $c_{\pm}^{(N)}(n)=T^{(N)}c_{\pm}(n-1)=c_{\pm}^{(N)}(n-1)-Q^{(N)}(c_{\pm}^{(N)}(n-1)),$
@@
 (Assumptions (i) and (iv) imply that $T^{(N)}$ maps constant functions to constant functions, and $\mathcal{L}^{(N)}$ vanishes on constants, as explained in the discussion in Section 5.1.) By assumption (vi) and standard ODE arguments, for any $T>0$,
-(83) $\lim_{\delta\downarrow 0}\sup\left\{|c_{\pm}^{(N)}(n)-c_{\pm}(t)|\,\mid\,|N^{-1}n-t|+N^{-1}\leq\delta,\,\,t\in[0,T]\right\}=0.$
+(83) $\lim_{\delta\downarrow 0}\sup\left\{|c_{\pm}^{(N)}(n)-c_{\pm}(t)|\,\mid\,0\leq t\leq T,\,\,|N^{-1}n-t|+N^{-1}\leq\delta\right\}=0.$
 Further, by assumption (iv), for any $N$,
 (84) $c_{-}^{(N)}(n)=F^{(N)}_{n}(-\infty)\leq F^{(N)}_{n}\leq F^{(N)}_{n}(+\infty)=c_{+}^{(N)}(n)\quad\text{for each}\,\,n\geq 0.$
-Step 1: Lipschitz Estimate in Time. Since $F_{\text{in}}$ is currently assumed to have bounded, uniformly continuous first and second derivatives, by the consistency condition (v),
-$N\mathscr{L}^{(N)}F_{0}^{(N)}=N\mathscr{L}^{(N)}F_{\text{in}}=\mathcal{F}(\partial_{x}F_{\text{in}},\partial^{2}_{x}F_{\text{in}})+R^{(N)}$
+Step 1: Lipschitz Estimate in Time.
+Since $F_{\text{in}}$ is currently assumed to have bounded, uniformly continuous first and second derivatives, by the consistency condition (v), for each $y\in\mathbb{R}$,
+$\mathscr{L}^{(N)}F_{0}^{(N)}(y)=\mathscr{L}^{(N)}F_{\text{in}}(y)=N^{-1}\mathcal{F}(\partial_{x}F_{\text{in}}(\delta^{(N)}y),\partial^{2}_{x}F_{\text{in}}(\delta^{(N)}y))+N^{-1}R^{(N)}(\delta^{(N)}y).$
 where $R^{(N)}\to 0$ uniformly as $N\to+\infty$.
@@
 $\|F^{(N)}_{n+1}-F_{n}^{(N)}\|_{\text{sup}}=\|T^{(N)}F^{(N)}_{n}-T^{(N)}F_{n-1}^{(N)}\|_{\text{sup}}\leq C_{0}N^{-1}(1+N^{-1}L)^{n}\quad\text{for each}\,\,n\in\mathbb{N}.$
-Summing this estimate in $n$, one obtains a constant $C_{0}^{\prime\prime}>0$ such that that $|F^{(N)}_{n}(x)-F_{\text{in}}(x)|\leq C_{0}^{\prime\prime}N^{-1}(1+N^{-1}L)^{n}$ uniformly in $(x,n)$ as $N\to+\infty$.
-In view of the definition of $F_{N}$, this implies that, for each $T>0$,
-$\sup\left\{|F_{N}(x,t)-F_{\text{in}}(x)|\,\mid\,(x,t)\in\mathbb{R}\times[0,T]\right\}\leq C_{0}^{\prime\prime}N^{-1}(1+LN^{-1})^{[NT]+1}.$
+Summing this estimate in $n$, one obtains a constant $C_{0}^{\prime\prime}>0$ such that that $|F^{(N)}_{n}(x)-F_{\text{in}}(x)|\leq C_{0}^{\prime\prime}L^{-1}\{(1+N^{-1}L)^{n}-1\}$ uniformly in $(x,n)$.
+In view of the definition of $F_{N}$, this implies that, for each $T>0$ and any $N\in\mathbb{N}$,
+$\sup\left\{|F_{N}(x,t)-F_{\text{in}}(x)|\,\mid\,(x,t)\in\mathbb{R}\times[0,T]\right\}\leq C_{0}^{\prime\prime}L^{-1}(e^{LT}-1).$
 Step 2: Convergence. The remainder of the proof uses the classical half-relaxed limit method. Let $F_{\star}$ and $F^{\star}$ be the half-relaxed limits of $\{F^{(N)}_{n}\}$ defined by
@@
 This will follow from the comparison principle, Theorem 4, as soon as it is established that $F^{\star}$ and $F_{\star}$ define respectively a viscosity sub- and a viscosity supersolution of the initial value problem (61).
-Step 3: $F^{\star}$ is a Subsolution.
-In view of what was proved in Step 1 and the uniform continuity of $F_{\text{in}}$, it is clear that $F^{\star}$ coincides with $F_{\text{in}}$ in the sense of (65).
-Therefore, to see that $F^{\star}$ is a viscosity subsolution of (61), it only remains to establish that $\partial_{t}F^{\star}-\mathcal{F}(\partial_{x}F^{\star},\partial_{x}^{2}F^{\star})+Q(F^{\star})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$.
-In order to avoid boundary issues that can arise if $F^{\star}$ is somewhere equal to one, it will be convenient to pull $F^{\star}$ down: To see that $F^{\star}$ is a subsolution, it suffices to prove that, for any $\delta\in(0,1)$, the function $G^{\star}:\mathbb{R}\times[0,\infty)\to[0,1-\delta]$ defined by $G^{\star}(x,t)=\max\{F^{\star}(x,t)-\delta e^{Lt},0\}$ satisfies
+Step 3: $F^{\star}$ is a Subsolution. In view of what was proved in Step 1 and the uniform continuity of $F_{\text{in}}$, it is clear that
+$\lim_{\delta\downarrow 0}\sup\left\{|F^{\star}(x,0)-F_{\text{in}}(y)|\,\mid\,x,y\in\mathbb{R}\,\,\text{such that}\,\,|x-y|\leq\delta\right\}=0,$
+so $F^{\star}$ satisfies (65).
+Therefore, to see that $F^{\star}$ is a viscosity subsolution of (61), it only remains to establish that $\partial_{t}F^{\star}-\mathcal{F}(\partial_{x}F^{\star},\partial_{x}^{2}F^{\star})+Q(F^{\star})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$.
+In order to avoid boundary issues that can arise if $F^{\star}$ is somewhere equal to one, it will be convenient to pull $F^{\star}$ down: To see that $F^{\star}$ is a subsolution, it suffices to prove that, for any $\delta\in(0,1)$, the function $G^{\star}:\mathbb{R}\times[0,+\infty)\to[0,1-\delta]$ defined by $G^{\star}(x,t)=\max\{F^{\star}(x,t)-\delta e^{Lt},0\}$ satisfies
 $\partial_{t}G^{\star}-\mathcal{F}(\partial_{x}G^{\star},\partial_{x}^{2}G^{\star})+Q(G^{\star})\leq 0\quad\text{in the viscosity sense in}\,\,\mathbb{R}\times(0,+\infty).$
@@
 Thus, in passing from $F$’s to $G$’s, nothing has changed except the stricter upper bound $G^{(N)}_{n}\leq 1-\delta(1+LN^{-1})^{n}$ holds for each $n$ and the evolution equation has been weakened to an inequality in (85).
-Suppose that $\varphi$ is a smooth function defined in some open subset of space-time that touches $G^{\star}$ from above at some point $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ in its domain.
+Suppose that $\varphi$ is a smooth function defined in some open subset of space-time that touches $G^{\star}$ from above at some point $(x_{0},t_{0})\in\mathbb{R}\times(0,+\infty)$ in its domain.
 The goal is to show that
@@
 First, consider case (I).
-Since $\varphi\geq G^{\star}$ holds in a neighborhood of $(x_{0},t_{0})$, $\varphi(x_{0},t_{0})=G^{*}(x_{0},t_{0})$, and $G^{\star}(x,t)\geq\max\{c_{-}(t)-\delta e^{Lt},0\}$ globally, in this case, $(x_{0},t_{0})$ is a local minimum of the function $(x,t)\mapsto\varphi(x,t)-\max\{c_{-}(t)-\delta e^{Lt},0\}$.
+Since $\varphi\geq G^{\star}$ holds in a neighborhood of $(x_{0},t_{0})$, $\varphi(x_{0},t_{0})=G^{\star}(x_{0},t_{0})$, and $G^{\star}(x,t)\geq\max\{c_{-}(t)-\delta e^{Lt},0\}$ globally, in this case, $(x_{0},t_{0})$ is a local minimum of the function $(x,t)\mapsto\varphi(x,t)-\max\{c_{-}(t)-\delta e^{Lt},0\}$.
 In particular, $x_{0}$ is a local minimum of $x\mapsto\varphi(x,t_{0})$, hence
@@
 $\partial_{t}\varphi(x_{0},t_{0})=0\leq-Q(0)=-Q(\varphi(x_{0},t_{0}))$
-if $c_{-}(t_{0})-\delta e^{Lt_{0}}\leq 0$.
-(That $Q(0)\leq 0$ is a consequence of the fact that the function $c\mapsto c-Q(c)$ maps $[0,1]$ into itself by assumptions (i) and (iv); see the discussion in Section 5.1.) Combining the inequality $\partial_{t}\varphi(x_{0},t_{0})\leq-Q(\varphi(x_{0},t_{0}))$ with (88), one obtains the desired inequality (87).
+if $c_{-}(t_{0})-\delta e^{Lt_{0}}\leq 0$. Combining the inequality $\partial_{t}\varphi(x_{0},t_{0})\leq-Q(\varphi(x_{0},t_{0}))$ with (88), one obtains the desired inequality (87).
 Step 3, Case (II). It only remains to consider case (II). That is, from now on, assume that $\varphi(x_{0},t_{0})>\max\{c_{-}(t_{0})-\delta e^{Lt_{0}},0\}$.
@@
 To avoid repetition, only case (b) will be considered in what follows.
-Therefore, assume henceforth that (b) holds. According to Lemma 3 with $A=1-\frac{\delta}{2}$, there is no loss of generality in fixing an $r>0$ and assuming that
-$\varphi(x,t)=\varphi_{1}(x)+\varphi_{2}(t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r],$
-where $\varphi_{1}$ is a nondecreasing $C^{2}$ function with $\varphi_{1}(x)=\varphi_{1}(x_{0})$ for $x\leq x_{0}$ and $\varphi_{1}(x)=1-\frac{\delta}{2}$ for $x\geq x_{0}+r$; $\varphi_{2}$ is $C^{2}$ in $t_{0}-r,t_{0}+r]$ with $\varphi_{2}(t_{0})=0$; and
-$G^{\star}(x,t)\leq\varphi(x,t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r],$
-$(x_{0},t_{0})\in\{(x,t)\,\mid\,G^{\star}(x,t)=\varphi(x,t)\}\subseteq(-\infty,x_{0}]\times\{t_{0}\}.$
-Since $\varphi_{2}$ is continuous in $t_{0}-r,t_{0}+r]$ and $\varphi_{1}(-\infty)=\varphi(x_{0},t_{0})>\max\{c_{-}(t_{0})-\delta e^{Lt_{0}},0\}$, up to shrinking $r$ if necessary, there is no loss of generality in assuming that there are constants $c_{*},c^{*}\in[0,1]$ such that, for each $t\in[t_{0}-r,t_{0}+r]$,
-(89) $\max\{c_{-}(t)-\delta e^{Lt},0\}<c_{*}\leq\varphi(-\infty,t)\quad\text{and}\quad 1-\delta<c_{*}\leq\varphi(+\infty,t)\leq 1-\frac{\delta}{4}.$
+Therefore, assume henceforth that (b) holds.
+According to Lemma 3 with $A=1-\frac{\delta}{2}$, it is possible to fix an $r>0$ and a function $\psi:\mathbb{R}\times[t_{0}-r,t_{0}+r]\to\mathbb{R}$ of the form
+$\psi(x,t)=\psi_{1}(x)+\psi_{2}(t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r],$
+where $\psi_{1}$ is a nondecreasing $C^{2}$ function with $\psi_{1}(x)=\psi_{1}(x_{0})$ for $x\leq x_{0}$ and $\psi_{1}(x)=1-\frac{\delta}{2}$ for $x\geq x_{0}+r$; $\psi_{2}$ is $C^{2}$ in $t_{0}-r,t_{0}+r]$ with $\psi_{2}(t_{0})=0$; and
+$\psi(x_{0},t_{0})=\varphi(x_{0},t_{0}),\quad\partial_{t}\psi(x_{0},t_{0})=\partial_{t}\varphi(x_{0},t_{0}),$
+$G^{\star}(x,t)\leq\psi(x,t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r],$
+$(x_{0},t_{0})\in\{(x,t)\,\mid\,G^{\star}(x,t)=\psi(x,t)\}\subseteq(-\infty,x_{0}]\times\{t_{0}\}.$
+Since $\psi_{2}$ is continuous in $t_{0}-r,t_{0}+r]$ and $\psi_{1}(-\infty)=\varphi(x_{0},t_{0})>\max\{c_{-}(t_{0})-\delta e^{Lt_{0}},0\}$, up to shrinking $r$ if necessary, there is no loss of generality in assuming that there are constants $c_{*},c^{*}\in(0,1)$ such that, for each $t\in[t_{0}-r,t_{0}+r]$,
+(89) $\max\{c_{-}(t)-\delta e^{Lt},0\}<c_{*}\leq\psi(-\infty,t)\quad\text{and}\quad 1-\delta<c^{*}\leq\psi(+\infty,t)\leq 1-\frac{\delta}{4}.$
 Finally, for each $N>0$, let $\nu(N)$ be the constant
-$\nu(N)=\sup\left\{G^{(N)}_{n}(y)-\varphi(\delta^{(N)}y,N^{-1}n)\,\mid\,y\in\mathbb{R},\,\,N^{-1}n\in[t_{0}-r,t_{0}+r]\right\}.$
-Notice that (89), (84), and (83) imply that there is an $R>0$ such that, for any $N$,
-$\sup\left\{G^{(N)}_{n}(y)-\varphi(\delta^{(N)}y,N^{-1}n)\,\mid\,\delta^{(N)}|y|\geq R,\,\,N^{-1}n\in[t_{0}-r,t_{0}+r]\right\}<0.$
-Therefore, by the equation (86) relating $\{G^{(N)}_{n}\}$ to $G^{\star}$, it follows that there is a sequence $\{(N_{k},y_{k},n_{k})\}$ such that $(N^{-1}_{k},\nu(N_{k}))\to(0,0)$ as $k\to\infty$; $\delta^{(N)}|y_{k}|\leq R$ and $|N^{-1}n_{k}-t_{0}|\leq r$ for each $k$; and
-(90) $G^{(N_{k})}_{n_{k}}(y_{k})-\varphi(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k})=\nu(N_{k})\quad\text{for each}\,\,k.$
-Further, since the set $\{(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r]\,\mid\,G^{\star}(x,t)=\varphi(x,t)\}$ is contained in $(-\infty,x_{0}]\times\{t_{0}\}$, there is no loss of generality in assuming that there is an $x^{\prime}_{0}\leq x_{0}$ such that $(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k})\to(x^{\prime}_{0},t_{0})$ and $G^{\star}(x^{\prime}_{0},t_{0})=\varphi(x^{\prime}_{0},t_{0})$.
-Note, in particular, that the inequality $x^{\prime}_{0}\leq x_{0}$ implies, by the constancy of $\varphi_{1}$ in $(-\infty,x_{0}]$,
-(91) $\lim_{k\to\infty}(\varphi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}\varphi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}^{2}\varphi_{1}(\delta^{(N_{k})}y_{k}))=(\varphi_{1}(x_{0}),0,0).$
-Finally, for notational ease, let $\Phi_{k}:\mathbb{R}\times[t_{0}-r,t_{0}+r]\to\mathbb{R}$ be the function $\Phi_{k}(y,t)=\varphi(\delta^{(N_{k})}y,t)+\nu(N_{k})$.
-Since $\nu(N_{k})\to 0$, the bounds (89) imply that $0\leq\Phi_{k}\leq 1$ pointwise for all $k$ large enough.
-That is, $\Phi_{k}\in CDF(\overline{\mathbb{R}})$ for large enough $k$.
+$\nu(N)=\sup\left\{G^{(N)}_{n}(y)-\psi(\delta^{(N)}y,N^{-1}n)\,\mid\,y\in\mathbb{R},\,\,N^{-1}n\in[t_{0}-r,t_{0}+r]\right\}.$
+Notice that (89), (84), and (83) imply that there is an $R>0$ such that, for any large enough $N$,
+$\sup\left\{G^{(N)}_{n}(y)-\psi(\delta^{(N)}y,N^{-1}n)\,\mid\,\delta^{(N)}|y|\geq R,\,\,N^{-1}n\in[t_{0}-r,t_{0}+r]\right\}<0.$
+Therefore, by the equation (86) relating $\{G^{(N)}_{n}\}$ to $G^{\star}$, it follows that there is a sequence $\{(N_{k},y_{k},n_{k})\}$ such that $(N^{-1}_{k},\nu(N_{k}))\to(0,0)$ as $k\to\infty$; $\delta^{(N_{k})}|y_{k}|\leq R$ and $|N_{k}^{-1}n_{k}-t_{0}|\leq r$ for each $k$; and
+(90) $G^{(N_{k})}_{n_{k}}(y_{k})-\psi(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k})=\nu(N_{k})\quad\text{for each}\,\,k.$
+Further, since the set $\{(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r]\,\mid\,G^{\star}(x,t)=\psi(x,t)\}$ is contained in $(-\infty,x_{0}]\times\{t_{0}\}$, there is no loss of generality in assuming that there is an $x^{\prime}_{0}\leq x_{0}$ such that $(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k})\to(x^{\prime}_{0},t_{0})$ and $G^{\star}(x^{\prime}_{0},t_{0})=\psi(x^{\prime}_{0},t_{0})$.
+Note, in particular, that the inequality $x^{\prime}_{0}\leq x_{0}$ implies, by the constancy of $\psi_{1}$ in $(-\infty,x_{0}]$,
+(91) $\lim_{k\to\infty}(\psi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}\psi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}^{2}\psi_{1}(\delta^{(N_{k})}y_{k}))=(\psi_{1}(x_{0}),0,0).$
+Finally, for notational ease, let $\Psi_{k}:\mathbb{R}\times[t_{0}-r,t_{0}+r]\to\mathbb{R}$ be the function $\Psi_{k}(y,t)=\psi(\delta^{(N_{k})}y,t)+\nu(N_{k})$.
+Since $\nu(N_{k})\to 0$, the bounds (89) imply that $0\leq\Psi_{k}\leq 1$ pointwise for all $k$ large enough.
+That is, $\Psi_{k}\in CDF(\overline{\mathbb{R}})$ for large enough $k$.
 By the choice of $\nu(N_{k})$,
-$G^{(N_{k})}_{n}\leq\Phi_{k}(\cdot,N_{k}^{-1}n)\quad\text{pointwise for each}\,\,n\in[N(t_{0}-r),N(t_{0}+r)].$
+$G^{(N_{k})}_{n}\leq\Psi_{k}(\cdot,N_{k}^{-1}n)\quad\text{pointwise for each integer}\,\,n\in[N(t_{0}-r),N(t_{0}+r)].$
 Therefore, by monotonicity (assumption (i)),
-$T^{(N_{k})}G^{(N_{k})}_{n_{k}-1}\leq T^{(N_{k})}\Phi_{k}(\cdot,N_{k}^{-1}(n_{k}-1)).$
-By (85), (90), and the definition of $\Phi_{k}$, this implies
-$\Phi_{k}(y_{k},N_{k}^{-1}n_{k})=G^{(N_{k})}_{n_{k}}(y_{k})\leq T^{(N_{k})}G^{(N_{k})}_{n_{k}-1}(y_{k})\leq T^{(N_{k})}\Phi_{k}(y_{k},N_{k}^{-1}n_{k}-N_{k}^{-1}).$
-Upon rearranging, this becomes
-$N_{k}\Big(\Phi_{k}(y_{k},N_{k}^{-1}n_{k})-\Phi_{k}(y_{k},N_{k}^{-1}n_{k}-N_{k}^{-1})\Big)$
-$\quad\quad\quad\leq N_{k}(\mathscr{L}^{(N_{k})}\Phi_{k})(y_{k},N_{k}^{-1}(n_{k}-1))+N_{k}Q^{(N_{k})}(\Phi_{k}(y_{k},N_{k}^{-1}(n_{k}-1)).$
-By assumption (iii), the definition of $\Phi_{k}$, and the form of $\varphi$,
-$(\mathscr{L}^{(N_{k})}\Phi_{k})(y_{k},N_{k}^{-1}(n_{k}-1))=(\mathscr{L}^{(N_{k})}(\varphi_{1})_{\delta^{(N_{k})}})(y_{k}),$
-hence, by assumptions (v) and (vi), (91), and the form of $\varphi$, sending $k\to\infty$ above yields
-$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(0,0)+Q(\varphi(x_{0},t_{0})))$ $=\partial_{t}\varphi_{2}(t_{0})-\lim_{k\to\infty}\mathcal{F}(\partial_{x}\varphi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}^{2}\varphi_{1}(\delta^{(N_{k})}y_{k}))$
-$\quad+\lim_{k\to\infty}N_{k}Q^{(N_{k})}(\varphi(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k}))$
+$T^{(N_{k})}G^{(N_{k})}_{n_{k}-1}\leq T^{(N_{k})}\Psi_{k}(\cdot,N_{k}^{-1}(n_{k}-1)).$
+By (85), (90), and the definition of $\Psi_{k}$, this implies
+$\Psi_{k}(y_{k},N_{k}^{-1}n_{k})=G^{(N_{k})}_{n_{k}}(y_{k})\leq T^{(N_{k})}G^{(N_{k})}_{n_{k}-1}(y_{k})\leq T^{(N_{k})}\Psi_{k}(y_{k},N_{k}^{-1}n_{k}-N_{k}^{-1}).$
+After rearranging terms, this becomes
+$N_{k}\Big(\Psi_{k}(y_{k},N_{k}^{-1}n_{k})-\Psi_{k}(y_{k},N_{k}^{-1}n_{k}-N_{k}^{-1})\Big)$
+$\quad\quad\quad\leq N_{k}(\mathscr{L}^{(N_{k})}\Psi_{k})(y_{k},N_{k}^{-1}(n_{k}-1))-N_{k}Q^{(N_{k})}(\Psi_{k}(y_{k},N_{k}^{-1}(n_{k}-1)).$
+By assumption (iii), the definition of $\Psi_{k}$, and the form of $\psi$,
+$(\mathscr{L}^{(N_{k})}\Psi_{k})(y_{k},N_{k}^{-1}(n_{k}-1))=(\mathscr{L}^{(N_{k})}(\psi_{1})_{\delta^{(N_{k})}})(y_{k}),$
+hence, by assumptions (v) and (vi) and (91), sending $k\to\infty$ above yields
+$\partial_{t}\psi(x_{0},t_{0})-\mathcal{F}(0,0)+Q(\psi(x_{0},t_{0})))$ $=\partial_{t}\psi_{2}(t_{0})-\lim_{k\to\infty}\mathcal{F}(\partial_{x}\psi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}^{2}\psi_{1}(\delta^{(N_{k})}y_{k}))$
+$\quad+\lim_{k\to\infty}N_{k}Q^{(N_{k})}(\psi(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k}))$
 $\leq 0.$
-Since $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$ by assumption, this yields (87) as desired.
+Since $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$ by assumption and $\partial^{m}_{t}\psi(x_{0},t_{0})=\partial_{t}^{m}\varphi(x_{0},t_{0})$ for $m\in\{0,1\}$, this yields (87) as desired.
 Step 4: $F_{\star}$ is a Supersolution.
@@
 B.2. Proof of Corollary 4
-In the proof that follows, it it convenient to work with the half-relaxed limits $\liminf_{*}F_{N}$ and $\limsup_{*}F_{N}$ of the rescaled CDF $F_{N}$ defined by
+In the proof that follows, it it convenient to work with the half-relaxed limits $\liminf_{*}F_{N}$ and $\limsup^{*}F_{N}$ of the rescaled CDF $F_{N}$ defined by
 $(\liminf\nolimits_{*}F_{N})(x,t)$ $=\lim_{\delta\downarrow 0}\inf\left\{F_{N}(x^{\prime},s)\,\mid\,|x^{\prime}-x|+|s-t|+N^{-1}\leq\delta\right\},$
-$(\limsup\nolimits_{*}F_{N})(x,t)$ $=\lim_{\delta\downarrow 0}\sup\left\{F_{\epsilon}(x^{\prime},s)\,\mid\,|x^{\prime}-x|+|s-t|+N^{-1}\leq\delta\right\}.$
+$(\limsup\nolimits^{*}F_{N})(x,t)$ $=\lim_{\delta\downarrow 0}\sup\left\{F_{\epsilon}(x^{\prime},s)\,\mid\,|x^{\prime}-x|+|s-t|+N^{-1}\leq\delta\right\}.$
 Proof of Corollary 4.
@@
 In view of Proposition 13 and well-known properties of half-relaxed limits (see [15, Remark 6.4] or [10, Lemma 6.2]), to see that $F_{N}$ converges locally uniformly in $\mathbb{R}\times(0,+\infty)$ to the discontinuous viscosity solution as $N\to+\infty$, it suffices to establish that
-$\underline{F}\leq\liminf\nolimits_{*}F_{\epsilon}\leq\limsup\nolimits^{*}F_{\epsilon}\leq\overline{F}\quad\text{in}\,\,\mathbb{R}\times[0,\infty).$
+$\underline{F}\leq\liminf\nolimits_{*}F_{N}\leq\limsup\nolimits^{*}F_{N}\leq\overline{F}\quad\text{in}\,\,\mathbb{R}\times[0,+\infty).$
 The details will be provided only for the lower bound as the upper bound follows similarly.
@@
 $F_{N,M}(x,0)=2^{M}\int_{0}^{\infty}F_{N}(x-y,0)\rho(2^{M}y)\,dy,$
-where $\rho:(0,\infty)\to(0,\infty)$ is a smooth function supported in $[1,2]$ with $\int_{0}^{\infty}\rho(y)\,dy=1$.
+where $\rho:(0,+\infty)\to(0,+\infty)$ is a smooth function supported in $[1,2]$ with $\int_{0}^{\infty}\rho(y)\,dy=1$.
 From the fact that $\rho$ is supported in $[1,2]$ and $F_{N}$ is nondecreasing in $x$, it follows that $F_{N,M}(x,0)\leq F_{N,M+1}(x,0)\leq F_{N}(x,0)$ for every $x\in\mathbb{R}$.
@@
 By Theorem 3, $F_{N,M}\to F_{M}$ locally uniformly in $\mathbb{R}\times[0,+\infty)$ as $N\to+\infty$, where $F_{M}$ is the solution of (61) with initial datum $F_{\text{in},M}$ given by $F_{\text{in},M}(x)=2^{M}\int_{0}^{\infty}F_{\text{in}}(x-y)\rho(2^{M}y)\,dy$.
-Further, by construction, the bound $F_{M}\leq F_{M+1}\leq\liminf\nolimits_{*}F_{N}$ holds pointwise in $\mathbb{R}\times[0,\infty)$ for any fixed $M$.
+Further, by construction, the bound $F_{M}\leq F_{M+1}\leq\liminf\nolimits_{*}F_{N}$ holds pointwise in $\mathbb{R}\times[0,+\infty)$ for any fixed $M$.
 Standard arguments (see [10, Theorem 6.2]) show that the function $F_{\infty}$ defined by
 $F_{\infty}(x,t)=\lim_{\delta\downarrow 0}\inf\left\{F_{M}(x^{\prime},s)\,\mid\,|x^{\prime}-x|+|t-s|+M^{-1}\leq\delta\right\}$
-satisfies $\partial_{t}F_{\infty}-\mathcal{F}(\partial_{x}F_{\infty},\partial_{x}^{2}F_{\infty})+Q(F_{\infty})\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and $F_{\infty}(\cdot,0)\geq F_{M}(\cdot,0)$ for any $M$.
-In particular, by the comparison principle, $F_{\infty}$ is at least as large as any viscosity subsolution of the initial-value problem (61).
-Therefore, $F_{\infty}\geq\underline{F}$ in $\mathbb{R}\times[0,\infty)$, and this implies $\liminf\nolimits_{*}F_{N}\geq F_{\infty}\geq\underline{F}$ as desired.
+satisfies $\partial_{t}F_{\infty}-\mathcal{F}(\partial_{x}F_{\infty},\partial_{x}^{2}F_{\infty})+Q(F_{\infty})\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,+\infty)$ and $F_{\infty}(\cdot,0)\geq F_{M}(\cdot,0)$ for any $M$.
+In the limit $M\to+\infty$, this last bound becomes
+$F_{\infty}(x,0)\geq\lim_{\delta\downarrow 0}F_{\text{in}}(x-\delta).$
+Therefore, by the comparison principle (Theorem 4), $F_{\infty}$ is at least as large as any viscosity subsolution of the initial-value problem (61).
+Therefore, $F_{\infty}\geq\underline{F}$ in $\mathbb{R}\times[0,+\infty)$, and this implies $\liminf\nolimits_{*}F_{N}\geq F_{\infty}\geq\underline{F}$ as desired.
 ∎
@@
 In particular, $F_{n}\leq G_{n}+c_{n}$ for any $n$, as desired. ∎
-Appendix C Effective Coefficients in the Series-Parallel Graph
-This appendix computes the constants $\sigma_{D}$ and $a_{R}$ for the logarithms of the distance $D^{(n)}$ and the resistance $R^{(n)}$ of the critical series-parallel graph.
-In either case, the relevant forcing $f$ takes the form
+Appendix C Effective Coefficients in the Series-Parallel Graph Case
+This appendix computes the constants $\sigma_{\mathcal{D}}$ and $a_{R}$ in the case of the logarithm of the distance and the resistance of the series-parallel graph.
+Recall that, in this setting, the relevant forcing $f$ takes the form
 $f(u)=\log(1+e^{-u}).$
-The left-inverse $g$ is given by
+The corresponding function $g$ from Section 1.4.1 is given by
 $g(s)=\log(e^{s}-1)\quad\text{if}\,\,s\geq\log 2,\quad g(s)=0,\quad\text{otherwise.}$
-Since $g(s)+f(g(s))=s$ for $s\geq\log 2$, the formulas from Section 1.4.1 become
-$\sigma_{D}$ $=-\int_{0}^{\log 2}s\,ds-\int_{\log 2}^{\infty}\log\left(\frac{1}{1-e^{-s}}\right)\,ds,$
-$a_{R}$ $=\int_{0}^{\log 2}s^{2}\,ds+\int_{\log 2}^{\infty}\log^{2}\left(\frac{1}{1-e^{-s}}\right)\,ds$
-$\quad+2\int_{0}^{\log 2}s^{2}\,ds+2\int_{\log 2}^{\infty}s\log\left(\frac{1}{1-e^{-s}}\right)\,ds$
+Since $g(s)+f(g(s))=s$ for $s\geq\log 2$, the formulae for $\sigma_{\mathcal{D}}$ and $a_{R}$ become
+$\sigma_{\mathcal{D}}$ $=-\int_{0}^{\log 2}s\,ds-\int_{\log 2}^{\infty}\log\left(\frac{1}{1-e^{-s}}\right)\,ds,$
+$a_{R}$ $=3\int_{0}^{\log 2}s^{2}\,ds+\int_{\log 2}^{\infty}\log^{2}\left(\frac{1}{1-e^{-s}}\right)\,ds+2\int_{\log 2}^{\infty}s\log\left(\frac{1}{1-e^{-s}}\right)\,ds$
 The integrals above can be computed using polylogarithms, a family of functions related to the Riemann zeta function $\zeta$.
-The dilogarithm and trilogarithm $Li_{2}$ and $Li_{3}$ are defined for $t\in[0,1]$ by the formulas
+The dilogarithm and trilogarithm $Li_{2}$ and $Li_{3}$ are defined for $t\in[0,1]$ via the integrals
 $Li_{2}(t)=-\int_{0}^{t}\frac{1}{x}\log(1-x)\,dt,\quad Li_{3}(x)=\int_{0}^{t}\frac{1}{x}Li_{2}(x)\,dx.$
@@
 The computations that follow seem to be related to Euler’s formula for $Li_{2}(\frac{1}{2})$ and Landen’s formula for $Li_{3}(\frac{1}{2})$ (see [22, Section 1.1]).
-To compute the integrals, use the change-of-variables $x=e^{-s}$, $ds=\frac{1}{x}dx$. In the case of $\sigma_{D}$, this leads to
-$\sigma_{D}$ $=\int_{\frac{1}{2}}^{1}\log(x)\frac{1}{x}\,dx+\int_{0}^{\frac{1}{2}}\log(1-x)\frac{1}{x}\,dx$
+To compute the integrals, use the change-of-variables $x=e^{-s}$, $ds=\frac{1}{x}dx$. In the case of $\sigma_{\mathcal{D}}$, this leads to
+$\sigma_{\mathcal{D}}$ $=\int_{\frac{1}{2}}^{1}\log(x)\frac{1}{x}\,dx+\int_{0}^{\frac{1}{2}}\log(1-x)\frac{1}{x}\,dx$
 $=-\frac{1}{2}\log^{2}(\frac{1}{2})+\frac{1}{2}\int_{\frac{1}{2}}^{1}\log(x)\frac{1}{1-x}\,dx+\frac{1}{2}\int_{0}^{\frac{1}{2}}\log(1-x)\frac{1}{x}\,dx$
@@
 $=2\int_{0}^{1}\log(x)\log(1-x)\frac{1}{x}\,dx.$
-Finally, integrating-by-parts again and invoking the definitions of $Li_{2}$ and $Li_{3}$ yields
-$a_{R}$ $=2\int_{0}^{1}Li_{2}(x)\frac{1}{x}\,dx-Li_{2}(1)\log(1)=2\int_{0}^{1}Li_{2}(x)\frac{1}{x}\,dx=2Li_{3}(1)=2\zeta(3).$
+Finally, integrating by parts and invoking the definitions of $Li_{2}$ and $Li_{3}$ yields
+$a_{R}$ $=2\int_{0}^{1}Li_{2}(x)\frac{1}{x}\,dx-2Li_{2}(1)\log(1)=2\int_{0}^{1}Li_{2}(x)\frac{1}{x}\,dx=2Li_{3}(1)=2\zeta(3).$